A Complete Bibliography of ACM Transactions on Computer-Human Interaction

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 November 2024 Version 1.100

Title word cross-reference

#ActuallyAutistic [Gub23].

+ [UIJ05]. 3 [IIY^{+}24, IW03, LBP17, MWVK21, NHGC16, QB05, SGL09, WLB09].

12 [VSS^{+}23]. 12-Year-Old [RSC^{+}24]. 19 [BHH^{+}24, DPL^{+}23, MBN^{+}23].

3D [CNG^{+}24, HWSB99, ZBM96].

AAC [NWH21]. Abbreviated [LQZ23]. Abilities [VJR24, SHCP08]. abstracted [TKH11]. abstraction [PSS09]. abstraction-level [PSS09]. Abundance [ERL^{+}23]. Abuse [AHCF18]. Academia [WMRW16]. Academia-Industry [WMRW16]. Academic [NSP^{+}18, NDSG06]. Accept [GSPZ24]. Acceptance [HH17, NEE^{+}24, CT07]. Access [DV18, HCH^{+}14, PRJ16, BA04, DS98, DHMV14, LEF^{+}00, POS^{+}01, RP96, SW09, WFD98]. Accessibility [FLST13, SWZ^{+}24, HB07, PCH^{+}06, SHM07, SHMA07]. Accessible [DPG22, FJM24, SED^{+}16, SHMA07]. accomplishment [LJY^{+}13]. Account [PTR23]. accountable [SSRW13]. Accuracy [GRG18, PKES22, DSG09]. Achieving [BBD^{+}24, MD23, SJZ^{+}98]. ACM [YHS96]. Acquisition [PD16, CD11]. across [VTS^{+}04, WIKW23]. Action [GMPS17, JRB^{+}24, KGYQ15, KKH^{+}13, LW23, Dou13, Hay11, RSJ02, MZL^{+}23]. Action-Transferred [KGYQ15]. Actions [FH23, Roo13, LJY^{+}13, PK94]. Active [BFC12, Hor16, CGS12]. Activism [ABC20]. activities [CGS12]. Activity [BMNH20, Bar09, BFC12, CVZBB20, GRKB16, HFP12, KHM20, TVH16, BFAM11, LDF12, MCSN03]. Activity-based [Bar09, BFAM11]. Actuated [DMOJ18]. Ad [BAAL^{+}16, RLP14, ENSS09]. Ad-Hoc [BAAL^{+}16]. Adapt [LAW18]. adaptable [MBB07]. Adaptation [KSCB21, PCR15, SXS^{+}06, Vic00]. Adapting [MKS19]. Adaptive [BVL^{+}19, MSM23, SIK^{+}12, SJU19, GR11, MCC^{+}04, OS04, ODC04, RB11]. Adding [KTBR15a, YHWK16]. Addressing [HFK^{+}22, MLC^{+}13, SLP22, XJS23, NPLB09]. ADHD [ACS^{+}23, GDW^{+}24]. Adopting [LJPS21]. Adoption [NWH21]. Ads [BHNG05]. Adult [FBE23, KH18, RBH24, SED^{+}16, SHCP08]. Adults [LCK24, MM21, OAV^{+}16, PLF20, RBH24, Bec04, HBDG04, MJV^{+}06]. Advance [LJPS21]. Advanced [FMP19, OWOZ17]. Advancing [ICM^{+}23, SBSS12]. aesthetic [BSW08]. Aesthetics [EDT^{+}23, WWM08, BT08, BBMT06, DH08, HR02, HSD08, PHJ08a, PHJ08b, RB11]. Affect [OSBB^{+}19, YS20, LDF12]. Affecting [FBE23]. Affective [PBBJ^{+}22]. Affects [GPPD^{+}22]. affordances [OCM^{+}12]. Africa [BSM^{+}13]. African [Ada23]. Again [USK^{+}23]. Age [LF14, LW15, OTV19, VPW^{+}15, MRF09]. Age-Old [VPW^{+}15]. age-related [MRF09]. Ageing [VPW^{+}15]. Agency [BKPH22, GC22, SB18, SFKF19, TNLK22, SLS94]. Agenda [VSS^{+}23, SAT^{+}23, TM05]. Agent [ACR^{+}16, CVB16, LWLL22, RRC^{+}22, RK22, TCD^{+}21]. Agent-Assisted [RRC^{+}22]. agents [CGA06, LB10, VBHK10]. Aging [LJPS21, HC06]. Agreement [Tsa18, VW22]. AI [ANW^{+}23, BP23, BBT^{+}23, CRL^{+}23, CWM^{+}23, DWL^{+}24, EDT^{+}23, GCC^{+}24, GLX^{+}23, ICM^{+}23, LVDA23, PKES22, PTR23, RM23, THL^{+}23, WWD^{+}22, XZL^{+}20, ZLD^{+}23, vBBSB23]. AI-assisted [GCC^{+}24]. AI-Based [EDT^{+}23]. AI-Enabled [vBBSB23]. AI-Generated [DWL^{+}24]. AI-Powered [XZL^{+}20]. Aids [MSH^{+}23]. Aim [VS14]. Air [NPCBL15, Mac99, SHS24]. Alert [LHF^{+}22]. Alerts [MWS18]. Alerts-Using [MWS18]. algebra [Thi04]. Algorithm [BGK^{+}22, BBS01a, MS94]. Algorithmic [BTBM24, ICM^{+}23, KKJ24]. Algorithms [KYZ23, vBBSB23]. Alibaba [WSWL23]. allocating [FPST99]. Alloys [MMH^{+}22]. Alone [SLM^{+}20]. Aloud [FLCT19, Her24]. also [YCVV23]. Altering [KIW16]. Alternative [KB23, LH08]. Amateur [RT18]. Amazon [DKT^{+}21]. Ambient [BVR15]. Ambiguity [SHSS19]. ambiguous [MF10]. Ambulatory [PWG18]. Ameliorating [RAT^{+}22]. Among [MM21, KSJB11, RZK24, WSO16]. Amplifiers [WSO16]. Amplifying [GBH^{+}18]. Anachronistic [Men22]. Analog [PF18]. Analogical [KQH^{+}22]. Analysing [HFK^{+}22]. Analysis [CFH^{+}20, DTP^{+}23, ERT23, GCC^{+}24, HMKV19, KPO18, LA17, TSFA07, TLY^{+}24, VML15, VW22, WWH19, XGA^{+}22, ZSC^{+}15, BG05, GSM99, HK99, JK96a, MKP05, MMHM23, NM09, PGG03, MD23, TCJ01]. Analytic [Her24]. Analytics [MMEFN^{+}24, SLY^{+}18]. Analyzing [PVSZ24, TLA^{+}19, WM06]. Anatomy [WBW^{+}23, LST08]. animated [ODC04, Wol97]. animation [Dee95, KGZ07, MS94, TC01]. ANISMA [MMH^{+}22]. annexation [AO11]. Annotation [MBHC17]. annotations [SGL09]. annoying [MRC^{+}22]. Anonymity [AHCF18]. Anticipating [SLBB19]. Anxiety [LRS19, MFPB24]. any [LEF^{+}00]. anytime [POS^{+}01]. anywhere [LEF^{+}00, POS^{+}01]. appeal [LDS^{+}11]. Appearing [ANW^{+}23]. Appendices [BYS02a]. application [BI08, BRS99, JN96, VM95]. application-sharing [BRS99]. Applications [AJSW12, BBT^{+}23, GMPS17, Har19, HGOZ19, MMH^{+}22, THL^{+}23, HBR^{+}94, BPR08, DM05, Dou98, Edw05, LMG^{+}11, LH08, PSS09, SXS^{+}06, TSFA07]. Applied [DS08]. Apply [KTN^{+}18, TH15]. Applying [ERT23, Pil18, PMM^{+}13, PJL^{+}16, TC01]. Appointments [HHG^{+}22]. Appraisal [BKJ18]. Approach [ALDR^{+}22, BMNH20, BPG^{+}22, BGK^{+}22, CDT^{+}21, GSK22, HM23, HSA^{+}23, KGYQ15, XJS23, CL08, CC13, EK00, GB05, LR13, MTDM14, PKHD09, RO97, Wex95]. Approachable [AWB^{+}23]. approaches [DRW13, PSSB13]. Approaching [Tan07]. Appropriate [HJS^{+}20]. Appropriateness [AATC22]. Appropriation [LBP17]. Apps [LWLL22, PAM^{+}22, SWZ^{+}24]. Archaeological [ACDL12]. Architects [Kir19]. Architecture [ACW^{+}19, SIK^{+}12, SJU19, HBR^{+}94, DELS99, KWM97, LFT06, SLS94, TNB^{+}95]. archives [WH01]. archiving [KS10]. Arguments [FMSS17]. arithmetic [BBS01a]. Around-the-Head [KRMS21]. arrangement [Coh97]. arrangements [SM11b]. Article [JRB^{+}24]. Artifacts [FH23]. Artificial [CWZL23, KWCS22, PPAA^{+}18, ZCX^{+}24]. Artisans [ZSNP14]. Artistic [AZS16]. Aspects [GC22, KSJB11]. Assessing [LVDA23, MCSN03, BA04]. Assessment [LRP15, RLP14, Tsa18, KWL^{+}13]. Assessments [GFK^{+}17]. Assist [WWD^{+}22]. Assistance [XRL^{+}22]. Assistant [CEH^{+}22, CGL24]. Assistants [AKTB19, PLF20]. Assisted [JKS18, KNK^{+}21, PWG18, RRC^{+}22, VS14, GCC^{+}24]. Assistive [ICC^{+}22, RCFR22].

Associated [ANK^{+}23]. Association

[XPL23]. Assumption [BMB^{+}13, OSPK23]. Asylum [AV19]. asymmetric [RSJ02]. At-a-Distance [MWB16]. Athletes [RT18]. Attention [KBB^{+}17, NSCR19, VAF17, MRF09]. Attitudes [NKDB23, RC22, WM15]. attitudinal [ZSE^{+}12]. Audience [HCS^{+}23, RSR21]. Audio [BVR15, TCD^{+}21, VAAB^{+}16, ASHM97, RV95, SS00]. Audio-Augmented [VAAB^{+}16]. audio-only [ASHM97]. Audio-Visual [TCD^{+}21]. Auditory [VAAB^{+}16, VAF17]. Augmentation [MBP^{+}22]. Augmentations [BHA18, HLSR24, SLP22]. Augmented [DVK18, KSB^{+}24, LCC23, PSD24, UK21, VAAB^{+}16, ZRT^{+}20, BBS01b]. Augmenting [BJC^{+}23, CZH^{+}15, KQH^{+}22, MBP^{+}22]. Authentication [MRC^{+}22, MWVK21]. Author [Ano96, Ano97]. Authors [DWL^{+}24]. Autism [RBH24]. Autistic [PPAA^{+}18, RBH24, RC22, SRE^{+}21, SFKF19]. Autocompletion [KB23]. Automata [DKA^{+}15]. Automated [Har19, SWZ^{+}24, VML15, BA04]. Automatic [CM03, DKA^{+}15, LA15]. Automatically [KKH^{+}13]. Automatics [LAW18]. Automation [BWR^{+}17, NKDB23, VP24]. Automoderator [JBGB19]. Autonomous [BM24]. availability [WW05]. Available [HGOZ19, TLA^{+}19]. Avatar [AATC22]. avatars [QB05]. Avoidance [YS20]. Award [Hin17b]. aware [CFLC23, DM05, PKHD09]. Awareness [BVR15, BHA18, MZL^{+}23, NVPE19, SP21, WM15, CMS^{+}11, GG99, MMS^{+}08, NBG09].

back [OHM^{+}13]. background

[HPHS05, HWSB99]. Backtracking [AJSW12]. Bad [TVH16]. Balancing [RBB15]. BAN [SYK23]. banner [BHNG05]. Barriers [DV18, JNR^{+}24, HFB09]. Based [BKJ18, BSR^{+}23, DS08, DVK18, EDT^{+}23, FSMP22, GFK^{+}17, GBBM12, GLZH20, GSS^{+}19, HGH^{+}24, HHG^{+}22, HLEG18, HFP12, KWS^{+}14, KNK^{+}21, LA15, LQZ23, LMN24, MKS19, OAV^{+}16, PCL^{+}20, RRC^{+}22, UB19, XPL23, AGZ10, Bar09, BFAM11, BSK^{+}05, Ber94, BPR08, CL08, CP10, CC13, DBT^{+}12, FS04, HT11, HLJ^{+}97, KWM97, LV20, LGHH08, LG12, LR13, LV09, NPLB09, NGB06, RM00, RO97, SP21, SHR07, SWLM22, TIG09, VM95, Vic00, WLB09, YTL^{+}23, ZB05]. basic [LV09]. bathing [KdJvE13]. Batman [BP23]. Bayesian [HSA^{+}23]. BCI [KTBR15b]. BCIs [KTBR15a]. Be [BFC12, LBGC24, OSBB^{+}19, Dou13]. Bearing [COFH16]. been [Dou13]. Beginning [OWOZ17]. Behavior [BTS^{+}13, CRL^{+}23, DLL24, GCB16, HSA^{+}23, PVSZ24, RTT19, RB23, VJR24, BFAM11, MJV^{+}06]. Behavioral [TGG24, HT11]. Behaviors [LMN24, RC22, YS20, VM95]. Behaviour [PVCB18, SRE^{+}21]. Behaviours [GPPD^{+}22]. Being [KIW16, NSP^{+}18]. Benefits [GMW05, SSC^{+}16, KSR14, LZB98, RL09]. Best [Hin17b, RHK^{+}24, XGA^{+}22]. better [DGK^{+}10]. Between [BFG^{+}23, JNR^{+}24, TNLK22, VBR20, BTS^{+}13, CALH^{+}19, GRR20, HBDG04, KH23, LDS^{+}11, PKES22, SHMA07, TRZ15]. Beyond [BM24, Bro12, HBJP24, MREN^{+}22, PS02, PDR17, YTL^{+}23]. Biases [TSTH17]. Big [BK18, BJC^{+}23]. Bigraphs [BCRS16]. Binocular [CZH^{+}15]. Biocentric [FKGW22]. Biodata [RMT^{+}15]. Biofeedback [VSW23]. Biological [BGK^{+}22]. biomedical [WFD98]. biometric [HT11]. Bisexual [TB24]. Blackness [CBRT23]. Blended [OKP11]. Blending [PPAA^{+}18]. blind [PRB^{+}11, QO13, TSFA07]. blindness [BHNG05]. Blockchain [MREN^{+}22].

Blockchains [MREN^{+}22]. Blocklists [JGBG18]. Blood [KLK^{+}23]. Blueprints [DLL24]. blues [RV95]. Blur [NGB06]. Bodily [MLH14, MGVE17]. Body [ABBH20, BBC^{+}22, KTN^{+}18, MYR^{+}20, MMHM23, RT18, SHG^{+}24, Sva13]. BodyWire [MYR^{+}20]. BodyWire-HCI [MYR^{+}20]. Books [GCS23]. Boolean [WFD98]. BotMap [DMOJ18]. Bots [VP24]. Boundaries [VBR20, BGR^{+}98]. Boundary [AFP13, WIKW23, CT07]. Bounds [MKFB20]. Bowl [WOB22]. Box [COFH16, ORRH99]. Brain [ACM18, CKS16, KTBR15a, PPS^{+}22, SZM23, DBT^{+}12]. Brain-Computer [SZM23, DBT^{+}12]. Brazilian [DKT^{+}21]. Breached [ZLM^{+}24]. Breaches [MZL^{+}23]. breadth [KB03]. Break [PVCB18]. Breakdown [FLvA24]. Breakdowns [LMMBL19, SM11b]. Breaking [CAP24]. Breathing [KSG^{+}22, MFU^{+}20]. Bridge [MBHC17]. Brief [KOP^{+}10]. broadcasting [BGC^{+}00]. brokers [AO12]. browser [PCV03]. Browsing [HF96, OC03]. BubbleView [KBB^{+}17]. Budgeting [PMAN24]. Build [RZK24]. builder [Wol97]. Building [ACW^{+}19, GB19, LW23, LKS19, RG96]. Buildings [BVL^{+}19]. Business [YRA20]. buttons [FH14]. ByteDance [WSWL23].

C [LQZ23]. C-PAK [LQZ23]. Calculations [VW22]. calendar [BCCR04, NBG09]. calendars [PCH^{+}06]. Call [CDC15, LBGC24, RPFMP17]. Calls [BAAL^{+}16]. Calming [MFPB24]. Camera [VS14]. Campaigns [LBO^{+}15]. Can [BCF^{+}06, BD22, DKA^{+}15, EBNM22, HFB09, LRS19, OSBB^{+}19, SSS^{+}22]. Cancellation [HLSR24]. Cannot [SLM^{+}20]. capabilities [MS94]. Capacity [PBBJS22]. Capital [AV19, LVG^{+}14]. Capitalism [DKT^{+}21]. Capture [HCH^{+}14, STKB12, BA04]. Captured [CDT^{+}21, CCO14]. Capturing [SGL09]. Care [GPH^{+}23, HJS^{+}20, MPS22, NVF^{+}15, WMRW16]. Caregivers [HCH^{+}14]. Carers [LGW^{+}19]. Cartography [TLA^{+}19]. cartoon [TC01]. Case [BQDB13, BBT^{+}23, DLL24, DBDK18, HBT^{+}21, ICC^{+}22, JGBG18, JBGB19, LJPS21, RCFR22, SHSS19, WSWL23, ZPSL08, ZSNP14, ABL05, FPST99, GEC^{+}09, NDSG06]. Cassowary [BBS01a]. Catwalk [AFP13]. causality [SJZ^{+}98]. CAVE [SGFT06]. CBT [THL^{+}23]. cell [SHR07]. Censorship [WM15]. Centered [ABC20, BMNH20, CKK^{+}21, FMP19, PJL^{+}16, TJLS21, WWD^{+}22, JNBF23, Rob05, THL^{+}23, WWM08]. Centers [RPFMP17]. Centred [ANW^{+}23]. centric [TE12]. Chair [SED^{+}16]. Challenges [ANW^{+}23, GBH^{+}18, JPM^{+}20, LPMST22, PWG18, PTR23, PPS^{+}22, RMT^{+}15, SEA22, TDKS19, WMRW16, WIKW23, XJS23, ZLD^{+}23, DHMV14]. Change [BTS^{+}13, DLL24, KSG^{+}22, OMV17, PVCB18, RTT19, RB23, ZLM^{+}24]. Changes [CKS16, BI08]. Changing [NHGC16, SYK23, BB09]. character [WH01, WM06]. character-level [WM06]. Characteristics [ANK^{+}23, BHH^{+}24, SRE^{+}21]. charge [VTS^{+}04]. charging [BSM^{+}13]. Charting [AM00]. Charts [FLST13]. Chat [BHA18, GPE06]. Chatbot [FLvA24, XZL^{+}20, YTL^{+}23]. Chatbot-based [YTL^{+}23]. Chatbots [CEH^{+}22]. Chatting [GPE06, KCL^{+}16]. Cheating [CGGN^{+}15]. Check [MBN^{+}23]. Checker [YTL^{+}23]. Checklists [KSCB21]. CheerOn [HLEG18]. cherished [KS10]. CHI [HMKV19, RHK^{+}24]. Chief [Zha15]. Child [SRE^{+}21]. Child-Robot [SRE^{+}21]. Children [ACS^{+}23, ACM18, BWTR12, GPPD^{+}22, JMP^{+}17, LRS19, PPAA^{+}18, RSC^{+}24, SRE^{+}21, SP21, SFKF19, HBDG04, Ink01, PRB^{+}11]. Children-Variance [SRE^{+}21]. China [WM15, Wan21]. Chiron [TNB^{+}95]. Chiron-1 [TNB^{+}95]. Choices [BWLW23, DLBBS24]. Choreography [HCS^{+}23]. Christmas [PL14]. Chronic [HLN^{+}14]. CIDER [OSPK23]. Citizen [LRP15]. Citizen-Generated [LRP15]. Citizens [MM21]. City [PCL^{+}20, Wan21]. Civic [Pil18]. Clarifying [VW22]. classes [KWL^{+}13]. Classification [KFG15]. Classifying [TLY^{+}24]. Classroom [LA17, SDBK24, BA04, SDBK24]. Classrooms [LPMST22]. CLERA [DTP^{+}23]. cliché [SKB^{+}23]. click [Ink01]. Clinical [BP23, JPM^{+}20, NKDB23]. Clinically [THL^{+}23]. Clinician [ZLD^{+}23]. Clinician-Facing [ZLD^{+}23]. Close [MBHC17, JH14]. clustered [SZG^{+}96]. Clustering [BPOW15]. clutter [CMH12, STH08]. Co [BT08, BPG^{+}22, BOMM18, GR11, KSB^{+}24, RBH24, RM23, SP21, WOB22, ZSE^{+}12]. Co-adaptive [GR11]. Co-constructed [BT08]. Co-Creative [RM23]. Co-Design [BPG^{+}22, KSB^{+}24, SP21]. Co-lead [RBH24]. Co-narrating [ZSE^{+}12]. Co-Present [BOMM18]. Co-speculation [WOB22]. Coactivation [BPOW15]. CoAIcoder [GCC^{+}24]. Code [BSR^{+}23, WWD^{+}22]. Codesign [USK^{+}23]. COFI [RM23]. Cognition [BAZ24, HHK00, Kir13, SKW01]. Cognitive [ARK^{+}21, CRH12, DTP^{+}23, MSH^{+}23, SP24, VML15, VJR24, CDF^{+}05, CL08, HK99, KSJB11, LZB98, RBJY00, RVB11, SHMA07, SHCP08, PW06]. Cognitive-Modeling [VML15]. cognitively [SHMA07]. CogTool [YLR21]. coincidence [LVH12]. Collaborating [JMP^{+}17, VP24]. Collaboration [BEJM14, BP23, GCC^{+}24, JBGB19, MM21, NKDB23, RZK24, RLP14, WKK^{+}23, BRS99, NDSG06, OKP11, PGG03, SXS^{+}06, YR12]. Collaborative [ARK^{+}21, ASJB22, AO12, CWM^{+}23, CALH^{+}19, DPG22, EZC24, JH14, RBK19, RAT^{+}22, SSC^{+}16, AEF^{+}00, BDR00, BGC^{+}00, Dou95, Edw05, GB95, HDM11, HFH^{+}00, JGH07, JMJ03, JH03, PK94, SRGS00, SC02]. Collaboratively [HJS^{+}20]. collaboratorium [BB02]. collaboratory [SWM03]. Collapse [TBN^{+}13]. Collecting [JSJ^{+}21]. Collections [CLY^{+}23, THA99]. Collective [CKKL18]. Collectively [FMEB20]. Color [GSX^{+}23, JKVA17, XPL23]. Color2Vec [XPL23]. ColorBless [CZH^{+}15]. Colorblind [CZH^{+}15]. Colors [JKVA17]. Colour [SLP22]. Colours [SLP22]. Com [GCS23]. combined [WZ97]. Combining [CJV16, ZCX^{+}24]. Comfort [MFPB24]. Comic [BKJ18]. Command [FSMP22, Ber94, GMW05, LMG^{+}11, LGHH08, LG12]. Commanding [GLZH20]. Commensality [FPD^{+}16]. Commentaries [DRW13]. commerce [KB03, QB05, WSWL23]. commercial [SHMA07]. Commitment [PFGE24]. Common [HH23, RSR21, CMS^{+}11]. Communicating [MYR^{+}20, VTS^{+}04]. Communication [BWTR12, HJS^{+}20, IIY^{+}24, MYR^{+}20, PDR17, PPAA^{+}18, SHD^{+}24, FPST99, KLMC12, KSJB11, Tan07, WJN^{+}04]. communications [PS02]. Communicative [PBBJS22]. Communities [CALH^{+}19, DV18, DVHZ^{+}21, GRKB16, HWC^{+}16, PRJ16, SLM^{+}20, XJS23]. Community [Bid21, BAHU24, CJBG22, MLC^{+}13, TJLS21, MKP05]. Community-Supporting [MLC^{+}13]. Community-Wide [CJBG22]. Companion [CNG^{+}24]. Comparative [LA17, SHMA07]. Comparing [CVZBB20, FPST99, WMMS08]. Comparison [DLL24, RRC^{+}22, RSR21, JK96a, WZ97, WFD98]. comparisons [PW06]. Compassion [MLH21]. Competence [GFK^{+}17]. Competence-Based [GFK^{+}17]. competition [CC13]. Compilation [HBJP24]. Compiler [PKRR15]. Complementarity [ICM^{+}23]. complementary [IA08]. Complementing [VK14]. Complete [KKH^{+}13]. Completing [LQZ23]. Complex [BWTR12, JS10, SN17, WKK^{+}23]. complexity [MJV^{+}06]. Complicated [PKES22]. Component [HH17]. composing [YLZ14]. Composition [DAM17, VK14]. comprehension [YLZ14]. Compromised [PFGE24]. Computational [BKG23, BFG^{+}23, OFLK17, SSHM^{+}23, SNT^{+}23, SLP22, VML15, WWD^{+}22, WL21, HR02, LFT06]. Computer [ASL^{+}22, ACM18, BMNH20, DFH^{+}15, FFKM22, KWCS22, KTBR15a, LMK21, MSA^{+}23, PKRR15, RSR21, SZM23, SKN24, SW09, TNLK22, WL21, XRL^{+}22, BDR00, BPR08, BP05, DBT^{+}12, Hay11, HHK00, HFB09, KWM97, KP10, SC03, Tan07, VTS^{+}04, ZBM96]. Computer-Driven [TNLK22]. computer-mediated [Tan07]. Computer-supported [SW09]. Computers [BD22, SHM07]. Computing [DFH^{+}15, GB19, RBH24, SAP^{+}15, SHVH^{+}21, AM00, Bar09, EJS02, Edw05, ENSS09]. concept [SKW01, WLB09]. Concept-Oriented [WLB09]. Conception [ALDR^{+}22]. Concepts [ALR20, JKVA17, Pil18, HL12]. Conceptual [KTBR15b, SSH22, ZCX^{+}24]. Conceptualization [CGL24, RBH24]. Conceptualizing [SBSS12]. Concern [CRH12]. Concerns [JSJ^{+}21, MM21]. Concurrent [FLCT19, Her24, Sun02]. conditions [CT07]. Conduct [XZL^{+}20]. Conductance [SHSS19]. Conferences [NSP^{+}18]. Conferencing [HHG^{+}22, GM03, NGB06]. confidence [FS04]. Confident [LK20]. Configurations [PMAN24]. Conflict [Jun16, ZSE^{+}12]. Conflicts [GC22]. Confrontation [NKDB23]. connecting [HLN04, BFL^{+}23]. Connections [KH23, NPF^{+}15].

Connectivity [DVHZ^{+}21]. Consensus [Tsa18]. Consent [GSPZ24, SKDM24]. Consequences [RAT^{+}22]. Considerations [LW23]. Considering [KC23, GPE06]. Consistency [SC02]. Constrained [PRJ16, CG10]. constrained-input [CG10]. constraint [BBS01a, UIJ05, VMGS94, VM95]. constraint-based [VM95]. Constraints [LKS19, VH01, WMMS08]. Construct [DKA^{+}15]. constructed [BT08]. Constructing [CNE^{+}07, THA99, HBR^{+}94, BGR^{+}98]. Construction [TB24, JH03, LV09]. Consumers [AWB^{+}23]. Consumption [KBJ^{+}13, MWB16, SSRW13]. contact [GM03, Tan07]. ContactMap [WJN^{+}04]. Contemplative [MLH21]. Contemporary [Men22]. Content [GRKB16, JGBG18, JBGB19, JNBF23, LCE^{+}19, LRP15, JBH13, KGZ07]. Context [AYK23, ASL^{+}22, BTS^{+}13, BHA18, FMSS17, GMPS17, KFG15, KSCB21, MBN^{+}23, NWH21, SSH22, DM05, Edw05, KP10, LG04, LDF12, PKHD09, LG04]. context-aware [DM05, PKHD09]. Context-Dependent [GMPS17]. Context-Driven [KSCB21]. context-enhanced [Edw05]. Context-sensitive [LG04]. Contexts [GC22, KSCB21, LAW18, XPL23, FPST99]. Contextual [ACDL12, FBE23, MM17, SS00]. Contingent [VAF17]. Continuance [XLC12]. Continue [GSPZ24]. Continuous [TPM^{+}21, FS04]. Contradiction [DAK20]. contrast [JK96a]. contributions [HWSB99]. Control [FFKM22, GCB16, KFF^{+}23, KTBR15a, MGMMS21, MOMS17, AO12, FKK07, LH08, Mac99, MD23, SW09]. controlled [DBT^{+}12, SWM03]. controller [NM09]. Controlling [DS98]. Convergence [MHL^{+}23, ODC04, SJZ^{+}98]. Conversational [FLvA24, PVSZ24, RK22, XZL^{+}20, LG04, MCC^{+}04, OS04, ODC04]. convey [Coh97]. Cooperative [BNS02, TCD^{+}21, FPST99, MS04, RSJ02, SJZ^{+}98]. Coordinated [MWVK21]. Coordination [CE20, RZK24, SG20, TKH11, NBG09]. Copilot [PRD^{+}24]. Coping [CLG20]. copresence [KOP^{+}10]. Correct [KLK^{+}23]. Correcting [LQZ23]. Correction [PD16, SMW01]. corrective [LG04]. Correlation [VCN^{+}17]. Corrigendum [YHS96]. Cost [JSJ^{+}21, PRJ16, BHNG05]. costs [PW06]. Could [JPR^{+}19]. Countdown [GRR20]. Counterfactual [FH23]. Counterintuitive [GPPD^{+}22]. Counterintuitive-Problem [GPPD^{+}22]. Coupling [DC95, Jun16]. Course [OWOZ17]. Covert [PVSZ24]. COVID [BHH^{+}24, MBN^{+}23, DPL^{+}23]. COVID-19 [BHH^{+}24, MBN^{+}23, DPL^{+}23]. Cracks [HDF21]. Craft [PF21]. Crafting [BKQ^{+}17, BPW12, GHB^{+}22]. Crane [SNT^{+}23]. Create [HGOZ19]. Creating [BCF^{+}11, CNG^{+}24, CALH^{+}19, JN96, MBP^{+}22, NM09, Shn00, YCVV23, AEF^{+}00]. Creation [AJSW12]. Creation-Oriented [AJSW12]. Creative [CLY^{+}23, GSX^{+}23, RM23, SBSG23]. Creativity [CL14, KQH^{+}22, ZCX^{+}24, Shn00]. Creatures [AFP13]. credibility [LF14]. Crip [Gub23]. Crisis [MHL^{+}23, HDM11]. Criteria [HG24]. Critical [CFH^{+}20, GC22, SFKF19, Tsa18, FPST99, GSM99, GPP99]. Criticality [FLvA24]. Criticism [RBB15]. Cross [CL17, CKKL18]. Cross-Cultural [CKKL18]. Cross-Device [CL17]. Crossing [Gru04, UK21, YS20, AGZ10]. Crowd [BGK^{+}22, SB18]. Crowd-Algorithm [BGK^{+}22]. Crowdfunding [GH13]. Crowdsourced [HRB14]. Crowdsourcing [GFK^{+}17, GCB16, KBB^{+}17]. Crowdworker [GCB16]. Crowdworkers [LBGC24]. crucial [NBG09]. CSCW [BEJM14, Dou98]. Cued [BLNH21].

Cued-Recall [BLNH21]. cues [Bre98]. Cultivation [MLH21]. Cultural [AATC22, BGA^{+}15, CKKL18, KFG15]. culturally [RB11]. Culture [TB24]. cultures [BPW12]. Curation [KWS^{+}14]. Current [HHE^{+}12]. cursor [WL97]. Custom [LMMBL19, TPM^{+}21]. Customer [FLvA24]. Customization [KKJ24]. Customize [DAM17]. CyberGlove [KHW95]. Cybersecurity [ZR21].

D [BYS02a, IIY^{+}24, IW03, LBP17, MWVK21, NHGC16, QB05, SGL09, WLB09]. Daily [BBC^{+}22, CCO14]. DANTE [YSHG07]. Dark [KSB^{+}24, RSC^{+}24]. DASS [NVPE19]. Data [ACS^{+}23, BTS^{+}13, CRL^{+}23, CFLC23, HHHV24, HSA^{+}23, JPR^{+}19, JSJ^{+}21, KHM20, KS15, MZL^{+}23, PSD24, SSH22, SJU19, SKDM24, WWD^{+}22, ZPSL08, ATH^{+}03, FKGB10, MJV^{+}06, SY97, CMLS10]. Data-Driven [KS15]. Data-enabled [CFLC23]. dataflow [VH01]. DateLens [BCCR04]. Dealing [POS^{+}01]. Debriefing [BLNH21]. debugger [MS94]. debugging [GBW^{+}12, MS94]. Dec [YHS96]. Decade [SEA22]. Decades [TLY^{+}24]. Deceptions [HBL22]. Decision [AF18, DLBBS24, FMSS17, ICM^{+}23, KB23, LK20, MSH^{+}23, WKK^{+}23, WWH19, ZSC^{+}15, vBBSB23, Rob05]. Decision-Making [AF18, LK20, WKK^{+}23, DLBBS24, KB23]. Decisions [FBE23, HBL22, ZR21, LB10]. declarative [PSS09]. Declare [DWL^{+}24]. Decoding [TGG24]. Decoration [BKQ^{+}17]. Dedicated [NSP^{+}18, NPLB09]. deductive [BYS02a, BYS02b]. Deep [EDT^{+}23, MRC^{+}22, SVCB21]. Defamiliarization [BBS05]. Default [WLB15]. Deficiency [SLP22]. deficits [SHMA07]. definitions [BG98]. Deformation [MMH^{+}22]. degradation [WWHW97]. Delay [GRR20]. delayed [JGH07]. Delays [LV20, PH23]. delegation [MCC^{+}04]. Deliberate [GSPZ24]. Delivery [PTR23, IB10]. demands [MRF09]. Dementia [FMP19, HFLH^{+}23, HJS^{+}20, MPS22, WMRW16]. Demographics [LCE^{+}19]. Demonstrating [RCFR22]. Demonstration [CL17]. Demonstrational [VM95, MCM97]. Department [PCR15]. Dependent [GMPS17]. Deployment [KH18, MMEFN^{+}24]. Deployments [MLC^{+}13]. Depression [CDT^{+}21]. depth [HWSB99]. Deriving [KSCB21]. Describe [SSH22]. description [NPLB09, NM09]. descriptions [MTDM14]. Descriptive [CMH12]. Design [ANO19, ACW^{+}19, AATC22, ABC20, ASL^{+}22, ACM18, AZS16, BPOW15, BBB18, Bar18, BPG^{+}22, BPL^{+}23, BK18, BWLW23, BBT^{+}23, BT15, CEH^{+}22, CWM^{+}23, CNG^{+}24, CFLC23, CKK^{+}21, DPG22, DS08, DBDK18, EYK^{+}16, EBNM22, ERT23, FKGW22, FMP19, GC22, GPPD^{+}22, GSK22, GDW^{+}24, Gil19, GBH^{+}18, GCS23, GLX^{+}23, GSX^{+}23, HL21, HWC^{+}16, Hin17b, HBT^{+}21, HHHV24, HDF21, ICC^{+}22, JKVA17, JBD^{+}22, KGYQ15, KS15, KTN^{+}18, KSB^{+}24, LGW^{+}19, LMG^{+}11, LWA^{+}23, LCK24, LA17, MCD^{+}20, MSH^{+}23, MSA^{+}23, OSPK23, OFLK17, PRR^{+}19, PAM^{+}22, PL14, Pil18, PMS24, RSR21, RTR^{+}16, RAT^{+}22, SWZ16, SZM23, SSHM^{+}23, Søn20, SVDM17, STB21, SWLM22, SNT^{+}23, TLA^{+}19, VL07, VBR20, WMRW16, WBFDK21, XRL^{+}22, YCVV23, YRA20, ZCX^{+}24, AEF^{+}00, BBS05, Bla06, BB02, BG05, CMS^{+}11, DRD^{+}00, DGK^{+}10, FPST99, FH08, GSM99, GFC13, HK99, HL12]. design [JK96b, KLS95, Kir13, KdJvE13, LB10, LST08, LR13, MAVR13, MS04, MBB07, PKHD09, PG94, Rob05, SP21, SJ09, SGL09, Sva13, Thi04, TBN^{+}13, WDHM13, WWM08, YR12, WLB09]. Design-Emerging [BBB18]. Design-Sketching [STB21]. Design-Space [HL21]. Designer [LMMBL19]. Designer-Developer [LMMBL19]. Designerly [BPG^{+}22]. Designers [ID20, Kir19, WOB22]. Designing [BSW17, BCF^{+}11, BMM20, CGS12, DM05, FJM24, HRW^{+}23, KHM20, KB23, KSB^{+}20, LWG^{+}14, LRS19, MPB^{+}11, MGVE17, NDSG06, NVPE19, NSCR19, OTV19, PF21, RM23, RK22, RBK19, SVCB21, SAT^{+}23, SAP^{+}15, Søn20, SC03, SHC05, TJLS21, THL^{+}23, TSGK14, VAAB^{+}16, XGA^{+}22, AGZ10, BSK^{+}05, EK00, PS02, Vic00, Wol97]. Designs [AGM23, DLL24, NBK24, GPP99, MB05]. desired [BSK^{+}05]. desk [BBS01b]. desktop [GM03, RSK04, WSKS97, WJN^{+}04]. detail [WWHW97]. Detected [OSBB^{+}19]. Detecting [BD22, CDT^{+}21, PVSZ24]. Determination [TM24]. deterrents [GH13]. Deterring [CGGN^{+}15]. Developer [LMMBL19]. Developing [Dou95, GLZH20, KTN^{+}18, SF15, THL^{+}23]. Development [EBNM22, MNPP17, TBD20, YHS95, YHS96, KLS95, PG94, TNB^{+}95]. Device [CL17, LPMST22, NWH21, HB07, KHW95]. Devices [FSMP22, JWS12, TRZ15, HPHS05, JSM^{+}94, KHA11, MRF09]. Diagnosis [GLX^{+}23, YTL^{+}23]. diagram [CM03, WMMS08]. diagrammatic [HK99]. Diagramming [IW03]. Diagrams [RSC15, HK99]. dialogs [MCC^{+}04]. Dialogue [RSR21]. Dictated [GLZH20]. Dictation [GLZH20, FS04]. Did [CRL^{+}23, MA23, RZK24]. Differences [HBDG04, TRZ15, LF14, MRF09, ZS06]. Different [WBD^{+}22, KSJB11]. Differential [KYZ23]. Differently [Kir19]. difficulty [CD11]. Diffractive [LJPS21]. Digital [AGM23, BMM20, CGL24, CL14, CFH^{+}24, DKT^{+}21, HL18, HBJP24, HFK^{+}22, JDV^{+}21, JMP^{+}17, JNR^{+}24, NEE^{+}24, OTV19, PMAN24, PCL^{+}20, PF18, PVCB18, PBBJS22, MD23, RTR^{+}16, RAT^{+}22, VJR24, ABL05, BBS01b, ENSS09, LVH12, SGL09, UIJ05]. dimensional [GB05, Hor01, WST14]. Dimensions [CRH12]. Diminished [GCB16]. Direct [AGB14, BG98, GMW05, HFB09, MRF09]. directed [BI08, IB10]. Direction [SHG^{+}24]. Directions [SEA22, WIKW23]. disabilities [CDF^{+}05]. Disability [GDW^{+}24]. disabled [YSHG07]. Disambiguation [WST14]. Disclosure [And20, KB23]. Disclosures [AF18, AHCF18]. Discomfort [HLSR24]. Discourages [BBC20]. Discourse [RBB15, QMB^{+}02]. Discourses [VPW^{+}15]. Discussion [Roo13]. Display [MLC^{+}13, MSD^{+}21, AO11, AO12, JH14, KSK02, WWHW97]. Displays [BVL^{+}19, HFK^{+}22, VAAB^{+}16, VAF17, MMS^{+}08, MWW06, TGSP06, WB94]. Disposal [SWZ16]. disruption [SM11b]. Disruptive [PVSZ24]. Dissecting [HBL22]. Distance [BEJM14, MWB16]. Distant [MBHC17]. distortion [LA94]. distortion-oriented [LA94]. Distract [LRS19]. Distraction [AOB^{+}20]. Distractions [BBC20]. Distractor [YS20]. Distress [BAAL^{+}16]. Distressing [And20]. Distributed [BEJM14, HHK00, GG99, OKP11]. Distributions [EDT^{+}23]. Distrust [KH18]. Dive [MRC^{+}22]. divergence [DGK^{+}10]. Diverse [CRM17]. Diversifying [PRR^{+}19]. divide [Gru04]. DIY [SN17]. Do [BLNH21, DMG16, HMKV19, JPR^{+}19, Kir19, LCK24, SLM^{+}20, DWL^{+}24, MO94]. Doctors [HHG^{+}22]. document [DELS99, LEF^{+}00, RP96]. Documentation [BP23, WWD^{+}22, HF96]. documents [HF03]. Does [BEJM14, GBBM12, TH15]. Dolls [SLBB19]. Domain [DAM17, TVH16]. Domain-Specific [DAM17]. Domestic [NPF^{+}15, BBS05, GEC^{+}09]. Done [PVMK24]. Don’t [GSPZ24]. Dot [GCS23]. Dot-Com [GCS23]. Downside [CVZBB20].

Drag [ACPL15, Ink01]. Drag-and-Drop [ACPL15, Ink01]. Drawing [Coh97]. Driven [KS15, KSCB21, SIK^{+}12, TNLK22]. Drop [ACPL15, Ink01]. Dualistic [HBT^{+}21]. Duration [GRR20]. Durations [BBC20]. During [MHL^{+}23, MWS18, MYR^{+}20, SG20, WBW^{+}23, AGB14, BI08, FPD^{+}16, LVH12, LRP15, MJV^{+}06, RSJ02, SM11b, XHM^{+}13]. Dyadic [PBBJ^{+}22]. Dynamic [KSCB21, LRP15, PD16, JBH13, MWW06]. Dynamically [LAW18]. Dynamics [USK^{+}23, MKP05].

E-commerce [WSWL23, KB03]. E-government [DHMV14]. E-Textile [PMS24]. Early [JNR^{+}24, KYZ23, LVDA23]. EarthShake [YHWK16]. easier [HK99]. Easy [TPM^{+}21]. Eating [GSK22]. eClass [BA04]. Economic [BBMT06]. Economy [TSTH17, Vic00]. EcoSanté [KH23]. Edge [DKT^{+}21]. Editing [GLZH20, LMW^{+}20, SJZ^{+}98, SC02, WMMS08]. Editor [Hin16c, Hin16a, Hin16d, Hin16e, Hin16f, Hin16g, Hin17e, Hin17a, Hin17b, Hin17c, Hin17d, Hin18a, Hin18b, Hin18c, Zha15]. Editor-in-Chief [Zha15]. Editorial [CMLS10, Hin16b, MNPP17, Zha14]. editors [CM03, Sun02, RZK24]. edits [SGL09]. Education [HFP12, VSS^{+}23, NWH21, TDKS19, YHS95, YHS96]. Educational [HL18, LBO^{+}15, LA17, VWKL^{+}24]. Effect [AYK23, BHH^{+}24, CZH^{+}15, DWL^{+}24, EZC24, KWCS22, SSS^{+}22, TCD^{+}21, ZBM96]. Effective [TBD20, BPR08, Sut00]. effectively [SS94]. Effectiveness [GCC^{+}24, MMHM23, KLMC12, WFD98]. Effects [ARK^{+}21, BSG18, CJBG22, CD11, GWZC23, HK99, KB03, LCE^{+}19, MSM23, NHGC16, SHG^{+}24, VBHK10, XLC12, FH14, GG99, JGH07, LBT96, PCV03, QB05, WWHW97]. Efficacy [OMV17, DGK^{+}10]. efficiency [KLMC12, WFD98]. Efficient [FSMP22, TPM^{+}21, BPR08, MF10]. Electricity [JKS18, KBJ^{+}13, KTN^{+}18]. Electro [MYR^{+}20]. Electro-Quasistatic [MYR^{+}20]. electronic [HF03, QB05]. Electronics [BSR^{+}23, BPW12]. Elephant [SBSG23]. Elicitation [OSPK23, Tsa18, VW22]. Eliciting [HM23]. elision [HLN04]. Elite [RT18]. Email [LCE^{+}19]. Emails [LCE^{+}19]. Embedded [BSR^{+}23, LJY^{+}13]. Embodied [Kir13, SP21, LR13, MAVR13, VBHK10]. Embodiment [KPWS20, SWZ16]. Embracing [LMK21]. emergence [BT08]. Emergencies [LRP15, RLP14]. Emergency [PCR15, CMS^{+}11, TKH11]. Emergent [FH14]. Emerging [BBB18, KSB^{+}24, VSS^{+}23, Pil18, PF21, PMK02]. Emoji [PDR17]. Emotion [FLvA24, ICC^{+}22, SAT^{+}23]. Emotional [SF15]. Emotionally [MFPB24]. Emotions [BBC^{+}22, CLG20, GBBM12]. Empirical [BB18, BAZ24, CKS16, DFH^{+}21, HFP12, JBD^{+}22, PJL^{+}16, RPFMP17, YRA20, KSR14, MCM97]. Empowering [ABY17, DAM17]. empowerment [ABL05]. EMR [PCR15]. Enabled [KB21, SP24, CFLC23, vBBSB23]. Enabling [LV09, MYR^{+}20, QO13]. Enact [LMMBL19]. Enclosure [Bid21]. encountering [LCHD11]. Encounters [CVB16, KOP^{+}10]. Encourage [KH23]. Encouraging [ZLM^{+}24]. Encyclopedia [RZK24]. End [ABY17, BWR^{+}17, DAM17, GBW^{+}12, Hin17b, MNPP17, MM17, VW22, KSR14]. End-User [Hin17b, MNPP17, VW22, GBW^{+}12]. End-Users [ABY17]. ended [XZL^{+}20]. Endurance [DAK20]. Energy [ACR^{+}16, SKDM24, SSRW13]. Engagement [BBT^{+}23, JNR^{+}24, LCC23, MSM23, MPS22, MREN^{+}22, Pil18, SRE^{+}21, SDBK24, YRA20]. Engine [KQH^{+}22].

Engineer [BBT^{+}23]. Engineering [BAZ24, WBW^{+}23, KWM97]. Enhance [KB23, SSS^{+}22]. enhanced [Edw05, HPHS05]. EnhancedDesk [BBS01b]. Enhances [JWS12]. Enhancing [PAM^{+}22, CC13, FKK07]. Enjoyment [YHWK16]. Enriching [ACDL12]. Entanglement [Fra20]. Enterprise [GRKB16]. Entity [JDV^{+}21]. Entrepreneur [NBB20]. Entrepreneurs [BBR18]. Entry [DVK18, MF10, VK14, WST14, WM06]. Environment [BKJ18, KH23, CGS12, GB95, JGH07, TkWSR99, WL97]. Environments [AGWF19, CGGN^{+}15, DAM17, LK20, BhHSS00, BDR00, BGC^{+}00, CDF^{+}05, FKK07, HFH^{+}00, KSJB11, LJY^{+}13, MMS^{+}08, PRM00, PSS09, RSJ02, RL09, RVB11, SRGS00, SS00, SGFT06, TSGK14, Wex95]. EPIC [KWM97]. Epilogue [Dou13]. Epistemic [Gub23, TB24]. equivalent [SHMA07]. Era [GCS23, Hin16b]. Err [KYZ23]. error [GPP99, SMW01]. Errors [HBL22, KYZ23, SMB12, WM06]. Essentialised [KPWS20]. Establishing [BP05, NWH21]. estimation [HBE96]. Estrellita [HCH^{+}14]. eTextile [PF21]. Ethical [BGA^{+}15]. Ethics [SEA22]. ethnographic [ORRH99]. ethnography [PMM^{+}13]. ethnomethodology [MS04]. Evaluate [CRH12, KWS^{+}14]. Evaluating [CT07, FLST13, HL18, LG12, MPS22, SWM03, VWKL^{+}24, YSHG07, AGZ10, LWG^{+}14, SHC05]. Evaluation [BWLW23, BMB^{+}13, BLNH21, CFH^{+}24, DPG22, EBNM22, GLX^{+}23, Har19, KHW95, KWB^{+}15, MFPB24, RTR^{+}16, SVDM17, UB19, JK96b, LMG^{+}11, MCSN03, MBB07, MCM97, PGG03, Sal09, SHR07, YHS95, YHS96, YR12, IA08]. evaluations [VK14, WM06]. Even [KLK^{+}23]. Events [AJSW12]. Everyday [ASJB22, COFH16, JDV^{+}21, RBK19, SED^{+}16, HR02, LJY^{+}13, PMK02]. Evidence [RSR21, WKK^{+}23, YHWK16]. Evoke [EBNM22]. Evolution [ALR20, CWZL23, RSK04]. evolving [ENSS09]. Examining [And20, BAHU24, GCC^{+}24, MWB16, TDKS19, VPW^{+}15, WM15, WBFDK21, CJBG22, KSB^{+}24]. exclusion [ABL05]. executable [WLB09]. execution [BI08]. Exercise [SHG^{+}24, WZ97]. Exergame [UB19]. Exertion [ASL^{+}22, MGVE17]. Exertion-Understanding [ASL^{+}22]. Existential [RB23]. existing [SHMA07]. ExoBuilding [SIK^{+}12]. Expanding [MKFB20, BG98, MB05, RV95]. Expected [BSK^{+}05]. Experience [BM24, BKPH22, BLNH21, CEH^{+}22, CGL24, FMP19, GDW^{+}24, HH23, HH17, KKJ24, MZR^{+}21, MKS19, MLH14, RB23, SHG^{+}24, TH15, TVH16, vSHL12, BSW08, WWM08]. Experience-Centered [FMP19, WWM08]. Experiences [And20, BPG^{+}22, BMM20, CHAN20, DKT^{+}21, ENSS09, GSK22, HGOZ19, WBD^{+}22, Wan21]. Experiencing [LVH12, SKB^{+}23]. Experiential [HH17]. Experiment [PPS^{+}22, SSS^{+}22, SWM03]. Experimental [TSGK14, BhHSS00, DELS99, HFB09, LZB98, LBT96, MB05]. Experiments [ATH^{+}03, CP10]. Expert [BKG23, KSB^{+}24, LGC17, LV20, CG10, SHR07]. Expertise [CKS16, ICM^{+}23]. Explanation [KB21]. Explanations [PKES22]. Exploiting [DMG16, DRD^{+}00, PSD24]. Exploration [CFH^{+}20, SVCB21, UK21, WBW^{+}23, KdJvE13, LDS^{+}11, WZ97]. exploratory [CT07, Rie96]. Explore [BBC^{+}22, MMH^{+}22]. Exploring [AGB14, AGWF19, BWR^{+}17, CWM^{+}23, CFLC23, GC22, GSK22, GSPZ24, GDW^{+}24, HFLH^{+}23, HLSR24, HJS^{+}20, KBJ^{+}13, MSM23, PRJ16, RB23, SN17, SLM^{+}20, SKDM24, SYS19, TGG24, ENSS09]. Expression [MSM23]. expressions [HR02].

Expressive [AZS16, PD16, PDR17, SHS24]. Expressiveness [BHA18, Red08]. Expressivity [CDT15]. Extended [MBP^{+}22]. Extending [BMB^{+}13, SLS94, LH08]. extensible [Edw05]. extent [RVB11]. Extravaganza [Hin17b]. Extravaganza-Special [Hin17b]. Extruding [HBJP24]. Eye [CVC12, DTP^{+}23, GSX^{+}23, HFK^{+}22, MKS19, SSC^{+}16, SG20, WKK^{+}23, GM03]. Eye-tracked [GSX^{+}23]. Eye-Trackers [SSC^{+}16]. eyed [WL97]. Eyes [GLZH20]. Eyes-Free [GLZH20].

Fabricable [SNT^{+}23]. Fabrication [AGWF19, HBJP24, LAW18]. Facebook [LVG^{+}14, YRA20]. facial [VBHK10]. facilitate [SGL09, ZSE^{+}12]. Facilitates [IIY^{+}24]. Facilitating [HLEG18]. Facilitators [DV18]. Facing [BK18, MLC^{+}13, ZLD^{+}23]. Factor [MRC^{+}22]. Factorial [LVDA23]. Factors [CWZL23, FBE23, LCC23, WBW^{+}23, LDF12]. fading [PMK02]. Fails [SG21, NGB06]. Failure [HDF21, HLJ^{+}97]. Fall [OAV^{+}16, UB19]. Fallacies [Tsa18]. familiarity [LBT96]. Familie [SKDM24]. Family [FPD^{+}16, PL14, JK96a, NBG09, PCH^{+}06]. Farewell [Zha15]. Fast [DVK18, MWVK21]. Feature [CDT^{+}21, WWH19, YRA20, MBB07]. feature-rich [MBB07]. Features [ANO19, GPPD^{+}22, PVSZ24, PSD24]. Feedback [BBC20, CNG^{+}24, CLY^{+}23, DKA^{+}15, DLBBS24, FFKM22, LMDT22, MWS18, TGG24, FH14, JGH07, SRGS00]. Feedforward [MMH23]. Feel [BLNH21, DMG16]. Feeling [CFH^{+}24]. Feminism [Bar18]. few [TIG09]. Fiction [Men22, WDHM13]. fidelity [TKH11]. Field [CFH^{+}24, OAV^{+}16, ASHM97, MBB07, Rie96]. fieldwork [PRM00]. Filmmaking [RMT^{+}15]. Filter [BJC^{+}23].

Filtering [HL21]. filters [LST08]. filtration [NGB06]. Finding [DAK20]. Findings [JBD^{+}22, Ree19]. Fine [SSHM^{+}23]. Fine-grained [SSHM^{+}23]. Finger [SSHM^{+}23, TRZ15, BBS01b]. Finland [JMP^{+}17]. fire [TKH11]. First [CVB16, DTL^{+}21, GWZC23, WBFDK21]. First-Person [DTL^{+}21, WBFDK21]. Fish [DT23, WL97]. fisheye [BCCR04, HH07, SZG^{+}96]. fit [TkWSR99]. FITTS [GRG18, MB05]. Five [ALR20, BMB^{+}13]. Five-User [BMB^{+}13]. Flexibility [LBGC24]. Flexible [BCGVP24, BRS99, Dou98]. Flicking [JWS12]. flight [Mac99]. Flow [BSR^{+}23, QB05]. Flow-Based [BSR^{+}23]. Flowboard [BSR^{+}23]. Flud [BGK^{+}22]. fluid [DELS99, MMS^{+}08]. fNIRS [SAP^{+}15]. focus [DSG09]. focused [HFH^{+}00]. Foldable [SNT^{+}23]. Folding [EZC24]. Fonts [WBD^{+}22]. Food [BWLW23, COFH16, ERL^{+}23, GSK22, GFC13, SHC05]. Foot [EBNM22]. Foot-Tickling [EBNM22]. force [SRGS00]. Forced [DBDK18]. Foreground [HPHS05]. forgotten [BHNG05]. Form [KB23, PVCB18, RV95]. Form-Autocompletion [KB23]. Formal [CFH^{+}20, GRG18]. Formalization [JBD^{+}22]. formats [WZ97]. foundation [HHK00]. Foundations [AGZ10]. Four [LMK21]. fragments [LG04]. frame [WB94]. Framework [AF18, CL17, DS08, GSS^{+}19, JNBF23, KFG15, KHM20, LWA^{+}23, MLH21, MLC^{+}13, NVPE19, NSCR19, RMP^{+}21, RM23, SP24, SSH22, SZM23, VS14, BSK^{+}05, DRD^{+}00, IB10, LCHD11, PK94, SGL09, TM05, YR12]. Framing [BSG18, GPH^{+}23, vBBSB23]. Free [GLZH20, PRJ16, RZK24, FS04]. Freehand [UK21, SGL09]. frequency [SS94]. Friction [SB18]. Friendship [LVG^{+}14]. friendsourcing [BTS^{+}10]. Fro [Hor16]. Front [LBGC24]. Frustration [HH23]. full [SZG^{+}96]. full-zoom [SZG^{+}96]. Fully [LMW^{+}20]. Functional [SNT^{+}23]. Functionality [OFLK17]. Functions [CVC12]. Fundamental [DAK20]. Fundamentals [BEJM14]. Future [ACR^{+}16, BT15, FMEB20, GC22, VSS^{+}23, PRR^{+}19, SEA22, VPW^{+}15, AM00, Kir13, MHP00]. Future-Making [PRR^{+}19]. Futures [Bar18, FH23]. Futuring [Ada23]. Futurity [Gub23].

Gain [STKB12]. Game [GPPD^{+}22, GDW^{+}24, HL18, LA17, TLA^{+}19, YHWK16, ZRT^{+}20, CC13, KTBR15b, TKH11]. game-based [CC13]. Gameplay [ARK^{+}21, GBBM12]. Games [BFC12, HL18, JPM^{+}20, MGVE17, OMV17, SVDM17, SG21, TM24]. Gaming [SED^{+}16, XJS23]. Gap [VBR20]. GAVIN [KNK^{+}21]. Gaze [IIY^{+}24, KNK^{+}21, SHD^{+}24, SG20, VAF17]. Gaze-Assisted [KNK^{+}21]. Gaze-Contingent [VAF17]. Gazing [SHD^{+}24]. GenAI [CNG^{+}24]. GenAI-Powered [CNG^{+}24]. Gender [KPWS20, USK^{+}23, KLMC12]. General [NKDB23]. generalization [BYS02a, BYS02b]. Generated [DWL^{+}24, LRP15]. Generating [LAW18]. Generation [SWZ^{+}24, Wan21, CM03, SJGL09]. Generative [BLBM21, ZCX^{+}24, MTDM14]. generic [JN96, MO94]. Genie [DLBBS24]. Geographic [TSTH17]. Geometry [NHGC16]. Georeferenced [ZPSL08]. Geriatric [BBT^{+}23]. Gesture [CDT15, SWLM22, TPM^{+}21, Tsa18, XGA^{+}22, LGHH08, QMB^{+}02, Wex95]. Gesture-based [SWLM22, LGHH08]. Gestures [HGH^{+}24, LMW^{+}20, TRZ15, BG98, QO13]. Get [BFC12, PVMK24]. Getting [DV18]. Ghostwriter [DWL^{+}24]. Gifting [SKB^{+}23]. Gifts [KSB^{+}20]. Gigapixel [RTR^{+}16]. girl [TDKS19]. Giving [KLK^{+}23]. Glasses [SLP22]. Global [BJC^{+}23, Vic00]. Go [HBJP24, SWZ16, DPL^{+}23]. Goal [VML15, BI08, IB10, TSGK14]. goal-directed [BI08, IB10]. Goal-Specific [VML15]. goals [MCSN03]. Going [ANO19]. GOMS [JK96a, JK96b]. Good [RRC^{+}22, TVH16]. Got [DLBBS24]. Governance [TJLS21]. government [DHMV14]. GPS [WMRW16]. Graduate [YCVV23]. grained [SSHM^{+}23]. grammars [KZZ06]. graph [KZZ06]. Graphical [AZS16, BG98, MCM97, Ber94, HLN04, KZZ06, TC01]. graphics [SC02]. Graphs [FLST13, Coh97, VH01]. Grassroots [GB19]. Green [WDHM13]. Greenhouse [CHAN20]. Grid [CAP24]. Grief [SWZ16]. Ground [RSR21, CMS^{+}11]. Grounded [BMB^{+}13]. Grounds [CBRT23]. Group [EZC24, Sun02, TkWSR99]. GroupKit [RG96]. groups [MO94]. groupware [GG99, MO94, PGG03, RG96]. Growers [CHAN20]. Grows [DVHZ^{+}21]. growth [SKN24]. GSR [ZSC^{+}15]. GUI [LMN24]. guide [PKHD09]. Guideline [KTN^{+}18]. Guidelines [LVDA23, MSH^{+}23, SAP^{+}15, KdJvE13]. Guiding [LCK24, RHK^{+}24]. Guilty [DMG16]. Gullible [RSC^{+}24].

Habits [PVCB18]. Habituation [ASJB22, NVR^{+}14]. Hackathon [HFLH^{+}23]. hand [KHW95]. handed [HPPK98, KHA11, LZB98]. handheld [MWW06]. handle [WW05]. handles [WR99]. Handmaid [MA23]. Hands [WZ97, CNE^{+}07, FS04]. hands-free [FS04]. Hands-on [WZ97, CNE^{+}07]. handset [PS02]. handwriting [PRB^{+}11]. Hanging [ASHM97]. Haptic [MMH^{+}22, PBBJ^{+}22, SYK23, JGH07, SRGS00]. Haptics [SHS24]. Harassment [JGBG18]. Hardware [WBW^{+}23]. Harm [XJS23]. Having [LBGC24]. Haystack [LRP15]. HBI [ACW^{+}19]. HCI [Ada23, ACW^{+}19, ALR20, ABBH20, BGA^{+}15, BFL^{+}23, BV20, CR13, CWZL23, COFH16, CALH^{+}19, CBRT23, DTL^{+}21, DRW13, DBDK18, ERT23, FKGW22, FKGB10, Fra20, HH17, HFP12, JRB^{+}24, KPWS20, LJPS21, MLH21, MYR^{+}20, MAVR13, VSS^{+}23, Men22, MJV^{+}06, NVF^{+}15, PRM00, PSSB13, PMM^{+}13, PPS^{+}22, RB23, RBC^{+}21, RBB15, RHK^{+}24, SEA22, SAT^{+}23, SG21, Sut00, TBD20, TM24, VBR20, Vic00, VPW^{+}15, XGA^{+}22, ZLD^{+}23]. Head [DVK18, HFK^{+}22, KRMS21, SG20, UK21, KSK02, WWHW97]. Head-Mounted [DVK18, HFK^{+}22, UK21, KSK02, WWHW97]. Heads [SHD^{+}24]. Health [ABBH20, ACS^{+}23, BTS^{+}13, HWC^{+}16, HLN^{+}14, JNR^{+}24, KPWS20, KH23, NBB20, SP24, SLM^{+}20, Søn20, TBD20, THL^{+}23, TDKS19, Bec04, LWG^{+}14, LF14, MKP05]. Healthcare [ANW^{+}23, AWB^{+}23, BP23, BBT^{+}23, MHL^{+}23, PTR23]. Heating [JKS18]. Hedonic [XLC12]. Help [DKA^{+}15, PVMK24, VS14, QB05]. Helps [CLY^{+}23, KKH^{+}13]. Herbal [PKRR15]. here [VTS^{+}04]. Heritage [KFG15]. Herzberg [TH15]. Heuristic [EYK^{+}16]. heuristics [IA08]. Hey [DLBBS24]. Hidden [PWG18]. Hide [SBSG23]. hierarchically [SZG^{+}96]. hierarchies [Hor01]. High [BHNG05, GR11, HBE96, JH14, KWM97]. High-cost [BHNG05]. high-performance [KWM97]. high-resolution [JH14]. High-speed [HBE96]. Hinckley [Zha15]. hinder [BHNG05]. Hiring [SB18]. Historically [Men22]. Histories [FH23, HM23]. History [BFL^{+}23]. HMDs [MWB16, RMP^{+}21]. Hoc [BAAL^{+}16, RLP14, ENSS09]. Holding [PTR23]. Holistic [KKIT20]. HoloSketch [Dee95]. Home [ACR^{+}16, BWR^{+}17, CR13, GC22, HHG^{+}22, LW15, SN17, UB19, GEC^{+}09, KS10, NGB06, ORRH99, SM11a, STH08]. Home-Based [UB19, NGB06]. homepages [LDS^{+}11]. Homogeneous [HT11]. Honor [DAK20]. Hopeful [KC23]. Hospital [LRS19]. hospitals [Bar09]. Hotkeys [FSMP22]. Household [SKDM24, SKDM24]. Human [ACW^{+}19, ANW^{+}23, ASL^{+}22, CVB16, CWM^{+}23, CWZL23, DFH^{+}15, FFKM22, GCC^{+}24, GSK22, HSA^{+}23, JBGB19, KGYQ15, KTN^{+}18, KWCS22, KTBR15a, LMK21, LFT06, LKS19, MSM23, MYR^{+}20, MSA^{+}23, PKRR15, PH23, PPAA^{+}18, RTT19, RSR21, SVCB21, SKN24, THL^{+}23, TCD^{+}21, WWD^{+}22, WBW^{+}23, YHS96, YLR21, AEF^{+}00, BDR00, BP05, FHA^{+}05, FH08, Hay11, HHK00, KWM97, KS10, KP10, OCM^{+}12, QMB^{+}02, SC03, ZBM96, BP23, GLX^{+}23, ICM^{+}23, LVDA23, RM23]. Human-Agent [TCD^{+}21]. Human-AI [BP23, GLX^{+}23, ICM^{+}23, LVDA23, RM23]. Human-Building [ACW^{+}19]. Human-Centered [WWD^{+}22, THL^{+}23]. Human-Centred [ANW^{+}23]. Human-Computer [ASL^{+}22, DFH^{+}15, FFKM22, KWCS22, LMK21, MSA^{+}23, SKN24, BDR00, BP05, Hay11, HHK00, KWM97, KP10, SC03, ZBM96]. Human-Food-Technology [GSK22]. Human-Machine [JBGB19, LFT06]. Human-Robot [PH23]. Human-to-Human [GCC^{+}24]. Humanity [ABC20]. Humans [BD22, YCVV23]. Humor [BKJ18]. Hybrid [BGK^{+}22, BBD^{+}24, EYK^{+}16, KSB^{+}20, KWB^{+}15, LCC23, SKB^{+}23, ZCX^{+}24, ZSNP14]. Hygienes [TH15]. HyperActive [SLS94]. hyperbolic [PCV03]. hypermedia [SLS94, YHS95, YHS96]. Hypertext [WFD98]. Hypotension [GPH^{+}23]. Hypothesis [JBD^{+}22].

ICOs [NPLB09]. ICT [ERL^{+}23, OAV^{+}16]. ICT-Based [OAV^{+}16]. Ideas [ID20, LST08, STH08]. Ideation [KWS^{+}14, OCM^{+}12]. Identification [KPO18]. Identifying [FJM24]. Identity [BT15, HM23, PCL^{+}20]. Ideology [BHH^{+}24]. Idiosyncrasies [SRE^{+}21]. Igroup [TLY^{+}24]. Illiteracy [RAT^{+}22]. Illness [HLN^{+}14, LGW^{+}19]. Illusions [MMHM23]. Illustrated [AOB^{+}20]. Image [BBC^{+}22, KBB^{+}17, LCK24]. Image-schematic [LCK24]. imagery [DBT^{+}12]. Images [BJC^{+}23, CNG^{+}24, RTR^{+}16]. Imaginaries [BBT^{+}23]. imitation [FKKH10]. immersive [FKK07]. Impact [Bro12, FLvA24, GSS^{+}19, HL18, ICM^{+}23, KYZ23, SEA22, STKB12, USK^{+}23, GSM99, OCM^{+}12, DLBBS24]. Impacts [BSG18, BSR^{+}23, DPL^{+}23]. Impaired [VS14]. Impairment [ZPSL08, MJV^{+}06]. Impairments [DPG22, KRMS21]. Impala [BCRS16]. Implementable [TBD20]. Implementation [PCR15, SJ09]. Implications [BGA^{+}15, BPL^{+}23, FMP19, JBD^{+}22, Sva13, BD22]. Implicit [CJV16, KNK^{+}21, MBHC17, SAP^{+}15]. Importance [AATC22, KBB^{+}17, EJS02]. Impossible [BBD^{+}24]. imprecise [WST14]. Impressions [CVB16]. Improv [CL17]. Improve [CVZBB20, FS04, TGSP06]. Improved [VAF17]. Improves [YHWK16, RVB11]. Improving [FLST13, GLX^{+}23, HLN04, MKS19, OMV17, RB11, RM00, TSFA07]. improvisation [XHM^{+}13]. Improvising [CL17]. IMU [SSHM^{+}23]. In-game [KTBR15b]. In-Person [WKK^{+}23]. In-the-Wild [MMEFN^{+}24]. in-vehicle [Sal09, TSGK14]. Inaccessibility [BBR18]. Inaccurate [TGG24]. Incloodle [SDBK24]. Incloodle-Classroom [SDBK24]. Include [WL21]. Inclusion [ABC20]. Inclusive [BT15, LCK24, OSPK23, SDBK24]. Income [PAM^{+}22]. Inconspicuous [MFU^{+}20]. Incorporating [FMSS17]. Increase [LV20, WBD^{+}22, BHNG05]. increased [DGK^{+}10]. Increasing [CJV16, GRKB16, LMDT22, PAM^{+}22, SB18]. incremental [Coh97]. independence [BRS99, HB07]. Independent [BVR15]. Index [CL14, JWS12, Ano96, Ano97]. Indexicality [KP10]. Indexing [Hor16]. India [JMP^{+}17, KLK^{+}23, TDKS19]. Indicators [AJSW12]. indirect [MRF09]. Individual [ARK^{+}21, BCGVP24, MZL^{+}23, AEF^{+}00, JH03]. Individually [MZR^{+}21]. Individuals [WBD^{+}22, SHMA07]. Individuated [WBD^{+}22]. industrial [NDSG06]. industrial/academic [NDSG06]. Industry [WMRW16]. ineffable [BSW08]. Inequality [WSO16]. Infants [HCH^{+}14]. Inference [vSHL12]. inferential [OCM^{+}12]. Influence [CEH^{+}22, DCO13, TCD^{+}21, SUS95]. Influencing [LCC23]. Inform [GDW^{+}24, SHSS19, GFC13]. Informal [IIY^{+}24, DSG09]. Informatics [RT18, SP24, TBN^{+}13]. Information [CZH^{+}15, DS08, FBE23, GRG18, HJS^{+}20, KWS^{+}14, PRJ16, XLC12, BBS01b, CCO14, DSG09, HC06, HLN04, IW03, LF14, PCV03, SHCP08, UIJ05, WFD98, YLZ14]. Information-Based [KWS^{+}14]. Information-Sharing [FBE23]. Information-Theoretic [GRG18]. Informed [BWLW23, Men22, SWZ16]. Informing [BPOW15, HL21]. Infrastructural [BBR18]. infrastructure [Edw05, SM11b]. Infrastructures [LBP17]. Infrastructuring [CALH^{+}19]. Ingestible [LWA^{+}23]. Inhabited [BGC^{+}00]. Initiation [IIY^{+}24]. Initiative [BBT^{+}23]. Initiatives [Pil18]. Injustice [TB24]. Ink [SB18]. innovation [Shn00]. Input [BPOW15, CL17, PSD24, CG10, JSM^{+}94, KHW95, LZB98, MRF09, PT01, WST14, WM06]. inquiry [SSRW13]. ins [GEC^{+}09]. Insights [RPFMP17, SYK23, TLY^{+}24, TSGK14]. inspection [FH08]. Inspired [CFH^{+}24]. Institutions [CALH^{+}19]. instruction [BPR08, CL08]. instructions [LBT96]. Instrumental [SLM^{+}20]. Integrality [JSM^{+}94, GEF98]. Integrated [ASL^{+}22, SNT^{+}23, CL08]. Integrates [CKK^{+}21]. Integrating [BBS01b, CWZL23, VMGS94, MS94, OC03]. Integration [ASL^{+}22, GLX^{+}23, SZM23, HK99, RSJ02]. Integrations [MSA^{+}23, SZM23]. Integrity [BP23]. Intelligence [CWZL23, KWCS22, PPAA^{+}18, SBSG23, ZCX^{+}24]. Intelligent [BD22, LA17, PLF20, SLY^{+}18, ZSC^{+}15, CM03]. Intent [RBK19]. Intention [MZL^{+}23, SJZ^{+}98]. Interact [YHS96]. Interacting [ACR^{+}16]. Interaction [ACW^{+}19, AGB14, AZS16, BKG23, BMNH20, BLBM21, BBD^{+}24, BMB^{+}13, BOMM18, CL17, CJV16, DFH^{+}15, DS08, EZC24, FFKM22, GSK22, GSPZ24, GLZH20, Gil19, GSS^{+}19, HMKV19, ID20, KFF^{+}23, KWCS22, LGC17, LB10, LLZ14, LVDA23, LMK21, LMDT22, LMN24, LKS19, MYR^{+}20, MBHC17, MKS19, MPS22, MMH23, OFLK17, PVSZ24, PKRR15, PH23, PF18, PL14, RCFR22, RM23, SKN24, SWLM22, Sva13, TGG24, TCD^{+}21, WBFDK21, BDR00, BGC^{+}00, BSK^{+}05, DH08, FKK07, Hay11, HF96, HPHS05, HFH^{+}00, HHK00, HL12, HLJ^{+}97, Ink01, JS10, KWM97, Kir13, KP10, LR13, LJY^{+}13, MAVR13, MS04, MTDM14, MCC^{+}04, OKP11, OHM^{+}13, PHJ08a, PHJ08b, Red08, RM00, Rul08, SS00, SHR07, SGFT06, UIJ05, ZBM96, ZB05]. interactional [BCF^{+}11]. Interactions [AGM23, BCRS16, Jun16, KCL^{+}16, LMMBL19, MCD^{+}20, NVPE19, PRD^{+}24, RRC^{+}22, VAAB^{+}16, vBBSB23, BT08, BMDD00, WSKS97]. Interactive [BVL^{+}19, BKQ^{+}17, CAP24, Gil19, HWC^{+}16, KKJ24, LCK24, LRS19, SLY^{+}18, SVDM17, WWH19, YHWK16, DRD^{+}00, DELS99, GPP99, LGHH08, LG12, LBT96, NPLB09, OM11, ZSE^{+}12]. interactively [Aro97]. Interactivity [Kir19]. Intercorporeal [VSW23]. Interdependent [ANK^{+}23]. Interest [KFG15]. Interface [ANO19, AATC22, DMOJ18, GPPD^{+}22, KBB^{+}17, SSC^{+}16, SLY^{+}18, ZSC^{+}15, BCCR04, CP10, DBT^{+}12, GSM99, GPP99, HSD08, JK96a, JK96b, LBT96, MBB07, Mye95, MHP00, NPLB09, NM09, OCM^{+}12, PG94, PT01, RV95, RL09, SKW01, TIG09, TNB^{+}95, Thi04, Wol97, XHM^{+}13]. Interfaces [ACM18, BSW08, BD22, CFH^{+}20, CRH12, DT23, KTBR15a, LA15, LV20, MKS19, RTR^{+}16, SZM23, SAP^{+}15, SYK23, TLA^{+}19, AGZ10, BCF^{+}11, Ber94, CG10, CGA06, DC95, DS98, GEF98, HF96, HBP02, JDM99, KZZ06, LG04, LH08, MTDM14, MB05, MPB^{+}11, OS04, ODC04, PW06, Red08, RB11, Sal09, SJGL09, SJ09, Shn00, SC03, SMW01, TC01, TSGK14, WLB09, Wol97, YR12]. Interfacing [ACW^{+}19]. Interference [BJC^{+}23, KSK02]. Interferences [LMN24]. Interfering [KIW16]. Interior [GSX^{+}23]. interleaving [MCC^{+}04]. intermediaries [DHMV14]. Intermediate [HL12]. Intermediate-level [HL12]. Intermittent [MGMMS21]. Internal [KLS95]. International [SEA22]. Internationalization [LA15]. Internet [ABY17, DVHZ^{+}21, Hin17b, KB21, KGZ07, LCE^{+}19, LBP17, MNPP17, MZR^{+}21, OPL10, PRJ16, WM15]. Internet-Enabled [KB21]. Internet-Free [PRJ16]. interoperability [ENSS09, KSJB11]. Interpersonal [PBBJS22, XJS23]. Interplay [MGVE17]. Interpret [CLY^{+}23]. interpretation [LG04]. Interpretive [BSG18]. Interrogating [FMEB20]. interruptibility [FHA^{+}05]. Interruption [BBC20, BI08]. Interruptions [BBC20, WW05]. Intersectional [ALDR^{+}22, ERT23, KKIT20]. Intertwined [MSA^{+}23]. Intervention [CJBG22, CFH^{+}24, KH23, MFPB24].

Interventions [JNR^{+}24, PVCB18, SAT^{+}23, ZR21]. Intimacy [SLBB19]. Intimate [HHE^{+}12, ICC^{+}22]. Intimate-Space [ICC^{+}22]. Intonation [PD16]. Intrinsic [BM24]. Introducing [RPFMP17]. Introduction [ACW^{+}19, ABBH20, ANW^{+}23, BBB18, BRK15, BDR00, CCG^{+}13, DFH^{+}15, DTL^{+}21, GPP99, HC06, HDM11, JMJ03, JNM05, MAVR13, OSF95, OS04, PHJ08a, PHJ08b, PSSB13, PBBJS22, RD05, SY97, SHM07, SJGL09, SHVH^{+}21, TM02, ZB05]. Introspecting [MKS19]. Inverse [HSA^{+}23, Sun02]. Investigating [ARK^{+}21, AZS16, LGC17, MZR^{+}21, SHD^{+}24, SHCP08, SGFT06]. Investigation [GCB16, OWOZ17, WWH19, CT07, LBT96, QB05]. Invisible [GDW^{+}24]. Invitation [RBH24]. Involving [FKGW22]. IoT [AKTB19, LK20]. IPQ [TLY^{+}24]. ISIS [MCC^{+}04]. Islandness [RBC^{+}21]. Isolation [HRW^{+}23]. Isopleth [HGOZ19]. Issue [ABBH20, ANW^{+}23, CCG^{+}13, DFH^{+}15, DTL^{+}21, Hin16c, Hin16a, Hin16d, Hin16e, Hin16f, Hin16g, Hin17e, Hin17a, Hin17b, Hin17c, Hin17d, Hin18a, Hin18b, Hin18c, PBBJS22, BDR00, DRW13, GPP99, HC06, MAVR13, OSF95, PHJ08a, PHJ08b, PSSB13, SY97, SHM07, SJGL09]. Issues [BK18, MLC^{+}13, GPP99, PRM00]. Iterative [XRL^{+}22]. Iteratively [XGA^{+}22]. Itinerative [PRR^{+}19].

jam [XHM^{+}13]. Jets [SHS24]. jigsaws [JH03]. Joint [DTP^{+}23, ICM^{+}23, SDBK24]. judgment [HSD08]. judgments [LF14]. Just [FH23, LA15, MBN^{+}23, PDR17, MRC^{+}22]. Justice [XJS23].

Kalas [SHC05]. Kansuke [TIG09]. Ken [Zha15]. Kenyan [WSO16]. Kettle [ASJB22]. key [MF10]. Keyboard [FSMP22, MBP^{+}22, SYS19, MF10]. Keyboards [MBP^{+}22]. keyphrases [CMH12]. Keystrokes [LQZ23]. Kickstarting [NBB20]. Kindergarten [SDBK24]. Kinect [NVR^{+}14]. Knowledge [BP23, CFH^{+}24, LW23, HL12, JH03, SHCP08, Sut00, Vic00]. knowledge-based [Vic00]. Knows [PRD^{+}24].

Lab [ALR20]. labeled [Hor01]. laboratories [CNE^{+}07]. Laboratory [CR13, Roo13]. lag [WB94]. Landmarks [SYS19]. Language [AWB^{+}23, BTS^{+}13, CEH^{+}22, CKK^{+}21, NKDB23, HBR^{+}94, BG05, JDM99, NM09, PSS09]. languages [BG98]. Large [CLY^{+}23, KWB^{+}15, LBO^{+}15, VML15, YLR21, CCO14, RP96, RVB11, TGSP06]. Large-Scale [LBO^{+}15, CCO14, RVB11]. Larp [JRB^{+}24]. Late [KYZ23]. Later [LW15]. Laugh [BKJ18]. Laughter [EBNM22]. law [MB05]. Layer [PMS24]. Layered [MLC^{+}13]. LAYERS [MLC^{+}13]. Layouts [SSHM^{+}23]. lead [RBH24]. leads [DGK^{+}10]. Lean [JSJ^{+}21]. Leaps [SBSG23]. Learn [CRL^{+}23]. Learned [KWB^{+}15, SAP^{+}15, BPR08, BA04, TE12]. Learners [KS15]. Learning [BRK15, BSR^{+}23, CDT^{+}21, EDT^{+}23, GPH^{+}23, Gil19, HLEG18, HGOZ19, HSA^{+}23, JKVA17, KGYQ15, KTBR15a, LW23, MSM23, MMEFN^{+}24, PKRR15, SSC^{+}16, SVCB21, TBD20, VSW23, WWH19, YHWK16, CL08, CC13, Rie96, SKW01, WZ97, XHM^{+}13, BG05]. Learnt [MMEFN^{+}24]. learnware [SKW01]. leave [Tan07]. leave-taking [Tan07]. Led [BGC^{+}13]. Leisure [TVH16]. Length [LQZ23]. Lens [BA24, GSK22, KS15, RBC^{+}21, GFC13, PMM^{+}13]. Lessons [BA04, KWB^{+}15, MMEFN^{+}24, SAP^{+}15, BPR08, TE12]. Let [XHM^{+}13]. Letting [SWZ16]. level [HL12, MS94, PSS09, WWHW97, WM06].

Leveraging [MJV^{+}06, PVMK24]. LGBTQ [ALDR^{+}22]. libraries [ABL05]. Life [Bid21, BBC^{+}22, LW15, NPF^{+}15, Wan21, WIKW23, BSM^{+}13, GPE06]. Lifestyle [KH23]. lighting [HWSB99]. Lightweight [BOMM18, NM09, WSKS97, WW05]. Like [HBJP24, LBGC24, GCS23]. Limitations [JBD^{+}22]. Limits [LBGC24]. line [PRB^{+}11]. Linear [RSC15, BBS01a]. lines [HLN04]. Linguistic [BHH^{+}24]. Link [BKPH22]. Linking [HCS^{+}23, IB10, MS04]. Lions [BCRS16]. list [RV95]. listening [LVH12]. literacy [MPB^{+}11]. Literature [ALR20, BAZ24, DFH^{+}21, RZK24, SFKF19, TBD20, XGA^{+}22]. Live [BSR^{+}23, JRB^{+}24, QB05]. Lived [GDW^{+}24, RB23, Sva13]. Livestream [WSWL23]. Living [ALR20, BVR15, CR13, MPS22, PWG18, Wan21]. Load [DTP^{+}23]. Local [BJC^{+}23, Bro12, WWH19]. Localization [LA15]. Location [PCL^{+}20, DRD^{+}00]. Location-Based [PCL^{+}20]. logograph [TIG09]. Loneliness [HRW^{+}23]. Long [ASJB22, JKS18, KKIT20, NPF^{+}15, OAV^{+}16, BP05]. Long-Term [ASJB22, JKS18, NPF^{+}15, OAV^{+}16, BP05]. Longer [LV20]. Longitudinal [CDT^{+}21, MSD^{+}21]. look [TIG09]. look-up [TIG09]. looking [GM03]. Loss [ALDR^{+}22]. Lost [ICC^{+}22]. Loud [BKJ18]. Love [HHE^{+}12]. Low [JSJ^{+}21, PRJ16, PAM^{+}22, PLF20, SYS19, MPB^{+}11]. Low-Cost [PRJ16]. low-literacy [MPB^{+}11]. Low-Occlusion [SYS19]. lower [HFB09]. Luster [CZH^{+}15].

Machete [TPM^{+}21]. Machine [CDT^{+}21, GPH^{+}23, Gil19, JBGB19, LW23, SBSG23, TBD20, WWH19, LFT06]. Machines [DMG16, HBJP24]. macrotheory [BMDD00]. Madness [ERT23]. magical [Kir13]. maintaining [BP05]. maintenance [SC02, TNB^{+}95]. Make [CRL^{+}23, CVZBB20]. Makes [KSB^{+}24]. Making [AF18, AWB^{+}23, BBS05, FMSS17, Hor16, ICM^{+}23, LK20, MHL^{+}23, PRR^{+}19, PBBJ^{+}22, RBB15, STH08, TLA^{+}19, WKK^{+}23, ZSC^{+}15, DLBBS24, KB23, LR13, SSRW13]. Malicious [WSWL23]. Manage [ID20]. Management [CKKL18, HLN^{+}14, KLK^{+}23, LPMST22, BI08, CMS^{+}11, HDM11, WH01]. manager [TE12]. Managing [BAAL^{+}16, JKS18, PVMK24, PH23, WWHW97, WW05, WIKW23, YCVV23, WSKS97]. manifestations [LST08]. Manipulating [GPPD^{+}22]. Manipulation [MWVK21, BNS02, BG98, GMW05, HPPK98, HFB09, RSJ02, SKW01]. Manual [LZB98]. Manyfesto [Ada23]. map [RVB11]. Mapping [CVC12, MCD^{+}20, WLB15]. Maps [KBB^{+}17, TLA^{+}19]. Marginalized [LMK21]. MARIA [PSS09]. marking [KHA11]. Markov [TCJ01]. Marvista [CWM^{+}23]. Masculinity [GCS23]. Mashups [BTS^{+}13]. Mass [KCL^{+}16]. massive [KWL^{+}13, GB95]. Matching [VCN^{+}17, ZR21, TM05]. Material [KHM20]. materials [BPW12]. matrix [Thi04]. Matter [ANO19, BEJM14, PF21]. Matters [BK18, BHH^{+}24, WWD^{+}22]. Matters-Facing [BK18]. Me [ACS^{+}23, DLBBS24, PFGE24, XZL^{+}20, BCF^{+}06, GM03]. Mealtime [FPD^{+}16]. Mean [HMKV19]. Meaning [PBBJ^{+}22, RT18]. Meaningful [OTV19]. Meanings [SWLM22]. Measurable [ZSC^{+}15]. Measurement [BB18, CGL24]. Measurements [vBBSB23]. Measures [CE20, MWS18, BBMT06]. Measuring [BFAM11]. mechanics [PGG03]. Mechanism [EBNM22, FSMP22, Ber94]. Mechanisms [YR12]. Media [ANK^{+}23, AF18, AHCF18, And20, BA24, BFG^{+}23, CKKL18, CLG20, MWB16, SDBK24, ASHM97, BG05, FPST99, HDM11, LCHD11, SM11a]. Mediate [KLK^{+}23]. Mediated [HM23, HRW^{+}23, KIW16, LGW^{+}19, SHD^{+}24, XRL^{+}22, LJY^{+}13, SM11b, Tan07]. Mediating [HHE^{+}12]. mediation [DM05]. Medical [AWB^{+}23, Bar09]. Medium [EYK^{+}16, TkWSR99]. Meeting [BSW08]. melody [PT01]. Memories [BBC^{+}22, KIW16]. Memory [KIW16, MM21, MMH^{+}22, HLJ^{+}97]. MemoVis [CNG^{+}24]. Men [SLBB19]. Menstrual [FMEB20, TDKS19]. Menstruating [NBB20]. Mental [JNR^{+}24, LGW^{+}19, MWS18, RSC^{+}24, TBD20, THL^{+}23, BI08, LWG^{+}14]. Menu [DLBBS24, RV95, SHR07]. Menus [JWS12, HH07, KHA11, SS94]. merging [GMW05]. Merleau [Sva13]. Merleau-Ponty [Sva13]. Messages [BOMM18, CVZBB20]. Messaging [ASJB22, BPL^{+}23, BHA18, SS00]. Messengers [JKVA17]. Meta [Her24, KPO18, MMHM23, MD23]. Meta-Analysis [KPO18, MMHM23, MD23]. Meta-Analytic [Her24]. metalevel [Dou98]. Metaphor [AGM23, Bla06]. Metaphoraction [SWLM22]. Metaphorical [SWLM22]. Metaphors [DT23, FH08, LCK24]. Metatation [MBHC17]. Metering [JPR^{+}19]. Method [CRH12, ERT23, ZCX^{+}24, BBS01b, BYS02a, BYS02b, Coh97, FPST99, GSM99, TBN^{+}13]. Methodologies [HL18]. Methods [BPOW15, CKK^{+}21, DTL^{+}21, DFH^{+}21, NEE^{+}24, Tsa18, WBFDK21, IA08, SZG^{+}96]. Metrics [KWS^{+}14, BFAM11, YHS95, YHS96]. MHP [LFT06]. mice [HBDG04]. Micro [HBL22, KRMS21]. Micro-Decisions [HBL22]. Microgestures [SSHM^{+}23]. Microtasks [GFK^{+}17]. Mid [NPCBL15, SHS24]. Mid-Air [NPCBL15, SHS24]. Middle [WL21]. might [Dou13]. Migrants [DBDK18]. Migration [SEA22]. Milling [HBJP24]. mind [AEF^{+}00]. Mindfulness [NSCR19]. Minecraft [TJLS21]. Mingongs [Wan21]. Minimalist [BCGVP24]. Minimize [LMN24]. Mining [FKGB10, JKVA17, ATH^{+}03, CMLS10]. Mirror [KSG^{+}22]. Mismatched [AGM23]. Mitigating [BSG18, TB24]. Mixed [NEE^{+}24, SED^{+}16, BGR^{+}98]. Mixed-Methods [NEE^{+}24]. mixed-reality [BGR^{+}98]. ML [TBD20]. Mobile [ACDL12, BQDB13, BHA18, HCH^{+}14, LMW^{+}20, LWLL22, RLP14, SSC^{+}16, SWZ^{+}24, TRZ15, VAAB^{+}16, WSO16, DSG09, DCO13, DRD^{+}00, EJS02, HPHS05, KP10, LB10, MPB^{+}11, OS04]. mobility [KOP^{+}10, POS^{+}01]. Modalities [MYR^{+}20]. Modality [KTBR15a]. Model [BKG23, CRL^{+}23, CLG20, DAM17, DTP^{+}23, FBE23, KPO18, KFF^{+}23, LFT06, MGMMS21, PKES22, PJL^{+}16, RB23, SHR07, CT07, CG10, Dou95, HLN04, JDM99, MCSN03, NPLB09, SGL09, VH01]. Model-based [SHR07, NPLB09]. ModelCraft [SGL09]. Modeling [ANK^{+}23, FFKM22, HSA^{+}23, JGH07, LVG^{+}14, MSM23, MA23, PDR17, RM23, TVH16, VML15, WL08, XPL23, YLR21, CL08, GB05, JH03, KOP^{+}10, ODC04, PGG03]. Modelling [BCRS16]. Models [FMEB20, MOMS17, OMV17, RSC^{+}24, RSR21, STKB12, WL21, ZCX^{+}24, KWM97, RBJY00, TCJ01, VMGS94]. Moderated [LWG^{+}14]. Moderating [XLC12]. Moderation [CJBG22, JGBG18, JNBF23, TJLS21]. modern [RSK04]. modified [TIG09]. Moments [CRM17]. Money [PF18]. Moneywork [PF18]. Monitoring [NEE^{+}24, WMRW16, RO97]. Monolingual [HRB14]. Monster [KSG^{+}22]. MOOC [KWB^{+}15]. Mood [RPFMP17]. Mothering [DAK20]. Motion [PD16, VCN^{+}17]. Motivating [BFC12]. Motivation [BM24, CJV16, KS15, PVMK24, ZLM^{+}24]. Motivational [CVZBB20]. Motivations [AGM23, GH13]. Motivators [TH15]. Motives [KCL^{+}16]. motor [CD11, DBT^{+}12]. Mounted [DVK18, HFK^{+}22, UK21, KSK02, WWHW97]. Mouse [MGMMS21, HT11, Hor01, Ink01]. mouse-based [HT11]. mouse-pointing [Hor01]. Movement [GB19, Gil19, SHG^{+}24, VCN^{+}17, VSW23, LR13]. movement-based [LR13]. Movements [KFF^{+}23, MGMMS21, YS20]. Moving [LR13, PRM00]. Multi [KPO18, LPMST22, PMS24, CGS12, SXS^{+}06, WMMS08, YR12]. Multi-Device [LPMST22]. Multi-Layer [PMS24]. multi-slate [CGS12]. Multi-Touch [KPO18]. multi-user [SXS^{+}06, YR12]. multi-way [WMMS08]. Multidimensional [PSD24]. Multifaceted [MLC^{+}13]. Multilevel [VAAB^{+}16, MKP05]. multimedia [RP96]. Multimodal [CE20, MMEFN^{+}24, QMB^{+}02, SBSG23, SMW01, FH14, GEF98, KWM97, PRB^{+}11]. multiple [JMJ03, PSS09, PW06, TSGK14]. multiple-goal [TSGK14]. Multiples [FMEB20]. multitask [LFT06]. Multitasking [GCB16, BFAM11, MCSN03]. Multithreaded [LG04]. multitouch [JH14, KHA11]. Multiuser [FMSS17, HBR^{+}94, DC95, DS98]. Muscle [BPOW15, CDT15]. museum [YHS95, YHS96]. Museums [Hor16]. Music [AKTB19, BNS02, LVH12]. musical [XHM^{+}13]. Must [WL21]. My [ACS^{+}23, CRL^{+}23, DLBBS24, MLH14].

Nah [MRC^{+}22]. Name [USK^{+}23]. Nanites [RO97]. narrating [ZSE^{+}12]. Narrative [BWTR12, HDF21, WBFDK21]. Natural [AWB^{+}23, BTS^{+}13, MM17, NKDB23, Wex95, WLB09]. Naturalistic [VJR24]. naturalness [OHM^{+}13]. navigability [TSFA07]. Navigate [ALDR^{+}22, RL09]. Navigating [MWW06, PVMK24, RTR^{+}16, SZG^{+}96]. Navigation [HBP02, KRMS21, KGYQ15, SYK23, Bre98, CG10, FS04, MWW06, SHC05]. navigational [KLMC12]. Necessary [YS20]. Need [BKJ18, HHE^{+}12, PVMK24, MO94]. Needs [BWTR12, BWR^{+}17, CMLS10, CCO14]. Negotiating [GFC13, Tan07]. Negotiation [MBN^{+}23]. negotiations [WW05]. Neighborhood [CR13]. Network [LFT06, WL08]. Network-Model [LFT06]. Networking [PCL^{+}20, RC22, GEC^{+}09, SBSS12]. Networks [BGK^{+}22, Bid21, ENSS09, KOP^{+}10, SZG^{+}96]. Neurodivergent [SG21]. Neurodiverse [SDBK24]. Neurofeedback [ACM18]. Never [LHF^{+}22]. News [CWM^{+}23, HG24]. newsgroups [ZS06]. Next [Fra20, SJGL09]. next-generation [SJGL09]. No [Ada23, BKJ18, GPP99]. Noise [HLSR24]. Nomadic [SS00]. Non [And20, DMOJ18, HBT^{+}21, RC22, JDM99]. Non-Autistic [RC22]. Non-Disclosure [And20]. Non-Dualistic [HBT^{+}21]. Non-Visual [DMOJ18]. non-WIMP [JDM99]. nonexpert [LV09]. Nonprofit [BV20]. nonspeech [Bre98]. Nonvisual [BMNH20]. Norm [PVSZ24]. Norman [PJL^{+}16]. Norms [HSI^{+}20]. notation [BNS02]. Note [KNK^{+}21, DSG09]. Note-taking [KNK^{+}21, DSG09]. Notebooks [WWD^{+}22]. Nothing [HBJP24]. Notices [GSPZ24]. Notification [LHF^{+}22, IB10, MCSN03]. Notion [TH15]. Novel [BPOW15, EBNM22, MCD^{+}20]. Novice [BKG23, HLEG18, PRD^{+}24, CG10, MPB^{+}11]. Nudge [ZR21]. Nudges [GSPZ24, PFGE24]. NUI [OHM^{+}13].

Oasis [IB10]. Object

[HFH^{+}00, KLS95, RSJ02]. Object-focused [HFH^{+}00]. object-oriented [KLS95].

Objects [ABY17, BA24, CFLC23, MFPB24, NHGC16, RBK19, TGG24, VCN^{+}17, YHWK16, Ber94, KS10, VM95, WB94, WR99]. Observation [MLH21]. observer [PG94]. obstacles [PG94]. Occlusion [DVK18, SYS19, ZBM96]. Occupied [BBR18]. Ocular [TGG24]. Oculography [CVC12]. Off [TNLK22, GRR20, JNBF23]. Old [RSC^{+}24, VPW^{+}15]. Older [FBE23, KH18, LCK24, MM21, NVR^{+}14, OAV^{+}16, PLF20, Bec04, MJV^{+}06, NDSG06, SHCP08]. OLEI [KS15]. Olfactory [MCD^{+}20]. On-Skin [HGH^{+}24]. One [BPL^{+}23, PVMK24, MF10, VMGS94, WST14, WL97, WMMS08]. one-dimensional [WST14]. one-eyed [WL97]. one-key [MF10]. One-Way [BPL^{+}23, VMGS94, WMMS08]. ongoing [SM11b]. Online [ALDR^{+}22, BRK15, CLG20, CGGN^{+}15, GWZC23, GRKB16, HLEG18, HWC^{+}16, HBL22, HSI^{+}20, JGBG18, KS15, KCL^{+}16, PVSZ24, SLM^{+}20, TB24, THL^{+}23, XJS23, Bec04, HF96, KWL^{+}13, LWG^{+}14, LF14, MKP05, ZS06]. only [ASHM97]. Onset [CDT^{+}21]. Ontologies [CBRT23]. Open [COFH16, XZL^{+}20, SLS94]. Open-ended [XZL^{+}20]. Opening [ACM18]. OpenStreetMap [VP24]. Opera [GHB^{+}22]. Operations [YS20]. Opinions [CLY^{+}23]. Opportunities [GPH^{+}23, GBH^{+}18, HLSR24, JPM^{+}20, LPMST22, NVF^{+}15, PMS24, SKDM24, TDKS19, WIKW23, XJS23, ZLD^{+}23, PMK02]. Oppression [ERT23]. Optimal [FFKM22]. Optimisation [LMDT22]. Optimism [RBB15]. Optimistic [LLZ14]. Optional [HBL22]. options [FPST99]. Order [FLvA24, GWZC23]. Ordinary [CHAN20]. Organic [KSG^{+}22]. organization [CP10, SM11a]. Organizational [PCR15, PG94, GR11]. Organizations [BV20, Pil18, GR11, KSJB11]. organize [SS94]. organizing [THA99, WJN^{+}04]. Oriented [AJSW12, DRW13, KLS95, LA94, PSS09, PSSB13, WLB09]. Origami [SNT^{+}23]. Other [KSG^{+}22, STH08]. Our [KSG^{+}22, KIW16]. Outcomes [MHL^{+}23, PJL^{+}16]. Outdoor [SVDM17]. Outperform [BD22]. Output [USK^{+}23]. Outs [BD22, GEC^{+}09]. OverCode [GSS^{+}15]. Overview [SEA22, HBP02, NM09]. Own [BFC12, BSW08]. Owned [RMP^{+}21]. Ownership [DWL^{+}24, MMHM23].

P [MLC^{+}13]. P-LAYERS [MLC^{+}13]. Pacer [MFU^{+}20]. Pages [YRA20]. Pair [RK22]. PairBuddy [RK22]. PAK [LQZ23]. Palestine [BBR18]. Pandora [COFH16]. Panning [DMOJ18]. Paper [AWB^{+}23, Hin17b, BBS01b, LGHH08, LG12, Mac99, WH01]. Papers [AWB^{+}23]. PapierCraft [LGHH08]. paradigm [AGZ10, BYS02a, BYS02b, SJ09]. Paradigms [DAM17, TM24]. Paradoxing [KC23]. Parallel [DGK^{+}10, LH08]. Parameter [SVCB21]. parameters [VL07]. Park [ZRT^{+}20]. Parks [ACDL12]. Part [STKB12]. partial [TkWSR99, ZBM96]. partial-occlusion [ZBM96]. Participant [HFLH^{+}23, RMP^{+}21, PG94]. participant-observer [PG94]. Participant-Owned [RMP^{+}21]. Participation [Bar18, RLP14, GH13]. Participatory [BBB18, BK18, DBDK18, GBH^{+}18, PMAN24, PAM^{+}22, Pil18]. Partners [HJS^{+}20, RM23, RBK19]. Passive [CDT^{+}21]. Passwords [PFGE24, ZLM^{+}24]. Past [FMEB20, MHP00, AM00]. Pathology [GLX^{+}23, RTR^{+}16]. Patient [NEE^{+}24]. Patients [HHG^{+}22]. Pattern [MFU^{+}20, RSC^{+}24]. Patterns [BTS^{+}13, KCL^{+}16, KSB^{+}24, LVG^{+}14, MS04, CL08, HBP02, HF03]. Pavlov [Wol97]. Paying [CVZBB20]. PD [Pil18].

PDAs [BCCR04]. Pedestrian [SYK23]. peephole [MWW06]. Peer [CLY^{+}23, HRW^{+}23, KWL^{+}13, XHM^{+}13]. Pen [AGB14, CAP24, TRZ15, LG12, RM00, SGL09]. pen-based [LG12, RM00]. People [AKTB19, BJC^{+}23, CVZBB20, CZH^{+}15, DPG22, DMG16, KRMS21, KKH^{+}13, LGW^{+}19, LMK21, MPS22, NVR^{+}14, OWOZ17, CDF^{+}05, SC03]. Perceive [DWL^{+}24]. Perceived [BB18, ZRT^{+}20, BBMT06, BHNG05, LDS^{+}11, RB11]. Perception [DLBBS24, GRR20, KKJ24, NHGC16, SHD^{+}24, SHG^{+}24, SYK23, DH08, QB05]. Perceptions [ANK^{+}23, HBL22, MRC^{+}22]. Perceptual [SSC^{+}16, IB10, IW03]. Performance [AYK23, BGC^{+}13, Jun16, PKRR15, STKB12, YLR21, CG10, HBDG04, KWM97, LBT96, LFT06, RB11, RM00, TGSP06, WWHW97]. Performance-Led [BGC^{+}13]. Performing [DH08, PCL^{+}20, SHVH^{+}21]. Perioperative [GPH^{+}23]. peripheral [WWHW97]. Persistence [LGC17]. Person [ALDR^{+}22, DTL^{+}21, WBFDK21, WKK^{+}23]. Personal [AGWF19, BWTR12, BHA18, HM23, RT18, SP24, SJU19, JH14, NM09, WH01]. Personalisation [MSD^{+}21]. Personality [LVG^{+}14, DCO13]. Personalization [BTS^{+}10, GMPS17, MFU^{+}20, OMV17, TDH10]. Personas [SSS^{+}22, ODC04]. Persons [HJS^{+}20]. Perspective [BT15, FKGW22, GPH^{+}23, JPR^{+}19, LWLL22, SWZ16, SZM23, vSHL12, CL08, GBW^{+}12]. perspectives [GPH^{+}23]. Pervasive [MSD^{+}21]. Phasic [CKS16]. phenomenological [SSRW13]. phenomenology [Sva13]. Philosophers [WOB22]. Phishing [LCE^{+}19]. Phone [GHB^{+}22, DCO13, HLJ^{+}97, SHR07]. phone-based [HLJ^{+}97]. Phones [WSO16]. Photography [VS14]. Photos [HSI^{+}20]. Physecology [SSH22]. Physical [BCRS16, BFC12, CVZBB20, JMP^{+}17, KHM20, OSBB^{+}19, PSD24, SP21, YHWK16, ZCX^{+}24, LDF12, PKHD09]. Physical/Virtual [BCRS16]. Physicalization [HHHV24]. Physicalizations [SSH22]. Physically [TGSP06]. physio [HT11]. physio-behavioral [HT11]. Physiological [BKJ18, CJV16, DFH^{+}15, MWS18, SAP^{+}15]. Physiologically [SIK^{+}12]. Physiology [HCS^{+}23]. Physiopucks [CJV16]. pictorially [VM95]. Picture [AGM23, BJC^{+}23]. Pictures [SSS^{+}22]. Pitch [PD16]. Pitfalls [WWH19]. PIV [MFU^{+}20]. place [BB02, GPE06, STH08]. Placebo [KWCS22]. Placement [MFU^{+}20]. Placing [LW15]. Plain [AWB^{+}23]. planning [CMS^{+}11]. Plans [KKH^{+}13, Roo13]. Plant [DVHZ^{+}21]. Plants [FKGW22]. Platform [SVDM17, SNT^{+}23]. Platforms [MSH^{+}23, VTS^{+}04]. Play [BSW17, LWA^{+}23, NBK24, SED^{+}16, SLY^{+}18, SG21, TLA^{+}19]. Players [DPL^{+}23]. PLIERS [CKK^{+}21]. point [Ink01]. point-and-click [Ink01]. pointer [VMGS94]. Pointing [LMN24, MWVK21, MOMS17, NPCBL15, STKB12, GB05, Hor01, HBDG04]. Pointing-Based [LMN24]. Pokémon [DPL^{+}23]. Policies [FMSS17]. Political [BSG18, BHH^{+}24]. Politics [NBB20]. Polylogues [WOB22]. Polynomial [CVC12]. Ponty [Sva13]. Populations [SG21]. portal [NDSG06]. Possessions [OTV19]. Possibilities [MMH23]. Post [LGC17, MA23, PJL^{+}16, SKN24]. Post-growth [SKN24]. Post-Roe [MA23]. Post-Training [LGC17]. Post-WIMP [PJL^{+}16]. Posted [HSI^{+}20]. Postgraduate [HRW^{+}23]. Potential [PL14, TDH10]. Power [BBC^{+}22, ERT23, NBK24]. Powered [CNG^{+}24, SED^{+}16, XZL^{+}20]. PPE [MHL^{+}23]. Practical [KTBR15a].

Practice [JRB^{+}24, NKDB23, RAT^{+}22, SP21, VBR20, DRW13, GFC13, KS10, MAVR13, PSSB13, TBN^{+}13, TKH11, WZ97]. practice-oriented [DRW13, PSSB13]. Practices [ERL^{+}23, JSJ^{+}21, JPM^{+}20, KBJ^{+}13, KdJvE13, PF18, PF21, PBBJS22, PPS^{+}22, RHK^{+}24, SHSS19, SWZ16, SN17, XGA^{+}22, BB09, RSK04, SSRW13, WDHM13]. Practitioner [NKDB23]. Practitioners [Ree19]. Pragmatism [BFG^{+}23]. Praxis [BFL^{+}23]. Prayer [SLM^{+}20]. Pre [GFK^{+}17]. Pre-Selection [GFK^{+}17]. preattentive [HBE96]. Precise [DVK18, TPM^{+}21]. Predictable [SRE^{+}21]. Predicting [BKJ18, CDT^{+}21, EDT^{+}23, FHA^{+}05, JBH13, Jun16, LHF^{+}22, PKRR15, RP96]. Prediction [GPH^{+}23]. Predictions [PJL^{+}16]. Predictive [KWM97, KFF^{+}23, DBT^{+}12]. Predicts [VJR24]. Preference [GSX^{+}23, WLB15, CP10]. preference-based [CP10]. Preferences [FMSS17, HSI^{+}20, LHF^{+}22, JN96, WFD98, YLZ14]. Prefix [LQZ23]. Prefix-Based [LQZ23]. Pregnancy [ALDR^{+}22]. Preliminary [SP24]. preschool [HBDG04]. Presence [TLY^{+}24, HR02, QB05, SRGS00, SUS95]. Present [BOMM18, AM00, MHP00]. Presentation [EYK^{+}16, JN96, LA94]. Presentations [vBBSB23]. Presenting [BTS^{+}13]. preservation [SJZ^{+}98]. preserve [NGB06]. Pressure [KLK^{+}23]. Presto [DELS99]. Preterm [HCH^{+}14]. Prevent [KB23]. Prevention [OAV^{+}16]. Priming [KTBR15b]. primitives [IW03]. Principal [LWLL22]. Principal-Agent [LWLL22]. Principle [BFG^{+}23]. Principles [AOB^{+}20]. Printers [LBP17]. Priorities [RBH24]. PRISM [FKK07]. Privacy [ANK^{+}23, And20, BB18, CFLC23, CKKL18, CLG20, DFH^{+}21, FMSS17, HSI^{+}20, JPR^{+}19, JSJ^{+}21, KB23, KC23, LK20, LWLL22, MA23, RC22, WLB15, BG05, CGA06, IA08, NGB06]. Privacy-aware [CFLC23]. Proactive [DLBBS24, MMS^{+}08]. probabilistic [GB05]. Probe [BBC^{+}22]. Problem [GPPD^{+}22, VPW^{+}15, WBW^{+}23, FKGB10, OCM^{+}12]. problem-solving [FKGB10]. Problems [AJSW12, FLCT19, GSS^{+}15, RSK04]. Procedure [BMB^{+}13]. Process [CKK^{+}21, EZC24, HK99, HF96, SGL09]. Processes [AGWF19, USK^{+}23, BPW12, EK00]. Processing [AWB^{+}23, HBE96]. Processor [LFT06]. Produce [Ree19]. Product [GWZC23, LB10]. Production [WBFDK21]. Productive [CRM17]. Products [OTV19, SNT^{+}23]. Professional [HGOZ19, HBJP24]. Profiles [HWC^{+}16]. Program [Søn20, DC95]. Programmers [PRD^{+}24]. Programming [BWR^{+}17, BSR^{+}23, CKK^{+}21, CKS16, GSS^{+}15, LLZ14, RK22, HFB09, RC96]. programs [MCM97]. Progressive [GSX^{+}23]. Project [HLEG18, OWOZ17, WMRW16, CMS^{+}11, NM09]. Project-Based [HLEG18]. Projection [NHGC16]. Projects [CLY^{+}23]. Promises [TM24]. Promote [BTS^{+}13, HFB09, SP21]. promoting [HWSB99, PCH^{+}06]. Promotional [RMT^{+}15]. Properly [VS14]. proposed [MO94]. proposition [TkWSR99]. Proprioceptive [HGH^{+}24]. Protect [PFGE24]. Protecting [LWLL22]. Protection [ZLM^{+}24]. Proto [SHSS19]. Proto-Practices [SHSS19]. Prototype [HFP12, ZCX^{+}24]. Prototypes [MREN^{+}22, LST08, TIG09, LST08]. Prototyping [BFG^{+}23, LMMBL19, MFPB24, MMH^{+}22, SVDM17, XRL^{+}22, DGK^{+}10, Sal09]. Provide [MWS18, Bre98]. Providing [KGZ07, KKH^{+}13, KSR14, LEF^{+}00]. proximity [Coh97]. Psychological [BPL^{+}23]. Psychophysiological [KFG15]. Public [CALH^{+}19, MLC^{+}13, MREN^{+}22, Pil18, CDF^{+}05, ZS06]. publics [LCHD11]. Pupillary [ZSC^{+}15, MJV^{+}06]. Purpose [SG21]. pursuit [RSK04]. Pushed [SHG^{+}24]. Putting [Edw05, LBGC24, OHM^{+}13]. Puzzle [ZR21].

QN [LFT06]. QN-MHP [LFT06]. quadriplegic [SC03]. Qualitative [GCC^{+}24, OAV^{+}16]. qualities [Rul08]. Quality [LRP15, RHK^{+}24, SHG^{+}24, HSD08]. Quantifying [CL14, HSA^{+}23]. quantization [CD11]. Quarantined [CJBG22]. Quasistatic [MYR^{+}20]. querying [HF96]. Question [CDC15]. Questionnaire [TLY^{+}24]. Questions [RHK^{+}24, XZL^{+}20]. Queueing [LFT06]. Queues [ZRT^{+}20]. Queuing [WL08]. Quiet [GBH^{+}18]. QuintEssence [BBC^{+}22]. Qwerty [SYS19].

R [UB19]. Race [USK^{+}23]. Radar [HGH^{+}24]. RadarHand [HGH^{+}24]. Radio [WBFDK21, SS00]. Rainforest [DKT^{+}21]. Rapid [MHL^{+}23, Sal09, SVDM17]. RaPIDO [SVDM17]. rate [WB94]. Rating [EDT^{+}23, VL07]. Re [BFL^{+}23, GLZH20, LCHD11]. Re-Dictation [GLZH20]. Re-encountering [LCHD11]. Reaching [WB94]. reactable [XHM^{+}13]. Reactions [MZL^{+}23, TGG24, YRA20]. Readily [HGOZ19]. Reading [CWM^{+}23, HF03, KLK^{+}23, LJPS21, MBHC17, WBD^{+}22, CGS12]. Real [JNR^{+}24, JPM^{+}20, SSH22, GG99, RG96, SJZ^{+}98, SC02, SXS^{+}06, WR99]. real-time [GG99, RG96, SJZ^{+}98, SC02, SXS^{+}06]. Real-World [JPM^{+}20, SSH22, JNR^{+}24]. Realism [TGG24]. Realistic [BKJ18].

Realities [VWKL^{+}24]. Reality [DVK18, FJM24, GSS^{+}19, Har19, HFK^{+}22, JNR^{+}24, KSB^{+}24, LCC23, MWVK21, MBP^{+}22, MMHM23, MMH23, PSD24, RMP^{+}21, SED^{+}16, SG20, UK21, VAAB^{+}16, VAF17, ZRT^{+}20, vBBSB23, BGR^{+}98, CNE^{+}07, Dee95, OSF95, SUS95]. Reality-Based [GSS^{+}19]. Really [AKTB19, GCS23, PVMK24]. realtime [BBS01b]. Reasoning [MM17, OCM^{+}12]. Rebuilding [AV19]. Recall [BLNH21, CDC15, GRR20]. recallability [DSG09]. recipes [SHC05]. Reciprocal [ACPL15, BVR15, NVR^{+}14]. Reciprocity [AHCF18]. Recognition [FMP19, LG04]. recombinant [ENSS09]. Recommendation [GRKB16, JDV^{+}21, LB10, LMG^{+}11]. Recommendations [Bro12]. Recommender [GWZC23, KWB^{+}15, CP10, RD05]. Reconciling [COFH16]. recorded [Aro97]. Recourse [KKJ24]. Recov [UB19]. Recov-R [UB19]. Reddit [CJBG22, JBGB19]. Reduce [LRS19, SB18, UB19, VH01]. Reducing [LMMBL19, ZRT^{+}20]. Reference [CNG^{+}24]. Referent [PSD24]. Reflecting [Men22, PF21]. Reflection [BCGVP24, DLBBS24, HM23]. Reflections [BFG^{+}23, KH23]. reflective [Dou95]. Refugees [AV19]. Region [DTP^{+}23]. Regions [KB21]. Regulated [NSCR19]. Regulation [ICC^{+}22, JBGB19, NSCR19, PAM^{+}22, SAT^{+}23]. Rehabilitation [JPM^{+}20, OSBB^{+}19]. Rehearsal [LV20]. Rehearsal-based [LV20]. reification [Bla06]. Reimagining [ABBH20, BBB18, BPW12, FMEB20, KPWS20]. Reinforcement [HSA^{+}23, SVCB21]. related [MRF09, THA99]. Relation [JRB^{+}24]. Relations [RT18, LDS^{+}11]. Relationship [PKES22, Hay11]. Relationships [HHE^{+}12, KLK^{+}23, BP05]. relative [HWSB99]. Relevant [THL^{+}23]. Reliability [HG24, GR11, NPLB09]. Reliance [KYZ23]. Relief [RV95]. remains [KS10]. Remedy [PFGE24]. Remote [BPG^{+}22, RMP^{+}21, WKK^{+}23, XRL^{+}22, CNE^{+}07]. Remotely [RMP^{+}21]. Rendezvous [HBR^{+}94]. replicated [BRS99]. Report [SWZ^{+}24]. Reporting [PPS^{+}22, RHK^{+}24]. repositories [RP96]. Representation [DFH^{+}21, GDW^{+}24, HK99, KLS95, MCM97]. Representations [AZS16, HFLH^{+}23, KHM20]. Represented [OSBB^{+}19]. Reproducibility [PPS^{+}22]. Requirements [KSCB21, MSH^{+}23, MO94, VH01]. Research [AWB^{+}23, BGC^{+}13, BV20, BAZ24, CDC15, CALH^{+}19, DFH^{+}21, DBDK18, EDT^{+}23, FKGW22, GSS^{+}19, HRW^{+}23, HDF21, JRB^{+}24, LJPS21, LMK21, MLH21, VSS^{+}23, RBH24, SEA22, SAT^{+}23, SFKF19, SG21, TM24, VPW^{+}15, AM00, CMS^{+}11, Hay11, HHK00, HL12, SBSS12, TM05]. Researchers [YCVV23]. Reshaping [PBBJS22]. Resharing [BHH^{+}24]. Residential [JKS18, MPS22]. Resistance [Gub23]. resolution [JH14]. resolve [SM11b]. Resonant [MFPB24]. Resource [ERL^{+}23, PRJ16, Pil18, SHSS19]. Resource-Constrained [PRJ16]. Resource-Scarce [Pil18]. resources [Bec04, THA99]. Respond [YTL^{+}23]. Responding [AF18]. response [TKH11]. Responses [AHCF18, VBHK10]. Restorative [XJS23]. Results [OAV^{+}16, DGK^{+}10, SWM03]. Resuscitation [SMB12]. retail [LB10]. Rethinking [HDF21, RTT19]. Retrieval [DS08, SLY^{+}18]. Retrospection [WBFDK21]. Retrospective [Her24, HDF21]. Reuse [PPS^{+}22, RC96, Sut00]. Reveal [LVG^{+}14, MMH23, LDF12]. Revealing [RSC^{+}24]. Reverse [WBW^{+}23]. Review [BV20, Bro12, BAZ24, DFH^{+}21, Her24, HH17, JSJ^{+}21, MMHM23, RZK24, MD23, RHK^{+}24, SFKF19, TBD20, Tsa18, LA94]. reviewers [Ano08, Tra04, ACM03]. Reviewing [BMB^{+}13]. Reviews [RHK^{+}24]. Revisiting [BEJM14]. Reward [BFC12]. Rewards [CVZBB20]. Rhetoric [GCS23, JNR^{+}24]. rich [MBB07]. Rigour [STB21]. Risk [BB18, DFH^{+}21, HRW^{+}23, MBN^{+}23, UB19]. Risks [CLG20]. Rituals [PL14, SWZ16]. Rivalry [KSK02]. Robin [BP23]. Robot [ANO19, ICC^{+}22, MSM23, NSP^{+}18, PH23, SRE^{+}21, KLMC12]. Robotics [BBT^{+}23]. Robots [SRE^{+}21, SLBB19]. Robust [CDT^{+}21, LG04]. Robustness [BLNH21]. Roe [MA23]. Role [AGWF19, BM24, EZC24, GB19, Gil19, MWB16, SKW01, TVH16, BhHSS00, KLMC12, Mac99]. Roleplay [JRB^{+}24]. roles [SHCP08]. room [GPP99]. Rotating [WR99]. rotational [WST14]. Routine [NEE^{+}24]. rule [KLS95]. Rules [GMPS17, WL21]. run [TNB^{+}95]. run-time [TNB^{+}95]. Running [RMP^{+}21]. Rural [Bid21, DVHZ^{+}21, KB21, RBC^{+}21, Wan21, WSO16, BSM^{+}13]. Rurality [SHVH^{+}21]. Rushing [LBGC24].

S [SYK23]. S-BAN [SYK23]. SADIe [HB07]. Safe [KTN^{+}18]. safer [Mac99]. Safety [CFH^{+}20, MM21, RC22, FPST99, GSM99, GPP99]. safety-critical [FPST99, GSM99, GPP99]. SAK [MF10]. Same [BLNH21, LCK24]. Satchel [LEF^{+}00]. satisfaction [DCO13]. Saviourism [Ada23]. scaffolding [SKW01]. Scalability [PWG18, NPLB09]. Scalable [SHS24, BYS02a, BYS02b]. Scale [BRK15, GSS^{+}15, GBH^{+}18, KS15, KWB^{+}15, LBO^{+}15, YLR21, CD11, CCO14, RVB11]. scales [VL07]. Scanning [MF10]. Scarce [Pil18]. Scarcity [ERL^{+}23, HSA^{+}23]. Scenario [HHG^{+}22]. Scenario-Based [HHG^{+}22]. Scenarios [FMSS17, LH08]. Scene [FJM24]. Scene-Viewing [FJM24]. scenes [HWSB99]. scent [KB03, PCV03]. ScentTrails [OC03]. schematic [LCK24]. Scheme [GRG18]. Schemes [MZR^{+}21]. School [ANO19, VWKL^{+}24, WL21]. Science [WWD^{+}22, WL21]. Scientific [KQH^{+}22, SWM03]. Scientists [CRL^{+}23]. Scores [TLY^{+}24, FS04]. Scottish [RSC^{+}24]. Screen [JWS12, RRC^{+}22]. Screen-Based [RRC^{+}22]. Seamless [AO11, BSR^{+}23]. Search [AKTB19, KQH^{+}22, BHNG05, Hor01, KB03, PCV03, WWHW97]. searching [OC03]. Seated [MKFB20]. Second [Wan21]. section [RD05]. Sector [BV20, Pil18]. Secure [MWVK21]. Security [DFH^{+}21]. see [BCF^{+}06, QO13]. Seeing [BJC^{+}23, CFLC23, SLP22]. Seekers [AV19]. seeking [Bec04, SHCP08]. Segmentation [TPM^{+}21]. Selecting [VCN^{+}17]. Selection [CDT^{+}21, FSMP22, GFK^{+}17, OFLK17, PSD24, UK21, WL97, WWH19, GMW05, RM00, SS94]. selective [Ber94]. Self [BCGVP24, BA24, DWL^{+}24, GFK^{+}17, MLH21, MLH14, NEE^{+}24, NSCR19, NVF^{+}15, PCL^{+}20, PAM^{+}22, RPFMP17, MD23, TM24, YTL^{+}23, DGK^{+}10, KWL^{+}13]. Self-Assessments [GFK^{+}17]. Self-Care [NVF^{+}15]. Self-control [MD23]. Self-Declare [DWL^{+}24]. Self-Determination [TM24]. Self-diagnosis [YTL^{+}23]. self-efficacy [DGK^{+}10]. Self-Monitoring [NEE^{+}24]. Self-Observation [MLH21]. Self-Regulated [NSCR19]. Self-Regulation [PAM^{+}22]. Self-Tracking [BCGVP24, RPFMP17]. Selling [WSWL23]. Semantic [FKKH10, PDR17, MTDM14, YSHG07]. Semantics [JKVA17, HB07]. semitransparency [ZBM96]. Sending [CVZBB20]. senior [ZS06]. Seniors [UB19]. Sense [BKPH22, CRL^{+}23, Hor16]. sensed [BSK^{+}05]. Sensemaking [FMEB20, HGOZ19, GBW^{+}12]. Sensing [CDT15, CDT^{+}21, SSHM^{+}23, BSK^{+}05, ZB05, KOP^{+}10]. sensing-based [BSK^{+}05, ZB05]. Sensitive [AF18, BT15, LMW^{+}20, LG04]. Sensitivities [HFLH^{+}23]. sensor [HPHS05, LV09]. sensor-based [LV09]. sensor-enhanced [HPHS05]. sensors [FHA^{+}05]. sensory [PMM^{+}13]. Sentiment [WWH19]. separability [JSM^{+}94]. Serendipitous [DS08]. Service [CALH^{+}19, FLvA24, PSS09]. service-oriented [PSS09]. Services [BPL^{+}23, DCO13]. set [MO94, ORRH99]. set-top-box [ORRH99]. Sets [RSC15]. Setting [LRS19, MM21, NWH21]. Settings [JPM^{+}20, KBJ^{+}13, VWKL^{+}24, WLB15]. setup [LH08]. Sexual [AHCF18, SLBB19]. Shape [BBT^{+}23, KSG^{+}22, MMH^{+}22, SYK23]. Shape-Change [KSG^{+}22]. Shape-Changing [SYK23]. Shape-Memory [MMH^{+}22]. Shared [NSP^{+}18, NVPE19, PCH^{+}06, AEF^{+}00, BhHSS00, BGR^{+}98, PGG03]. shared-workspace [PGG03]. Sharing [BQDB13, ERL^{+}23, FMSS17, FBE23, MLH14, NPF^{+}15, TSTH17, BRS99, LCHD11]. Shelfie [KHM20]. shell [MCM97]. Shifting [DSG09, HFLH^{+}23, JKS18]. Shifts [SG20, ZSE^{+}12]. Shortcut [LMW^{+}20]. Shortens [BBC20]. Should [YTL^{+}23, Dou13]. shows [IA08]. Shumin [Zha15]. Sighted [XRL^{+}22]. Signals [BKJ18, PPS^{+}22]. signed [KGZ07]. Signing [PRB^{+}11]. Similarities [TRZ15, ZS06]. Similarity [PDR17, VBHK10]. Simple [ABY17]. simulated [CNE^{+}07]. Simulating [KFF^{+}23]. simulation [TKH11]. single [SXS^{+}06]. single-user [SXS^{+}06]. Site [Gub23, KB03]. site-specific [KB03]. Sites [Bro12, RC22, KB03, VL07]. situ [CCO14]. Situated [BPG^{+}22, Roo13]. Situation [RLP14]. sketchify [OM11]. Sketching [OM11, STB21, Dee95]. sketching/animation [Dee95]. Skills [OSPK23, PPAA^{+}18, SF15]. skimming [Aro97]. Skin [HGH^{+}24, MMH^{+}22, SHSS19]. slate [CGS12]. Slices [Jun16]. slides [JN96]. SlideSpace [EYK^{+}16]. Slippery [DT23]. small [CD11]. Smalltalk [RC96]. Smart [ACR^{+}16, DAM17, GC22, GHB^{+}22, JPR^{+}19, MWB16, ZSNP14, MTDM14]. Smartphone [AYK23, LHF^{+}22]. Smartphones [LMW^{+}20]. Smartwatch [ACS^{+}23, NEE^{+}24]. Smell [BBC^{+}22, MCD^{+}20]. Smiles [TCD^{+}21]. sociability [MKP05]. Sociable [LBP17]. Social [ABL05, AV19, ANK^{+}23, AF18, AHCF18, And20, BA24, BV20, BAZ24, CKKL18, CLG20, DLL24, DVHZ^{+}21, EK00, EZC24, FPD^{+}16, GC22, GB19, HLEG18, HLN^{+}14, LVG^{+}14, MFPB24, MGVE17, MPS22, PCL^{+}20, PF18, PPAA^{+}18, RT18, RC22, SF15, TM05, TCD^{+}21, WSO16, ATH^{+}03, BSM^{+}13, FKKH10, LWG^{+}14, MMS^{+}08, SM11b, SHC05, WJN^{+}04, HDM11, LRP15]. Socially [ICC^{+}22, SP24]. Socially-Enabled [SP24]. society [HC06]. Socio [CDF^{+}05, GPH^{+}23, KSJB11, SP24, PKHD09]. Socio-Cognitive [SP24, KSJB11]. socio-physical [PKHD09]. Socio-technical [CDF^{+}05, GPH^{+}23]. Sociocultural [KB21, XPL23]. Sociotechnical [ZLD^{+}23]. Soft [FSMP22, SYS19, Rul08]. Software [JBD^{+}22, JDM99, JN96, LMG^{+}11, MBB07, Mye95, MHP00, OSF95, TNB^{+}95, WZ97]. solar [BSM^{+}13]. Solutionism [Ada23, CBRT23]. Solutions [GSS^{+}15]. Solving [WBW^{+}23, BBS01a, FKGB10, OCM^{+}12]. Soma [HBT^{+}21, STB21]. Some [Sva13, HK99]. Sometimes [LBGC24].

SonicAIR [BVR15]. Sonification

[WBFDK21, ZPSL08]. SOS [BAAL^{+}16]. Sound [SHG^{+}24]. Sounds [ACDL12, Bre98]. source [MS94]. source-level [MS94]. Space [ACM18, CFLC23, GSK22, HL21, ICC^{+}22, MCD^{+}20, PCL^{+}20, YCVV23, ASHM97, AO12, BG05, DRD^{+}00, MCD^{+}20]. Spaces [ALDR^{+}22, BCRS16, CALH^{+}19, IIY^{+}24, SED^{+}16, BGR^{+}98, DELS99, OKP11]. Sparse [DVHZ^{+}21, SSHM^{+}23]. SparseIMU [SSHM^{+}23]. Spatial [KGYQ15, KZZ06, SYS19, VAF17, TGSP06]. Spear [LCE^{+}19]. Spear-Phishing [LCE^{+}19]. Special [ABBH20, ANW^{+}23, CCG^{+}13, DFH^{+}15, DTL^{+}21, Hin17b, NWH21, PBBJS22, BDR00, DRW13, GPP99, HC06, MAVR13, OSF95, PHJ08a, PHJ08b, PSSB13, RD05, SY97, SHM07, SJGL09]. Special-Education [NWH21]. Specialists [BP23]. Specific [DAM17, VML15, JN96, KB03]. specification [JDM99, SJ09]. specifying [VM95]. spectacle [BCF^{+}11]. spectator [BCF^{+}11]. spectrum [RV95]. speculation [WOB22]. Speculative [FH23, RAT^{+}22]. Speech [GCS23, JWS12, Aro97, FS04, GEF98, LG04, ODC04, QMB^{+}02, SS00, SY97, SMW01, TSGK14, QB05]. Speech-like [GCS23]. SpeechSkimmer [Aro97]. Speed [GRR20, GRG18, WBD^{+}22, HBE96]. Speed-Accuracy [GRG18]. speller [DBT^{+}12]. Spindex [JWS12]. Split [SS94]. Sport [RT18]. Sports [KCL^{+}16, SLY^{+}18, OPL10]. Spotlight [Hin16c, Hin16a, Hin16d, Hin16e, Hin16f, Hin16g, Hin17e, Hin17a, Hin17b, Hin17c, Hin17d, Hin18a, Hin18b, Hin18c]. spreadsheet [BG98, BYS02a, BYS02b]. Spreadsheets [CAP24]. Squash [DVHZ^{+}21]. staging [DH08]. Stakeholder [BV20]. States [MA23]. static [MWW06].

Stationary [TRZ15]. Statistical [BTS^{+}13]. steps [SUS95]. stereo [HWSB99]. Stigma [And20]. Still [BPG^{+}22, BEJM14, HH23]. Stock [ZLD^{+}23]. Stolen [SBSG23]. Stop [GPPD^{+}22]. Stop-And-Think [GPPD^{+}22]. storage [VH01]. stores [LB10]. Stories [BQDB13]. StoryKit [BQDB13]. Storytelling [BQDB13]. Strain [HFK^{+}22]. Strands [ALR20]. strange [BBS05, LR13]. Strategies [HHE^{+}12, JMP^{+}17, VPW^{+}15, WSWL23, FKGB10, GBW^{+}12, Rie96]. Strategy [BPR08]. Strategy-based [BPR08]. stream [WM06]. Stress [YCVV23]. Stressed [CFH^{+}24]. Strictly [MYR^{+}20]. string [PDR17]. Strips [BKJ18, Mac99]. Stroke [TRZ15, TIG09]. Strong [HL12]. Structural [ERT23, HB07]. structure [IB10, RO97]. structure-based [RO97]. Structures [PMS24, GR11, IW03, JH03]. Structuring [EZC24]. Struggle [DAK20, FMP19]. Student [GSS^{+}15, WL21, YCVV23]. Students [CLY^{+}23, DKA^{+}15, HRW^{+}23, CL08]. Studies [RMP^{+}21, VW22, MB05, PG94]. Studio [HFP12]. Studio-Based [HFP12]. Study [ALR20, AATC22, ASJB22, BQDB13, BBT^{+}23, BBC^{+}22, CVC12, CHAN20, CKS16, DBDK18, HHG^{+}22, HFP12, ICC^{+}22, JKS18, LJPS21, LVDA23, LA17, MLH21, MSD^{+}21, NEE^{+}24, NVR^{+}14, OAV^{+}16, PL14, WKK^{+}23, WSWL23, YRA20, ZPSL08, ZSNP14, ASHM97, ABL05, BhHSS00, Bec04, CCO14, CNE^{+}07, DSG09, FPST99, HF96, HFB09, JN96, KdJvE13, KSR14, LZB98, NDSG06, ORRH99, RM00, Rie96, SHMA07]. style [CL08, LBT96, SKW01]. Styles [ARK^{+}21, Ink01, RV95]. subjective [BBMT06]. Subjectivities [BTBM24]. Subjunctive [LH08]. Success [HDF21]. Support [AHCF18, BCGVP24, BPL^{+}23, BJC^{+}23, CL14, HLEG18, HRW^{+}23, HCH^{+}14, HLN^{+}14, LGW^{+}19, MSH^{+}23, NWH21, OSBB^{+}19, OFLK17, PDR17, PPAA^{+}18, SHG^{+}24, SLM^{+}20, SWLM22, TBD20, ZCX^{+}24, vBBSB23, CGS12, Edw05, EK00, GG99, KSR14, LH08, Rob05, SLS94, TkWSR99, TNB^{+}95, Vic00]. Supported [MPS22, SW09]. Supporting [BVR15, BWTR12, CFH^{+}20, CMS^{+}11, HGOZ19, Hor16, KRMS21, KGYQ15, MLC^{+}13, MBN^{+}23, RLP14, RBJY00, SRGS00, WKK^{+}23, BRS99, CDF^{+}05, Shn00, LB10, MMS^{+}08, WW05]. Surfaces [CAP24, KPO18]. Surplus [ERL^{+}23]. surrogate [CC13]. Surveillance [SKDM24]. Survey [KPO18, XGA^{+}22]. Surveys [LVDA23, XZL^{+}20]. Susceptibility [LCE^{+}19]. Sustainability [COFH16, PWG18]. Sustainable [KBJ^{+}13, RBB15, DRW13, PSSB13, PMM^{+}13, WDHM13]. Symmetric [RSJ02]. Symmetry [BBD^{+}24, PCH^{+}06]. Symptom [YTL^{+}23]. Symptoms [CDT^{+}21]. Synchronous [BBD^{+}24, MWB16]. Synthesis [JRB^{+}24]. synthesized [KGZ07]. Synthetic [SHS24]. System [ACR^{+}16, BAAL^{+}16, GLX^{+}23, KRMS21, LA17, OAV^{+}16, PVMK24, UB19, WMRW16, WWD^{+}22, ATH^{+}03, Aro97, BBS01b, HLN04, LMG^{+}11, LGHH08, LG12, MCC^{+}04, PRB^{+}11, Rob05, SHC05, TkWSR99]. Systematic [BV20, BAZ24, DFH^{+}21, MMHM23, MD23, RHK^{+}24, TBD20]. Systematizing [DLL24]. Systemic [TSTH17]. Systems [BMDD00, GWZC23, KTN^{+}18, KWB^{+}15, LCK24, NVPE19, RM23, TBD20, BRS99, CP10, DRD^{+}00, Dou95, EJS02, EK00, FS04, GR11, GPP99, HK99, HC06, HDM11, LFT06, LV09, MCSN03, NPLB09, OM11, PK94, RSK04, RM00, RD05, SBSS12, SJZ^{+}98, SC02, UIJ05].

Tabletop [DMOJ18, XHM^{+}13, ZSE^{+}12]. Tabletops [JMP^{+}17, NHGC16]. Tactile [AZS16, KRMS21, PRB^{+}11]. tagging [FKKH10]. Tailored [UB19]. Taking [KKIT20, LBP17, SUS95, ZLD^{+}23, DSG09, KNK^{+}21, Tan07]. Tale [MA23]. Talk [BOMM18, RRC^{+}22]. Tangible [CVZBB20, CJV16, DMOJ18, SSC^{+}16, Red08, SJ09, UIJ05, XHM^{+}13]. Tangled [Red08]. tank [WL97]. Taobao [WSWL23]. Tapping [JWS12]. Target [PD16, CD11]. targets [MB05]. Tariff [ACR^{+}16]. Task [FLvA24, PGG03, TCD^{+}21, HBDG04, JN96, JH03, KWM97, KSJB11, LBT96, TkWSR99, TE12, VK14]. task-centric [TE12]. task-medium [TkWSR99]. task-specific [JN96]. TaskGenies [KKH^{+}13]. TaskRabbit [TSTH17]. Tasks [AOB^{+}20, CE20, DLBBS24, JDV^{+}21, KKH^{+}13, LAW18, MWS18, RRC^{+}22, WIKW23, BI08, IB10, JMJ03, KLMC12, PGG03, RM00, TGSP06]. Taxonomy [FJM24, LA94]. Teachers [MSM23]. Teaching [KWB^{+}15, OSPK23, SF15, BPR08, PRB^{+}11]. Team [Jun16, OKP11, TKH11]. Teams [HLEG18]. Teamwork [SMB12]. Tech [BBR18]. technical [CDF^{+}05, GPH^{+}23]. Technique [KGYQ15, OSPK23, JK96b, NPLB09, SUS95]. Techniques [BKG23, DPG22, FJM24, LGC17, SP21, Dou98, JK96a, LA94, TC01, VM95]. Techno [GCS23]. Techno-Masculinity [GCS23]. Technological [OSBB^{+}19, OTV19, PCR15]. Technologies [AFP13, BWLW23, BT15, GC22, GB19, HJS^{+}20, LRS19, LBP17, VSS^{+}23, NSCR19, NVF^{+}15, RTT19, RB23, SP24, SN17, Wan21, BBS05, DSG09, GPE06]. Technology [ACDL12, DKT^{+}21, FPD^{+}16, GSK22, HM23, HRW^{+}23, HHE^{+}12, HH17, KIW16, LJPS21, LGW^{+}19, Men22, MPS22, PLF20, RCFR22, SM11b, SF15, SDBK24, SFKF19, VWKL^{+}24, XLC12, BB09, BPW12, CT07, GR11, LWG^{+}14, ORRH99, OSF95, PMK02]. Technology-Mediated [HM23, HRW^{+}23, KIW16, LGW^{+}19, SM11b].

Technology-Supported [MPS22]. teen [GPE06, ZS06]. Teenager [RC22]. Teenagers [GDW^{+}24, GPE06]. Tele [BPG^{+}22]. Tele-Experiences [BPG^{+}22]. teleconferencing [GB95]. TeleNotes [WSKS97]. Telepresence [NSP^{+}18]. television [BB09, BGC^{+}00, OPL10]. Tell [GBBM12, XZL^{+}20]. Temporal [LBGC24, LKS19]. Tension [MLH14]. Tensions [LMK21, NVF^{+}15]. Term [ASJB22, JKS18, NPF^{+}15, OAV^{+}16, BP05]. terms [BSW08]. Terrain [PVMK24]. Territory [PCL^{+}20]. test [TkWSR99]. Testing [Ree19]. Tests [Her24]. Text [BSG18, BPL^{+}23, BOMM18, DWL^{+}24, DVK18, GLZH20, LMW^{+}20, NKDB23, PDR17, CMH12, MF10, VK14, WST14, WM06, QB05]. Text-To-Speech [QB05]. Textile [PMS24]. Textures [DVHZ^{+}21]. Theatre [SP21]. Their [Bro12, HCH^{+}14, OTV19, VCN^{+}17, DAM17, SSH22]. Them [CVZBB20]. Theme [ZRT^{+}20]. Theoretic [GRG18, MOMS17]. Theoretical [PJL^{+}16, KdJvE13]. Theories [BLBM21]. Theorizing [BA24]. Theory [BFL^{+}23, DS08, DLL24, SAT^{+}23, TBN^{+}13, TM24, VBR20, ZLM^{+}24, HSD08, MAVR13]. Theory-Based [DS08]. Therapy [CFH^{+}24, LGW^{+}19, NEE^{+}24, SWZ16, LWG^{+}14]. Therapy-Inspired [CFH^{+}24]. There [BKPH22, DV18, NSP^{+}18, GPP99, KLMC12]. Thin [Jun16]. thing [TDKS19]. Things [CFLC23, GWZC23, Hin17b, PVMK24, RBK19, HR02, MTDM14, ABY17, LBP17, MNPP17]. Think [FLCT19, GPPD^{+}22, Kir19]. Think-Aloud [FLCT19]. Thinking [DLBBS24, HL21, Her24, WL21, FH08]. Threat [MA23]. Three [CFH^{+}20]. TickleFoot [EBNM22]. Tickling [EBNM22]. tiered [WLB09]. TikTok [WSWL23]. Tilting [WOB22]. Time [LA15, OPL10, ZRT^{+}20, GG99, LEF^{+}00, RG96, SJZ^{+}98, SC02, SXS^{+}06, TNB^{+}95].

Timer [GRR20]. TimeToFocus [BBC20]. Timing [AOB^{+}20]. TKS [JH03]. To-and-Fro [Hor16]. TOCHI [Hin16c, Hin16a, Hin16d, Hin16e, Hin16f, Hin16g, Hin16b, Hin17e, Hin17a, Hin17b, Hin17c, Hin17d, Hin18a, Hin18b, Hin18c, Zha14, Zha15]. Together [OWOZ17]. Togetherness [ASJB22]. Token [UIJ05]. Tomorrow [PFGE24]. Too [MM21]. Tool [CWM^{+}23, CNG^{+}24, FLST13, GSS^{+}19, HCH^{+}14, Bla06, Dee95]. toolbox [IA08]. Toolkit [MBP^{+}22, MMH^{+}22, AO11, Dou98, RG96]. Tools [CFH^{+}20, CL14, DAM17, ID20, KB23, PF21, MD23, ZSNP14, Mye95, MHP00]. top [ORRH99]. topically [THA99]. Topics [CLY^{+}23]. TopicShop [ATH^{+}03]. Torso [SG20]. Touch [AGB14, DVK18, FSMP22, GBBM12, HGH^{+}24, JWS12, KPO18, LMW^{+}20, LLZ14, MYR^{+}20, PBBJS22, PBBJ^{+}22, RAT^{+}22, TNLK22, BhHSS00]. Touch-Based [DVK18, FSMP22, HGH^{+}24]. Touching [KSG^{+}22]. touchless [OHM^{+}13]. Touchscreen [GBBM12]. Touchscreen-Based [GBBM12]. Tourist [CEH^{+}22]. town [PKHD09]. Track [LBGC24]. tracked [GSX^{+}23]. Trackers [CVC12, SSC^{+}16]. Tracking [BCGVP24, GSPZ24, KBB^{+}17, MKS19, RPFMP17, WKK^{+}23, BBS01b]. Tracking-Based [MKS19]. Trade [GRR20, JNBF23, TNLK22]. Trade-Off [TNLK22, GRR20]. Trade-off-centered [JNBF23]. Tradeoff [GRG18]. traffic [Mac99]. Train [LBGC24]. Training [KTBR15b, LGC17, MLH21, HFB09]. Trajectories [GHB^{+}22, SLY^{+}18, BCF^{+}11]. Trans [YHS96]. Transactions [HBL22]. Transcending [AEF^{+}00]. transcoding [YSHG07]. Transcription [WL08]. transfer [HFB09]. Transferred [KGYQ15]. Transformation [CALH^{+}19]. Transformations [ABY17]. Transformative [BSW17]. Transient [AOB^{+}20]. Transition [BKG23, Zha15]. Transitioning [LW15, MSH^{+}23]. Transitions [HM23]. Translate [WBFDK21]. Translation [HRB14, ICC^{+}22]. Translations [VBR20]. translucence [EJS02, EK00]. Transmission [GRG18]. transparency [BRS99]. Transparent [SXS^{+}06]. Transportation [DV18, CDF^{+}05]. Trauma [SMB12]. Treatment [THL^{+}23]. tree [PCV03]. Trends [NVF^{+}15, ZS06]. Trial [LGW^{+}19, ORRH99]. Tribal [DVHZ^{+}21]. Tricky [PVMK24]. Trigger [GMPS17]. Trigger-Action [GMPS17]. trilingual [MCC^{+}04]. Trioethnography [HDF21]. Triple [MBN^{+}23]. Troubling [Søn20]. Trust [FLvA24, KLK^{+}23, KC23, LMDT22, PKES22, DHMV14]. trustworthiness [LDS^{+}11]. Trustworthy [PTR23]. tsunami [GR11]. TUI [NBK24]. Tuning [ICM^{+}23]. Turn [BGA^{+}15, CCG^{+}13]. turns [Zha14]. Tutoring [LA17]. TVs [MWB16]. Tweets [BHH^{+}24]. twenty [Zha14]. Twisted [KSG^{+}22]. Twitter [Gub23]. Two [HPPK98, KHA11, MRC^{+}22, STKB12, TLY^{+}24, WOB22, GB05, Hor01, LZB98, MBB07, PG94]. two-dimensional [GB05, Hor01]. Two-Factor [MRC^{+}22]. Two-handed [HPPK98, KHA11, LZB98]. two-interface [MBB07]. Two-Part [STKB12]. Type [HL18]. Typing [AYK23, WL08].

U.S. [FBE23]. UberX [TSTH17]. Ubiquitous [BVL^{+}19, AM00, PSS09]. UIDL [SJGL09, WLB09]. UIs [PJL^{+}16]. UK [JMP^{+}17]. Ultra [NPCBL15]. Ultra-Walls [NPCBL15]. Umbrella [RHK^{+}24]. Un-Paradoxing [KC23]. Uncertainties [DAK20]. Uncommoning [Bid21]. unconstrained [WM06]. Uncovering [HM23, SSRW13]. Underdeveloped [AGM23]. Undergraduates [OWOZ17].

Underserved [DV18]. Understand [ACS^{+}23, CE20, KS15, YLZ14]. Understanding [AATC22, ASL^{+}22, AGB14, AGWF19, BI08, BGR^{+}98, CDT15, CKS16, ERL^{+}23, Gil19, GSS^{+}19, HFK^{+}22, KCL^{+}16, KKJ24, KSCB21, KSJB11, LHF^{+}22, MSA^{+}23, MREN^{+}22, NKDB23, PCL^{+}20, PPS^{+}22, TSTH17, Wan21, AEF^{+}00, CMLS10, LG12, POS^{+}01, KP10, TNLK22]. Undo [Sun02, Ber94]. undoing [PK94]. unequal [VTS^{+}04]. Unexpected [CVZBB20]. Unfulfilled [TM24]. Unhappy [SSS^{+}22]. Unified [DTP^{+}23, FSMP22, KPO18]. unimportant [CMH12]. Unintended [AGB14, KB23]. unit [KdJvE13]. United [MA23]. universal [NM09, PSS09]. unlabeled [Hor01]. Unmasking [NBK24]. Unnecessary [YS20]. Unpack [SSC^{+}16]. Unpacking [BB09, HBT^{+}21]. Unproductive [CFH^{+}24]. Unquestioned [TM24]. Untangling [HH07]. Uptake [JNR^{+}24]. Urban [ACW^{+}19, GBH^{+}18, KOP^{+}10, PKHD09]. Usability [AJSW12, FLCT19, Har19, HWC^{+}16, Her24, MZR^{+}21, PRD^{+}24, Ree19, TCJ01, Bec04, BBMT06, BB02, FH08, GG99, HBP02, HF03, HH07, LG12, LDS^{+}11, MKP05, NPLB09, PS02, PMK02, PGG03, RB11, TSFA07, TM02, WMMS08]. Usable [DFH^{+}21, GEC^{+}09]. Usage [KCL^{+}16, NSP^{+}18, PAM^{+}22, ZS06]. Use [ASJB22, AKTB19, ALDR^{+}22, BAHU24, CRM17, CRH12, ERL^{+}23, FPD^{+}16, GHB^{+}22, ID20, JKS18, KSCB21, LV20, LCK24, PF18, PLF20, SKDM24, BPR08, HR02, HLN04, KB03, Sut00, WZ97]. useful [GEC^{+}09]. User [ANO19, AYK23, AATC22, ANK^{+}23, BM24, BKPH22, BMB^{+}13, BD22, BBT^{+}23, BWR^{+}17, BLNH21, CFH^{+}20, CEH^{+}22, CGL24, CJV16, CHAN20, CKK^{+}21, CGA06, DT23, DLBBS24, FLvA24, GSX^{+}23, HWC^{+}16, HH23, Hin17b, HH17, Hor16, ICM^{+}23, JPR^{+}19, JSJ^{+}21, KPO18, KYZ23, KKJ24, LA15, LCE^{+}19, LMDT22, MNPP17, MZR^{+}21, MRC^{+}22, MTDM14, MKS19, Mye95, PKRR15, PKES22, PJL^{+}16, SSS^{+}22, SSC^{+}16, TNLK22, Thi04, TH15, TVH16, VW22, WM15, WLB15, YS20, ZSC^{+}15, vSHL12, AGZ10, Ber94, CMLS10, DC95, GSM99, GPP99, GBW^{+}12, HSD08, HF96, HBP02, JDM99, JK96a, JK96b, JN96, KZZ06, KLMC12, LB10, LBT96, LCHD11, MCSN03, MB05, MHP00, NPLB09, NM09, Red08, RB11, SJGL09, SJ09, Shn00, SMW01, SXS^{+}06, TNB^{+}95, TC01, VBHK10, WFD98, WLB09, YLZ14, YR12, BB09, RSK04]. User-Centered [CKK^{+}21, PJL^{+}16]. User-Driven [TNLK22]. User-Experience [vSHL12]. user-robot [KLMC12]. Users [ABY17, CE20, DAM17, DWL^{+}24, LHF^{+}22, MM17, OTV19, SED^{+}16, VS14, ZPSL08, ZLM^{+}24, JBH13, KSR14, MPB^{+}11, NDSG06, RBJY00, YSHG07]. uses [RC96]. Using [BQDB13, Bre98, CDT^{+}21, Dou98, FS04, FSMP22, GFK^{+}17, GFC13, GRKB16, JK96b, KFG15, KWS^{+}14, LDF12, MWS18, MYR^{+}20, MWVK21, MRF09, MMH^{+}22, MMH23, OMV17, PRM00, PF21, RMT^{+}15, RRC^{+}22, SSC^{+}16, SAP^{+}15, SYS19, VH01, WBFDK21, XZL^{+}20, ZLM^{+}24, CDF^{+}05, HBE96, HBDG04, IW03, JH03, RL09, SS94, SGL09, SLP22, WL97]. Utilitarian [XLC12]. Utilizing [SLY^{+}18, ZBM96]. Utopias [Bar18]. UX [Ree19].

Vaccinated [MBN^{+}23]. Vaccine [BHH^{+}24]. Valence [TVH16]. Validation [CKKL18, YHS95, YHS96]. Validity [BAAL^{+}16, STB21]. Value [BT15, OTV19, BBMT06, WH01]. Values [XLC12, KS10]. Variable [DVK18, LQZ23]. Variable-Length [LQZ23]. variables [VMGS94]. Variance [SRE^{+}21]. Variation [CEH^{+}22, GSS^{+}15]. Varying [LAW18]. vehicle [Sal09, TSGK14]. Verbalizations [FLCT19]. Verifiable [MZR^{+}21]. versioning [KSR14]. Versus [RRC^{+}22, Hor01, Ink01, JN96, MWW06, PW06, WFD98]. Vertigo [BMM20]. Via [RMP^{+}21, BTS^{+}10, HSA^{+}23, KFF^{+}23, SYK23, ZRT^{+}20]. Vibrotactile [AOB^{+}20, MFPB24, MFU^{+}20]. Video [CVC12, HHG^{+}22, NPF^{+}15, SHD^{+}24, BG05, GM03, LJY^{+}13, NGB06]. Video-Mediated [SHD^{+}24, LJY^{+}13]. Video-Oculography [CVC12]. View [KKIT20, RTT19, JBH13]. Viewing [FJM24, KCL^{+}16, LH08]. Violations [PVSZ24]. Virtual [BCRS16, CVB16, FJM24, Har19, HFK^{+}22, IIY^{+}24, KSB^{+}24, LCC23, MWVK21, MKFB20, MMHM23, MMH23, RMP^{+}21, SG20, SYK23, TGG24, VAF17, BhHSS00, BDR00, BGC^{+}00, Dee95, FH14, FKK07, GR11, GB95, HPPK98, HFH^{+}00, JGH07, OSF95, RSJ02, RL09, SUS95, SGFT06, WR99, Wex95]. Visibility [BA24, BAAL^{+}16]. Vision [DPG22, HLSR24, SLP22, XRL^{+}22]. Visitor [LRS19]. Visual [ABY17, CZH^{+}15, DMOJ18, HLSR24, Hor01, JKVA17, KRMS21, KBB^{+}17, TCD^{+}21, ZPSL08, BNS02, BHNG05, CD11, HT11, HBE96, JGH07, LDS^{+}11, MCM97, MJV^{+}06, MS94, PCV03, PW06]. Visualization [BSG18, IIY^{+}24, CMH12, HWSB99]. visualizations [HF03, YLZ14]. Visualizing [BGK^{+}22, CLY^{+}23, GSS^{+}15, KBJ^{+}13, RSC15, KOP^{+}10, THA99]. Visually [VS14, YSHG07]. Vlogs [HLN^{+}14]. Vocabularies [XGA^{+}22]. Vocabulary [HHHV24]. Voice [AKTB19, BV20, BAHU24, CGL24, DT23, GLZH20, KNK^{+}21, PLF20, RRC^{+}22, QB05]. Voice-Based [GLZH20, KNK^{+}21]. Voices [BBB18, BV20, GBH^{+}18]. Voter [Rob05]. Voter-centered [Rob05]. Voting [MZR^{+}21]. VR [MWB16, RMP^{+}21, WB94, WL97]. vs [KYZ23, PD16, TNLK22].

Waiting [CRM17, ZRT^{+}20]. WaitSuite [CRM17]. Walk [CVZBB20]. Walking [BSM^{+}13, RVB11, RL09, SUS95]. Walkthrough [HFP12]. wall [JH14]. Walls [NPCBL15]. Want [PRD^{+}24]. Wanted [MBN^{+}23]. warning [GR11]. warp [OPL10]. Warping [PD16]. Was [GCS23, Dou13]. wasn’t [Dou13]. waste [GFC13]. Watch [ACS^{+}23, MA23]. Watching [NSP^{+}18]. Wattch [SKDM24]. Wave [CKS16, Fra20]. Way [BPL^{+}23, VMGS94, WMMS08]. Ways [DAK20]. Wearable [PAM^{+}22]. Weave [PMS24]. Web [Bec04, Bro12, CL08, GCS23, HGOZ19, JBH13, KB03, LA15, OC03, SHMA07, SHCP08, TSFA07, THA99, VL07, XPL23]. Web-Based [LA15, XPL23, CL08]. Website [EDT^{+}23]. Websites [HG24, VML15]. Webstrates [BFG^{+}23]. Weird [PRD^{+}24]. Welcome [Hin16b, Zha15]. Well [KIW16]. Well-Being [KIW16]. Wellbeing [BTS^{+}13, BPL^{+}23, KKIT20, MD23]. We’re [RSC^{+}24, MBN^{+}23]. Wheeling [JWS12]. Where [Dou13, SBSG23]. whether [JBH13]. which [JK96b]. while [PRM00]. Who [VTS^{+}04]. whole [KHW95]. whole-hand [KHW95]. Whose [TNLK22]. Wide [CJBG22, JBH13]. Wikipedia [RZK24]. Wild [AYK23, ANW^{+}23, BGC^{+}13, BQDB13, CR13, CCG^{+}13, DTP^{+}23, MMEFN^{+}24, Roo13, SKB^{+}23, ZLD^{+}23]. WIMP [JDM99, PJL^{+}16]. Window [AO12, PW06, TE12]. windowing [AO11]. Windows [BVL^{+}19]. WindowScape [TE12]. WindowWall [BVL^{+}19]. Wire [ABY17, ASHM97]. wireless [PS02]. within [BGC^{+}00, KFG15, MS94]. Without [CMH12, HBP02]. Woman [ABC20]. Woman-Centered [ABC20]. Women [ABBH20, Bid21, KPWS20, KKIT20, NBB20, Søn20, WSO16]. Word [XPL23].

Word-Color [XPL23]. Wording [BHH^{+}24]. words [CMH12]. Work [BBD^{+}24, BP23, CHAN20, EZC24, KB21, LBGC24, PWG18, RPFMP17, TVH16, WIKW23, PVMK24, Bar09, HK99, JH14]. Work-Life [WIKW23]. Worker [GFK^{+}17, SB18, BRS99, Vic00]. Workers [CFH^{+}24, SB18]. Workflow [GLX^{+}23]. Workflows [HBJP24]. Working [HLJ^{+}97]. Working-memory [HLJ^{+}97]. Workload [MWS18, BI08, BHNG05]. Workplace [KBJ^{+}13, LW23]. Workshop [AGWF19]. workspace [GG99, PGG03]. Workspaces [MKFB20]. World [JPM^{+}20, SSH22, JNR^{+}24, JBH13]. Worn [HGH^{+}24]. Woven [PMS24]. Wrist [HGH^{+}24]. Wrist-Worn [HGH^{+}24]. Write [OWOZ17]. Writing [DPG22, SBSG23].

XICE [AO11]. xPath [GLX^{+}23]. XR [KSB^{+}24].

Year [RSC^{+}24]. Years [HMKV19, WOB22]. Young [DBDK18, LGW^{+}19]. Younger [LCK24]. Yourself [XZL^{+}20]. Youth [PAM^{+}22, TJLS21]. Youth-Centered [TJLS21].

Zero [TKH11]. Zero-fidelity [TKH11]. Zhai [Zha15]. Zone [BD22]. Zone-Outs [BD22]. zoom [SZG^{+}96]. zoomable [HBP02]. Zooming [DMOJ18, PW06].

References

Aljaroodi:2022:UIC

[AATC22] Hussain M. Aljaroodi, Marc T. P. Adam, Timm Teubner, and Raymond Chiong. Understanding the importance of cultural appropriateness for user interface design: an avatar study. ACM Transactions on Computer-Human Interaction, 29(6):52:1–52:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3517138.

Almeida:2020:ISI

[ABBH20] Teresa Almeida, Madeline Balaam, Shaowen Bardzell, and Lone Koefoed Hansen. Introduction to the special issue on HCI and the body: Reimagining women’s health. ACM Transactions on Computer-Human Interaction, 27(4): 20:1–20:32, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3406091.

Almeida:2020:WCD

[ABC20] Teresa Almeida, Madeline Balaam, and Rob Comber. Woman-centered design through humanity, activism, and inclusion. ACM Transactions on Computer-Human Interaction, 27(4):27:1–27:30, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3397176.

Adams:2005:SEE

[ABL05] Anne Adams, Ann Blandford, and Peter Lunt. Social empowerment and exclusion: a case study on digital libraries. ACM Transactions on Computer-Human Interaction, 12(2):174–200, June 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Akiki:2017:VST

[ABY17] Pierre A. Akiki, Arosha K.

Bandara, and Yijun Yu. Visual simple transformations: Empowering end-users to wire Internet of Things objects. ACM Transactions on Computer-Human Interaction, 24(2): 10:1–10:??, May 2017. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ardito:2012:EAP

[ACDL12] Carmelo Ardito, Maria F.

Costabile, Antonella De Angeli, and Rosa Lanzilotti. Enriching archaeological parks with contextual sounds and mobile technology. ACM Transactions on Computer-Human Interaction, 19(4): 29:1–29:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Staff:2003:R

[ACM03] ACM Transactions on Computer-Human Interaction staff. 2002 Reviewers. ACM Transactions on Computer-Human Interaction, 10(1):86, 2003. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Antle:2018:ODS

[ACM18] Alissa N. Antle, Leslie Chesick, and Elgin-Skye Mclaren. Opening up the design space of neurofeedback brain–computer interfaces for children. ACM Transactions on Computer-Human Interaction, 24(6): 38:1–38:??, January 2018. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Appert:2015:RDD

[ACPL15] Caroline Appert, Olivier Chapuis, Emmanuel Pietriga, and María-Jesús Lobo. Reciprocal drag-and-drop. ACM Transactions on Computer-Human Interaction, 22(6):29:1–29:??, December 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Alan:2016:TAI

[ACR^{+}16] Alper T. Alan, Enrico Costanza, Sarvapali D. Ramchurn, Joel Fischer, Tom Rodden, and Nicholas R. Jennings. Tariff agent: Interacting with a future smart energy system at home. ACM Transactions on Computer-Human Interaction, 23(4):25:1–25:??, September 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ankrah:2023:MMH

[ACS^{+}23] Elizabeth A. Ankrah, Franceli L. Cibrian, Lucas M. Silva, Arya Tavakoulnia, Jesus A. Beltran, Sabrina E.b. Schuck, Kimberley D. Lakes, and Gillian R. Hayes. Me, my health, and my watch: How children with ADHD understand smartwatch health data. ACM Transactions on Computer-Human Interaction, 30(4): 59:1–59:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3577008.

Alavi:2019:IHB

[ACW^{+}19] Hamed S. Alavi, Elizabeth F.

Churchill, Mikael Wiberg, Denis Lalanne, Peter Dalsgaard, Ava Fatah Gen Schieck, and Yvonne Rogers. Introduction to human-building interaction (HBI): Interfacing HCI with architecture and urban design. ACM Transactions on Computer-Human Interaction, 26(2):6:1–6:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309714.

Adamu:2023:NMS

[Ada23] Muhammad Sadi Adamu.

No more “Solutionism” or “Saviourism” in futuring African HCI: a manyfesto. ACM Transactions on Computer-Human Interaction, 30(2): 21:1–21:??, April 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571811.

Arias:2000:TIH

[AEF^{+}00] Ernesto Arias, Hal Eden, Gerhard Fischer, Andrew Gorman, and Eric Scharff. Transcending the individual human mind — creating shared understanding through collaborative design. ACM Transactions on Computer-Human Interaction, 7(1):84–113, March 2000. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tochi/2000-7-1/p84-arias/

.

Andalibi:2018:RSD

[AF18] Nazanin Andalibi and Andrea Forte. Responding to sensitive disclosures on social media: a decision-making framework. ACM Transactions on Computer-Human Interaction, 25(6):31:1–31:??, December 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Adams:2013:CTB

[AFP13] Anne Adams, Elizabeth Fitzgerald, and Gary Priestnall. Of catwalk technologies and boundary creatures. ACM Transactions on Computer-Human Interaction, 20(3): 15:1–15:??, July 2013. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Annett:2014:EUU

[AGB14] Michelle Annett, Anoop Gupta, and Walter F. Bischof. Exploring and understanding unintended touch during direct pen interaction. ACM Transactions on Computer-Human Interaction, 21(5):28:1–28:??, November 2014. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Axtell:2023:UMM

[AGM23] Benett Axtell, Eleen Gong, and Cosmin Munteanu. An underdeveloped metaphor: The mismatched designs and motivations of digital picture interactions. ACM Transactions on Computer-Human Interaction, 30(2):19:1–19:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569887.

Annett:2019:EUR

[AGWF19] Michelle Annett, Tovi Grossman, Daniel Wigdor, and George Fitzmaurice. Exploring and understanding the role of workshop environments in personal fabrication processes. ACM Transactions on Computer-Human Interaction, 26(2):10:1–10:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301420.

Apitz:2010:FDE

[AGZ10] Georg Apitz, François Guimbretière, and Shumin Zhai. Foundations for designing and evaluating user interfaces based on the crossing paradigm. ACM Transactions on Computer-Human Interaction, 17(2):9:1–9:??, May 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Andalibi:2018:SSR

[AHCF18] Nazanin Andalibi, Oliver L. Haimson, Munmun De Choudhury, and Andrea Forte. Social support, reciprocity, and anonymity in responses to sexual abuse disclosures on social media. ACM Transactions on Computer-Human Interaction, 25(5):28:1–28:??, October 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3234942.

Akers:2012:BEI

[AJSW12] David Akers, Robin Jeffries, Matthew Simpson, and Terry Winograd. Backtracking events as indicators of usability problems in creation-oriented applications. ACM Transactions on Computer-Human Interaction, 19(2):16:1–16:??, July 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ammari:2019:MSI

[AKTB19] Tawfiq Ammari, Jofish Kaye, Janice Y. Tsai, and Frank Bentley. Music, search, and IoT: How people (really) use voice assistants. ACM Transactions on Computer-Human Interaction, 26(3):17:1–17:??, June 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3311956.

Andalibi:2022:LPU

[ALDR^{+}22] Nazanin Andalibi, Ashley Lacombe-Duncan, Lee Roosevelt, Kylie Wojciechowski, and Cameron Giniel. LGBTQ persons’ use of online spaces to navigate conception, pregnancy, and pregnancy loss: an intersectional approach. ACM Transactions on Computer-Human Interaction, 29(1):2:1– 2:46, February 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3474362.

Alavi:2020:FSL

[ALR20] Hamed S. Alavi, Denis Lalanne, and Yvonne Rogers. The five strands of living lab: a literature study of the evolution of living lab concepts in HCI. ACM Transactions on Computer-Human Interaction, 27(2):10:1–10:26, April 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3380958.

Abowd:2000:CPP

[AM00] Gregory D. Abowd and Elizabeth D. Mynatt. Charting past, present, and future research in ubiquitous computing. ACM Transactions on Computer-Human Interaction, 7(1):29–58, March 2000. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tochi/2000-7-1/p29-abowd/

.

Andalibi:2020:DPS

[And20] Nazanin Andalibi. Disclosure, privacy, and stigma on social media: Examining nondisclosure of distressing experiences. ACM Transactions on Computer-Human Interaction, 27(3):18:1–18:43, June 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386600.

Amon:2023:MUC

[ANK^{+}23] Mary Jean Amon, Aaron Necaise, Nika Kartvelishvili, Aneka Williams, Yan Solihin, and Apu Kapadia. Modeling user characteristics associated with interdependent privacy perceptions on social media. ACM Transactions on Computer-Human Interaction, 30(3):40:1–40:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3577014.

Anonymous:1996:AI

[Ano96]  Anonymous.  Author in-

dex. ACM Transactions on Computer-Human Interaction, 3(4):376–377, December 1996. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-4/p376-author_
index/.

Anonymous:1997:AI

[Ano97]  Anonymous.  Author in-

dex. ACM Transactions on Computer-Human Interaction, 4(4):387–388, December 1997. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-4/p387-author_
index/.

Anonymous:2008:R

[Ano08] Anonymous. 2007 reviewers. ACM Transactions on Computer-Human Interaction, 14(4):22:1–22:??, January 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ahumada-Newhart:2019:GSR

[ANO19] Veronica Ahumada-Newhart and Judith S. Olson. Going to school on a robot: Robot and user interface design features that matter. ACM Transactions on Computer-Human Interaction, 26(4):25:1–25:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3325210.

Andersen:2023:ISI

[ANW^{+}23] Tariq Osman Andersen, Francisco Nunes, Lauren Wilcox, Enrico Coiera, and Yvonne Rogers. Introduction to the special issue on human-centred AI in healthcare: Challenges appearing in the wild. ACM Transactions on Computer-Human Interaction, 30(2): 25:1–25:??, April 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3589961.

Arthur:2011:XWT

[AO11] Richard Arthur and Dan R. Olsen, Jr. XICE windowing toolkit: Seamless display annexation. ACM Transactions on Computer-Human Interaction, 18(3):14:1–14:??, July 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Arthur:2012:WBC

[AO12] Richard Arthur and Dan R. Olsen, Jr. Window brokers: Collaborative display space control. ACM Transactions on Computer-Human Interaction, 19(3):17:1–17:??, October 2012. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Asplund:2020:ATP

[AOB^{+}20] Christopher L. Asplund, Takashi Obana, Parag Bhatnagar, Xun Quan Koh, and Simon T.

Perrault. It’s all in the timing: Principles of transient distraction illustrated with vibrotactile tasks. ACM Transactions on Computer-Human Interaction, 27(3):17:1–17:29, June 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3386358.

Alharthi:2021:IEI

[ARK^{+}21] Sultan A. Alharthi, George E.

Raptis, Christina Katsini, Igor Dolgov, Lennart E. Nacke, and Z. O. Toups. Investigating the effects of individual cognitive styles on collaborative gameplay. ACM Transactions on Computer-Human Interaction, 28(4):23:1–23:49, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3445792.

Arons:1997:SSI

[Aro97] Barry Arons. SpeechSkimmer:

a system for interactively skimming recorded speech. ACM Transactions on Computer-Human Interaction, 4(1):3– 38, March 1997. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-1/p3-arons/.

Ackerman:1997:HLF

[ASHM97] Mark S. Ackerman, Brian Starr, Debby Hindus, and Scott D. Mainwaring. Hanging on the ‘wire: a field study of an audio-only media space. ACM Transactions on Computer-Human Interaction, 4(1):39–66, March 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-1/p39-ackerman/

.

Ambe:2022:CHE

[ASJB22] Aloha Hufana Ambe, Alessandro Soro, Daniel Johnson, and Margot Brereton. From collaborative habituation to everyday togetherness: a long-term study of use of the messaging kettle. ACM Transactions on Computer-Human Interaction, 29(1):3:1–3:47, February 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3470973.

Andres:2022:IEU

[ASL^{+}22] Josh Andres, Nathan Semertzidis, Zhuying Li, Yan Wang, and Florian Floyd Mueller. Integrated exertionunderstanding the design of human-computer integration in an exertion context. ACM Transactions on Computer-Human Interaction, 29(6): 55:1–55:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3528352.

Amento:2003:ESD

[ATH^{+}03] Brian Amento, Loren Terveen, Will Hill, Deborah Hix, and Robert Schulman. Experiments in social data mining: The TopicShop system. ACM Transactions on Computer-Human Interaction, 10(1):54– 85, 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Almohamed:2019:RSC

[AV19] Asam Almohamed and Dhaval Vyas. Rebuilding social capital in refugees and asylum seekers. ACM Transactions on Computer-Human Interaction, 26(6):41:1–41:??, December 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3364996.

August:2023:PPM

[AWB^{+}23] Tal August, Lucy Lu Wang, Jonathan Bragg, Marti A. Hearst, Andrew Head, and Kyle Lo. Paper plain: Making medical research papers approachable to healthcare consumers with natural language processing. ACM Transactions on Computer-Human Interaction, 30(5):74:1–74:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589955.

Akpinar:2023:ECS

[AYK23] Elgin Akpinar, Yeliz Yesilada, and Pinar Karagöz. Effect of context on smartphone users’ typing performance in the wild. ACM Transactions on Computer-Human Interaction, 30(3):36:1–36:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3577013.

Azh:2016:IET

[AZS16] Maryam Azh, Shengdong Zhao, and Sriram Subramanian. Investigating expressive tactile interaction design in artistic graphical representations. ACM Transactions on Computer-Human Interaction, 23(5):32:1–32:??, November 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Brotherton:2004:LLE

[BA04] Jason A. Brotherton and Gregory D. Abowd. Lessons learned from eClass: Assessing automated capture and access in the classroom. ACM Transactions on Computer-Human Interaction, 11(2):121– 155, June 2004. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Barta:2024:TSV

[BA24] Kristen Barta and Nazanin Andalibi. Theorizing self visibility on social media: a visibility objects lens. ACM Transactions on Computer-Human Interaction, 31(3):31:1–31:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660337.

Boden:2016:MVV

[BAAL^{+}16] Alexander Boden, Amro Al-Akkad, Michael Liegl, Monika Buscher, Martin Stein, David Randall, and Volker Wulf. Managing visibility and validity of distress calls with an adhoc SOS system. ACM Transactions on Computer-Human Interaction, 23(6):38:1–38:??, December 2016. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Brewer:2024:EVC

[BAHU24] Robin Brewer, Sam Ankenbauer, Manahil Hashmi, and Pooja Upadhyay. Examining voice community use. ACM Transactions on Computer-Human Interaction, 31(2): 24:1–24:??, April 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635151.

Bardram:2009:ABC

[Bar09] Jakob E. Bardram. Activitybased computing for medical work in hospitals. ACM Transactions on Computer-Human Interaction, 16(2):10:1–10:??, June 2009. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

Bardzell:2018:UPF

[Bar18] Shaowen Bardzell. Utopias of participation: Feminism, design, and the futures. ACM Transactions on Computer-Human Interaction, 25(1):6:1– 6:??, February 2018. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Burda:2024:CSE

[BAZ24] Pavlo Burda, Luca Allodi, and Nicola Zannone. Cognition in social engineering empirical research: a systematic literature review. ACM Transactions on Computer-Human Interaction, 31(2):19:1–19:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3635149.

Bodker:2002:DCP

[BB02] Susanne Bødker and Jacob Buur. The design collaboratorium: a place for usability design. ACM Transactions on Computer-Human Interaction, 9(2):152–169, June 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Barkhuus:2009:UTU

[BB09] Louise Barkhuus and Barry Brown. Unpacking the television: User practices around a changing technology. ACM Transactions on Computer-Human Interaction, 16(3):

15:1–15:??, September 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bhatia:2018:EMP

[BB18] Jaspreet Bhatia and Travis D.

Breaux. Empirical measurement of perceived privacy risk. ACM Transactions on Computer-Human Interaction, 25(6):34:1–34:??, December 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bannon:2018:IRP

[BBB18] Liam Bannon, Jeffrey Bardzell, and Susanne Bødker. Introduction: Reimagining participatory design-emerging voices. ACM Transactions on Computer-Human Interaction, 25(1):1:1–1:??, February 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Borghouts:2020:TFI

[BBC20] Judith Borghouts, Duncan P.

Brumby, and Anna L. Cox. TimeToFocus: Feedback on interruption durations discourages distractions and shortens interruptions. ACM Transactions on Computer-Human Interaction, 27(5):32:1–32:31, October 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3396044.

Brianza:2022:QPS

[BBC^{+}22] Giada Brianza, Jesse Benjamin, Patricia Cornelio, Emanuela Maggioni, and Marianna Obrist. QuintEssence: a probe study to explore the power of smell on emotions, memories, and body image in daily life. ACM Transactions on Computer-Human Interaction, 29(6):58:1–58:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3526950.

Bjorn:2024:ASS

[BBD^{+}24] Pernille Bjørn, Juliane Busboom, Melanie Duckert, Susanne Bødker, Irina Shklovski, Eve Hoggan, Kellie Dunn, Qianqian Mu, Louise Barkhuus, and Nina Boulus-Rødje. Achieving symmetry in synchronous interaction in hybrid work is impossible. ACM Transactions on Computer-Human Interaction, 31(4):49:1–49:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3648617.

Ben-Bassat:2006:ESM

[BBMT06] Tamar Ben-Bassat, Joachim Meyer, and Noam Tractinsky. Economic and subjective measures of the perceived value of aesthetics and usability. ACM Transactions on Computer-Human Interaction, 13(2):210–234, June 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bjorn:2018:IIT

[BBR18] Pernille Bjørn and Nina Boulus-Rødje. Infrastructural inaccessibility: Tech entrepreneurs in occupied Palestine. ACM Transactions on Computer-Human Interaction, 25(5):26:1–26:??, October 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Badros:2001:CLA

[BBS01a] Greg J. Badros, Alan Borning, and Peter J. Stuckey. The Cassowary linear arithmetic constraint solving algorithm. ACM Transactions on Computer-Human Interaction, 8(4):267–306, December 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Badros:2001:IPD

[BBS01b] Greg J. Badros, Alan Borning, and Peter J. Stuckey. Integrating paper and digital information on EnhancedDesk: a method for realtime finger tracking on an augmented desk system. ACM Transactions on Computer-Human Interaction, 8(4):307–322, December 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bell:2005:MMS

[BBS05] Genevieve Bell, Mark Blythe, and Phoebe Sengers. Making by making strange: Defamiliarization and the design of domestic technologies. ACM Transactions on Computer-Human Interaction, 12(2):149– 173, June 2005. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Breuer:2023:HEI

[BBT^{+}23] Svenja Breuer, Maximilian Braun, Daniel Tigard, Alena Buyx, and Ruth Müller. How engineers’ imaginaries of healthcare shape design and user engagement: a case study of a robotics initiative for geriatric healthcare AI applications. ACM Transactions on Computer-Human Interaction, 30(2):30:1–30:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3577010.

Bederson:2004:DFC

[BCCR04] Benjamin B. Bederson, Aaron Clamage, Mary P. Czerwinski, and George G. Robertson. DateLens: a fisheye calendar interface for PDAs. ACM Transactions on Computer-Human Interaction, 11(1):90– 119, March 2004. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Benford:2006:CYS

[BCF^{+}06] Steve Benford, Andy Crabtree, Martin Flintham, Adam Drozd, Rob Anastasi, Mark Paxton, Nick Tandavanitj, Matt Adams, and Ju Row-Farr. Can you see me now? ACM Transactions on Computer-Human Interaction, 13(1):100–133, March 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Benford:2011:CSD

[BCF^{+}11] Steve Benford, Andy Crabtree, Martin Flintham, Chris Greenhalgh, Boriana Koleva, Matt Adams, Nick Tandavanitj, Ju Row Farr, Gabriella Giannachi, and Irma Lindt. Creating the spectacle: Designing interactional trajectories through spectator interfaces. ACM Transactions on Computer-Human Interaction, 18(3):11:1–11:??, July 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Barker-Canler:2024:FMS

[BCGVP24] Matthew Barker-Canler, Daniel Gooch, Janet Van Der Linden, and Marian Petre. Flexible minimalist self-tracking to support individual reflection. ACM Transactions on Computer-Human Interaction, 31(3):30:1–30:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3660339.

Benford:2016:LIB

[BCRS16] Steve Benford, Muffy Calder, Tom Rodden, and Michele Sevegnani. On lions, impala, and bigraphs: Modelling interactions in physical/virtual spaces. ACM Transactions on Computer-Human Interaction, 23(2):9:1–9:??, May 2016. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bosch:2022:CCO

[BD22] Nigel Bosch and Sidney K. D’Mello. Can computers outperform humans in detecting user zone-outs? Implications for intelligent interfaces. ACM Transactions on Computer-Human Interaction, 29(2):10:1–10:33, April 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3481889.

Benford:2000:ISI

[BDR00] Steve Benford, Paul Dourish, and Tom Rodden. Introduction to the special issue on human-computer interaction and collaborative virtual environments. ACM Transactions on Computer-Human Interaction, 7(4):439– 441, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-4/p439-benford/p439-
benford.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-4/p439-benford/

.

Becker:2004:SWU

[Bec04] Shirley Ann Becker. A study of Web usability for older adults seeking online health resources. ACM Transactions on Computer-Human Interaction, 11(4):387–406, December 2004. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bjorn:2014:DDS

[BEJM14] Pernille Bjørn, Morten Esbensen, Rasmus Eskild Jensen, and Stina Matthiesen. Does distance still matter? Revisiting the CSCW fundamentals on distributed collaboration. ACM Transactions on Computer-Human Interaction, 21(5):27:1–27:??, November 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Berlage:1994:SUM

[Ber94] Thomas Berlage. A selective undo mechanism for graphical user interfaces based on command objects. ACM Transactions on Computer-Human Interaction, 1(3):269– 294, September 1994. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1994-1-3/p269-berlage/

.

Benbunan-Fich:2011:MMB

[BFAM11] Raquel Benbunan-Fich, Rachel F. Adler, and Tamilla Mavlanova. Measuring multitasking behavior with activity-based metrics. ACM Transactions on Computer-Human Interaction, 18(2):7:1–7:??, June 2011. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Berkovsky:2012:PAM

[BFC12] Shlomo Berkovsky, Jill Freyne, and Mac Coombe. Physical activity motivating games: Be active and get your own reward. ACM Transactions on Computer-Human Interaction, 19(4):32:1–32:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Borowski:2023:BPP

[BFG^{+}23] Marcel Borowski, Bjarke V. Fog, Carla F. Griggio, James R. Eagan, and Clemens N. Klokmose. Between principle and pragmatism: Reflections on prototyping computational media with webstrates. ACM Transactions on Computer-Human Interaction, 30(4): 61:1–61:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3569895.

Bodker:2023:RCH

[BFL^{+}23] Susanne Bødker, Sarah Fox, Nicolas Lalone, Megh Marathe, and Robert Soden. (Re)Connecting history to the theory and praxis of HCI. ACM Transactions on Computer-Human Interaction, 30(2):16:1–16:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589804.

Burnett:1998:GDE

[BG98] Margaret M. Burnett and Herkimer J. Gottfried. Graphical definitions: expanding spreadsheet languages through direct manipulation and gestures. ACM Transactions on Computer-Human Interaction, 5(1):1–33, March 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1998-5-1/p1-burnett/

.

Boyle:2005:LPL

[BG05] Michael Boyle and Saul Greenberg. The language of privacy: Learning from video media space analysis and design. ACM Transactions on Computer-Human Interaction, 12(2):328–370, June 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Benford:2015:EIH

[BGA^{+}15] Steve Benford, Chris Greenhalgh, Bob Anderson, Rachel Jacobs, Mike Golembewski, Marina Jirotka, Bernd Carsten Stahl, Job Timmermans, Gabriella Giannachi, Matt Adams, Ju Row Farr, Nick Tandavanitj, and Kirsty Jennings. The ethical implications of HCI’s turn to the cultural. ACM Transactions on Computer-Human Interaction, 22(5):24:1–24:??, October 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Benford:2000:ITB

[BGC^{+}00] Steve Benford, Chris Greenhalgh, Mike Craven, Graham Walker, Tim Regan, Jason Morphett, and John Wyver. Inhabited television: broadcasting interaction from within collaborative virtual environments. ACM Transactions on Computer-Human Interaction, 7(4):510– 547, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-4/p510-benford/p510-
benford.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-4/p510-benford/

.

Benford:2013:PLR

[BGC^{+}13] Steve Benford, Chris Greenhalgh, Andy Crabtree, Martin Flintham, Brendan Walker, Joe Marshall, Boriana Koleva, Stefan Rennick Egglestone, Gabriella Giannachi, Matt Adams, Nick Tandavanitj, and Ju Row Farr. Performance-led research in the wild. ACM Transactions on Computer-Human Interaction, 20(3):14:1–14:??, July 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bharadwaj:2022:FHC

[BGK^{+}22] Aditya Bharadwaj, David Gwizdala, Yoonjin Kim, Kurt Luther, and T. M. Murali. Flud: a hybrid crowdalgorithm approach for visualizing biological networks. ACM Transactions on Computer-Human Interaction, 29(1):8:1– 8:53, February 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3479196.

Benford:1998:UCS

[BGR^{+}98] Steve Benford, Chris Greenhalgh, Gail Reynard, Chris Brown, and Boriana Koleva. Understanding and constructing shared spaces with mixedreality boundaries. ACM Transactions on Computer-Human Interaction, 5(3):185– 223, September 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1998-5-3/p185-benford/

.

Buschek:2018:PMM

[BHA18] Daniel Buschek, Mariam Hassib, and Florian Alt. Personal mobile messaging in context: Chat augmentations for expressiveness and awareness. ACM Transactions on Computer-Human Interaction, 25(4):23:1–23:??, September 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3201404.

Borghouts:2024:WME

[BHH^{+}24] Judith Borghouts, Yicong Huang, Suellen Hopfer, Chen Li, and Gloria Mark. Wording matters: The effect of linguistic characteristics and political ideology on resharing of COVID-19 vaccine tweets. ACM Transactions on Computer-Human Interaction, 31(4):44:1–44:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3637876.

Basdogan:2000:ESR

[BhHSS00] Cagatay Basdogan, Chih hao Ho, Mandayam A. Srinivasan, and Mel Slater. An experimental study on the role of touch in shared virtual environments. ACM Transactions on Computer-Human Interaction, 7(4):443– 460, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-4/p443-basdogan/p443-
basdogan.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-4/p443-basdogan/

.

Burke:2005:HCB

[BHNG05] Moira Burke, Anthony Hornof, Erik Nilsen, and Nicholas Gorman. High-cost banner blindness: Ads increase perceived workload, hinder visual search, and are forgotten. ACM Transactions on Computer-Human Interaction, 12(4):423– 445, December 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bailey:2008:UCM

[BI08] Brian P. Bailey and Shamsi T.

Iqbal. Understanding changes in mental workload during execution of goal-directed tasks and its application for interruption management. ACM Transactions on Computer-Human Interaction, 14(4): 21:1–21:??, January 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bidwell:2021:RUW

[Bid21] Nicola J. Bidwell. Rural uncommoning: Women, community networks and the enclosure of life. ACM Transactions on Computer-Human Interaction, 28(3):19:1–19:50, July 2021. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3445793.

Boyd:2023:GFA

[BJC^{+}23] LouAnne E. Boyd, Jazette Johnson, Franceli Cibrian, Deanna Hughes, Eliza Delpizzo-Cheng, Karen Lotich, Sara Jones, Hollis Pass, Viseth Sean, and Gillian Hayes. Global filter: Augmenting images to support seeing the “Big picture” for people with local interference. ACM Transactions on Computer-Human Interaction, 30(3):45:1–45:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571812.

Bodker:2018:PDM

[BK18] Susanne Bødker and Morten Kyng. Participatory design that matters-facing the big issues. ACM Transactions on Computer-Human Interaction, 25(1):4:1–4:??, February 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bailly:2023:CMT

[BKG23] Gilles Bailly, Mehdi Khamassi, and Benoît Girard. Computational model of the transition from novice to expert interaction techniques. ACM Transactions on Computer-Human Interaction, 30(5):

66:1–66:??, October 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3505557.

Barral:2018:NNL

[BKJ18] Oswald Barral, Ilkka Kosunen, and Giulio Jacucci. No need to laugh out loud: Predicting humor appraisal of comic strips based on physiological signals in a realistic environment. ACM Transactions on Computer-Human Interaction, 24(6):40:1–40:??, January 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bergstrom:2022:SAU

[BKPH22] Joanna Bergström, Jarrod Knibbe, Henning Pohl, and Kasper Hornbæk. Sense of agency and user experience: Is there a link? ACM Transactions on Computer-Human Interaction, 29(4): 28:1–28:22, August 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3490493.

Benford:2017:CID

[BKQ^{+}17] Steve Benford, Boriana Koleva, Anthony Quinn, Emily-Clare Thorn, Kevin Glover, William Preston, Adrian Hazzard, Stefan Rennick-Egglestone, Chris Greenhalgh, and Richard Mortier. Crafting interactive decoration. ACM Transactions on Computer-Human Interaction, 24(4): 26:1–26:??, September 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Blackwell:2006:RMD

[Bla06] Alan F. Blackwell. The reification of metaphor as a design tool. ACM Transactions on Computer-Human Interaction, 13(4):490–530, December 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Beaudouin-Lafon:2021:GTI

[BLBM21] Michel Beaudouin-Lafon, Susanne Bødker, and Wendy E. Mackay. Generative theories of interaction. ACM Transactions on Computer-Human Interaction, 28(6):45:1–45:54, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3468505.

Bruun:2021:DYF

[BLNH21] Anders Bruun, Effie Lai-Chong Law, Thomas Dyhre Nielsen, and Matthias Heintz. Do you feel the same? On the robustness of cued-recall debriefing for user experience evaluation. ACM Transactions on Computer-Human Interaction, 28(4):25:1–25:45, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3453479.

Bennett:2024:BIM

[BM24] Daniel Bennett and Elisa D.

Mekler. Beyond intrinsic motivation: The role of autonomous motivation in user experience. ACM Transactions on Computer-Human Interaction, 31(5):60:1–60:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3689044.

Borsci:2013:REF

[BMB^{+}13] Simone Borsci, Robert D. Macredie, Julie Barnett, Jennifer Martin, Jasna Kuljis, and Terry Young. Reviewing and extending the fiveuser assumption: a grounded procedure for interaction evaluation. ACM Transactions on Computer-Human Interaction, 20(5):29:1–29:??, November 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Barnard:2000:SIM

[BMDD00] Philip Barnard, Jon May, David Duke, and David Duce. Systems, interactions, and macrotheory. ACM Transactions on Computer-Human Interaction, 7(2):222–262, June 2000. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/

                                        citations/journals/tochi/
                                        2000-7-2/p222-barnard/.

Byrne:2020:DDV

[BMM20] Richard Byrne, Joe Marshall, and Florian ‘Floyd’ Mueller. Designing digital vertigo experiences. ACM Transactions on Computer-Human Interaction, 27(3):19:1–19:30, June 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3387167.

Baldwin:2020:ACA

[BMNH20] Mark S. Baldwin, Jennifer Mankoff, Bonnie Nardi, and Gillian Hayes. An activity centered approach to nonvisual computer interaction. ACM Transactions on Computer-Human Interaction, 27(2): 12:1–12:27, April 2020. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3374211.

Bellini:2002:CVM

[BNS02] P. Bellini, P. Nesi, and M. B. Spinu. Cooperative visual manipulation of music notation. ACM Transactions on Computer-Human Interaction, 9(3):194–237, September 2002. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Brown:2018:TTL

[BOMM18] Barry Brown, Kenton O’Hara, Moira McGregor, and Donald Mcmillan. Text in talk: Lightweight messages in copresent interaction. ACM Transactions on Computer-Human Interaction, 24(6): 42:1–42:??, January 2018. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bickmore:2005:EML

[BP05] Timothy W. Bickmore and Rosalind W. Picard. Establishing and maintaining longterm human-computer relationships. ACM Transactions on Computer-Human Interaction, 12(2):293–327, June 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bossen:2023:BRH

[BP23] Claus Bossen and Kathleen H.

Pine. Batman and Robin in healthcare knowledge work: Human-AI collaboration by clinical documentation integrity specialists. ACM Transactions on Computer-Human Interaction, 30(2): 26:1–26:??, April 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3569892.

Bertran:2022:DTE

[BPG^{+}22] Ferran Altarriba Bertran, Alexandra Pometko, Muskan Gupta, Lauren Wilcox, Reeta Banerjee, and Katherine Isbister. Designerly teleexperiences: a new approach to remote yet still situated codesign. ACM Transactions on Computer-Human Interaction, 29(5):44:1–44:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3506698.

Bhattacharjee:2023:DIO

[BPL^{+}23] Ananya Bhattacharjee, Jiyau Pang, Angelina Liu, Alex Mariakakis, and Joseph Jay Williams. Design implications for one-way text messaging services that support psychological wellbeing. ACM Transactions on Computer-Human Interaction, 30(3):34:1–34:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569888.

Bachynskyi:2015:IDN

[BPOW15] Myroslav Bachynskyi, Gregorio Palmas, Antti Oulasvirta, and Tino Weinkauf. Informing the design of novel input methods with muscle coactivation clustering. ACM Transactions on Computer-Human Interaction, 21(6):30:1–30:??, January 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bhavnani:2008:SBI

[BPR08] Suresh K. Bhavnani, Frederick A. Peck, and Frederick Reif. Strategy-based instruction: Lessons learned in teaching the effective and efficient use of computer applications. ACM Transactions on Computer-Human Interaction, 15(1):2:1–2:??, May 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Buechley:2012:CTR

[BPW12] Leah Buechley and Hannah Perner-Wilson. Crafting technology: Reimagining the processes, materials, and cultures of electronics. ACM Transactions on Computer-Human Interaction, 19(3):21:1–21:??, October 2012. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bonsignore:2013:SSW

[BQDB13] Elizabeth Bonsignore, Alexander J. Quinn, Allison Druin, and Benjamin B. Bederson. Sharing stories “in the wild”: a mobile storytelling case study using StoryKit. ACM Transactions on Computer-Human Interaction, 20(3):18:1–18:??, July 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Brewster:1998:UNS

[Bre98] Stephen A. Brewster. Using nonspeech sounds to provide navigation cues. ACM Transactions on Computer-Human Interaction, 5(3):224– 259, September 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:

                                        80/pubs/citations/journals/
                                        tochi/1998-5-3/p224-brewster/

.

Bederson:2015:IOL

[BRK15] Benjamin B. Bederson, Daniel M. Russell, and Scott Klemmer. Introduction to online learning at scale. ACM Transactions on Computer-Human Interaction, 22(2):5:1–5:??, April 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Brown:2012:BRL

[Bro12] Barry Brown. Beyond recommendations: Local review Web sites and their impact. ACM Transactions on Computer-Human Interaction, 19(4): 27:1–27:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Begole:1999:FCT

[BRS99] James Begole, Mary Beth Rosson, and Clifford A. Shaffer. Flexible collaboration transparency: supporting worker independence in replicated application-sharing systems. ACM Transactions on Computer-Human Interaction, 6(2):95–132, June 1999. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tochi/
                                        1999-6-2/p95-begole/.

Baumer:2018:IIT

[BSG18] Eric P. S. Baumer, Jaime Snyder, and Geri K. Gay. Interpretive impacts of text visualization: Mitigating political framing effects. ACM Transactions on Computer-Human Interaction, 25(4): 20:1–20:??, September 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3214353.

Benford:2005:ESD

[BSK^{+}05] Steve Benford, Holger Schnädelbach, Boriana Koleva, Rob Anastasi, Chris Greenhalgh, Tom Rodden, Jonathan Green, Ahmed Ghali, Tony Pridmore, Bill Gaver, Andy Boucher, Brendan Walker, Sarah Pennington, Albrecht Schmidt, Hans Gellersen, and Anthony Steed. Expected, sensed, and desired: a framework for designing sensing-based interaction. ACM Transactions on Computer-Human Interaction, 12(1):3–30, March 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bidwell:2013:WSL

[BSM^{+}13] Nicola J. Bidwell, Masbulele Siya, Gary Marsden, William D. Tucker, M. Tshemese, N. Gaven, S. Ntlangano, Simon Robinson, and Kristen Ali Eglinton. Walking and the social life of solar charging in rural Africa. ACM Transactions on Computer-Human Interaction, 20(4): 22:1–22:??, September 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Brocker:2023:FHS

[BSR^{+}23] Anke Brocker, René Schäfer, Christian Remy, Simon Voelker, and Jan Borchers. Flowboard: How seamless, live, flow-based programming impacts learning to code for embedded electronics. ACM Transactions on Computer-Human Interaction, 30(1):2:1–2:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3533015.

Boehner:2008:IIM

[BSW08] Kirsten Boehner, Phoebe Sengers, and Simeon Warner. Interfaces with the ineffable: Meeting aesthetic experience on its own terms. ACM Transactions on Computer-Human Interaction, 15(3):12:1–12:??, November 2008. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Back:2017:DTP

[BSW17] Jon Back, Elena Márquez Segura, and Annika Waern. Designing for transformative play. ACM Transactions on Computer-Human Interaction, 24(3):18:1–18:??, July 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Baljko:2008:AEC

[BT08] Melanie Baljko and Nell Tenhaaf. The aesthetics of emergence: Co-constructed interactions. ACM Transactions on Computer-Human Interaction, 15(3):11:1–11:??, November 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Briggs:2015:IVS

[BT15] Pam Briggs and Lisa Thomas.

An inclusive, value sensitive design perspective on future identity technologies. ACM Transactions on Computer-Human Interaction, 22(5): 23:1–23:??, October 2015. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Baumer:2024:AS

[BTBM24] Eric P. S. Baumer, Alex S.

Taylor, Jed R. Brubaker, and Micki McGee. Algorithmic subjectivities. ACM Transactions on Computer-Human Interaction, 31(3):35:1–35:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660344.

Bernstein:2010:PF

[BTS^{+}10] Michael S. Bernstein, Desney Tan, Greg Smith, Mary Czerwinski, and Eric Horvitz. Personalization via friendsourcing. ACM Transactions on Computer-Human Interaction, 17(2):6:1–6:??, May 2010. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bentley:2013:HMP

[BTS^{+}13] Frank Bentley, Konrad Tollmar, Peter Stephenson, Laura Levy, Brian Jones, Scott Robertson, Ed Price, Richard Catrambone, and Jeff Wilson. Health mashups: Presenting statistical patterns between wellbeing data and context in natural language to promote behavior change. ACM Transactions on Computer-Human Interaction, 20(5):30:1–30:??, November 2013. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bopp:2020:VSS

[BV20] Chris Bopp and Amy Voida. Voices of the social sector: a systematic review of stakeholder voice in HCI research with nonprofit organizations. ACM Transactions on Computer-Human Interaction, 27(2):9:1–9:26, April 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3368368.

Bader:2019:WTA

[BVL^{+}19] Patrick Bader, Alexandra Voit, Huy Viet Le, Paweł W. Woźniak, Niels Henze, and Albrecht Schmidt. WindowWall: Towards adaptive buildings with interactive windows as ubiquitous displays. ACM Transactions on Computer-Human Interaction, 26(2):11:1–11:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3310275.

Baharin:2015:SSI

[BVR15] Hanif Baharin, Stephen Viller, and Sean Rintel. SonicAIR: Supporting independent living with reciprocal ambient audio awareness. ACM Transactions on Computer-Human Interaction, 22(4):18:1–18:??, July 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bomfim:2023:DET

[BWLW23] Marcela Bomfim, Erin Wong, Paige Liang, and James Wallace. Design and evaluation of technologies for informed food choices. ACM Transactions on Computer-Human Interaction, 30(4):54:1–54:??, August 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565482.

Brich:2017:EEU

[BWR^{+}17] Julia Brich, Marcel Walch, Michael Rietzler, Michael Weber, and Florian Schaub. Exploring end user programming needs in home automation. ACM Transactions on Computer-Human Interaction, 24(2):11:1–11:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Black:2012:SPN

[BWTR12] Rolf Black, Annalu Waller, Ross Turner, and Ehud Reiter. Supporting personal narrative for children with complex communication needs. ACM Transactions on Computer-Human Interaction, 19(2): 15:1–15:??, July 2012. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Burnett:2002:ADS

[BYS02a] Margaret Burnett, Sherry Yang, and Jay Summet. Appendices A–D: a scalable method for deductive generalization in the spreadsheet paradigm. ACM Transactions on Computer-Human Interaction, 9(4):1–5, December 2002. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Burnett:2002:SMD

[BYS02b] Margaret Burnett, Sherry Yang, and Jay Summet. A scalable method for deductive generalization in the spreadsheet paradigm. ACM Transactions on Computer-Human Interaction, 9(4):253–284, December 2002. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Crivellaro:2019:IPS

[CALH^{+}19] Clara Crivellaro, Rob Anderson, Daniel Lambton-Howard, Tom Nappey, Patrick Olivier, Vasilis Vlachokyriakos, Alexander Wilson, and Pete Wright. Infrastructuring public service transformation: Creating collaborative spaces between communities and institutions through HCI research. ACM Transactions on Computer-Human Interaction, 26(3):15:1–15:??, June 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3310284.

Cavez:2024:SIS

[CAP24] Vincent Cavez, Caroline Appert, and Emmanuel Pietriga. Spreadsheets on interactive surfaces: Breaking through the grid with the pen. ACM Transactions on Computer-Human Interaction, 31(2):16:1–16:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630097.

Cunningham:2023:GSO

[CBRT23] Jay Cunningham, Gabrielle Benabdallah, Daniela Rosner, and Alex Taylor. On the grounds of solutionism: Ontologies of blackness and HCI. ACM Transactions on Computer-Human Interaction, 30(2):20:1–20:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3557890.

Chen:2013:SCA

[CC13] Zhi-Hong Chen and Sherry Y. Chen. A surrogate competition approach to enhancing gamebased learning. ACM Transactions on Computer-Human Interaction, 20(6):35:1–35:??, December 2013. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Crabtree:2013:ISI

[CCG^{+}13] Andy Crabtree, Alan Chamberlain, Rebecca E. Grinter, Matt Jones, Tom Rodden, and Yvonne Rogers. Introduction to the special issue of “The Turn to The Wild”. ACM Transactions on Computer-Human Interaction, 20(3):13:1–13:??, July 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Church:2014:LSS

[CCO14] Karen Church, Mauro Cherubini, and Nuria Oliver. A large-scale study of daily information needs captured in situ. ACM Transactions on Computer-Human Interaction, 21(2):10:1–10:??, April 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chapuis:2011:EMS

[CD11] Olivier Chapuis and Pierre Dragicevic. Effects of motor scale, visual scale, and quantization on small target acquisition difficulty. ACM Transactions on Computer-Human Interaction, 18(3):13:1–13:??, July 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Carrascal:2015:CRT

[CDC15] Juan Pablo Carrascal, Rodrigo De Oliveira, and Mauro Cherubini. To call or to recall? That’s the research question. ACM Transactions on Computer-Human Interaction, 22(1):4:1–4:??, March 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Carmien:2005:STE

[CDF^{+}05] Stefan Carmien, Melissa Dawe, Gerhard Fischer, Andrew Gorman, Anja Kintsch, and James F. Sullivan, Jr. Sociotechnical environments supporting people with cognitive disabilities using public transportation. ACM Transactions on Computer-Human Interaction, 12(2):233–262, June 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Caramiaux:2015:UGE

[CDT15] Baptiste Caramiaux, Marco Donnarumma, and Atau Tanaka.

Understanding gesture expressivity through muscle sensing. ACM Transactions on Computer-Human Interaction, 21(6):31:1–31:??, January 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chikersal:2021:DDP

[CDT^{+}21] Prerna Chikersal, Afsaneh Doryab, Michael Tumminia, Daniella K. Villalba, Janine M. Dutcher, Xinwen Liu, Sheldon Cohen, Kasey G. Creswell, Jennifer Mankoff, J. David Creswell, Mayank Goel, and Anind K. Dey. Detecting depression and predicting its onset using longitudinal symptoms captured by passive sensing: a machine learning approach with robust feature selection. ACM Transactions on Computer-Human Interaction, 28(1):3:1–3:41, February 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3422821.

Chen:2020:MCM

[CE20] Siyuan Chen and Julien Epps. Multimodal coordination measures to understand users and tasks. ACM Transactions on Computer-Human Interaction, 27(6):42:1–42:26, November 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3412365.

Chaves:2022:CLD

[CEH^{+}22] Ana Paula Chaves, Jesse Egbert, Toby Hocking, Eck Doerry, and Marco Aurelio Gerosa. Chatbots language design: The influence of language variation on user experience with tourist assistant chatbots. ACM Transactions on Computer-Human Interaction, 29(2):13:1–13:38, April 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487193.

Campos:2020:SAS

[CFH^{+}20] José Creissac Campos, Camille Fayollas, Michael D. Harrison, Célia Martinie, Paolo Masci, and Philippe Palanque. Supporting the analysis of safety critical user interfaces: an exploration of three formal tools. ACM Transactions on Computer-Human Interaction, 27(5):35:1–35:48, October 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3404199.

Chow:2024:FSU

[CFH^{+}24] Kevin Chow, Thomas Fritz, Liisa Holsti, Skye Barbic, and Joanna McGrenere. Feeling stressed and unproductive? A field evaluation of a therapy-inspired digital intervention for knowledge workers. ACM Transactions on Computer-Human Interaction, 31(1):12:1–12:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3609330.

Cheng:2023:STT

[CFLC23] Yu-Ting Cheng, Mathias Funk, Rung-Huei Liang, and Lin-Lin Chen. Seeing through things: Exploring the design space of privacy-aware data-enabled objects. ACM Transactions on Computer-Human Interaction, 30(4): 56:1–56:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3577012.

Cockburn:2010:MNE

[CG10] Andy Cockburn and Carl Gutwin. A model of novice and expert navigation performance in constrained-input interfaces. ACM Transactions on Computer-Human Interaction, 17(3):13:1–13:??, July 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cranor:2006:UIP

[CGA06] Lorrie Faith Cranor, Praveen Guduru, and Manjula Arjula. User interfaces for privacy agents. ACM Transactions on Computer-Human Interaction, 13(2):135–178, June 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Corrigan-Gibbs:2015:DCO

[CGGN^{+}15] Henry Corrigan-Gibbs, Nakull Gupta, Curtis Northcutt, Edward Cutrell, and William Thies. Deterring cheating in online environments. ACM Transactions on Computer-Human Interaction, 22(6): 28:1–28:??, December 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chen:2024:UED

[CGL24] Qian Chen, Yeming Gong, and Yaobin Lu. User experience of digital voice assistant: Conceptualization and measurement. ACM Transactions on Computer-Human Interaction, 31(1):10:1–10:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3622782.

Chen:2012:DMS

[CGS12] Nicholas Chen, Francois Guimbretiere, and Abigail Sellen. Designing a multi-slate reading environment to support active reading activities. ACM Transactions on Computer-Human Interaction, 19(3):18:1–18:??, October 2012. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Clemmensen:2020:OUE

[CHAN20] Torkil Clemmensen, Morten Hertzum, and Jose Abdelnour-Nocera. Ordinary user experiences at work: a study of greenhouse growers. ACM Transactions on Computer-Human Interaction, 27(3): 16:1–16:31, June 2020. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386089.

Chandrasekharan:2022:QEE

[CJBG22] Eshwar Chandrasekharan, Shagun Jhaver, Amy Bruckman, and Eric Gilbert. Quarantined! Examining the effects of a community-wide moderation intervention on Reddit. ACM Transactions on Computer-Human Interaction, 29(4):29:1–29:26, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490499.

Cincuegrani:2016:PIU

[CJV16] S. Mealla Cincuegrani, S. Jordà, and A. Väljamäe. Physiopucks: Increasing user motivation by combining tangible and implicit physiological interaction. ACM Transactions on Computer-Human Interaction, 23(1):4:1–4:??, February 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Coblenz:2021:PPI

[CKK^{+}21] Michael Coblenz, Gauri Kambhatla, Paulette Koronkevich, Jenna L. Wise, Celeste Barnaby, Joshua Sunshine, Jonathan Aldrich, and Brad A. Myers. PLIERS: a process that integrates user-centered methods into programming language design. ACM Transactions on Computer-Human Interaction, 28(4):28:1–28:53, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3452379.

Cho:2018:CPM

[CKKL18] Hichang Cho, Bart Knijnenburg, Alfred Kobsa, and Yao Li. Collective privacy management in social media: a cross-cultural validation. ACM Transactions on Computer-Human Interaction, 25(3): 17:1–17:??, June 2018. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Crk:2016:UPE

[CKS16] Igor Crk, Timothy Kluthe, and Andreas Stefik. Understanding programming expertise: an empirical study of phasic brain wave changes. ACM Transactions on Computer-Human Interaction, 23(1):2:1– 2:??, February 2016. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chen:2008:IAM

[CL08] Sherry Y. Chen and Xiaohui Liu. An integrated approach for modeling learning patterns of students in Web-based instruction: a cognitive style perspective. ACM Transactions on Computer-Human Interaction, 15(1):1:1–1:??, May 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cherry:2014:QCS

[CL14] Erin Cherry and Celine Latulipe. Quantifying the creativity support of digital tools through the creativity support index. ACM Transactions on Computer-Human Interaction, 21(4):21:1–21:??, August 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chen:2017:IIF

[CL17] Xiang ‘Anthony’ Chen and Yang Li. Improv: an input framework for improvising cross-device interaction by demonstration. ACM Transactions on Computer-Human Interaction, 24(2):15:1–15:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cho:2020:PRE

[CLG20] Hichang Cho, Pengxiang Li, and Zhang Hao Goh. Privacy risks, emotions, and social media: a coping model of online privacy. ACM Transactions on Computer-Human Interaction, 27(6):40:1–40:28, December 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3412367.

Crain:2023:VTO

[CLY^{+}23] Patrick Crain, Jaewook Lee, Yu-Chun Yen, Joy Kim, Alyssa Aiello, and Brian Bailey. Visualizing topics and opinions helps students interpret large collections of peer feedback for creative projects. ACM Transactions on Computer-Human Interaction, 30(3):49:1–49:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571817.

Chok:2003:AGI

[CM03] Sitt Sen Chok and Kim Marriott. Automatic generation of intelligent diagram editors. ACM Transactions on Computer-Human Interaction, 10(3):244–276, September 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chuang:2012:CUW

[CMH12] Jason Chuang, Christopher D.

Manning, and Jeffrey Heer. “without the clutter of unimportant words”: Descriptive keyphrases for text visualization. ACM Transactions on Computer-Human Interaction, 19(3):19:1–19:??, October 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chen:2010:EDM

[CMLS10] Sherry Y. Chen, Robert D. Macredie, Xiaohui Liu, and Alistair Sutcliffe. Editorial: Data mining for understanding user needs. ACM Transactions on Computer-Human Interaction, 17(1):1:1–1:??, March 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Convertino:2011:SCG

[CMS^{+}11] Gregorio Convertino, Helena M. Mentis, Aleksandra Slavkovic, Mary Beth Rosson, and John M. Carroll. Supporting common ground and awareness in emergency management planning: a design research project. ACM Transactions on Computer-Human Interaction, 18(4):22:1–22:??, December 2011. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Corter:2007:CRS

[CNE^{+}07] James E. Corter, Jeffrey V. Nickerson, Sven K. Esche, Constantin Chassapis, Seongah Im, and Jing Ma. Constructing reality: a study of remote, hands-on, and simulated laboratories. ACM Transactions on Computer-Human Interaction, 14(2):7:1–7:??, August 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chen:2024:MGP

[CNG^{+}24] Chen Chen, Cuong Nguyen, Thibault Groueix, Vladimir G. Kim, and Nadir Weibel. MemoVis: a GenAI-powered tool for creating companion reference images for 3D design feedback. ACM Transactions on Computer-Human Interaction, 31(5):67:1–67:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3694681.

Clear:2016:BOP

[COFH16] Adrian K. Clear, Kirstie O’Neill, Adrian Friday, and Mike Hazas. Bearing an open “Pandora’s Box”: HCI for reconciling everyday food and sustainability. ACM Transactions on Computer-Human Interaction, 23(5):28:1–28:??, November 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cohen:1997:DGC

[Coh97] Jonathan D. Cohen. Drawing graphs to convey proximity: an incremental arrangement method. ACM Transactions on Computer-Human Interaction, 4(3):197– 229, September 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-3/p197-cohen/

.

Chen:2010:EPB

[CP10] Li Chen and Pearl Pu. Experiments on the preferencebased organization interface in recommender systems. ACM Transactions on Computer-Human Interaction, 17(1):5:1– 5:??, March 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Carroll:2013:WHN

[CR13] John M. Carroll and Mary Beth Rosson. Wild at home: The neighborhood as a living laboratory for HCI. ACM Transactions on Computer-Human Interaction, 20(3):16:1–16:??, July 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cohen:2012:DCM

[CRH12] Mark A. Cohen, Frank E. Ritter, and Steven R. Haynes. Dimensions of concern: a method to use cognitive dimensions to evaluate interfaces. ACM Transactions on Computer-Human Interaction, 19(2):9:1–9:??, July 2012. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cabrera:2023:WDM

[CRL^{+}23] Ángel Alexander Cabrera, Marco Tulio Ribeiro, Bongshin Lee, Robert Deline, Adam Perer, and Steven M. Drucker. What did my AI learn? How data scientists make sense of model behavior. ACM Transactions on Computer-Human Interaction, 30(1):1:1– 1:??, February 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3542921.

Cai:2017:WPU

[CRM17] Carrie J. Cai, Anji Ren, and Robert C. Miller. Wait-Suite: Productive use of diverse waiting moments. ACM Transactions on Computer-Human Interaction, 24(1):7:1– 7:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Chan:2007:EBC

[CT07] Hock Chuan Chan and Hock-Hai Teo. Evaluating the boundary conditions of the technology acceptance model: an exploratory investigation. ACM Transactions on Computer-Human Interaction, 14(2):9:1– 9:??, August 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cafaro:2016:FIH

[CVB16] Angelo Cafaro, Hannes Högni Vilhjálmsson, and Timothy Bickmore. First impressions in human–agent virtual encounters. ACM Transactions on Computer-Human Interaction, 23(4):24:1–24:??, September 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cerrolaza:2012:SPM

[CVC12] Juan J. Cerrolaza, Arantxa Villanueva, and Rafael Cabeza. Study of polynomial mapping functions in video-oculography eye trackers. ACM Transactions on Computer-Human Interaction, 19(2):10:1–10:??, July 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Cherubini:2020:UDP

[CVZBB20] Mauro Cherubini, Gabriela Villalobos-Zuñiga, Marc-Olivier Boldi, and Riccardo Bonazzi. The unexpected downside of paying or sending messages to people to make them walk: Comparing tangible rewards and motivational messages to improve physical activity. ACM Transactions on Computer-Human Interaction, 27(2):8:1–8:44, April 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3365665.

Chen:2023:MED

[CWM^{+}23] Xiang “Anthony” Chen, Chien-Sheng Wu, Lidiya Murakhovs’ka, Philippe Laban, Tong Niu, Wenhao Liu, and Caiming Xiong. Marvista: Exploring the design of a human– AI collaborative news reading tool. ACM Transactions on Computer-Human Interaction, 30(6):92:1–92:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3609331.

Chignell:2023:EHH

[CWZL23] Mark Chignell, Lu Wang, Atefeh Zare, and Jamy Li. The evolution of HCI and human factors: Integrating human and artificial intelligence. ACM Transactions on Computer-Human Interaction, 30(2):17:1–17:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557891.

Chua:2015:CAV

[CZH^{+}15] Soon Hau Chua, Haimo Zhang, Muhammad Hammad, Shengdong Zhao, Sahil Goyal, and Karan Singh. ColorBless: Augmenting visual information for colorblind people with binocular luster effect. ACM Transactions on Computer-Human Interaction, 21(6):32:1–32:??, January 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Devendorf:2020:FUM

[DAK20] Laura Devendorf, Kristina Andersen, and Aisling Kelliher. The fundamental uncertainties of mothering: Finding ways to honor endurance, struggle, and contradiction. ACM Transactions on Computer-Human Interaction, 27(4):26:1–26:24, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3397177.

Desolda:2017:EEU

[DAM17] Giuseppe Desolda, Carmelo Ardito, and Maristella Matera. Empowering end users to customize their smart environments: Model, composition paradigms, and domainspecific tools. ACM Transactions on Computer-Human Interaction, 24(2):12:1–12:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Duarte:2018:PDP

[DBDK18] Ana Maria Bustamante Duarte, Nina Brendel, Auriol Degbelo, and Christian Kray. Participatory design and participatory research: an HCI case study with young forced migrants. ACM Transactions on Computer-Human Interaction, 25(1):3:1–3:??, February 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

DAlbis:2012:PSC

[DBT^{+}12] Tiziano D’Albis, Rossella Blatt, Roberto Tedesco, Licia Sbattella, and Matteo Matteucci. A predictive speller controlled by a brain-computer interface based on motor imagery. ACM Transactions on Computer-Human Interaction, 19(3):20:1–20:??, October 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Dewan:1995:CUI

[DC95] Prasun Dewan and Rajiv Choudhary. Coupling the user interfaces of a multiuser program. ACM Transactions on Computer-Human Interaction, 2(1):1–39, March 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-1/p1-dewan/.

DeOliveira:2013:IPS

[DCO13] Rodrigo De Oliveira, Mauro Cherubini, and Nuria Oliver. Influence of personality on satisfaction with mobile phone services. ACM Transactions on Computer-Human Interaction, 20(2):10:1–10:??, May 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Deering:1995:HVR

[Dee95] Michael F. Deering. HoloSketch: a virtual reality sketching/animation tool. ACM Transactions on Computer-Human Interaction, 2(3):220– 238, September 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-3/p220-deering/

.

Dourish:1999:PEA

[DELS99] Paul Dourish, W. Keith Edwards, Anthony LaMarca, and Michael Salisbury. Presto: an experimental architecture for fluid interactive document spaces. ACM Transactions on Computer-Human Interaction, 6(2):133–161, June 1999. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-2/p133-dourish/.

DaSilva:2015:ISI

[DFH^{+}15] Hugo Plácido Da Silva, Stephen Fairclough, Andreas Holzinger, Robert Jacob, and Desney Tan. Introduction to the special issue on physiological computing for human-computer interaction. ACM Transactions on Computer-Human Interaction, 21(6):29:1–29:??, January 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Distler:2021:SLR

[DFH^{+}21] Verena Distler, Matthias Fassl, Hana Habib, Katharina Krombholz, Gabriele Lenzini, Carine Lallemand, Lorrie Faith Cranor, and Vincent Koenig. A systematic literature review of empirical methods and risk representation in usable privacy and security research. ACM Transactions on Computer-Human Interaction, 28(6):43:1–43:50, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469845.

Dow:2010:PPL

[DGK^{+}10] Steven P. Dow, Alana Glassco, Jonathan Kass, Melissa Schwarz, Daniel L. Schwartz, and Scott R. Klemmer. Parallel prototyping leads to better design results, more divergence, and increased selfefficacy. ACM Transactions on Computer-Human Interaction, 17(4):18:1–18:??, December 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Dalsgaard:2008:PPS

[DH08] Peter Dalsgaard and Lone Koefoed Hansen. Performing perception — staging aesthetics of interaction. ACM Transactions on Computer-Human Interaction, 15(3):13:1–13:??, November 2008. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Dombrowski:2014:GIC

[DHMV14] Lynn Dombrowski, Gillian R.

Hayes, Melissa Mazmanian, and Amy Voida. E-government intermediaries and the challenges of access and trust. ACM Transactions on Computer-Human Interaction, 21(2): 13:1–13:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

DAntoni:2015:HCA

[DKA^{+}15] Loris D’Antoni, Dileep Kini, Rajeev Alur, Sumit Gulwani, Mahesh Viswanathan, and Björn Hartmann. How can automatic feedback help students construct automata? ACM Transactions on Computer-Human Interaction, 22(2):9:1– 9:??, April 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

DeCastroLeal:2021:DTE

[DKT^{+}21] Débora De Castro Leal, Max Krüger, Vanessa Teles E.Teles, Carlos Antônio Teles E.Teles, Denise Machado Cardoso, Dave Randall, and Volker Wulf. Digital technology at the edge of capitalism: Experiences from the Brazilian Amazon rainforest. ACM Transactions on Computer-Human Interaction, 28(3):18:1–18:39, July 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3448072.

Dubiel:2024:HGY

[DLBBS24] Mateusz Dubiel, Luis A. Leiva, Kerstin Bongard-Blanchy, and Anastasia Sergeeva. “Hey genie, you got me thinking about my menu choices!” Impact of proactive feedback on user perception and reflection in decision-making tasks. ACM Transactions on Computer-Human Interaction, 31(5): 56:1–56:??, October 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3685274.

DeVries:2024:BSB

[DLL24] Roelof A. J. De Vries, Mailin Lemke, and Geke D. S. Ludden. Blueprints: Systematizing behavior change designs — the case of social comparison theory. ACM Transactions on Computer-Human Interaction, 31(1):11:1–11:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617364.

Dey:2005:DMC

[DM05] Anind K. Dey and Jennifer Mankoff. Designing mediation for context-aware applications. ACM Transactions on Computer-Human Interaction, 12(1):53–80, March 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

DeMelo:2016:PDF

[DMG16] Celso De Melo, Stacy Marsella, and Jonathan Gratch. People do not feel guilty about exploiting machines. ACM Transactions on Computer-Human Interaction, 23(2):8:1– 8:??, May 2016. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ducasse:2018:BNV

[DMOJ18] Julie Ducasse, Marc Macé, Bernard Oriola, and Christophe Jouffrais. BotMap: Non-visual panning and zooming with an actuated tabletop tangible interface. ACM Transactions on Computer-Human Interaction, 25(4):24:1–24:??, September 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3204460.

Dourish:1995:DRM

[Dou95] Paul Dourish. Developing a reflective model of collaborative systems. ACM Transactions on Computer-Human Interaction, 2(1):40– 63, March 1995. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-1/p40-dourish/

.

Dourish:1998:UMT

[Dou98] Paul Dourish. Using metalevel techniques in a flexible toolkit for CSCW applications. ACM Transactions on Computer-Human Interaction, 5(2):109–155, June 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1998-5-2/p109-dourish/

.

Dourish:2013:EWA

[Dou13]  Paul Dourish.  Epilogue:

Where the action was, wasn’t, should have been, and might yet be. ACM Transactions on Computer-Human Interaction, 20(1):2:1–2:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Das:2022:DEA

[DPG22] Maitraye Das, Anne Marie Piper, and Darren Gergle. Design and evaluation of accessible collaborative writing techniques for people with vision impairments. ACM Transactions on Computer-Human Interaction, 29(2):9:1–9:42, April 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480169.

Dunham:2023:ICP

[DPL^{+}23] John Dunham, Konstantinos Papangelis, Samuli Laato, Nicolas Lalone, Jin Lee, and Michael Saker. The impacts of Covid-19 on players of Pokémon GO. ACM Transactions on Computer-Human Interaction, 30(4): 57:1–57:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3569896.

Dix:2000:ESL

[DRD^{+}00] Alan Dix, Tom Rodden, Nigel Davies, Jonathan Trevor, Adrian Friday, and Kevin Palfreyman. Exploiting space and location as a design framework for interactive mobile systems. ACM Transactions on Computer-Human Interaction, 7(3):285–321, 2000. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tochi/2000-7-3/p285-dix/.

Disalvo:2013:CSI

[DRW13] Carl Disalvo, Johan Redström, and Matt Watson. Commentaries on the special issue on practice-oriented approaches to sustainable HCI. ACM Transactions on Computer-Human Interaction, 20(4): 26:1–26:??, September 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Dewan:1998:CAM

[DS98] Prasun Dewan and Honghai Shen. Controlling access in multiuser interfaces. ACM Transactions on Computer-Human Interaction, 5(1):34– 62, March 1998. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1998-5-1/p34-dewan/

.

DeBruijn:2008:NFT

[DS08] Oscar De Bruijn and Robert Spence. A new framework for theory-based interaction design applied to serendipitous information retrieval. ACM Transactions on Computer-Human Interaction, 15(1):5:1– 5:??, May 2008. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Dai:2009:SFA

[DSG09] Liwei Dai, Andrew Sears, and Rich Goldman. Shifting the focus from accuracy to recallability: a study of informal notetaking on mobile information technologies. ACM Transactions on Computer-Human Interaction, 16(1):4:1–4:??, April 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Desai:2023:MVU

[DT23] Smit Desai and Michael Twidale. Metaphors in voice user interfaces: a slippery fish. ACM Transactions on Computer-Human Interaction, 30(6):89:1–89:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3609326.

Desjardins:2021:ISI

[DTL^{+}21] Audrey Desjardins, Oscar Tomico, Andrés Lucero, Marta E. Cecchinato, and Carman Neustaedter. Introduction to the special issue on first-person methods in HCI. ACM Transactions on Computer-Human Interaction, 28(6):37:1–37:12, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3492342.

Ding:2023:CUM

[DTP^{+}23] Li Ding, Jack Terwilliger, Aishni Parab, Meng Wang, Lex Fridman, Bruce Mehler, and Bryan Reimer. CLERA: a unified model for joint cognitive load and eye region analysis in the wild. ACM Transactions on Computer-Human Interaction, 30(6):84:1–84:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3603622.

Dillahunt:2018:GTB

[DV18] Tawanna R. Dillahunt and Tiffany C. Veinot. Getting there: Barriers and facilitators to transportation access in underserved communities. ACM Transactions on Computer-Human Interaction, 25(5):29:1–29:??, October 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Duarte:2021:SPG

[DVHZ^{+}21] Marisa Elena Duarte, Morgan Vigil-Hayes, Ellen Zegura, Elizabeth Belding, Ivone Masara, and Jennifer Case Nevarez. As a squash plant grows: Social textures of sparse Internet connectivity in rural and tribal communities. ACM Transactions on Computer-Human Interaction, 28(3):16:1–16:16, July 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3453862.

Dudley:2018:FPT

[DVK18] John J. Dudley, Keith Vertanen, and Per Ola Kristensson. Fast and precise touch-based text entry for head-mounted augmented reality with variable occlusion. ACM Transactions on Computer-Human Interaction, 25(6):30:1–30:??, December 2018. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Draxler:2024:AGE

[DWL^{+}24] Fiona Draxler, Anna Werner, Florian Lehmann, Matthias Hoppe, Albrecht Schmidt, Daniel Buschek, and Robin Welsch. The AI ghostwriter effect: When users do not perceive ownership of AIgenerated text but self-declare as authors. ACM Transactions on Computer-Human Interaction, 31(2):25:1–25:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3637875.

Elvitigala:2022:TDD

[EBNM22] Don Samitha Elvitigala, Roger Boldu, Suranga Nanayakkara, and Denys J. C. Matthies. TickleFoot: Design, development and evaluation of a novel foot-tickling mechanism that can evoke laughter. ACM Transactions on Computer-Human Interaction, 29(3): 20:1–20:23, June 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3490496.

Eisbach:2023:PRD

[EDT^{+}23] Simon Eisbach, Fabian Daugs, Meinald T. Thielsch, Matthias Böhmer, and Guido Hertel. Predicting rating distributions of Website aesthetics with deep learning for AI-based research. ACM Transactions on Computer-Human Interaction, 30(3):37:1–37:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3569889.

Edwards:2005:PCC

[Edw05] W. Keith Edwards. Putting computing in context: an infrastructure to support extensible context-enhanced collaborative applications. ACM Transactions on Computer-Human Interaction, 12(4):446– 474, December 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ebling:2002:ITM

[EJS02] Maria R. Ebling, Bonnie E. John, and M. Satyanarayanan. The importance of translucence in mobile computing systems. ACM Transactions on Computer-Human Interaction, 9(1):42–67, March 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Erickson:2000:STA

[EK00] Thomas Erickson and Wendy A.

Kellogg. Social translucence: an approach to designing systems that support social processes. ACM Transactions on Computer-Human Interaction, 7(1):59– 83, March 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
2000-7-1/p59-erickson/.

Edwards:2009:ERC

[ENSS09] W. Keith Edwards, Mark W.

Newman, Jana Z. Sedivy, and Trevor F. Smith. Experiences with recombinant computing: Exploring ad hoc interoperability in evolving digital networks. ACM Transactions on Computer-Human Interaction, 16(1):3:1–3:??, April 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Engelbutzeder:2023:SST

[ERL^{+}23] Philip Engelbutzeder, Dave Randell, Marvin Landwehr, Konstantin Aal, Gunnar Stevens, and Volker Wulf. From surplus and scarcity toward abundance: Understanding the use of ICT in food resource sharing practices. ACM Transactions on Computer-Human Interaction, 30(5):80:1–80:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589957.

Erete:2023:MMA

[ERT23] Sheena Erete, Yolanda Rankin, and Jakita Thomas. A method to the madness: Applying an intersectional analysis of structural oppression and power in HCI and design. ACM Transactions on Computer-Human Interaction, 30(2):24:1–24:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3507695.

Edge:2016:SHD

[EYK^{+}16] Darren Edge, Xi Yang, Yasmine Kotturi, Shuoping Wang, Dan Feng, Bongshin Lee, and Steven Drucker. SlideSpace: Heuristic design of a hybrid presentation medium. ACM Transactions on Computer-Human Interaction, 23(3): 16:1–16:??, July 2016. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ez-Zaouia:2024:GFE

[EZC24] Mohamed Ez-Zaouia and Rubiela Carrillo. The group folding effect: The role of collaborative process structuring and social interaction in group work. ACM Transactions on Computer-Human Interaction, 31(2):15:1–15:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3622783.

Frik:2023:MCF

[FBE23] Alisa Frik, Julia Bernd, and Serge Egelman. A model of contextual factors affecting older adults’ informationsharing decisions in the U.S. ACM Transactions on Computer-Human Interaction, 30(1):12:1–12:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557888.

Fischer:2022:OFC

[FFKM22] Florian Fischer, Arthur Fleig, Markus Klar, and Jörg Müller. Optimal feedback control for modeling human-computer interaction. ACM Transactions on Computer-Human Interaction, 29(6):51:1–51:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3524122.

Frokjaer:2008:MHT

[FH08] Erik Frøkjær and Kasper Hornbæk. Metaphors of human thinking for usability inspection and design. ACM Transactions on Computer-Human Interaction, 14(4): 20:1–20:??, January 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Faeth:2014:EEM

[FH14] Adam Faeth and Chris Harding. Emergent effects in multimodal feedback from virtual buttons. ACM Transactions on Computer-Human Interaction, 21(1):3:1–3:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Forlano:2023:SHJ

[FH23] Laura E. Forlano and Megan K. Halpern. Speculative histories, just futures: From counterfactual artifacts to counterfactual actions. ACM Transactions on Computer-Human Interaction, 30(2):22:1–22:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3577212.

Fogarty:2005:PHI

[FHA^{+}05] James Fogarty, Scott E. Hudson, Christopher G. Atkeson, Daniel Avrahami, Jodi Forlizzi, Sara Kiesler, Johnny C. Lee, and Jie Yang. Predicting human interruptibility with sensors. ACM Transactions on Computer-Human Interaction, 12(1):119–146, March 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Franz:2024:VRS

[FJM24] Rachel L. Franz, Sasa Junuzovic, and Martez Mott. A virtual reality scene taxonomy: Identifying and designing accessible scene-viewing techniques. ACM Transactions on Computer-Human Interaction, 31(2):23:1–23:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635142.

Fern:2010:MPS

[FKGB10] Xiaoli Fern, Chaitanya Komireddy, Valentina Grigoreanu, and Margaret Burnett. Mining problem-solving strategies from HCI data. ACM Transactions on Computer-Human Interaction, 17(1):3:1– 3:??, March 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Fell:2022:BPH

[FKGW22] Jan Fell, Pei-Yi Kuo, Travis Greene, and Jyun-Cheng Wang. A biocentric perspective on HCI design research involving plants. ACM Transactions on Computer-Human Interaction, 29(5): 46:1–46:??, October 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3512887.

Frees:2007:PIE

[FKK07] Scott Frees, G. Drew Kessler, and Edwin Kay. PRISM interaction for enhancing control in immersive virtual environments. ACM Transactions on Computer-Human Interaction, 14(1):2:1–2:??, May 2007. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Fu:2010:SIS

[FKKH10] Wai-Tat Fu, Thomas Kannampallil, Ruogu Kang, and Jibo He. Semantic imitation in social tagging. ACM Transactions on Computer-Human Interaction, 17(3):12:1–12:??, July 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Fan:2019:CTA

[FLCT19] Mingming Fan, Jinglan Lin, Christina Chung, and Khai N. Truong. Concurrent thinkaloud verbalizations and usability problems. ACM Transactions on Computer-Human Interaction, 26(5): 28:1–28:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3325281.

Ferres:2013:ETI

[FLST13] Leo Ferres, Gitte Lindgaard, Livia Sumegi, and Bruce Tsuji. Evaluating a tool for improving accessibility to charts and graphs. ACM Transactions on Computer-Human Interaction, 20(5):28:1–28:??, November 2013. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

Folstad:2024:CBC

[FLvA24] Asbjørn Følstad, Effie L.-C.

Law, and Nena van As. Conversational breakdown in a customer service chatbot: Impact of task order and criticality on user trust and emotion. ACM Transactions on Computer-Human Interaction, 31(5):66:1–66:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3690383.

Fox:2020:MMI

[FMEB20] Sarah E. Fox, Amanda Menking, Jordan Eschler, and Uba Backonja. Multiples over models: Interrogating the past and collectively reimagining the future of menstrual sensemaking. ACM Transactions on Computer-Human Interaction, 27(4):22:1–22:24, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3397178.

Foley:2019:SRA

[FMP19] Sarah Foley, John McCarthy, and Nadia Pantidi. The struggle for recognition in advanced dementia: Implications for experience-centered design. ACM Transactions on Computer-Human Interaction, 26(6):40:1–40:??, December 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3359594.

Fogues:2017:SPM

[FMSS17] Ricard L. Fogues, Pradeep K. Murukannaiah, Jose M. Such, and Munindar P. Singh. Sharing policies in multiuser privacy scenarios: Incorporating context, preferences, and arguments in decision making. ACM Transactions on Computer-Human Interaction, 24(1):5:1–5:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ferdous:2016:CSU

[FPD^{+}16] Hasan Shahid Ferdous, Bernd Ploderer, Hilary Davis, Frank Vetere, and Kenton O’Hara. Commensality and the social use of technology during family mealtime. ACM Transactions on Computer-Human Interaction, 23(6):37:1–37:??, December 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Fields:1999:CDO

[FPST99] Robert Fields, Fabio Paternò, Carmen Santoro, and Sophie Tahmassebi. Comparing design options for allocating communication media in cooperative safety-critical contexts: a method and a case study. ACM Transactions on Computer-Human Interaction, 6(4):370–398, December 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-4/p370-fields/.

Frauenberger:2020:EHN

[Fra20] Christopher Frauenberger. Entanglement HCI: The next wave? ACM Transactions on Computer-Human Interaction, 27(1):2:1–2:27, January 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3364998.

Feng:2004:UCS

[FS04] Jinjuan Feng and Andrew Sears. Using confidence scores to improve hands-free speech based navigation in continuous dictation systems. ACM Transactions on Computer-Human Interaction, 11(4):329– 356, December 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Fennedy:2022:TUE

[FSMP22] Katherine Fennedy, Angad Srivastava, Sylvain Malacria, and Simon T. Perrault. Towards a unified and efficient command selection mechanism for touch-based devices using soft keyboard hotkeys. ACM Transactions on Computer-Human Interaction, 29(1):4:1– 4:39, February 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3476510.

Greenhalgh:1995:MCV

[GB95] Chris Greenhalgh and Steven Benford. MASSIVE: a collaborative virtual environment for teleconferencing. ACM Transactions on Computer-Human Interaction, 2(3):239– 261, September 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1995-2-3/p239-greenhalgh/

.

Grossman:2005:PAM

[GB05] Tovi Grossman and Ravin Balakrishnan. A probabilistic approach to modeling twodimensional pointing. ACM Transactions on Computer-Human Interaction, 12(3):435– 459, September 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ghoshal:2019:RSC

[GB19] Sucheta Ghoshal and Amy Bruckman. The role of social computing technologies in grassroots movement building. ACM Transactions on Computer-Human Interaction, 26(3):18:1–18:??, June 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3318140.

Gao:2012:WDT

[GBBM12] Yuan Gao, Nadia Bianchi-Berthouze, and Hongying Meng. What does touch tell us about emotions in touchscreenbased gameplay? ACM Transactions on Computer-Human Interaction, 19(4):31:1–31:??, December 2012. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Gooch:2018:AQV

[GBH^{+}18] Daniel Gooch, Matthew Barker, Lorraine Hudson, Ryan Kelly, Gerd Kortuem, Janet Van Der Linden, Marian Petre, Rebecca Brown, Anna Klis-Davies, Hannah Forbes, Jessica Mackinnon, Robbie Macpherson, and Clare Walton. Amplifying quiet voices: Challenges and opportunities for participatory design at an urban scale. ACM Transactions on Computer-Human Interaction, 25(1):2:1–2:??, February 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Grigoreanu:2012:EUD

[GBW^{+}12] Valentina Grigoreanu, Margaret Burnett, Susan Wiedenbeck, Jill Cao, Kyle Rector, and Irwin Kwan. End-user debugging strategies: a sensemaking perspective. ACM Transactions on Computer-Human Interaction, 19(1):5:1– 5:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Garg:2022:SCA

[GC22] Radhika Garg and Hua Cui. Social contexts, agency, and conflicts: Exploring critical aspects of design for future smart home technologies. ACM Transactions on Computer-Human Interaction, 29(2): 11:1–11:30, April 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3485058.

Gould:2016:DCC

[GCB16] Sandy J. J. Gould, Anna L. Cox, and Duncan P. Brumby. Diminished control in crowdsourcing: an investigation of crowdworker multitasking behavior. ACM Transactions on Computer-Human Interaction, 23(3):19:1–19:??, July 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Gao:2024:CEE

[GCC^{+}24] Jie Gao, Kenny Tsu Wei Choo, Junming Cao, Roy Ka-Wei Lee, and Simon Perrault. CoAIcoder: Examining the effectiveness of AIassisted human-to-human collaboration in qualitative analysis. ACM Transactions on Computer-Human Interaction, 31(1):6:1–6:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617362.

Goree:2023:IWR

[GCS23] Samuel Goree, David Crandall, and Norman Makoto Su. “It Was Really All About Books:” speech-like technomasculinity in the rhetoric of dot-com era Web design books. ACM Transactions on Computer-Human Interaction, 30(2):18:1–18:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3508067.

Gerling:2024:RID

[GDW^{+}24] Kathrin Gerling, Arno Depoortere, Jeroen Wauters, Katta Spiel, Dmitry Alexandrovsky, Marina Danckaerts, Dieter Baeyens, and Saskia Van der Oord. Representation of invisible disability: Exploring the lived experience of teenagers with ADHD to inform game design. ACM Transactions on Computer-Human Interaction, 31(5): 58:1–58:??, October 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3685276.

Grinter:2009:IOH

[GEC^{+}09] Rebecca E. Grinter, W. Keith Edwards, Marshini Chetty, Erika S. Poole, Ja-Young Sung, Jeonghwa Yang, Andy Crabtree, Peter Tolmie, Tom Rodden, Chris Greenhalgh, and Steve Benford. The ins and outs of home networking: The case for useful and usable domestic networking. ACM Transactions on Computer-Human Interaction, 16(2):8:1– 8:??, June 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Grasso:1998:ISM

[GEF98] Michael A. Grasso, David S. Ebert, and Timothy W. Finin. The integrality of speech in multimodal interfaces. ACM Transactions on Computer-Human Interaction, 5(4):303– 325, December 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1998-5-4/p303-grasso/

.

Ganglbauer:2013:NFW

[GFC13] Eva Ganglbauer, Geraldine Fitzpatrick, and Rob Comber. Negotiating food waste: Using a practice lens to inform design. ACM Transactions on Computer-Human Interaction, 20(2):11:1–11:??, May 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Gadiraju:2017:UWS

[GFK^{+}17] Ujwal Gadiraju, Besnik Fetahu, Ricardo Kawase, Patrick Siehndel, and Stefan Dietze. Using worker self-assessments for competence-based preselection in crowdsourcing microtasks. ACM Transactions on Computer-Human Interaction, 24(4):30:1–30:??, September 2017. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Gutwin:1999:EWA

[GG99] Carl Gutwin and Saul Greenberg. The effects of workspace awareness support on the usability of real-time distributed groupware. ACM Transactions on Computer-Human Interaction, 6(3):243–281, September 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-3/p243-gutwin/.

Gerber:2013:CMD

[GH13] Elizabeth M. Gerber and Julie Hui. Crowdfunding: Motivations and deterrents for participation. ACM Transactions on Computer-Human Interaction, 20(6):34:1–34:??, December 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Greenhalgh:2022:CTS

[GHB^{+}22] Chris Greenhalgh, Adrian Hazzard, Steve Benford, Laurence Cliffe, and Elizabeth Kelly. Crafting trajectories of smart phone use at the opera. ACM Transactions on Computer-Human Interaction, 29(6):59:1–59:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531007.

Gillies:2019:URI

[Gil19] Marco Gillies. Understanding the role of interactive machine learning in movement interaction design. ACM Transactions on Computer-Human Interaction, 26(1):5:1–5:??, February 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3287307.

Gu:2023:IWI

[GLX^{+}23] Hongyan Gu, Yuan Liang, Yifan Xu, Christopher Kazu Williams, Shino Magaki, Negar Khanlou, Harry Vinters, Zesheng Chen, Shuo Ni, Chunxu Yang, Wenzhong Yan, Xinhai Robert Zhang, Yang Li, Mohammad Haeri, and Xiang ‘Anthony’ Chen. Improving workflow integration with xPath: Design and evaluation of a Human-AI diagnosis system in pathology. ACM Transactions on Computer-Human Interaction, 30(2):28:1–28:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3577011.

Ghosh:2020:CRD

[GLZH20] Debjyoti Ghosh, Can Liu, Shengdong Zhao, and Kotaro Hara. Commanding and re-dictation: Developing eyes-free voice-based interaction for editing dictated text. ACM Transactions on Computer-Human Interaction, 27(4):28:1–28:31, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3390889.

Grayson:2003:YLM

[GM03] David M. Grayson and Andrew F. Monk. Are you looking at me? Eye contact and desktop video conferencing. ACM Transactions on Computer-Human Interaction, 10(3):221–243, September 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ghiani:2017:PCD

[GMPS17] Giuseppe Ghiani, Marco Manca, Fabio Paternò, and Carmen Santoro. Personalization of context-dependent applications through trigger-action rules. ACM Transactions on Computer-Human Interaction, 24(2):14:1–14:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Guimbretiere:2005:BMC

[GMW05] François Guimbretière, Andrew Martin, and Terry Winograd. Benefits of merging command selection and direct manipulation. ACM Transactions on Computer-Human Interaction, 12(3):460–476, September 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Grinter:2006:CTC

[GPE06] Rebecca E. Grinter, Leysia Palen, and Margery Eldridge. Chatting with teenagers: Considering the place of chat technologies in teen life. ACM Transactions on Computer-Human Interaction, 13(4):423– 447, December 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ghosh:2023:FML

[GPH^{+}23] Pratik Ghosh, Karen L. Posner, Stephanie L. Hyland, Wil van Cleve, Melissa Bristow, Dustin R. Long, Konstantina Palla, Bala Nair, Christine Fong, Ronald Pauldine, Monica S. Vavilala, and Kenton O’Hara. Framing machine learning opportunities for hypotension prediction in perioperative care: a socio-technical perspective: Socio-technical perspectives on hypotension prediction. ACM Transactions on Computer-Human Interaction, 30(5):79:1–79:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589953.

Gray:1999:ISI

[GPP99] Wayne D. Gray, Philippe Palanque, and Fabio Paternó. Introduction to the special issue on interface issues and designs for safetycritical interactive systems: when there is no room for user error. ACM Transactions on Computer-Human Interaction, 6(4):309–310, December 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
1999-6-4/p309-gray/p309-
gray.pdf;   http://www.acm.
org/pubs/citations/journals/
tochi/1999-6-4/p309-gray/

.

Gauthier:2022:MID

[GPPD^{+}22] Andrea Gauthier, Kaska Porayska-Pomsta, Iroise Dumontheil, Sveta Mayer, and Denis Mareschal. Manipulating interface design features affects children’s stop-and-think behaviours in a counterintuitiveproblem game. ACM Transactions on Computer-Human Interaction, 29(2):12:1–12:22, April 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485168.

Grabowski:2011:HRV

[GR11] Martha Grabowski and Karlene Roberts. High reliability virtual organizations: Co-adaptive technology and organizational structures in tsunami warning systems. ACM Transactions on Computer-Human Interaction, 18(4):19:1–19:??, December 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Gori:2018:SAT

[GRG18] Julien Gori, Olivier Rioul, and Yves Guiard. Speed-accuracy tradeoff: a formal informationtheoretic transmission scheme (FITTS). ACM Transactions on Computer-Human Interaction, 25(5):27:1–27:??, October 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3231595.

Guy:2016:IAE

[GRKB16] Ido Guy, Inbal Ronen, Elad Kravi, and Maya Barnea. Increasing activity in enterprise online communities using content recommendation. ACM Transactions on Computer-Human Interaction, 23(4): 22:1–22:??, September 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ghafurian:2020:CTS

[GRR20] Moojan Ghafurian, David Reitter, and Frank E. Ritter. Countdown timer speed: a trade-off between delay duration perception and recall. ACM Transactions on Computer-Human Interaction, 27(2):11:1–11:25, April 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3380961.

Grudin:2004:CD

[Gru04] Jonathan Grudin. Crossing the divide. ACM Transactions on Computer-Human Interaction, 11(1):1–25, March 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Gayler:2022:EDS

[GSK22] Tom Gayler, Corina Sas, and Vaiva Kalnikaite. Exploring the design space for humanfood-technology interaction: an approach from the lens of eating experiences. ACM Transactions on Computer-Human Interaction, 29(2): 16:1–16:52, April 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3484439.

Galliers:1999:IAM

[GSM99] Julia Galliers, Alistair Sutcliffe, and Shailey Minocha. An impact analysis method for safety-critical user interface design. ACM Transactions on Computer-Human Interaction, 6(4):341–369, December 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tochi/
                                        1999-6-4/p341-galliers/.

Gerber:2024:DAA

[GSPZ24] Nina Gerber, Alina Stöver, Justin Peschke, and Verena Zimmermann. Don’t accept all and continue: Exploring nudges for more deliberate interaction with tracking consent notices. ACM Transactions on Computer-Human Interaction, 31(1):1:1–1:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617363.

Glassman:2015:OVV

[GSS^{+}15] Elena L. Glassman, Jeremy Scott, Rishabh Singh, Philip J. Guo, and Robert C. Miller. OverCode: Visualizing variation in student solutions to programming problems at scale. ACM Transactions on Computer-Human Interaction, 22(2):7:1–7:??, April 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Girouard:2019:RRB

[GSS^{+}19] Audrey Girouard, Orit Shaer, Erin T. Solovey, G. Michael Poor, and Robert J. K. Jacob. The reality of reality-based interaction: Understanding the impact of a framework as a research tool. ACM Transactions on Computer-Human Interaction, 26(5): 35:1–35:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3319617.

Guo:2023:CPI

[GSX^{+}23] Shihui Guo, Yubin Shi, Pintong Xiao, Yinan Fu, Juncong Lin, Wei Zeng, and Tong-Yee Lee. Creative and progressive interior color design with eye-tracked user preference. ACM Transactions on Computer-Human Interaction, 30(1):5:1–5:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542922.

Guberman:2023:ATS

[Gub23]  Josh Guberman.  #Actu-

allyAutistic Twitter as a site for epistemic resistance and crip futurity. ACM Transactions on Computer-Human Interaction, 30(3):38:1–38:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569891.

Guo:2023:FTF

[GWZC23] Xunhua Guo, Lingli Wang, Mingyue Zhang, and Guoqing Chen. First things first? Order effects in online product recommender systems. ACM Transactions on Computer-Human Interaction, 30(1):15:1–15:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3557886.

Harms:2019:AUE

[Har19]  Patrick Harms.  Auto-

mated usability evaluation of virtual reality applications. ACM Transactions on Computer-Human Interaction, 26(3):14:1–14:??, June 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301423.

Hayes:2011:RAR

[Hay11] Gillian R. Hayes. The relationship of action research to human-computer interaction. ACM Transactions on Computer-Human Interaction, 18(3):15:1–15:??, July 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Harper:2007:SSS

[HB07] Simon Harper and Sean Bechhofer. SADIe: Structural semantics for accessibility and device independence. ACM Transactions on Computer-Human Interaction, 14(2): 10:1–10:??, August 2007. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hourcade:2004:DPT

[HBDG04] Juan Pablo Hourcade, Benjamin B. Bederson, Allison Druin, and François Guimbretière. Differences in pointing task performance between preschool children and adults using mice. ACM Transactions on Computer-Human Interaction, 11(4):357–386, December 2004. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Healey:1996:HVE

[HBE96] Christopher G. Healey, Kellogg S. Booth, and James T. Enns. High-speed visual estimation using preattentive processing. ACM Transactions on Computer-Human Interaction, 3(2):107–135, June 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-2/p107-healey/

.

Hirsch:2024:NLC

[HBJP24] Mare Hirsch, Gabrielle Benabdallah, Jennifer Jacobs, and Nadya Peek. Nothing like compilation: How professional digital fabrication workflows go beyond extruding, milling, and machines. ACM Transactions on Computer-Human Interaction, 31(1):13:1–13:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3609328.

Hogan:2022:DOM

[HBL22] Mairéad Hogan, Chris Barry, and Michael Lang. Dissecting optional micro-decisions in online transactions: Perceptions, deceptions, and errors. ACM Transactions on Computer-Human Interaction, 29(6):53:1–53:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3531005.

Hornbaek:2002:NPU

[HBP02] Kasper Hornbæk, Benjamin B. Bederson, and Catherine Plaisant. Navigation patterns and usability of zoomable user interfaces with and without an overview. ACM Transactions on Computer-Human Interaction, 9(4):362–389, December 2002. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Anonymous:1994:I

[HBR^{+}94] Ralph D. Hill, Tom Brinck, Steven L. Rohall, John F. Patterson, and Wayne Wilner. The Rendezvous architecture and language for constructing multiuser applications. ACM Transactions on Computer-Human Interaction, 1(2):81– 125, June 1994. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-2/p81-hill/.

Hook:2021:UND

[HBT^{+}21] Kristina Höök, Steve Benford, Paul Tennent, Vasiliki Tsaknaki, Miquel Alfaras, Juan Martinez Avila, Christine Li, Joseph Marshall, Claudia Daudén Roquet, Pedro Sanches, Anna Ståhl, Muhammad Umair, Charles Windlin, and Feng Zhou. Unpacking non-dualistic design: The soma design case. ACM Transactions on Computer-Human Interaction, 28(6):40:1–40:36, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3462448.

Hiltz:2006:ISI

[HC06] Starr Roxanne Hiltz and Sara J. Czaja. Introduction to the special issue on information systems for an aging society. ACM Transactions on Computer-Human Interaction, 13(3):309–312, September 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hayes:2014:EMC

[HCH^{+}14] Gillian R. Hayes, Karen G.

Cheng, Sen H. Hirano, Karen P. Tang, Marni S. Nagel, and Dianne E. Baker. Estrellita: a mobile capture and access tool for the support of preterm infants and their caregivers. ACM Transactions on Computer-Human Interaction, 21(3):19:1–19:??, June 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Han:2023:LAP

[HCS^{+}23] Jiawen Han, George Chernyshov, Moe Sugawa, Dingding Zheng, Danny Hynds, Taichi Furukawa, Marcelo Padovani Macieira, Karola Marky, Kouta Minamizawa, Jamie A. Ward, and Kai Kunze. Linking audience physiology to choreography. ACM Transactions on Computer-Human Interaction, 30(1):9:1–9:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557887.

Howell:2021:CSN

[HDF21] Noura Howell, Audrey Desjardins, and Sarah Fox. Cracks in the success narrative: Rethinking failure in design research through a retrospective trioethnography. ACM Transactions on Computer-Human Interaction, 28(6):42:1–42:31, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3462447.

Hiltz:2011:ISM

[HDM11] Starr Roxanne Hiltz, Paloma Diaz, and Gloria Mark. Introduction: Social media and collaborative systems for crisis management. ACM Transactions on Computer-Human Interaction, 18(4):18:1–18:??, December 2011. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hertzum:2024:CRT

[Her24] Morten Hertzum. Concurrent or retrospective thinking aloud in usability tests: a metaanalytic review. ACM Transactions on Computer-Human Interaction, 31(3):37:1–37:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665327.

Hertzum:1996:BQO

[HF96] Morten Hertzum and Erik Frøkjær. Browsing and querying in online documentation: a study of user interfaces and the interaction process. ACM Transactions on Computer-Human Interaction, 3(2):136– 161, June 1996. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-2/p136-hertzum/

.

Hornbaek:2003:RPU

[HF03] Kasper Hornbæk and Erik Frøkjær. Reading patterns and usability in visualizations of electronic documents. ACM Transactions on Computer-Human Interaction, 10(2):119– 149, 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hundhausen:2009:CDM

[HFB09] Christopher D. Hundhausen, Sean F. Farley, and Jonathan L. Brown. Can direct manipulation lower the barriers to computer programming and promote transfer of training? An experimental study. ACM Transactions on Computer-Human Interaction, 16(3): 13:1–13:??, September 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hindmarsh:2000:OFI

[HFH^{+}00] Jon Hindmarsh, Mike Fraser, Christian Heath, Steve Benford, and Chris Greenhalgh. Object-focused interaction in collaborative virtual environments. ACM Transactions on Computer-Human Interaction, 7(4):477–509, 2000. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org/
                                        pubs/articles/journals/tochi/
                                        2000-7-4/p477-hindmarsh/
                                        p477-hindmarsh.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tochi/2000-7-4/
                                        p477-hindmarsh/.

Hirzle:2022:UAA

[HFK^{+}22] Teresa Hirzle, Fabian Fischbach, Julian Karlbauer, Pascal Jansen, Jan Gugenheimer, Enrico Rukzio, and Andreas Bulling. Understanding, addressing, and analysing digital eye strain in virtual reality head-mounted displays. ACM Transactions on Computer-Human Interaction, 29(4): 33:1–33:80, August 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3492802.

Hodge:2023:EPR

[HFLH^{+}23] James Hodge, Sarah Foley, Dan Lambton-Howard, Laura Booi, Kyle Montague, Sandra Coulter, David Kirk, and Kellie Morrissey. Exploring participants’ representations and shifting sensitivities in a hackathon for dementia. ACM Transactions on Computer-Human Interaction, 30(3):46:1–46:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571814.

Hundhausen:2012:ESP

[HFP12] C. D. Hundhausen, D. Fairbrother, and M. Petre. An empirical study of the “prototype walkthrough”: a studiobased activity for HCI education. ACM Transactions on Computer-Human Interaction, 19(4):26:1–26:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Heuer:2024:RCN

[HG24] Hendrik Heuer and Elena L. Glassman. Reliability criteria for news websites. ACM Transactions on Computer-Human Interaction, 31(2):21:1–21:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3635147.

Hajika:2024:RWW

[HGH^{+}24] Ryo Hajika, Tamil Selvan Gunasekaran, Chloe Dolma Si Ying Haigh, Yun Suen Pai, Eiji Hayashi, Jaime Lien, Danielle Lottridge, and Mark Billinghurst. RadarHand: a wrist-worn radar for onskin touch-based proprioceptive gestures. ACM Transactions on Computer-Human Interaction, 31(2):17:1–17:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617365.

Hibschman:2019:ISS

[HGOZ19] Joshua Hibschman, Darren Gergle, Eleanor O’Rourke, and Haoqi Zhang. Isopleth: Supporting sensemaking of professional Web applications to create readily available learning experiences. ACM Transactions on Computer-Human Interaction, 26(3):16:1–16:??, June 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3310274.

Hornbaek:2007:UUF

[HH07] Kasper Hornbæk and Morten Hertzum. Untangling the usability of fisheye menus. ACM Transactions on Computer-Human Interaction, 14(2):6:1– 6:??, August 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hornbaek:2017:TAU

[HH17] Kasper Hornbæk and Morten Hertzum. Technology acceptance and user experience: a review of the experiential component in HCI. ACM Transactions on Computer-Human Interaction, 24(5):33:1–33:??, November 2017. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hertzum:2023:FSC

[HH23] Morten Hertzum and Kasper Hornbæk. Frustration: Still a common user experience. ACM Transactions on Computer-Human Interaction, 30(3): 42:1–42:??, June 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3582432.

Hassenzahl:2012:AYN

[HHE^{+}12] Marc Hassenzahl, Stephanie Heidecker, Kai Eckoldt, Sarah Diefenbach, and Uwe Hillmann. All you need is love: Current strategies of mediating intimate relationships through technology. ACM Transactions on Computer-Human Interaction, 19(4):30:1–30:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Han:2022:SBS

[HHG^{+}22] Dongqi Han, Yasamin Heshmat, Denise Y. Geiskkovitch, Zixuan Tan, and Carman Neustaedter. A scenario-based study of doctors and patients on video conferencing appointments from home. ACM Transactions on Computer-Human Interaction, 29(5): 49:1–49:??, October 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514234.

Hornecker:2024:DVD

[HHHV24] Eva Hornecker, Trevor Hogan, Uta Hinrichs, and Rosa Van Koningsbruggen. A design vocabulary for data physicalization. ACM Transactions on Computer-Human Interaction, 31(1):2:1–2:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617366.

Hollan:2000:DCT

[HHK00] James Hollan, Edwin Hutchins, and David Kirsh. Distributed cognition: toward a new foundation for human-computer interaction research. ACM Transactions on Computer-Human Interaction, 7(2):174– 196, June 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
2000-7-2/p174-hollan/.

Hinckley:2016:ESTb

[Hin16a] K. Hinckley. The Editor’s spotlight: TOCHI issue 23:2. ACM Transactions on Computer-Human Interaction, 23(2):7:1– 7:??, May 2016. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2016:EWN

[Hin16b]  Ken Hinckley.  Editorial:

Welcome to a new era for TOCHI. ACM Transactions on Computer-Human Interaction, 23(1):1:1–1:??, February 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2016:ESTa

[Hin16c]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 23:1. ACM Transactions on Computer-Human Interaction, 23(1):1:1–1:??, February 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2016:ESTc

[Hin16d]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 23:3. ACM Transactions on Computer-Human Interaction, 23(3):13:1–13:??, July 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2016:ESTd

[Hin16e]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 23:4. ACM Transactions on Computer-Human Interaction, 23(4):20:1–20:??, September 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2016:ESTe

[Hin16f]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 23:5. ACM Transactions on Computer-Human Interaction, 23(5):27:1–27:??, November 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2016:ESTf

[Hin16g]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 23:6. ACM Transactions on Computer-Human Interaction, 23(6):34:1–34:??, December 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2017:ESTa

[Hin17a]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 24:1. ACM Transactions on Computer-Human Interaction, 24(1):1:1–1:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2017:ESTb

[Hin17b]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 24:2 extravaganza-special issue on end-user design for the Internet of things, and the TOCHI best paper award 2016. ACM Transactions on Computer-Human Interaction, 24(2):8:1–8:??, May 2017. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2017:ESTc

[Hin17c]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 24:3. ACM Transactions on Computer-Human Interaction, 24(3):18:1–18:??, July 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2017:ESTd

[Hin17d]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 24:4. ACM Transactions on Computer-Human Interaction, 24(4):25:1–25:??, September 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2017:EST

[Hin17e]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 24:5. ACM Transactions on Computer-Human Interaction, 24(5):31:1–31:??, November 2017. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2018:ESTa

[Hin18a]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 24:6. ACM Transactions on Computer-Human Interaction, 24(6):37:1–37:??, January 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2018:ESTb

[Hin18b]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 25:2. ACM Transactions on Computer-Human Interaction, 25(2):7:1–7:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2018:ESTc

[Hin18c]  Ken Hinckley.  The Edi-

tor’s spotlight: TOCHI issue 25:3. ACM Transactions on Computer-Human Interaction, 25(3):14:1–14:??, June 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hwang:2020:EHP

[HJS^{+}20] Amy S. Hwang, Piper Jackson, Andrew Sixsmith, Louise Nygård, Arlene Astell, Khai N. Truong, and Alex Mihailidis. Exploring how persons with dementia and care partners collaboratively appropriate information and communication technologies. ACM Transactions on Computer-Human Interaction, 27(6):46:1–46:38, November 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3389377.

Hahn:1999:WSD

[HK99] Jungpil Hahn and Jinwoo Kim.

Why are some diagrams easier to work with? effects of diagrammatic representation on the cognitive integration process of systems analysis and design. ACM Transactions on Computer-Human Interaction, 6(3):181–213, September 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-3/p181-hahn/.

Hook:2012:SCI

[HL12] Kristina Höök and Jonas Löwgren. Strong concepts: Intermediate-level knowledge in interaction design research. ACM Transactions on Computer-Human Interaction, 19(3):23:1–23:??, October 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Heintz:2018:DEG

[HL18] Stephanie Heintz and Effie L.-C. Law. Digital educational games: Methodologies for evaluating the impact of game type. ACM Transactions on Computer-Human Interaction, 25(2):8:1–8:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Halskov:2021:FID

[HL21] Kim Halskov and Caroline Lundqvist. Filtering and informing the design space: Towards design-space thinking. ACM Transactions on Computer-Human Interaction, 28(1):8:1–8:28, February 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3434462.

Harburg:2018:CFO

[HLEG18] Emily Harburg, Daniel Rees Lewis, Matthew Easterday, and Elizabeth M. Gerber. CheerOn: Facilitating online social support for novice project-based learning teams. ACM Transactions on Computer-Human Interaction, 25(6): 32:1–32:??, December 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Huguenard:1997:WFP

[HLJ^{+}97] Brian R. Huguenard, F. Javier Lerch, Brian W. Junker, Richard J. Patz, and Robert E. Kass. Working-memory failure in phone-based interaction. ACM Transactions on Computer-Human Interaction, 4(2):67–102, June 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-2/p67-huguenard/

.

Huotari:2004:IGI

[HLN04] Jouni Huotari, Kalle Lyytinen, and Marketta Niemelä. Improving graphical information system model use with elision and connecting lines. ACM Transactions on Computer-Human Interaction, 11(1):26– 58, March 2004. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Huh:2014:HVS

[HLN^{+}14] Jina Huh, Leslie S. Liu, Tina Neogi, Kori Inkpen, and Wanda Pratt. Health vlogs as social support for chronic illness management. ACM Transactions on Computer-Human Interaction, 21(4): 23:1–23:??, August 2014. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hong:2024:VNC

[HLSR24] Junlei Hong, Tobias Langlotz, Jonathan Sutton, and Holger Regenbrecht. Visual noise cancellation: Exploring visual discomfort and opportunities for vision augmentations. ACM Transactions on Computer-Human Interaction, 31(2):22:1–22:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3634699.

Haimson:2023:UPH

[HM23] Oliver L. Haimson and Megh Marathe. Uncovering personal histories: a technologymediated approach to eliciting reflection on identity transitions. ACM Transactions on Computer-Human Interaction, 30(2):23:1–23:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3504004.

Hornbaek:2019:WDW

[HMKV19] Kasper Hornbæk, Aske Mottelson, Jarrod Knibbe, and Daniel Vogel. What do we mean by “interaction”? An analysis of 35 years of CHI. ACM Transactions on Computer-Human Interaction, 26(4):27:1–27:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3325285.

Hornof:2001:VSM

[Hor01] Anthony J. Hornof. Visual search and mouse-pointing in labeled versus unlabeled two-dimensional visual hierarchies. ACM Transactions on Computer-Human Interaction, 8(3):171–197, September 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hornecker:2016:FSM

[Hor16] Eva Hornecker. The to-and-fro of sense making: Supporting users’ active indexing in museums. ACM Transactions on Computer-Human Interaction, 23(2):10:1–10:??, May 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:2005:FBI

[HPHS05] Ken Hinckley, Jeff Pierce, Eric Horvitz, and Mike Sinclair. Foreground and background interaction with sensor-enhanced mobile devices. ACM Transactions on Computer-Human Interaction, 12(1):31–52, March 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hinckley:1998:TVM

[HPPK98] Ken Hinckley, Randy Pausch, Dennis Proffitt, and Neal F. Kassell. Two-handed virtual manipulation. ACM Transactions on Computer-Human Interaction, 5(3):260– 302, September 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1998-5-3/p260-hinckley/

.

Hallnas:2002:UPE

[HR02] Lars Hallnäs and Johan Redström. From use to presence: on the expressions and aesthetics of everyday computational things. ACM Transactions on Computer-Human Interaction, 9(2):106–124, June 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hu:2014:CMT

[HRB14] Chang Hu, Philip Resnik, and Benjamin B. Bederson. Crowdsourced monolingual translation. ACM Transactions on Computer-Human Interaction, 21(4):22:1–22:??, August 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Harrison:2023:DTM

[HRW^{+}23] Daniel Harrison, Scarlett Rowland, Gavin Wood, Lyndsey Bakewell, Ioannis Petridis, Kiel Long, Konstantina Vasileiou, Julie Barnett, Manuela Barreto, Michael Wilson, Shaun Lawson, and John Vines. Designing technology-mediated peer support for postgraduate research students at risk of loneliness and isolation. ACM Transactions on Computer-Human Interaction, 30(1):10:1–10:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534961.

Hossain:2023:BAQ

[HSA^{+}23] Tahera Hossain, Wanggang Shen, Anindya Antar, Snehal Prabhudesai, Sozo Inoue, Xun Huan, and Nikola Banovic. A Bayesian approach for quantifying data scarcity when modeling human behavior via inverse reinforcement learning. ACM Transactions on Computer-Human Interaction, 30(1):8:1–8:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551388.

Hartmann:2008:TTU

[HSD08] Jan Hartmann, Alistair Sutcliffe, and Antonella De Angeli. Towards a theory of user judgment of aesthetics and user interface quality. ACM Transactions on Computer-Human Interaction, 15(4):15:1–15:??, November 2008. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hoyle:2020:PNP

[HSI^{+}20] Roberto Hoyle, Luke Stark, Qatrunnada Ismail, David Crandall, Apu Kapadia, and Denise Anthony. Privacy norms and preferences for photos posted online. ACM Transactions on Computer-Human Interaction, 27(4):30:1–30:27, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3380960.

Hamdy:2011:HPB

[HT11] Omar Hamdy and Issa Traoré.

Homogeneous physio-behavioral visual and mouse-based biometric. ACM Transactions on Computer-Human Interaction, 18(3):12:1–12:??, July 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hartzler:2016:DUI

[HWC^{+}16] Andrea L. Hartzler, Bridget Weis, Carly Cahill, Wanda Pratt, Albert Park, Uba Backonja, and David W. Mc-Donald. Design and usability of interactive user profiles for online health communities. ACM Transactions on Computer-Human Interaction, 23(3):15:1–15:??, July 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Hubona:1999:RCS

[HWSB99] Geoffrey S. Hubona, Philip N. Wheeler, Gregory W. Shirah, and Matthew Brandt. The relative contributions of stereo, lighting, and background scenes in promoting 3D depth visualization. ACM Transactions on Computer-Human Interaction, 6(3):214– 242, September 1999. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tochi/1999-6-3/p214-hubona/

.

Iachello:2008:PMP

[IA08] Giovanni Iachello and Gregory D. Abowd. From privacy methods to a privacy toolbox: Evaluation shows that heuristics are complementary. ACM Transactions on Computer-Human Interaction, 15(2):8:1–8:??, July 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Iqbal:2010:OFL

[IB10] Shamsi T. Iqbal and Brian P.

Bailey. Oasis: a framework for linking notification delivery to the perceptual structure of goal-directed tasks. ACM Transactions on Computer-Human Interaction, 17(4): 15:1–15:??, December 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Isbister:2022:DLT

[ICC^{+}22] Katherine Isbister, Peter Cottrell, Alessia Cecchet, Ella Dagan, Nikki Theofanopoulou, Ferran Altarriba Bertran, Aaron J. Horowitz, Nick Mead, Joel B. Schwartz, and Petr Slovak. Design (not) lost in translation: a case study of an intimate-space socially assistive “robot” for emotion regulation. ACM Transactions on Computer-Human Interaction, 29(4):32:1–32:36, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491083.

Inkpen:2023:AHA

[ICM^{+}23] Kori Inkpen, Shreya Chappidi, Keri Mallari, Besmira Nushi, Divya Ramesh, Pietro Michelucci, Vani Mandava, Libuse Hannah Veprek, and Gabrielle Quinn. Advancing Human-AI complementarity: The impact of user expertise and algorithmic tuning on joint decision making. ACM Transactions on Computer-Human Interaction, 30(5): 71:1–71:??, October 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3534561.

Inie:2020:HID

[ID20] Nanna Inie and Peter Dalsgaard. How interaction designers use tools to manage ideas. ACM Transactions on Computer-Human Interaction, 27(2):7:1–7:26, April 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3365104.

Ichino:2024:HGV

[IIY^{+}24] Junko Ichino, Masahiro Ide, Takehito Yoshiki, Hitomi Yokoyama, Hirotoshi Asano, Hideo Miyachi, and Daisuke Okabe. How gaze visualization facilitates initiation of informal communication in 3D virtual spaces. ACM Transactions on Computer-Human Interaction, 31(1):5:1–5:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617368.

Inkpen:2001:DDV

[Ink01] Kori M. Inkpen. Drag-and-drop versus point-and-click mouse interaction styles for children. ACM Transactions on Computer-Human Interaction, 8(1):1–33, 2001. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
tochi/2001-8-1/p1-inkpen/

.

Irani:2003:DIS

[IW03] Pourang Irani and Colin Ware. Diagramming information structures using 3D perceptual primitives. ACM Transactions on Computer-Human Interaction, 10(1):1– 19, 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jun:2022:HFE

[JBD^{+}22] Eunice Jun, Melissa Birchfield, Nicole De Moura, Jeffrey Heer, and René Just. Hypothesis formalization: Empirical findings, software limitations, and design implications. ACM Transactions on Computer-Human Interaction, 29(1):6:1– 6:28, February 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3476980.

Jhaver:2019:HMC

[JBGB19] Shagun Jhaver, Iris Birman, Eric Gilbert, and Amy Bruckman. Human-machine collaboration for content regulation: The case of Reddit Automoderator. ACM Transactions on Computer-Human Interaction, 26(5):31:1–31:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3338243.

Jay:2013:PWU

[JBH13] Caroline Jay, Andy Brown, and Simon Harper. Predicting whether users view dynamic content on the World Wide Web. ACM Transactions on Computer-Human Interaction, 20(2):9:1–9:??, May 2013. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jacob:1999:SMS

[JDM99] Robert J. K. Jacob, Leonidas Deligiannidis, and Stephen Morrison. A software model and specification language for non-WIMP user interfaces. ACM Transactions on Computer-Human Interaction, 6(1):1–46, March 1999. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tochi/1999-6-1/p1-jacob/.

Jacucci:2021:ERE

[JDV^{+}21] Giulio Jacucci, Pedram Daee, Tung Vuong, Salvatore Andolina, Khalil Klouche, Mats SjÖberg, Tuukka Ruotsalo, and Samuel Kaski. Entity recommendation for everyday digital tasks. ACM Transactions on Computer-Human Interaction, 28(5):29:1–29:41, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458919.

Jhaver:2018:OHC

[JGBG18] Shagun Jhaver, Sucheta Ghoshal, Amy Bruckman, and Eric Gilbert. Online harassment and content moderation: The case of blocklists. ACM Transactions on Computer-Human Interaction, 25(2):12:1–12:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jay:2007:MED

[JGH07] Caroline Jay, Mashhuda Glencross, and Roger Hubbold. Modeling the effects of delayed haptic and visual feedback in a collaborative virtual environment. ACM Transactions on Computer-Human Interaction, 14(2):8:1–8:??, August 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Johnson:2003:TMI

[JH03] Hilary Johnson and Joanne Hyde. Towards modeling individual and collaborative construction of jigsaws using task knowledge structures (TKS). ACM Transactions on Computer-Human Interaction, 10(4):339–387, December 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jakobsen:2014:CPC

[JH14] Mikkel R. Jakobsen and Kasper Hornbæk. Up close and personal: Collaborative work on a high-resolution multitouch wall display. ACM Transactions on Computer-Human Interaction, 21(2): 11:1–11:??, April 2014. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

John:1996:GFU

[JK96a] Bonnie E. John and David E. Kieras. The GOMS family of user interface analysis techniques: comparison and contrast. ACM Transactions on Computer-Human Interaction, 3(4):320– 351, December 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-4/p320-john/

.

John:1996:UGU

[JK96b] Bonnie E. John and David E. Kieras. Using GOMS for user interface design and evaluation: which technique? ACM Transactions on Computer-Human Interaction, 3(4):287– 319, December 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-4/p287-john/

.

Jensen:2018:ASE

[JKS18] Rikke Hagensby Jensen, Jesper Kjeldskov, and Mikael B. Skov. Assisted shifting of electricity use: a long-term study of managing residential heating. ACM Transactions on Computer-Human Interaction, 25(5):25:1–25:??, October 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3210310.

Jahanian:2017:CMC

[JKVA17] Ali Jahanian, Shaiyan Keshvari, S. V. N. Vishwanathan, and Jan P. Allebach. Colors — messengers of concepts: Visual design mining for learning color semantics. ACM Transactions on Computer-Human Interaction, 24(1):2:1–2:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Johnson:2003:IMC

[JMJ03] Peter Johnson, Jon May, and Hilary Johnson. Introduction to multiple and collaborative tasks. ACM Transactions on Computer-Human Interaction, 10(4):277–280, December 2003. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jamil:2017:CAD

[JMP^{+}17] Izdihar Jamil, Calkin Suero Montero, Mark Perry, Kenton O’Hara, Abhijit Karnik, Kaisa Pihlainen, Mark T. Marshall, Swathi Jha, Sanjay Gupta, and Sriram Subramanian. Collaborating around digital tabletops: Children’s physical strategies from India, the UK and Finland. ACM Transactions on Computer-Human Interaction, 24(3): 23:1–23:??, July 2017. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Johnson:1996:CPS

[JN96] Jeff A. Johnson and Bonnie A. Nardi. Creating presentation slides: a study of user preferences for task-specific versus generic application software. ACM Transactions on Computer-Human Interaction, 3(1):38–65, March 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1996-3-1/p38-johnson/

.

Jiang:2023:TCF

[JNBF23] Jialun Aaron Jiang, Peipei Nie, Jed R. Brubaker, and Casey Fiesler. A trade-off-centered framework of content moderation. ACM Transactions on Computer-Human Interaction, 30(1):3:1–3:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534929.

Jones:2005:I

[JNM05] Matt Jones, Bonnie Nardi, and Elizabeth D. Mynatt. Introduction. ACM Transactions on Computer-Human Interaction, 12(2):147–148, June 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jardine:2024:BRR

[JNR^{+}24] Jacinta Jardine, Camille Nadal, Sarah Robinson, Angel Enrique, Marcus Hanratty, and Gavin Doherty. Between rhetoric and reality: Realworld barriers to uptake and early engagement in digital mental health interventions. ACM Transactions on Computer-Human Interaction, 31(2):27:1–27:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635472.

Jung:2020:RGR

[JPM^{+}20] Hee-Tae Jung, Taiwoo Park, Narges MAhyar, Sungji Park, Taekyeong Ryu, Yangsoo Kim, and Sunghoon Ivan Lee. Rehabilitation games in realworld clinical settings: Practices, challenges, and opportunities. ACM Transactions on Computer-Human Interaction, 27(6):41:1–41:43, December 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418197.

Jakobi:2019:IAW

[JPR^{+}19] Timo Jakobi, Sameer Patil, Dave Randall, Gunnar Stevens, and Volker Wulf. It is about what they could do with the data: a user perspective on privacy in smart metering. ACM Transactions on Computer-Human Interaction, 26(1):2:1–2:??, February 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3281444.

Johansson:2024:WLS

[JRB^{+}24] Karin Johansson, Raquel Robinson, Jon Back, Sarah Lynne Bowman, James Fey, Elena Márquez Segura, Annika Waern, and Katherine Isbister. Why Larp? A synthesis article on live action roleplay in relation to HCI research and practice. ACM Transactions on Computer-Human Interaction, 31(5):64:1–64:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3689045.

Janlert:2010:CI

[JS10] Lars-Erik Janlert and Erik Stolterman. Complex interaction. ACM Transactions on Computer-Human Interaction, 17(2):8:1–8:??, May 2010. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jin:2021:LPR

[JSJ^{+}21] Haojian Jin, Hong Shen, Mayank Jain, Swarun Kumar, and Jason I. Hong. Lean privacy review: Collecting users’ privacy concerns of data practices at a low cost. ACM Transactions on Computer-Human Interaction, 28(5): 34:1–34:55, October 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3463910.

Jacob:1994:ISI

[JSM^{+}94] Robert J. K. Jacob, Linda E.

Sibert, Daniel C. McFarlane, M. Preston, and J. R. Mullen. Integrality and separability of input devices. ACM Transactions on Computer-Human Interaction, 1(1):3– 26, March 1994. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-1/p3-jacob/.

Jung:2016:CIP

[Jun16] Malte F. Jung. Coupling interactions and performance: Predicting team performance from thin slices of conflict. ACM Transactions on Computer-Human Interaction, 23(3): 18:1–18:??, July 2016. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Jeon:2012:SSI

[JWS12] Myounghoon Jeon, Bruce N. Walker, and Abhishek Srivastava. “Spindex” (Speech Index) enhances menus on touch screen devices with tapping, wheeling, and flicking. ACM Transactions on Computer-Human Interaction, 19(2): 14:1–14:??, July 2012. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Katz:2003:ESB

[KB03] Michael A. Katz and Michael D. Byrne. Effects of scent and breadth on use of site-specific search on e-commerce Web sites. ACM Transactions on Computer-Human Interaction, 10(3):198–220, September 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kahn:2021:SEI

[KB21] Zoe Kahn and Jenna Burrell. A sociocultural explanation of Internet-enabled work in rural regions. ACM Transactions on Computer-Human Interaction, 28(3):17:1–17:22, July 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3443705.

Knijnenburg:2023:DAF

[KB23] Bart P. Knijnenburg and Burcu Bulgurcu. Designing alternative form-autocompletion tools to enhance privacy decision-making and prevent unintended disclosure. ACM Transactions on Computer-Human Interaction, 30(6): 91:1–91:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610366.

Kim:2017:BIC

[KBB^{+}17] Nam Wook Kim, Zoya Bylinskii, Michelle A. Borkin, Krzysztof Z. Gajos, Aude Oliva, Fredo Durand, and Hanspeter Pfister. Bubble-View: an interface for crowdsourcing image importance maps and tracking visual attention. ACM Transactions on Computer-Human Interaction, 24(5):36:1–36:??, November 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Katzeff:2013:ESP

[KBJ^{+}13] Cecilia Katzeff, Loove Broms, Li Jönsson, Ulrika Westholm, and Minna Räsänen. Exploring sustainable practices in workplace settings through visualizing electricity consumption. ACM Transactions on Computer-Human Interaction, 20(5):31:1–31:??, November 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Knowles:2023:PPC

[KC23] Bran Knowles and Stacey Conchie. Un-paradoxing privacy: Considering hopeful trust. ACM Transactions on Computer-Human Interaction, 30(6):87:1–87:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3609329.

Ko:2016:UMI

[KCL^{+}16] Minsam Ko, Seungwoo Choi, Joonwon Lee, Uichin Lee, and Aviv Segev. Understanding mass interactions in online sports viewing: Chatting motives and usage patterns. ACM Transactions on Computer-Human Interaction, 23(1):6:1–6:??, February 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kuijer:2013:PUD

[KdJvE13] Lenneke Kuijer, Annelise de Jong, and Daan van Eijk. Practices as a unit of design: an exploration of theoretical guidelines in a study on bathing. ACM Transactions on Computer-Human Interaction, 20(4):21:1–21:??, September 2013. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Klar:2023:SIM

[KFF^{+}23] Markus Klar, Florian Fischer, Arthur Fleig, Miroslav Bachinski, and Jörg Müller. Simulating interaction movements via model predictive control. ACM Transactions on Computer-Human Interaction, 30(3):44:1–44:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3577016.

Karran:2015:FPC

[KFG15] Alexander J. Karran, Stephen H.

Fairclough, and Kiel Gilleade. A framework for psychophysiological classification within a cultural heritage context using interest. ACM Transactions on Computer-Human Interaction, 21(6):34:1–34:??, January 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kim:2015:ATN

[KGYQ15] Ji-Sun Kim, Denis Gracanin, Taeyoung Yang, and Francis Quek. Action-transferred navigation technique design approach supporting human spatial learning. ACM Transactions on Computer-Human Interaction, 22(6):30:1–30:??, December 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kennaway:2007:PSC

[KGZ07] J. R. Kennaway, J. R. W.

Glauert, and I. Zwitserlood. Providing signed content on the Internet by synthesized animation. ACM Transactions on Computer-Human Interaction, 14(3):15:1–15:??, September 2007. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Knowles:2018:OAD

[KH18] Bran Knowles and Vicki L. Hanson. Older adults’ deployment of ‘distrust’. ACM Transactions on Computer-Human Interaction, 25(4): 21:1–21:??, September 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3196490.

Kuo:2023:ELI

[KH23] Pei-Yi (Patricia) Kuo and Michael S. Horn. EcoSanté lifestyle intervention: Encourage reflections on the connections between health and environment. ACM Transactions on Computer-Human Interaction, 30(6):88:1–88:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3609325.

Kin:2011:THM

[KHA11] Kenrick Kin, Björn Hartmann, and Maneesh Agrawala. Twohanded marking menus for multitouch devices. ACM Transactions on Computer-Human Interaction, 18(3): 16:1–16:??, July 2011. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Khot:2020:SFD

[KHM20] Rohit Ashok Khot, Larissa Hjorth, and Florian Mueller. Shelfie: a framework for designing material representations of physical activity data. ACM Transactions on Computer-Human Interaction, 27(3):14:1–14:52, June 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3379539.

Kessler:1995:ECW

[KHW95] G. Drew Kessler, Larry F.

Hodges, and Neff Walker. Evaluation of the Cyber-Glove as a whole-hand input device. ACM Transactions on Computer-Human Interaction, 2(4):263–283, December 1995. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-4/p263-kessler/

.

Kirsh:2013:ECM

[Kir13] David Kirsh. Embodied cognition and the magical future of interaction design. ACM Transactions on Computer-Human Interaction, 20(1):3:1– 3:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kirsh:2019:DAD

[Kir19] David Kirsh. Do architects and designers think about interactivity differently? ACM Transactions on Computer-Human Interaction, 26(2):7:1–7:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301425.

Konrad:2016:TMM

[KIW16] Artie Konrad, Ellen Isaacs, and Steve Whittaker. Technologymediated memory: Is technology altering our memories and interfering with wellbeing? ACM Transactions on Computer-Human Interaction, 23(4):23:1–23:??, September 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kokkalis:2013:TAP

[KKH^{+}13] Nicolas Kokkalis, Thomas Köhn, Johannes Huebner, Moontae Lee, Florian Schulze, and Scott R. Klemmer. TaskGenies: Automatically providing action plans helps people complete tasks. ACM Transactions on Computer-Human Interaction, 20(5):27:1–27:??, November 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kumar:2020:TLH

[KKIT20] Neha Kumar, Naveena Karusala, Azra Ismail, and Anupriya Tuli. Taking the long, holistic, and intersectional view to women’s wellbeing. ACM Transactions on Computer-Human Interaction, 27(4): 23:1–23:32, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3397159.

Koh:2024:UUP

[KKJ24] Seunghun Koh, Byung Hyung Kim, and Sungho Jo. Understanding the user perception and experience of interactive algorithmic recourse customization. ACM Transactions on Computer-Human Interaction, 31(3):43:1–43:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674503.

Karnatak:2023:IEG

[KLK^{+}23] Nimisha Karnatak, Brooke Loughrin, Tiffany Amy Kuo, Odeline Mateu-Silvernail, Indrani Medhi Thies, William Thies, and Mohit Jain. “Is it Even Giving the Correct Reading or Not?”: How trust and relationships mediate blood pressure management in India. ACM Transactions on Computer-Human Interaction, 30(6):90:1–90:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3609327.

Koulouri:2012:WTY

[KLMC12] Theodora Koulouri, Stanislao Lauria, Robert D. Macredie, and Sherry Chen. Are we there yet?: The role of gender on the effectiveness and efficiency of user-robot communication in navigational tasks. ACM Transactions on Computer-Human Interaction, 19(1):4:1–4:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kim:1995:IRR

[KLS95] Jinwoo Kim, F. Javier Lerch, and Herbert A. Simon. Internal representation and rule development in object-oriented design. ACM Transactions on Computer-Human Interaction, 2(4):357–390, December 1995. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1995-2-4/p357-kim/.

Khan:2021:GGA

[KNK^{+}21] Anam Ahmad Khan, Joshua Newn, Ryan M. Kelly, Namrata Srivastava, James Bailey, and Eduardo Velloso. GAVIN: Gaze-assisted voice-based implicit note-taking. ACM Transactions on Computer-Human Interaction, 28(4): 26:1–26:32, October 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3453988.

Kostakos:2010:BES

[KOP^{+}10] Vassilis Kostakos, Eamonn O’Neill, Alan Penn, George Roussos, and Dikaios Papadongonas. Brief encounters: Sensing, modeling and visualizing urban mobility and copresence networks. ACM Transactions on Computer-Human Interaction, 17(1):2:1– 2:??, March 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kjeldskov:2010:IUM

[KP10] Jesper Kjeldskov and Jeni Paay. Indexicality: Understanding mobile humancomputer interaction in context. ACM Transactions on Computer-Human Interaction, 17(4):14:1–14:??, December 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kharrufa:2018:UMU

[KPO18] Ahmed Kharrufa, Thomas Ploetz, and Patrick Olivier. A unified model for user identification on multi-touch surfaces: a survey and metaanalysis. ACM Transactions on Computer-Human Interaction, 24(6):39:1–39:??, January 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Keyes:2020:RWH

[KPWS20] Os Keyes, Burren Peil, Rua M. Williams, and Katta Spiel. Reimagining (women’s) health: HCI, gender and essentialised embodiment. ACM Transactions on Computer-Human Interaction, 27(4):25:1–25:42, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3404218.

Kang:2022:ASC

[KQH^{+}22] Hyeonsu B. Kang, Xin Qian, Tom Hope, Dafna Shahaf, Joel Chan, and Aniket Kittur. Augmenting scientific creativity with an analogical search engine. ACM Transactions on Computer-Human Interaction, 29(6):57:1–57:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3530013.

Kaul:2021:AHT

[KRMS21] Oliver Beren Kaul, Michael Rohs, Marc Mogalle, and Benjamin Simon. Around-the-head tactile system for supporting micro navigation of people with visual impairments. ACM Transactions on Computer-Human Interaction, 28(4):27:1–27:35, October 2021. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458021.

Kirk:2010:HRV

[KS10] David S. Kirk and Abigail Sellen. On human remains: Values and practice in the home archiving of cherished objects. ACM Transactions on Computer-Human Interaction, 17(3):10:1–10:??, July 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kizilcec:2015:MLU

[KS15] René F. Kizilcec and Emily Schneider. Motivation as a lens to understand online learners: Toward data-driven design with the OLEI scale. ACM Transactions on Computer-Human Interaction, 22(2):6:1– 6:??, April 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Koleva:2020:DHG

[KSB^{+}20] Boriana Koleva, Jocelyn Spence, Steve Benford, Hyosun Kwon, Holger Schnädelbach, Emily Thorn, William Preston, Adrian Hazzard, Chris Greenhalgh, Matt Adams, Ju Row Farr, Nick Tandavanitj, Alice Angus, and Giles Lane. Designing hybrid gifts. ACM Transactions on Computer-Human Interaction, 27(5): 37:1–37:33, October 2020. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3398193.

Krauss:2024:WMX

[KSB^{+}24] Veronika Krauß, Pejman Saeghe, Alexander Boden, Mohamed Khamis, Mark McGill, Jan Gugenheimer, and Michael Nebeling. What makes XR dark? Examining emerging dark patterns in augmented and virtual reality through expert co-design. ACM Transactions on Computer-Human Interaction, 31(3):32:1–32:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660340.

Kulp:2021:TDC

[KSCB21] Leah Kulp, Aleksandra Sarcevic, Megan Cheng, and Randall S. Burd. Towards dynamic checklists: Understanding contexts of use and deriving requirements for context-driven adaptation. ACM Transactions on Computer-Human Interaction, 28(2):12:1–12:33, April 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3444947.

Karpashevich:2022:TOB

[KSG^{+}22] Pavel Karpashevich, Pedro Sanches, Rachael Garrett, Yoav Luft, Kelsey Cotton, Vasiliki Tsaknaki, and Kristina Höök. Touching our breathing through shape-change: Monster, organic other, or twisted mirror. ACM Transactions on Computer-Human Interaction, 29(3):22:1–22:40, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490498.

Kwon:2011:SCA

[KSJB11] Gyu Hyun Kwon, Tonya L.

Smith-Jackson, and Charles W. Bostian. Socio-cognitive aspects of interoperability: Understanding communication task environments among different organizations. ACM Transactions on Computer-Human Interaction, 18(4): 20:1–20:??, December 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Koike:2002:RIH

[KSK02] Hideki Koike, Yoichi Sato, and Yoshinori Kobayashi. Rivalry and interference with a head-mounted display. ACM Transactions on Computer-Human Interaction, 9(3):238– 251, September 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kuttal:2014:BPV

[KSR14] Sandeep K. Kuttal, Anita Sarma, and Gregg Rothermel. On the benefits of providing versioning support for end users: an empirical study. ACM Transactions on Computer-Human Interaction, 21(2):9:1–9:??, April 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kosmyna:2015:AHL

[KTBR15a] Nataliya Kosmyna, Franck Tarpin-Bernard, and Bertrand Rivet. Adding human learning in brain–computer interfaces (BCIs): Towards a practical control modality. ACM Transactions on Computer-Human Interaction, 22(3):12:1–12:??, June 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kosmyna:2015:CPG

[KTBR15b] Nataliya Kosmyna, Franck Tarpin-Bernard, and Bertrand Rivet. Conceptual priming for in-game BCI training. ACM Transactions on Computer-Human Interaction, 22(5): 26:1–26:??, October 2015. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kono:2018:DGD

[KTN^{+}18] Michinari Kono, Takumi Takahashi, Hiromi Nakamura, Takashi Miyaki, and Jun Rekimoto. Design guideline for developing safe systems that apply electricity to the human body. ACM Transactions on Computer-Human Interaction, 25(3):19:1–19:??, June 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Konstan:2015:TRS

[KWB^{+}15] Joseph A. Konstan, J. D.

Walker, D. Christopher Brooks, Keith Brown, and Michael D. Ekstrand. Teaching recommender systems at large scale: Evaluation and lessons learned from a hybrid MOOC. ACM Transactions on Computer-Human Interaction, 22(2): 10:1–10:??, April 2015. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kosch:2022:PEA

[KWCS22] Thomas Kosch, Robin Welsch, Lewis Chuang, and Albrecht Schmidt. The placebo effect of artificial intelligence in human-computer interaction. ACM Transactions on Computer-Human Interaction, 29(6):56:1–56:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529225.

Kulkarni:2013:PSA

[KWL^{+}13] Chinmay Kulkarni, Koh Pang Wei, Huy Le, Daniel Chia, Kathryn Papadopoulos, Justin Cheng, Daphne Koller, and Scott R. Klemmer. Peer and self assessment in massive online classes. ACM Transactions on Computer-Human Interaction, 20(6):33:1–33:??, December 2013. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

Kieras:1997:PEM

[KWM97] David E. Kieras, Scott D. Wood, and David E. Meyer. Predictive engineering models based on the EPIC architecture for a multimodal highperformance human-computer interaction task. ACM Transactions on Computer-Human Interaction, 4(3):230– 275, September 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1997-4-3/p230-kieras/

.

Kerne:2014:UMC

[KWS^{+}14] Andruid Kerne, Andrew M. Webb, Steven M. Smith, Rhema Linder, Nic Lupfer, Yin Qu, Jon Moeller, and Sashikanth Damaraju. Using metrics of curation to evaluate information-based ideation. ACM Transactions on Computer-Human Interaction, 21(3):14:1–14:??, June 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Kim:2023:WAE

[KYZ23] Antino Kim, Mochen Yang, and Jingjing Zhang. When algorithms err: Differential impact of early vs. late errors on users’ reliance on algorithms. ACM Transactions on Computer-Human Interaction, 30(1):14:1–14:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557889.

Kong:2006:SGG

[KZZ06] Jun Kong, Kang Zhang, and Xiaoqin Zeng. Spatial graph grammars for graphical user interfaces. ACM Transactions on Computer-Human Interaction, 13(2):268–307, June 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Leung:1994:RTD

[LA94] Y. K. Leung and M. D. Aerley. A review and taxonomy of distortion-oriented presentation techniques. ACM Transactions on Computer-Human Interaction, 1(2):126– 160, June 1994. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-2/p126-leung/

.

Leiva:2015:AIJ

[LA15] Luis A. Leiva and Vicent Alabau. Automatic internationalization for just in time localization of Web-based user interfaces. ACM Transactions on Computer-Human Interaction, 22(3):13:1–13:??, June 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Long:2017:EGI

[LA17] Yanjin Long and Vincent Aleven. Educational game and intelligent tutoring system: a classroom study and comparative design analysis. ACM Transactions on Computer-Human Interaction, 24(3): 20:1–20:??, July 2017. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lakier:2018:ADG

[LAW18] Matthew Lakier, Michelle Annett, and Daniel Wigdor. Automatics: Dynamically generating fabrication tasks to adapt to varying contexts. ACM Transactions on Computer-Human Interaction, 25(4):22:1–22:??, September 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3185065.

Lee:2010:IDM

[LB10] Young Eun Lee and Izak Benbasat. Interaction design for mobile product recommendation agents: Supporting users’ decisions in retail stores. ACM Transactions on Computer-Human Interaction, 17(4):17:1–17:??, December 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lascau:2024:SLP

[LBGC24] Laura Lascau, Duncan P.

Brumby, Sandy J. J. Gould, and Anna L. Cox. “Sometimes It’s Like Putting the Track in Front of the Rushing Train”: Having to be ‘on call’ for work limits the temporal flexibility of crowdworkers. ACM Transactions on Computer-Human Interaction, 31(2):18:1–18:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635145.

Liu:2015:LSE

[LBO^{+}15] Yun-En Liu, Christy Ballweber, Eleanor O’Rourke, Eric Butler, Phonraphee Thummaphan, and Zoran Popović. Largescale educational campaigns. ACM Transactions on Computer-Human Interaction, 22(2):8:1– 8:??, April 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ludwig:2017:PST

[LBP17] Thomas Ludwig, Alexander Boden, and Volkmar Pipek. 3D printers as sociable technologies: Taking appropriation infrastructures to the Internet of Things. ACM Transactions on Computer-Human Interaction, 24(2):17:1–17:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lim:1996:EII

[LBT96] Kai H. Lim, Izak Benbasat, and Peter A. Todd. An experimental investigation of the interactive effects of interface style, instructions, and task familiarity on user performance. ACM Transactions on Computer-Human Interaction, 3(1):1–37, March 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-1/p1-lim/.

Li:2023:FIE

[LCC23] Yue Li, Eugene Ch’ng, and Sue Cobb. Factors influencing engagement in hybrid virtual and augmented reality. ACM Transactions on Computer-Human Interaction, 30(4): 65:1–65:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3589952.

Lin:2019:SSP

[LCE^{+}19] Tian Lin, Daniel E. Capecci, Donovan M. Ellis, Harold A. Rocha, Sandeep Dommaraju, Daniela S. Oliveira, and Natalie C. Ebner. Susceptibility to spear-phishing emails: Effects of Internet user demographics and email content. ACM Transactions on Computer-Human Interaction, 26(5):32:1–32:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3336141.

Lindtner:2011:TFP

[LCHD11] Silvia Lindtner, Judy Chen, Gillian R. Hayes, and Paul Dourish. Towards a framework of publics: Re-encountering media sharing and its user. ACM Transactions on Computer-Human Interaction, 18(2):5:1– 5:??, June 2011. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Li:2024:GDI

[LCK24] Jingyi Li, Nathan Crilly, and Per Ola Kristensson. Guiding the design of inclusive interactive systems: Do younger and older adults use the same image-schematic metaphors? ACM Transactions on Computer-Human Interaction, 31(4):47:1–47:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3648618.

Li:2012:UCR

[LDF12] Ian Li, Anind K. Dey, and Jodi Forlizzi. Using context to reveal factors that affect physical activity. ACM Transactions on Computer-Human Interaction, 19(1):7:1–7:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lindgaard:2011:ERB

[LDS^{+}11] Gitte Lindgaard, Cathy Dudek, Devjani Sen, Livia Sumegi, and Patrick Noonan. An exploration of relations between visual appeal, trustworthiness and perceived usability of homepages. ACM Transactions on Computer-Human Interaction, 18(1):1:1–1:??, April 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lamming:2000:SPA

[LEF^{+}00] Mik Lamming, Marge Eldridge, Mike Flynn, Chris Jones, and David Pendlebury. Satchel: providing access to any document, any time, anywhere. ACM Transactions on Computer-Human Interaction, 7(3):322– 352, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-3/p322-lamming/p322-
lamming.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-3/p322-lamming/

.

Liao:2014:ADC

[LF14] Q. Vera Liao and Wai-Tat Fu. Age differences in credibility judgments of online health information. ACM Transactions on Computer-Human Interaction, 21(1):2:1–2:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Liu:2006:QNM

[LFT06] Yili Liu, Robert Feyen, and Omer Tsimhoni. Queueing Network-Model Human Processor (QN-MHP): a computational architecture for multitask performance in humanmachine systems. ACM Transactions on Computer-Human Interaction, 13(1):37– 70, March 2006. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lemon:2004:MCR

[LG04] Oliver Lemon and Alexander Gruenstein. Multithreaded context for robust conversational interfaces: Contextsensitive speech recognition and interpretation of corrective fragments. ACM Transactions on Computer-Human Interaction, 11(3):241–267, September 2004. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Liao:2012:EUU

[LG12] Chunyuan Liao and Françcois Guimbretièere. Evaluating and understanding the usability of a pen-based command system for interactive paper. ACM Transactions on Computer-Human Interaction, 19(1):3:1–3:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lafreniere:2017:IPT

[LGC17] Benjamin Lafreniere, Carl Gutwin, and Andy Cockburn. Investigating the post-training persistence of expert interaction techniques. ACM Transactions on Computer-Human Interaction, 24(4):29:1–29:??, September 2017. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Liao:2008:PGB

[LGHH08] Chunyuan Liao, François Guimbretière, Ken Hinckley, and Jim Hollan. Papier-Craft: a gesture-based command system for interactive paper. ACM Transactions on Computer-Human Interaction, 14(4):18:1–18:??, January 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lederman:2019:SCY

[LGW^{+}19] Reeva Lederman, John Gleeson, Greg Wadley, Simon D’Alfonso, Simon Rice, Olga Santesteban-Echarri, and Mario Alvarez-Jimenez. Support for carers of young people with mental illness: Design and trial of a technology-mediated therapy. ACM Transactions on Computer-Human Interaction, 26(1):4:1–4:??, February 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301421.

Lunzer:2008:SIE

[LH08] Aran Lunzer and Kasper Hornbæk. Subjunctive interfaces: Extending applications to support parallel setup, viewing and control of alternative scenarios. ACM Transactions on Computer-Human Interaction, 14(4):17:1–17:??, January 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Li:2022:ANN

[LHF^{+}22] Tianshi Li, Julia Katherine Haines, Miguel Flores Ruiz De Eguino, Jason I. Hong, and Jeffrey Nichols. Alert now or never: Understanding and predicting notification preferences of smartphone users. ACM Transactions on Computer-Human Interaction, 29(5):39:1–39:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3478868.

Lazar:2021:ADR

[LJPS21] Amanda Lazar, Ben Jelen, Alisha Pradhan, and Katie A. Siek. Adopting diffractive reading to advance HCI research: a case study on technology for aging. ACM Transactions on Computer-Human Interaction, 28(5): 32:1–32:29, October 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3462326.

Luff:2013:EIA

[LJY^{+}13] Paul Luff, Marina Jirotka, Naomi Yamashita, Hideaki Kuzuoka, Christian Heath, and Grace Eden. Embedded interaction: The accomplishment of actions in everyday and video-mediated environments. ACM Transactions on Computer-Human Interaction, 20(1):6:1–6:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lee:2020:CPD

[LK20] Hosub Lee and Alfred Kobsa. Confident privacy decisionmaking in IoT environments. ACM Transactions on Computer-Human Interaction, 27(1):6:1– 6:39, January 2020. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3364223.

Lundgaard:2019:TCH

[LKS19] Stine S. Lundgaard, Jesper Kjeldskov, and Mikael B. Skov. Temporal constraints in human–building interaction. ACM Transactions on Computer-Human Interaction, 26(2):8:1–8:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301424.

Li:2014:OPT

[LLZ14] Yang Li, Hao Lu, and Haimo Zhang. Optimistic programming of touch interaction. ACM Transactions on Computer-Human Interaction, 21(4):24:1–24:??, August 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Liu:2022:IUT

[LMDT22] Jie Liu, Kim Marriott, Tim Dwyer, and Guido Tack. Increasing user trust in optimisation through feedback and interaction. ACM Transactions on Computer-Human Interaction, 29(5):42:1–42:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503461.

Li:2011:DEC

[LMG^{+}11] Wei Li, Justin Matejka, Tovi Grossman, Joseph A. Konstan, and George Fitzmaurice. Design and evaluation of a command recommendation system for software applications. ACM Transactions on Computer-Human Interaction, 18(2):6:1–6:??, June 2011. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Liang:2021:EFT

[LMK21] Calvin A. Liang, Sean A.

Munson, and Julie A. Kientz. Embracing four tensions in human-computer interaction research with marginalized people. ACM Transactions on Computer-Human Interaction, 28(2):14:1–14:47, April 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3443686.

Leiva:2019:ERD

[LMMBL19] Germán Leiva, Nolwenn Maudet, Wendy Mackay, and Michel Beaudouin-Lafon. Enact: Reducing designer-developer breakdowns when prototyping custom interactions. ACM Transactions on Computer-Human Interaction, 26(3): 19:1–19:??, June 2019. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3310276.

Loizeau:2024:GBM

[LMN24] Alice Loizeau, Sylvain Malacria, and Mathieu Nancel. GUI behaviors to minimize pointingbased interaction interferences. ACM Transactions on Computer-Human Interaction, 31(3):34:1–34:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660338.

Le:2020:SGM

[LMW^{+}20] Huy Viet Le, Sven Mayer, Maximilian Weiß, Jonas Vogelsang, Henrike Weingärtner, and Niels Henze. Shortcut gestures for mobile text editing on fully touch sensitive smartphones. ACM Transactions on Computer-Human Interaction, 27(5):33:1–33:38, October 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3396233.

Lachand-Pascal:2022:COM

[LPMST22] Valentin Lachand-Pascal, Christine Michel, Audrey Serna, and Aurélien Tabard. Challenges and opportunities for multidevice management in classrooms. ACM Transactions on Computer-Human Interaction, 29(6):54:1–54:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3519025.

Li:2023:CPC

[LQZ23] Tianshi Li, Philip Quinn, and Shumin Zhai. C-PAK: Correcting and completing variable-length prefix-based abbreviated keystrokes. ACM Transactions on Computer-Human Interaction, 30(1):7:1– 7:??, February 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3544101.

Loke:2013:MMS

[LR13] Lian Loke and Toni Robertson.

Moving and making strange: an embodied approach to movement-based interaction design. ACM Transactions on Computer-Human Interaction, 20(1):7:1–7:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ludwig:2015:SHD

[LRP15] Thomas Ludwig, Christian Reuter, and Volkmar Pipek. Social Haystack: Dynamic quality assessment of citizengenerated content during emergencies. ACM Transactions on Computer-Human Interaction, 22(4):17:1–17:??, July 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lim:2019:DDC

[LRS19] Bohyeon Lim, Yvonne Rogers, and Neil Sebire. Designing to distract: Can interactive technologies reduce visitor anxiety in a children’s hospital setting? ACM Transactions on Computer-Human Interaction, 26(2):9:1–9:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301427.

Lim:2008:APP

[LST08] Youn-Kyung Lim, Erik Stolterman, and Josh Tenenberg. The anatomy of prototypes: Prototypes as filters, prototypes as manifestations of design ideas. ACM Transactions on Computer-Human Interaction, 15(2):7:1–7:??, July 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lysecky:2009:ENC

[LV09] Susan Lysecky and Frank Vahid. Enabling nonexpert construction of basic sensorbased systems. ACM Transactions on Computer-Human Interaction, 16(1):1:1–1:??, April 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lewis:2020:LDR

[LV20] Blaine Lewis and Daniel Vogel. Longer delays in rehearsalbased interfaces increase expert use. ACM Transactions on Computer-Human Interaction, 27(6):45:1–45:41, November 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418196.

Li:2023:AHA

[LVDA23] Tianyi Li, Mihaela Vorvoreanu, Derek Debellis, and Saleema Amershi. Assessing Human-AI interaction early through factorial surveys: a study on the guidelines for Human-AI interaction. ACM Transactions on Computer-Human Interaction, 30(5): 69:1–69:??, October 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3511605.

Liu:2014:MWF

[LVG^{+}14] Yong Liu, Jayant Venkatanathan, Jorge Goncalves, Evangelos Karapanos, and Vassilis Kostakos. Modeling what friendship patterns on Facebook reveal about personality and social capital. ACM Transactions on Computer-Human Interaction, 21(3): 17:1–17:??, June 2014. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Leong:2012:ECD

[LVH12] Tuck W. Leong, Frank Vetere, and Steve Howard. Experiencing coincidence during digital music listening. ACM Transactions on Computer-Human Interaction, 19(1):6:1– 6:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lindley:2015:PAT

[LW15] Siân Lindley and Jayne Wallace. Placing in age: Transitioning to a new home in later life. ACM Transactions on Computer-Human Interaction, 22(4):20:1–20:??, July 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Lindley:2023:BKT

[LW23] Siân E. Lindley and Denise J. Wilkins. Building knowledge through action: Considerations for machine learning in the workplace. ACM Transactions on Computer-Human Interaction, 30(5): 72:1–72:??, October 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3584947.

Li:2023:DFI

[LWA^{+}23] Zhuying Li, Yan Wang, Josh Andres, Nathan Semertzidis, Stefan Greuter, and Florian Mueller. A design framework for ingestible play. ACM Transactions on Computer-Human Interaction, 30(4): 58:1–58:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3589954.

Lederman:2014:MOS

[LWG^{+}14] Reeva Lederman, Greg Wadley, John Gleeson, Sarah Bendall, and Mario Álvarez-Jiménez. Moderated online social therapy: Designing and evaluating technology for mental health. ACM Transactions on Computer-Human Interaction, 21(1):5:1–5:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Liu:2022:PPM

[LWLL22] Zilong Liu, Xuequn Wang, Xiaohan Li, and Jun Liu. Protecting privacy on mobile apps: a principal-agent perspective. ACM Transactions on Computer-Human Interaction, 29(1):7:1–7:32, February 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3475797.

Leganchuk:1998:MCB

[LZB98] Andrea Leganchuk, Shumin Zhai, and William Buxton. Manual and cognitive benefits of two-handed input: an experimental study. ACM Transactions on Computer-Human Interaction, 5(4):326– 359, December 1998. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1998-5-4/p326-leganchuk/

.

McDonald:2023:DWH

[MA23] Nora McDonald and Nazanin Andalibi. “I Did Watch ‘The Handmaid’s Tale”’: Threat modeling privacy post-Roe in the United States. ACM Transactions on Computer-Human Interaction, 30(4): 63:1–63:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3589960.

MacKay:1999:PSR

[Mac99] Wendy E. MacKay. Is paper safer? the role of paper flight strips in air traffic control. ACM Transactions on Computer-Human Interaction, 6(4):311–340, December 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-4/p311-mackay/.

Marshall:2013:ISI

[MAVR13] Paul Marshall, Alissa Antle, Elise Van Den Hoven, and Yvonne Rogers. Introduction to the special issue on the theory and practice of embodied interaction in HCI and interaction design. ACM Transactions on Computer-Human Interaction, 20(1):1:1–1:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

McGuffin:2005:FLE

[MB05] Michael J. McGuffin and Ravin Balakrishnan. Fitts’ law and expanding targets: Experimental studies and designs for user interfaces. ACM Transactions on Computer-Human Interaction, 12(4):388–422, December 2005. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

McGrenere:2007:FEA

[MBB07] Joanna McGrenere, Ronald M.

Baecker, and Kellogg S. Booth. A field evaluation of an adaptable two-interface design for feature-rich software. ACM Transactions on Computer-Human Interaction, 14(1):3:1– 3:??, May 2007. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mehta:2017:MAI

[MBHC17] Hrim Mehta, Adam Bradley, Mark Hancock, and Christopher Collins. Metatation: Annotation as implicit interaction to bridge close and distant reading. ACM Transactions on Computer-Human Interaction, 24(5):35:1–35:??, November 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Morris:2023:JWT

[MBN^{+}23] Margaret E. Morris, Jennifer Brown, Paula S. Nurius, Savanna Yee, Jennifer C. Mankoff, and Sunny Consolvo. “I Just Wanted to Triple Check… They were all Vaccinated”: Supporting risk negotiation in the context of COVID-19. ACM Transactions on Computer-Human Interaction, 30(4): 60:1–60:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3569938.

McGill:2022:CAK

[MBP^{+}22] Mark McGill, Stephen Brewster, Daniel Pires De Sa Medeiros, Sidney Bovet, Mario Gutierrez, and Aidan Kehoe. Creating and augmenting keyboards for extended reality with the keyboard augmentation toolkit. ACM Transactions on Computer-Human Interaction, 29(2):15:1–15:39, April 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490495.

Meng:2004:IAT

[MCC^{+}04] Helen Meng, P. C. Ching, Shuk Fong Chan, Yee Fong Wong, and Cheong Chat Chan. ISIS: an adaptive, trilingual conversational system with interleaving interaction and delegation dialogs. ACM Transactions on Computer-Human Interaction, 11(3):268– 299, September 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Maggioni:2020:SSM

[MCD^{+}20] Emanuela Maggioni, Robert Cobden, Dmitrijs Dmitrenko, Kasper Hornbæk, and Marianna Obrist. SMELL SPACE: Mapping out the olfactory design space for novel interactions. ACM Transactions on Computer-Human Interaction, 27(5):36:1–36:26, October 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3402449.

Modugno:1997:GRP

[MCM97] Francesmary Modugno, Albert T. Corbett, and Brad A. Myers. Graphical representation of programs in a demonstrational visual shell — an empirical evaluation. ACM Transactions on Computer-Human Interaction, 4(3):276– 308, September 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1997-4-3/p276-modugno/

.

McCrickard:2003:MNS

[MCSN03] D. Scott McCrickard, C. M. Chewar, Jacob P. Somervell, and Ali Ndiwalana. A model for notification systems evaluation—assessing user goals for multitasking activity. ACM Transactions on Computer-Human Interaction, 10(4):312–338, December 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Roffarello:2023:ADW

[MD23] Alberto Monge Roffarello and Luigi De Russis. Achieving digital wellbeing through digital self-control tools: a systematic review and metaanalysis. ACM Transactions on Computer-Human Interaction, 30(4):53:1–53:??, August 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571810.

Mensonge:2022:HIH

[Men22]  Kien Mensonge.  Histori-

cally informed HCI: Reflecting on contemporary technology through anachronistic fiction. ACM Transactions on Computer-Human Interaction, 29(6):60:1–60:??, December 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3517144.

Mackenzie:2010:SSA

[MF10] I. Scott Mackenzie and Torsten Felzer. SAK: Scanning ambiguous keyboard for efficient one-key text entry. ACM Transactions on Computer-Human Interaction, 17(3): 11:1–11:??, July 2010. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

MacDonald:2024:PEE

[MFPB24] Shaun MacDonald, Euan Freeman, Frank Pollick, and Stephen Brewster. Prototyping and evaluation of emotionally resonant vibrotactile comfort objects as a calming social anxiety intervention. ACM Transactions on Computer-Human Interaction, 31(4):46:1–46:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3648615.

Miri:2020:PPP

[MFU^{+}20] Pardis Miri, Robert Flory, Andero Uusberg, Heather Culbertson, Richard H. Harvey, Agata Kelman, Davis Erik Peper, James J. Gross, Katherine Isbister, and Keith Marzullo. PIV: Placement, pattern, and personalization of an inconspicuous vibrotactile breathing pacer. ACM Transactions on Computer-Human Interaction, 27(1):5:1–5:44, January 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3365107.

Martin:2021:ICM

[MGMMS21] J. Alberto Álvarez Martín, Henrik Gollee, Jörg Müller, and Roderick Murray-Smith. Intermittent control as a model of mouse movements. ACM Transactions on Computer-Human Interaction, 28(5): 35:1–35:46, October 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3461836.

Mueller:2017:DBI

[MGVE17] Florian ‘Floyd’ Mueller, Martin R. Gibbs, Frank Vetere, and Darren Edge. Designing for bodily interplay in social exertion games. ACM Transactions on Computer-Human Interaction, 24(3):24:1–24:??, July 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mack:2023:RCO

[MHL^{+}23] Kelly Mack, Megan Hofmann, Udaya Lakshmi, Jerry Cao, Nayha Auradkar, Rosa Arriaga, Scott Hudson, and Jen Mankoff. Rapid convergence: The outcomes of making PPE during a healthcare crisis. ACM Transactions on Computer-Human Interaction, 30(1):4:1–4:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542923.

Myers:2000:PPF

[MHP00] Brad Myers, Scott E. Hudson, and Randy Pausch. Past, present, and future of user interface software tools. ACM Transactions on Computer-Human Interaction, 7(1):3–28, March 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
2000-7-1/p3-myers/.

Moloney:2006:LDC

[MJV^{+}06] Kevin P. Moloney, Julie A.

Jacko, Brani Vidakovic, François Sainfort, V. Kathlene Leonard, and Bin Shi. Leveraging data complexity: Pupillary behavior of older adults with visual impairment during HCI. ACM Transactions on Computer-Human Interaction, 13(3):376– 402, September 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mcgill:2020:EBS

[MKFB20] Mark Mcgill, Aidan Kehoe, Euan Freeman, and Stephen Brewster. Expanding the bounds of seated virtual workspaces. ACM Transactions on Computer-Human Interaction, 27(3):13:1–13:40, June 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3380959.

Maloney-Krichmar:2005:MAS

[MKP05] Diane Maloney-Krichmar and Jenny Preece. A multilevel analysis of sociability, usability, and community dynamics in an online health community. ACM Transactions on Computer-Human Interaction, 12(2):201–232, June 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Menges:2019:IUE

[MKS19] Raphael Menges, Chandan Kumar, and Steffen Staab. Improving user experience of eye tracking-based interaction: Introspecting and adapting interfaces. ACM Transactions on Computer-Human Interaction, 26(6):37:1–37:??, December 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3338844.

Memarovic:2013:PLL

[MLC^{+}13] Nemanja Memarovic, Marc Langheinrich, Keith Cheverst, Nick Taylor, and Florian Alt. P-LAYERS — a layered framework addressing the multifaceted issues facing community-supporting public display deployments. ACM Transactions on Computer-Human Interaction, 20(3): 17:1–17:??, July 2013. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mentis:2014:MSY

[MLH14] Helena M. Mentis, Jarmo Laaksolahti, and Kristina Höök. My self and you: Tension in bodily sharing of experience. ACM Transactions on Computer-Human Interaction, 21(4):20:1–20:??, August 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mah:2021:TCR

[MLH21] Kristina Mah, Lian Loke, and Luke Hespanhol. Towards a contemplative research framework for training self-observation in HCI: a study of compassion cultivation. ACM Transactions on Computer-Human Interaction, 28(6):39:1–39:27, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3471932.

Metaxas:2017:NCR

[MM17] Georgios Metaxas and Panos Markopoulos. Natural contextual reasoning for end users. ACM Transactions on Computer-Human Interaction, 24(2):13:1–13:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mcdonald:2021:CTS

[MM21] Nora Mcdonald and Helena M. Mentis. “Citizens Too”: Safety setting collaboration among older adults with memory concerns. ACM Transactions on Computer-Human Interaction, 28(5):31:1–31:32, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3465217.

Martinez-Maldonado:2024:LLM

[MMEFN^{+}24] Roberto Martinez-Maldonado, Vanessa Echeverria, Gloria Fernandez-Nieto, Lixiang Yan, Linxuan Zhao, Riordan Alfredo, Xinyu Li, Samantha Dix, Hollie Jaggard, Rosie Wotherspoon, Abra Osborne, Simon Buckingham Shum, and Dragan Gasević. Lessons learnt from a multimodal learning analytics deployment in-the-wild. ACM Transactions on Computer-Human Interaction, 31(1):8:1–8:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3622784.

Messerschmidt:2022:APT

[MMH^{+}22] Moritz Alexander Messerschmidt, Sachith Muthukumarana, Nur Al-Huda Hamdan, Adrian Wagner, Haimo Zhang, Jan Borchers, and Suranga Chandima Nanayakkara. ANISMA: a prototyping toolkit to explore haptic skin deformation applications using shapememory alloys. ACM Transactions on Computer-Human Interaction, 29(3):19:1–19:34, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490497.

Muresan:2023:UFR

[MMH23] Andreea Muresan, Jess Mcintosh, and Kasper Hornbæk. Using feedforward to reveal interaction possibilities in virtual reality. ACM Transactions on Computer-Human Interaction, 30(6):82:1–82:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603623.

Mottelson:2023:SRM

[MMHM23] Aske Mottelson, Andreea Muresan, Kasper Hornbæk, and Guido Makransky. A systematic review and metaanalysis of the effectiveness of body ownership illusions in virtual reality. ACM Transactions on Computer-Human Interaction, 30(5):76:1–76:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3590767.

McDonald:2008:PDS

[MMS^{+}08] David W. McDonald, Joseph F. McCarthy, Suzanne Soroczak, David H. Nguyen, and Al M. Rashid. Proactive displays: Supporting awareness in fluid social environments. ACM Transactions on Computer-Human Interaction, 14(4): 16:1–16:??, January 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Markopoulos:2017:EEU

[MNPP17] Panos Markopoulos, Jeffrey Nichols, Fabio Paternò, and Volkmar Pipek. Editorial: End-user development for the Internet of Things. ACM Transactions on Computer-Human Interaction, 24(2):9:1– 9:??, May 2017. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mandviwalla:1994:WDG

[MO94] Munir Mandviwalla and Lorne Olfman. What do groups need? A proposed set of generic groupware requirements. ACM Transactions on Computer-Human Interaction, 1(3):245– 268, September 1994. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1994-1-3/p245-mandviwalla/

.

Muller:2017:CTM

[MOMS17] Jörg Müller, Antti Oulasvirta, and Roderick Murray-Smith. Control theoretic models of pointing. ACM Transactions on Computer-Human Interaction, 24(4):27:1–27:??, September 2017. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Medhi:2011:DMI

[MPB^{+}11] Indrani Medhi, Somani Patnaik, Emma Brunskill, S. N. Nagasena Gautama, William Thies, and Kentaro Toyama. Designing mobile interfaces for novice and low-literacy users. ACM Transactions on Computer-Human Interaction, 18(1):2:1–2:??, April 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Munoz:2022:EET

[MPS22] Diego Muñoz, Sonja Pedell, and Leon Sterling. Evaluating engagement in technologysupported social interaction by people living with dementia in residential care. ACM Transactions on Computer-Human Interaction, 29(5): 50:1–50:??, October 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514497.

Marky:2022:NJA

[MRC^{+}22] Karola Marky, Kirill Ragozin, George Chernyshov, Andrii Matviienko, Martin Schmitz, Max Mühlhäuser, Chloe Eghtebas, and Kai Kunze. “Nah, it’s just annoying!” a deep dive into user perceptions of twofactor authentication. ACM Transactions on Computer-Human Interaction, 29(5): 43:1–43:??, October 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3503514.

Murray-Rust:2022:BBU

[MREN^{+}22] Dave Murray-Rust, Chris Elsden, Bettina Nissen, Ella Tallyn, Larissa Pschetz, and Chris Speed. Blockchain and beyond: Understanding blockchains through prototypes and public engagement. ACM Transactions on Computer-Human Interaction, 29(5):41:1–41:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503462.

McLaughlin:2009:UDI

[MRF09] Anne Collins McLaughlin, Wendy A. Rogers, and Arthur D. Fisk. Using direct and indirect input devices: Attention demands and age-related differences. ACM Transactions on Computer-Human Interaction, 16(1):2:1–2:??, April 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mukherjea:1994:TVD

[MS94] Sougata Mukherjea and John T.

Stasko. Toward visual debugging: integrating algorithm animation capabilities within a source-level debugger. ACM Transactions on Computer-Human Interaction, 1(3):215– 244, September 1994. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-3/p215-mukherjea/

.

Martin:2004:PCI

[MS04] David Martin and Ian Sommerville. Patterns of cooperative interaction: Linking ethnomethodology and design. ACM Transactions on Computer-Human Interaction, 11(1):59–89, March 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mueller:2023:TUD

[MSA^{+}23] Florian ‘Floyd’ Mueller, Nathan Semertzidis, Josh Andres, Joe Marshall, Steve Benford, Xiang Li, Louise Matjeka, and Yash Mehta. Toward understanding the design of intertwined human-computer integrations. ACM Transactions on Computer-Human Interaction, 30(5):73:1–73:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3590766.

Mikusz:2021:LSP

[MSD^{+}21] Mateusz Mikusz, Peter Shaw, Nigel Davies, Petteri Nurmi, Sarah Clinch, Ludwig Trotter, Ivan Elhart, Marc Langheinrich, and Adrian Friday. A longitudinal study of pervasive display personalisation. ACM Transactions on Computer-Human Interaction, 28(1):2:1– 2:45, February 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3418352.

Mastrianni:2023:TCA

[MSH^{+}23] Angela Mastrianni, Aleksandra Sarcevic, Allison Hu, Lynn Almengor, Peyton Tempel, Sarah Gao, and Randall S. Burd. Transitioning cognitive aids into decision support platforms: Requirements and design guidelines. ACM Transactions on Computer-Human Interaction, 30(3):41:1–41:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3582431.

Ma:2023:MAE

[MSM23] Shuai Ma, Mingfei Sun, and Xiaojuan Ma. Modeling adaptive expression of robot learning engagement and exploring its effects on human teachers. ACM Transactions on Computer-Human Interaction, 30(5):70:1–70:??, October 2023. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571813.

Mayer:2014:UIS

[MTDM14] Simon Mayer, Andreas Tschofen, Anind K. Dey, and Friedemann Mattern. User interfaces for smart things — a generative approach with semantic interaction descriptions. ACM Transactions on Computer-Human Interaction, 21(2):12:1–12:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

McGill:2016:ERS

[MWB16] Mark McGill, John H. Williamson, and Stephen Brewster. Examining the role of smart TVs and VR HMDs in synchronous at-a-distance media consumption. ACM Transactions on Computer-Human Interaction, 23(5):33:1–33:??, November 2016. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Maior:2018:WAU

[MWS18] Horia A. Maior, Max L.

Wilson, and Sarah Sharples. Workload alerts-using physiological measures of mental workload to provide feedback during tasks. ACM Transactions on Computer-Human Interaction, 25(2):9:1–9:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mathis:2021:FSA

[MWVK21] Florian Mathis, John H. Williamson, Kami Vaniea, and Mohamed Khamis. Fast and secure authentication in virtual reality using coordinated 3D manipulation and pointing. ACM Transactions on Computer-Human Interaction, 28(1):6:1–6:44, February 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428121.

Mehra:2006:NHD

[MWW06] Sumit Mehra, Peter Werkhoven, and Marcel Worring. Navigating on handheld displays: Dynamic versus static peephole navigation. ACM Transactions on Computer-Human Interaction, 13(4):448–457, December 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Myers:1995:UIS

[Mye95] Brad A. Myers. User interface software tools. ACM Transactions on Computer-Human Interaction, 2(1):64– 103, March 1995. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1995-2-1/p64-myers/

.

Maity:2020:BHE

[MYR^{+}20] Shovan Maity, David Yang, Scott Stanton Redford, Debayan Das, Baibhab Chatterjee, and Shreyas Sen. BodyWire-HCI: Enabling new interaction modalities by communicating strictly during touch using electro-quasistatic human body communication. ACM Transactions on Computer-Human Interaction, 27(6):39:1–39:25, November 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3406238.

Mayer:2023:AIA

[MZL^{+}23] Peter Mayer, Yixin Zou, Byron M. Lowens, Hunter A. Dyer, Khue Le, Florian Schaub, and Adam J. Aviv. Awareness, intention, (in)Action: Individuals’ reactions to data breaches. ACM Transactions on Computer-Human Interaction, 30(5):77:1–77:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589958.

Marky:2021:IUU

[MZR^{+}21] Karola Marky, Marie-Laure Zollinger, Peter Roenne, Peter Y. A. Ryan, Tim Grube, and Kai Kunze. Investigating usability and user experience of individually verifiable Internet voting schemes. ACM Transactions on Computer-Human Interaction, 28(5): 30:1–30:36, October 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3459604.

Ng:2020:MEK

[NBB20] Sarah Ng, Shaowen Bardzell, and Jeffrey Bardzell. The menstruating entrepreneur kickstarting a new politics of women’s health. ACM Transactions on Computer-Human Interaction, 27(4):21:1–21:25, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3397158.

Neustaedter:2009:CCC

[NBG09] Carman Neustaedter, A. J. Bernheim Brush, and Saul Greenberg. The calendar is crucial: Coordination and awareness through the family calendar. ACM Transactions on Computer-Human Interaction, 16(1):6:1–6:??, April 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Nonnis:2024:UPP

[NBK24] A. Nonnis and N. Bryan-Kinns. Unmasking the power of play through TUI designs. ACM Transactions on Computer-Human Interaction, 31(4):52:1–52:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3648619.

Newell:2006:DPO

[NDSG06] Alan F. Newell, Anna Dickinson, Mick J. Smith, and Peter Gregor. Designing a portal for older users: a case study of an industrial/academic collaboration. ACM Transactions on Computer-Human Interaction, 13(3):347–375, September 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Nadal:2024:PAS

[NEE^{+}24] Camille Nadal, Caroline Earley, Angel Enrique, Corina Sas, Derek Richards, and Gavin Doherty. Patient acceptance of self-monitoring on a smartwatch in a routine digital therapy: a mixed-methods study. ACM Transactions on Computer-Human Interaction, 31(1):3:1–3:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617361.

Neustaedter:2006:BFF

[NGB06] Carman Neustaedter, Saul Greenberg, and Michael Boyle. Blur filtration fails to preserve privacy for home-based video conferencing. ACM Transactions on Computer-Human Interaction, 13(1):1–36, March 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Nacenta:2016:ECP

[NHGC16] Miguel A. Nacenta, Mark Hancock, Carl Gutwin, and Sheelagh Carpendale. The effects of changing projection geometry on perception of 3D objects on and around tabletops. ACM Transactions on Computer-Human Interaction, 23(2):11:1–11:??, May 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Navarro:2023:CCU

[NKDB23] David Fraile Navarro, A. Baki Kocaballi, Mark Dras, and Shlomo Berkovsky. Collaboration, not confrontation: Understanding general practitioners’ attitudes towards natural language and text automation in clinical practice. ACM Transactions on Computer-Human Interaction, 30(2): 29:1–29:??, April 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3569893.

Nichols:2009:CLU

[NM09] Jeffrey Nichols and Brad A. Myers. Creating a lightweight user interface description language: an overview and analysis of the personal universal controller project. ACM Transactions on Computer-Human Interaction, 16(4): 17:1–17:??, November 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Nancel:2015:MAP

[NPCBL15] Mathieu Nancel, Emmanuel Pietriga, Olivier Chapuis, and Michel Beaudouin-Lafon. Mid-air pointing on ultrawalls. ACM Transactions on Computer-Human Interaction, 22(5):21:1–21:??, October 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Neustaedter:2015:SDL

[NPF^{+}15] Carman Neustaedter, Carolyn Pang, Azadeh Forghani, Erick Oduor, Serena Hillman, Tejinder K. Judge, Michael Massimi, and Saul Greenberg. Sharing domestic life through long-term video connections. ACM Transactions on Computer-Human Interaction, 22(1):3:1–3:??, March 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Navarre:2009:IMB

[NPLB09] David Navarre, Philippe Palanque, Jean-Francois Ladry, and Eric Barboni. ICOs: a modelbased user interface description technique dedicated to interactive systems addressing usability, reliability and scalability. ACM Transactions on Computer-Human Interaction, 16(4):18:1–18:??, November 2009. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

Niksirat:2019:ARF

[NSCR19] Kavous Salehzadeh Niksirat, Chaklam Silpasuwanchai, Peng Cheng, and Xiangshi Ren. Attention regulation framework: Designing self-regulated mindfulness technologies. ACM Transactions on Computer-Human Interaction, 26(6): 39:1–39:??, December 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3359593.

Neustaedter:2018:BTW

[NSP^{+}18] Carman Neustaedter, Samarth Singhal, Rui Pan, Yasamin Heshmat, Azadeh Forghani, and John Tang. From being there to watching: Shared and dedicated telepresence robot usage at academic conferences. ACM Transactions on Computer-Human Interaction, 25(6):33:1–33:??, December 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Nunes:2015:SCT

[NVF^{+}15] Francisco Nunes, Nervo Verdezoto, Geraldine Fitzpatrick, Morten Kyng, Erik Grönvall, and Cristiano Storni. Selfcare technologies in HCI: Trends, tensions, and opportunities. ACM Transactions on Computer-Human Interaction, 22(6):33:1–33:??, December 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Niemantsverdriet:2019:DAI

[NVPE19] Karin Niemantsverdriet, Harm Van Essen, Minna Pakanen, and Berry Eggen. Designing for awareness in interactions with shared systems: The DASS framework. ACM Transactions on Computer-Human Interaction, 26(6): 36:1–36:??, December 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3338845.

Nansen:2014:RHS

[NVR^{+}14] Bjorn Nansen, Frank Vetere, Toni Robertson, John Downs, Margot Brereton, and Jeannette Durick. Reciprocal habituation: a study of older people and the Kinect. ACM Transactions on Computer-Human Interaction, 21(3):18:1–18:??, June 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Norrie:2021:ECA

[NWH21] Christopher S. Norrie, Annalu Waller, and Elizabeth F. S. Hannah. Establishing context: AAC device adoption and support in a special-education setting. ACM Transactions on Computer-Human Interaction, 28(2):13:1–13:30, April 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3446205.

Ogonowski:2016:IBF

[OAV^{+}16] Corinna Ogonowski, Konstantin Aal, Daryoush Vaziri, Thomas Von Rekowski, Dave Randall, Dirk Schreiber, Rainer Wieching, and Volker Wulf. ICT-based fall prevention system for older adults: Qualitative results from a long-term field study. ACM Transactions on Computer-Human Interaction, 23(5):29:1–29:??, November 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Olston:2003:SIB

[OC03] Christopher Olston and Ed H. Chi. ScentTrails: Integrating browsing and searching on the Web. ACM Transactions on Computer-Human Interaction, 10(3):177–197, September 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Oviatt:2012:IIA

[OCM^{+}12] Sharon Oviatt, Adrienne Cohen, Andrea Miller, Kumi Hodge, and Ariana Mann. The impact of interface affordances on human ideation, problem solving, and inferential reasoning. ACM Transactions on Computer-Human Interaction, 19(3):22:1–22:??, October 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Oviatt:2004:TAC

[ODC04] Sharon Oviatt, Courtney Darves, and Rachel Coulston. Toward adaptive conversational interfaces: Modeling speech convergence with animated personas. ACM Transactions on Computer-Human Interaction, 11(3):300– 328, September 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Oulasvirta:2017:CSF

[OFLK17] Antti Oulasvirta, Anna Feit, Perttu Lähteenlahti, and Andreas Karrenbauer. Computational support for functionality selection in interaction design. ACM Transactions on Computer-Human Interaction, 24(5):34:1–34:??, November 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ohara:2013:NTP

[OHM^{+}13] Kenton O’Hara, Richard Harper, Helena Mentis, Abigail Sellen, and Alex Taylor. On the naturalness of touchless: Putting the “interaction” back into NUI. ACM Transactions on Computer-Human Interaction, 20(1):5:1– 5:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ohara:2011:BIS

[OKP11] Kenton O’Hara, Jesper Kjeldskov, and Jeni Paay. Blended interaction spaces for distributed team collaboration. ACM Transactions on Computer-Human Interaction, 18(1):3:1– 3:??, April 2011. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Obrenovic:2011:SIS

[OM11] Željko Obrenovic and Jean-Bernard Martens. Sketching interactive systems with sketchify. ACM Transactions on Computer-Human Interaction, 18(1):4:1–4:??, April 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Orji:2017:IEG

[OMV17] Rita Orji, Regan L. Mandryk, and Julita Vassileva. Improving the efficacy of games for change using personalization models. ACM Transactions on Computer-Human Interaction, 24(5):32:1–32:??, November 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Olsen:2010:TWS

[OPL10] Dan R. Olsen, Brett Partridge, and Stephen Lynn. Time warp sports for Internet television. ACM Transactions on Computer-Human Interaction, 17(4):16:1–16:??, December 2010. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

OBrien:1999:HTE

[ORRH99] Jon O’Brien, Tom Rodden, Mark Rouncefield, and John Hughes. At home with the technology: an ethnographic study of a set-top-box trial. ACM Transactions on Computer-Human Interaction, 6(3):282–308, September 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-3/p282-o_brien/.

Oviatt:2004:IMA

[OS04] Sharon Oviatt and Stephanie Seneff. Introduction to mobile and adaptive conversational interfaces. ACM Transactions on Computer-Human Interaction, 11(3):237–240, September 2004. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Olugbade:2019:HCA

[OSBB^{+}19] Temitayo A. Olugbade, Aneesha Singh, Nadia Bianchi-Berthouze, Nicolai Marquardt, Min S. H. Aung, and Amanda C. De C. Williams. How can affect be detected and represented in technological support for physical rehabilitation? ACM Transactions on Computer-Human Interaction, 26(1):1:1–1:??, February 2019. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3299095.

Olsen:1995:ISI

[OSF95] Dan R. Olsen, Jr., Germinder Singh, and Steven K. Feiner. Introduction to the special issue on virtual reality software and technology. ACM Transactions on Computer-Human Interaction, 2(3):177– 178, September 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-3/p177-olsen/

.

Oleson:2023:TID

[OSPK23] Alannah Oleson, Meron Solomon, Christopher Perdriau, and Amy Ko. Teaching inclusive design skills with the CIDER assumption elicitation technique. ACM Transactions on Computer-Human Interaction, 30(1):6:1–6:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549074.

Orth:2019:DMP

[OTV19] Daniel Orth, Clementine Thurgood, and Elise Van Den Hoven. Designing meaningful products in the digital age: How users value their technological possessions. ACM Transactions on Computer-Human Interaction, 26(5): 34:1–34:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3341980.

Olson:2017:HPW

[OWOZ17] Judith S. Olson, Dakuo Wang, Gary M. Olson, and Jingwen Zhang. How people write together now: Beginning the investigation with advanced undergraduates in a project course. ACM Transactions on Computer-Human Interaction, 24(1):4:1–4:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Persa:2022:EYS

[PAM^{+}22] Nicholas Persa, Craig G. Anderson, Richard Martinez, Max Collins, Maria J. Anderson-Coto, and Kurt D. Squire. Enhancing youth self-regulation through wearable apps: Increasing usage through participatory design in low income youth. ACM Transactions on Computer-Human Interaction, 29(5):40:1–40:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490169.

Price:2022:MMT

[PBBJ^{+}22] Sara Price, Nadia Bianchi-Berthouze, Carey Jewitt, Nikoleta Yiannoutsou, Katerina Fotopoulou, Svetlana Dajic, Juspreet Virdee, Yixin Zhao, Douglas Atkinson, and Frederik Brudy. The making of meaning through dyadic haptic affective touch. ACM Transactions on Computer-Human Interaction, 29(3):21:1–21:42, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490494.

Price:2022:ISI

[PBBJS22] Sara Price, Nadia Bianchi-Berthouze, Carey Jewitt, and Jürgen Steimle. Introduction to the special issue on digital touch: Reshaping interpersonal communicative capacity and touch practices. ACM Transactions on Computer-Human Interaction, 29(3): 18:1–18:8, June 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3505591.

Plaisant:2006:SFC

[PCH^{+}06] Catherine Plaisant, Aaron Clamage, Hilary Browne Hutchinson, Benjamin B. Bederson, and Allison Druin. Shared family calendars: Promoting symmetry and accessibility. ACM Transactions on Computer-Human Interaction, 13(3):313–346, September 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Papangelis:2020:PDS

[PCL^{+}20] Konstantinos Papangelis, Alan Chamberlain, Ioanna Lykourentzou, Vassilis-Javed Khan, Michael Saker, Hai-Ning Liang, Irwyn Sadien, and Ting Cao. Performing the digital self: Understanding location-based social networking, territory, space, and identity in the city. ACM Transactions on Computer-Human Interaction, 27(1):1:1–1:26, January 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3364997.

Park:2015:TOA

[PCR15] Sun Young Park, Yunan Chen, and Scott Rudkin. Technological and organizational adaptation of EMR implementation in an emergency department. ACM Transactions on Computer-Human Interaction, 22(1):1:1–1:??, March 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pirolli:2003:EIS

[PCV03] Peter Pirolli, Stuart K. Card, and Mija M. Van Der Wege. The effects of information scent on visual search in the hyperbolic tree browser. ACM Transactions on Computer-Human Interaction, 10(1):20– 53, 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Perrotin:2016:TAV

[PD16] Olivier Perrotin and Christophe D’Alessandro. Target acquisition vs. expressive motion: Dynamic pitch warping for intonation correction. ACM Transactions on Computer-Human Interaction, 23(3):17:1–17:??, July 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pohl:2017:BJT

[PDR17] Henning Pohl, Christian Domin, and Michael Rohs. Beyond just text: Semantic emoji similarity modeling to support expressive communication [emoji string]. ACM Transactions on Computer-Human Interaction, 24(1):6:1–6:??, March 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Perry:2018:MPU

[PF18] Mark Perry and Jennifer Ferreira. Moneywork: Practices of use and social interaction around digital and analog money. ACM Transactions on Computer-Human Interaction, 24(6):41:1–41:??, January 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Posch:2021:MTD

[PF21] Irene Posch and Geraldine Fitzpatrick. The matter of tools: Designing, using and reflecting on new tools for emerging eTextile craft practices. ACM Transactions on Computer-Human Interaction, 28(1):4:1–4:38, February 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3426776.

Peer:2024:PMT

[PFGE24] Eyal Peer, Alisa Frik, Conor Gilsenan, and Serge Egelman. “Protect me tomorrow”: Commitment nudges to remedy compromised passwords. ACM Transactions on Computer-Human Interaction, 31(5):59:1–59:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3689038.

Poltrock:1994:OOI

[PG94] Steven E. Poltrock and Jonathan Grudin. Organizational obstacles to interface design and development: two participantobserver studies. ACM Transactions on Computer-Human Interaction, 1(1):52– 80, March 1994. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-1/p52-poltrock/

.

Pinelle:2003:TAG

[PGG03] David Pinelle, Carl Gutwin, and Saul Greenberg. Task analysis for groupware usability evaluation: Modeling shared-workspace tasks with the mechanics of collaboration. ACM Transactions on Computer-Human Interaction, 10(4):281–311, December 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pelikan:2023:MDH

[PH23] Hannah Pelikan and Emily Hofstetter. Managing delays in human-robot interaction. ACM Transactions on Computer-Human Interaction, 30(4):50:1–50:??, August 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3569890.

Petersen:2008:ISIa

[PHJ08a] Marianne Graves Petersen, Lars Hallnäs, and Robert J. K. Jacob. Introduction to special issue on the aesthetics of interaction. ACM Transactions on Computer-Human Interaction, 15(3):10:1–10:??, November 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Petersen:2008:ISIb

[PHJ08b] Marianne Graves Petersen, Lars Hallnäs, and Robert J. K. Jacob. Introduction to special issue on the aesthetics of interaction. ACM Transactions on Computer-Human Interaction, 15(4):14:1–14:??, November 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pilemalm:2018:PDE

[Pil18] Sofie Pilemalm. Participatory design in emerging civic engagement initiatives in the new public sector: Applying PD concepts in resourcescarce organizations. ACM Transactions on Computer-Human Interaction, 25(1):5:1– 5:??, February 2018. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Poor:2016:ANU

[PJL^{+}16] G. Michael Poor, Samuel D.

Jaffee, Laura Marie Leventhal, Jordan Ringenberg, Dale S. Klopfer, Guy Zimmerman, and Brandi A. Klein. Applying the Norman 1986 usercentered model to post-WIMP UIs: Theoretical predictions and empirical outcomes. ACM Transactions on Computer-Human Interaction, 23(5): 30:1–30:??, November 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Prakash:1994:FUA

[PK94] Atul Prakash and Michael J.

Knister. A framework for undoing actions in collaborative systems. ACM Transactions on Computer-Human Interaction, 1(4):295–330, December 1994. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-4/p295-prakash/

.

Papenmeier:2022:CRB

[PKES22] Andrea Papenmeier, Dagmar Kern, Gwenn Englebienne, and Christin Seifert. It’s complicated: The relationship between user trust, model accuracy and explanations in AI. ACM Transactions on Computer-Human Interaction, 29(4):35:1–35:33, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495013.

Paay:2009:TSP

[PKHD09] Jeni Paay, Jesper Kjeldskov, Steve Howard, and Bharat Dave. Out on the town: a socio-physical approach to the design of a context-aware urban guide. ACM Transactions on Computer-Human Interaction, 16(2):7:1–7:??, June 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Paik:2015:PUP

[PKRR15] Jaehyon Paik, Jong W. Kim, Frank E. Ritter, and David Reitter. Predicting user performance and learning in human– computer interaction with the Herbal compiler. ACM Transactions on Computer-Human Interaction, 22(5):25:1–25:??, October 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Petrelli:2014:FRP

[PL14] Daniela Petrelli and Ann Light. Family rituals and the potential for interaction design: a study of Christmas. ACM Transactions on Computer-Human Interaction, 21(3):16:1–16:??, June 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pradhan:2020:UIV

[PLF20] Alisha Pradhan, Amanda Lazar, and Leah Findlater. Use of intelligent voice assistants by older adults with low technology use. ACM Transactions on Computer-Human Interaction, 27(4):31:1–31:27, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3373759.

Palacin:2024:CDP

[PMAN24] Victoria Palacin, Samantha McDonald, Pablo Aragón, and Matti Nelimarkka. Configurations of digital participatory budgeting. ACM Transactions on Computer-Human Interaction, 31(2):28:1–28:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3635144.

Petersen:2002:UET

[PMK02] Marianne Graves Petersen, Kim Halskov Madsen, and Arne Kjær. The usability of everyday technology: emerging and fading opportunities. ACM Transactions on Computer-Human Interaction, 9(2):74–105, June 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pink:2013:ALS

[PMM^{+}13] Sarah Pink, Kerstin Leder Mackley, Val Mitchell, Marcus Hanratty, Carolina Escobar-Tello, Tracy Bhamra, and Roxana Morosanu. Applying the lens of sensory ethnography to sustainable HCI. ACM Transactions on Computer-Human Interaction, 20(4):25:1–25:??, September 2013. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pouta:2024:OML

[PMS24] Emmi Pouta, Jussi Ville Mikkonen, and Antti Salovaara. Opportunities with multi-layer weave structures in woven e-textile design. ACM Transactions on Computer-Human Interaction, 31(5): 62:1–62:??, October 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3689039.

Perry:2001:DMU

[POS^{+}01] Mark Perry, Kenton O’Hara, Abigail Sellen, Barry Brown, and Richard Harper. Dealing with mobility: understanding access anytime, anywhere. ACM Transactions on Computer-Human Interaction, 8(4):323–347, December 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Porayska-Pomsta:2018:BHA

[PPAA^{+}18] Kaśka Porayska-Pomsta, Alyssa M.

Alcorn, Katerina Avramides, Sandra Beale, Sara Bernardini, Mary Ellen Foster, Christopher Frauenberger, Judith Good, Karen Guldberg, Wendy Keay-Bright, Lila Kossyvaki, Oliver Lemon, Marilena Mademtzi, Rachel Menzies, Helen Pain, Gnanathusharan Rajendran, Annalu Waller, Sam Wass, and Tim J. Smith. Blending human and artificial intelligence to support autistic children’s social communication skills. ACM Transactions on Computer-Human Interaction, 25(6):35:1–35:??, December 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Putze:2022:UHP

[PPS^{+}22] Felix Putze, Susanne Putze, Merle Sagehorn, Christopher Micek, and Erin T. Solovey. Understanding HCI practices and challenges of experiment reporting with brain signals: Towards reproducibility and reuse. ACM Transactions on Computer-Human Interaction, 29(4):31:1–31:43, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490554.

Plimmer:2011:STL

[PRB^{+}11] Beryl Plimmer, Peter Reid, Rachel Blagojevic, Andrew Crossan, and Stephen Brewster. Signing on the tactile line: a multimodal system for teaching handwriting to blind children. ACM Transactions on Computer-Human Interaction, 18(3):17:1–17:??, July 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Prather:2024:WIK

[PRD^{+}24] James Prather, Brent N. Reeves, Paul Denny, Brett A. Becker, Juho Leinonen, Andrew Luxton-Reilly, Garrett Powell, James Finnie-Ansley, and Eddie Antonio Santos. “It’s Weird That it Knows What I Want”: Usability and interactions with Copilot for novice programmers. ACM Transactions on Computer-Human Interaction, 31(1):4:1– 4:??, February 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3617367.

Pearson:2016:ELC

[PRJ16] Jennifer Pearson, Simon Robinson, and Matt Jones. Exploring low-cost, Internetfree information access for resource-constrained communities. ACM Transactions on Computer-Human Interaction, 23(6):36:1–36:??, December 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pascoe:2000:UWM

[PRM00] Jason Pascoe, Nick Ryan, and David Morse. Using while moving: HCI issues in fieldwork environments. ACM Transactions on Computer-Human Interaction, 7(3):417– 437, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-3/p417-pascoe/p417-
pascoe.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-3/p417-pascoe/

.

Pearson:2019:DFM

[PRR^{+}19] Jennifer Pearson, Simon Robinson, Thomas Reitmaier, Matt Jones, and Anirudha Joshi. Diversifying future-making through itinerative design. ACM Transactions on Computer-Human Interaction, 26(5): 33:1–33:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3341727.

Palen:2002:BHD

[PS02] Leysia Palen and Marilyn Salzman. Beyond the handset: designing for wireless communications usability. ACM Transactions on Computer-Human Interaction, 9(2):125– 151, June 2002. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Perelman:2024:EPR

[PSD24] Gary Perelman, Marcos Serrano, and Emmanuel Dubois. Exploiting physical referent features as input for multidimensional data selection in augmented reality. ACM Transactions on Computer-Human Interaction, 31(4): 50:1–50:??, August 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3648613.

Paterno:2009:MUD

[PSS09] Fabio Paterno’, Carmen Santoro, and Lucio Davide Spano. MARIA: a universal, declarative, multiple abstraction-level language for service-oriented applications in ubiquitous environments. ACM Transactions on Computer-Human Interaction, 16(4):19:1–19:??, November 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Pierce:2013:ISI

[PSSB13] James Pierce, Yolande Strengers, Phoebe Sengers, and Susanne Bødker. Introduction to the special issue on practiceoriented approaches to sustainable HCI. ACM Transactions on Computer-Human Interaction, 20(4):20:1–20:??, September 2013. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Prechelt:2001:IMI

[PT01] Lutz Prechelt and Rainer Typke. An interface for melody input. ACM Transactions on Computer-Human Interaction, 8(2):133–149, 2001. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org/
pubs/articles/journals/tochi/
2001-8-2/p133-prechelt/p133-
prechelt.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2001-8-2/p133-prechelt/

.

Procter:2023:HAA

[PTR23] Rob Procter, Peter Tolmie, and Mark Rouncefield. Holding AI to account: Challenges for the delivery of trustworthy AI in healthcare. ACM Transactions on Computer-Human Interaction, 30(2):31:1–31:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3577009.

Pinder:2018:DBC

[PVCB18] Charlie Pinder, Jo Vermeulen, Benjamin R. Cowan, and Russell Beale. Digital behaviour change interventions to break and form habits. ACM Transactions on Computer-Human Interaction, 25(3):15:1–15:??, June 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Park:2024:RNY

[PVMK24] Soya Park, Stuti Vishwabhan, Michael Muller, and David R. Karger. “I Really Need Your Help with This Work… ”: a system for navigating the tricky terrain of managing up by leveraging one’s motivation to get things done. ACM Transactions on Computer-Human Interaction, 31(4):51:1–51:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3652603.

Paakki:2024:DCD

[PVSZ24] Henna Paakki, Heidi Vepsäläinen, Antti Salovaara, and Bushra Zafar. Detecting covert disruptive behavior in online interaction by analyzing conversational features and norm violations. ACM Transactions on Computer-Human Interaction, 31(2):20:1–20:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635143.

Plumlee:2006:ZVM

[PW06] Matthew D. Plumlee and Colin Ware. Zooming versus multiple window interfaces: Cognitive costs of visual comparisons. ACM Transactions on Computer-Human Interaction, 13(2):179–209, June 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Procter:2018:HWC

[PWG18] Rob Procter, Joe Wherton, and Trisha Greenhalgh. Hidden work and the challenges of scalability and sustainability in ambulatory assisted living. ACM Transactions on Computer-Human Interaction, 25(2):11:1–11:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Qiu:2005:IET

[QB05] Lingyun Qiu and Izak Benbasat. An investigation into the effects of Text-To-Speech voice and 3D avatars on the perception of presence and flow of live help in electronic commerce. ACM Transactions on Computer-Human Interaction, 12(4):329–355, December 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Quek:2002:MHD

[QMB^{+}02] Francis Quek, David McNeill, Robert Bryll, Susan Duncan, Xin-Feng Ma, Cemil Kirbas, Karl E. McCullough, and Rashid Ansari. Multimodal human discourse: gesture and speech. ACM Transactions on Computer-Human Interaction, 9(3):171–193, September 2002. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Quek:2013:EBS

[QO13] Francis Quek and Francisco Oliveira. Enabling the blind to see gestures. ACM Transactions on Computer-Human Interaction, 20(1):4:1–4:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ruller:2022:SDC

[RAT^{+}22] Sarah Rüller, Konstantin Aal, Peter Tolmie, Andrea Hartmann, Markus Rohde, and Volker Wulf. Speculative design as a collaborative practice: Ameliorating the consequences of illiteracy through digital touch. ACM Transactions on Computer-Human Interaction, 29(3):23:1–23:58, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487917.

Reinecke:2011:IPP

[RB11] Katharina Reinecke and Abraham Bernstein. Improving performance, perceived usability, and aesthetics with culturally adaptive user interfaces. ACM Transactions on Computer-Human Interaction, 18(2):8:1–8:??, June 2011. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Rapp:2023:ELE

[RB23] Amon Rapp and Arianna Boldi. Exploring the lived experience of behavior change technologies: Towards an existential model of behavior change for HCI. ACM Transactions on Computer-Human Interaction, 30(6):81:1–81:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3603497.

Roedl:2015:SMB

[RBB15] David Roedl, Shaowen Bardzell, and Jeffrey Bardzell. Sustainable making? Balancing optimism and criticism in HCI discourse. ACM Transactions on Computer-Human Interaction, 22(3):15:1–15:??, June 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Robinson:2021:RIL

[RBC^{+}21] Sarah Robinson, Nicola J. Bidwell, Roberto Cibin, Conor Linehan, Laura Maye, John Mccarthy, Nadia Pantidi, and Maurizio Teli. Rural islandness as a lens for (rural) HCI. ACM Transactions on Computer-Human Interaction, 28(3):20:1–20:32, July 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3443704.

Ramirez:2024:AAR

[RBH24] Dafne Zuleima Morgado Ramirez, Giulia Barbareschi, and Cathy Holloway. Adult autism research priorities and conceptualization in computing research: Invitation to co-lead with autistic adults. ACM Transactions on Computer-Human Interaction, 31(4): 45:1–45:??, August 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635148.

Ritter:2000:SCM

[RBJY00] Frank E. Ritter, Gordon D. Baxter, Gary Jones, and Richard M. Young. Supporting cognitive models as users. ACM Transactions on Computer-Human Interaction, 7(2):141–173, June 2000. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-2/p141-ritter/p141-
ritter.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-2/p141-ritter/

.

Rozendaal:2019:OID

[RBK19] Marco C. Rozendaal, Boudewijn Boon, and Victor Kaptelinin. Objects with intent: Designing everyday things as collaborative partners. ACM Transactions on Computer-Human Interaction, 26(4):26:1–26:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3325277.

Rosson:1996:RUS

[RC96] Mary Beth Rosson and John M.

Carroll. The reuse of uses in Smalltalk programming. ACM Transactions on Computer-Human Interaction, 3(3):219– 253, September 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-3/p219-rosson/

.

Rocheleau:2022:PSS

[RC22] Jessica N. Rocheleau and Sonia Chiasson. Privacy and safety on social networking sites: Autistic and non-autistic teenagers’ attitudes and behaviors. ACM Transactions on Computer-Human Interaction, 29(1):1:1–1:39, February 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469859.

Reyes-Cruz:2022:DIC

[RCFR22] Gisela Reyes-Cruz, Joel E. Fischer, and Stuart Reeves. Demonstrating interaction: The case of assistive technology. ACM Transactions on Computer-Human Interaction, 29(5):48:1–48:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3514236.

Riedl:2005:ISS

[RD05] John Riedl and Paul Dourish. Introduction to the special section on recommender systems. ACM Transactions on Computer-Human Interaction, 12(3):371–373, September 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Redstrom:2008:TIE

[Red08] Johan Redström. Tangled interaction: On the expressiveness of tangible user interfaces. ACM Transactions on Computer-Human Interaction, 15(4):16:1–16:??, November 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Reeves:2019:HUP

[Ree19]  Stuart Reeves.  How UX

practitioners produce findings in usability testing. ACM Transactions on Computer-Human Interaction, 26(1):3:1– 3:??, February 2019. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3299096.

Roseman:1996:BRG

[RG96] Mark Roseman and Saul Greenberg. Building realtime groupware with Group-Kit, a groupware toolkit. ACM Transactions on Computer-Human Interaction, 3(1):66– 106, March 1996. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-1/p66-roseman/

.

Rogers:2024:URR

[RHK^{+}24] Katja Rogers, Teresa Hirzle, Sukran Karaosmanoglu, Paula Toledo Palomino, Ekaterina Durmanova, Seiji Isotani, and Lennart E. Nacke. An umbrella review of reporting quality in CHI systematic reviews: Guiding questions and best practices for HCI. ACM Transactions on Computer-Human Interaction, 31(5):57:1–57:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3685266.

Rieman:1996:FSE

[Rie96] John Rieman. A field study of exploratory learning strategies. ACM Transactions on Computer-Human Interaction, 3(3):189–218, September 1996. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-3/p189-rieman/

.

Robe:2022:DPC

[RK22] Peter Robe and Sandeep Kaur Kuttal. Designing PairBuddy — a conversational agent for pair programming. ACM Transactions on Computer-Human Interaction, 29(4): 34:1–34:44, August 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3498326.

Ruddle:2009:BUW

[RL09] Roy A. Ruddle and Simon Lessels. The benefits of using a walking interface to navigate virtual environments. ACM Transactions on Computer-Human Interaction, 16(1):5:1– 5:??, April 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Reuter:2014:AHP

[RLP14] Christian Reuter, Thomas Ludwig, and Volkmar Pipek. Ad hoc participation in situation assessment: Supporting mobile collaboration in emergencies. ACM Transactions on Computer-Human Interaction, 21(5):26:1–26:??, November 2014. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic).

Ren:2000:ISP

[RM00] Xiangshi Ren and Shinju Moriya. Improving selection performance on pen-based systems: a study of penbased interaction for selection tasks. ACM Transactions on Computer-Human Interaction, 7(3):384–416, 2000. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org/
pubs/articles/journals/tochi/
2000-7-3/p384-ren/p384-ren.
pdf;   http://www.acm.org/
pubs/citations/journals/
tochi/2000-7-3/p384-ren/.

Rezwana:2023:DCA

[RM23] Jeba Rezwana and Mary Lou Maher. Designing creative AI partners with COFI: a framework for modeling interaction in Human-AI co-creative systems. ACM Transactions on Computer-Human Interaction, 30(5):67:1–67:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3519026.

Radiah:2021:RVS

[RMP^{+}21] Rivu Radiah, Ville Mäkelä, Sarah Prange, Sarah Delgado Rodriguez, Robin Piening, Yumeng Zhou, Kay Köhle, Ken Pfeuffer, Yomna Abdelrahman, Matthias Hoppe, Albrecht Schmidt, and Florian Alt. Remote VR studies: a framework for running virtual reality studies remotely via participant-owned HMDs. ACM Transactions on Computer-Human Interaction, 28(6):46:1–46:36, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472617.

Reeves:2015:CUB

[RMT^{+}15] Stuart Reeves, Sarah Martindale, Paul Tennent, Steve Benford, Joe Marshall, and Brendan Walker. The challenges of using biodata in promotional filmmaking. ACM Transactions on Computer-Human Interaction, 22(3):11:1–11:??, June 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Rodham:1997:NAS

[RO97] Kenneth J. Rodham and Dan R. Olsen, Jr. Nanites: an approach to structurebased monitoring. ACM Transactions on Computer-Human Interaction, 4(2):103– 136, June 1997. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-2/p103-rodham/

.

Robertson:2005:VCD

[Rob05] Scott P. Robertson. Votercentered design: Toward a voter decision support system. ACM Transactions on Computer-Human Interaction, 12(2):263–292, June 2005. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Rooksby:2013:WLD

[Roo13]  John Rooksby.  Wild in

the laboratory: a discussion of plans and situated actions. ACM Transactions on Computer-Human Interaction, 20(3):19:1–19:??, July 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Recker:1996:PDA

[RP96] Margaret M. Recker and James E. Pitkow. Predicting document access in large multimedia repositories. ACM Transactions on Computer-Human Interaction, 3(4):352– 375, December 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-4/p352-recker/

.

Rivera-Pelayo:2017:IMS

[RPFMP17] Verónica Rivera-Pelayo, Angela Fessl, Lars Müller, and Viktoria Pammer. Introducing mood self-tracking at work: Empirical insights from call centers. ACM Transactions on Computer-Human Interaction, 24(1):3:1–3:??, March 2017.

CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Reicherts:2022:GTC

[RRC^{+}22] Leon Reicherts, Yvonne Rogers, Licia Capra, Ethan Wood, Tu Dinh Duong, and Neil Sebire. It’s good to talk: a comparison of using voice versus screen-based interactions for agent-assisted tasks. ACM Transactions on Computer-Human Interaction, 29(3): 25:1–25:41, June 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3484221.

Rodgers:2015:VSL

[RSC15] Peter Rodgers, Gem Stapleton, and Peter Chapman. Visualizing sets with linear diagrams. ACM Transactions on Computer-Human Interaction, 22(6):27:1–27:??, December 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Renaud:2024:WGR

[RSC^{+}24] Karen Renaud, Cigdem Sengul, Kovila Coopamootoo, Bryan Clift, Jacqui Taylor, Mark Springett, and Ben Morrison. “We’re Not That Gullible!” revealing dark pattern mental models of 11–12-year-old Scottish children. ACM Transactions on Computer-Human Interaction, 31(3):33:1–33:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660342.

Ruddle:2002:SAA

[RSJ02] Roy A. Ruddle, Justin C. D.

Savage, and Dylan M. Jones. Symmetric and asymmetric action integration during cooperative object manipulation in virtual environments. ACM Transactions on Computer-Human Interaction, 9(4):285– 308, December 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ravasio:2004:PDE

[RSK04] Pamela Ravasio, Sissel Guttormsen Schär, and Helmut Krueger. In pursuit of desktop evolution: User problems and practices with modern desktop systems. ACM Transactions on Computer-Human Interaction, 11(2):156–180, June 2004. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Rothwell:2021:CCG

[RSR21] Clayton D. Rothwell, Valerie L. Shalin, and Griffin D. Romigh. Comparison of common ground models for human–computer dialogue: Evidence for audience design. ACM Transactions on Computer-Human Interaction, 28(2):9:1–9:35, April 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410876.

Rapp:2018:PIS

[RT18] Amon Rapp and Lia Tirabeni. Personal informatics for sport: Meaning, body, and social relations in amateur and elite athletes. ACM Transactions on Computer-Human Interaction, 25(3):16:1–16:??, June 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ruddle:2016:DEI

[RTR^{+}16] Roy A. Ruddle, Rhys G. Thomas, Rebecca Randell, Philip Quirke, and Darren Treanor. The design and evaluation of interfaces for navigating gigapixel images in digital pathology. ACM Transactions on Computer-Human Interaction, 23(1):5:1–5:??, February 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Rapp:2019:RTB

[RTT19] Amon Rapp, Maurizio Tirassa, and Lia Tirabeni. Rethinking technologies for behavior change: a view from the inside of human change. ACM Transactions on Computer-Human Interaction, 26(4):22:1–22:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3318142.

Rullo:2008:SQI

[Rul08] Alessia Rullo. The soft qualities of interaction. ACM Transactions on Computer-Human Interaction, 15(4): 17:1–17:??, November 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Resnick:1995:RAI

[RV95] Paul Resnick and Robert A.

Virzi. Relief from the audio interface blues: expanding the spectrum of menu, list, and form styles. ACM Transactions on Computer-Human Interaction, 2(2):145– 176, June 1995. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-2/p145-resnick/

.

Ruddle:2011:WIY

[RVB11] Roy A. Ruddle, Ekaterina Volkova, and Heinrich H. Bülthoff. Walking improves your cognitive map in environments that are large-scale and large in extent. ACM Transactions on Computer-Human Interaction, 18(2):10:1–10:??, June 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Ren:2024:HDT

[RZK24] Yuqing Ren, Haifeng Zhang, and Robert E. Kraut. How did they build the free encyclopedia? A literature review of collaboration and coordination among Wikipedia Editors. ACM Transactions on Computer-Human Interaction, 31(1):7:1–7:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617369.

Salvucci:2009:RPE

[Sal09] Dario D. Salvucci. Rapid prototyping and evaluation of in-vehicle interfaces. ACM Transactions on Computer-Human Interaction, 16(2):9:1– 9:??, June 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Solovey:2015:DII

[SAP^{+}15] Erin Treacy Solovey, Daniel Afergan, Evan M. Peck, Samuel W. Hincks, and Robert J. K. Jacob. Designing implicit interfaces for physiological computing: Guidelines and lessons learned using fNIRS. ACM Transactions on Computer-Human Interaction, 21(6):35:1–35:??, January 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Slovak:2023:DER

[SAT^{+}23] Petr Slovak, Alissa Antle, Nikki Theofanopoulou, Claudia Daudén Roquet, James Gross, and Katherine Isbister. Designing for emotion regulation interventions: an agenda for HCI theory and research. ACM Transactions on Computer-Human Interaction, 30(1):13:1–13:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569898.

Salehi:2018:IIW

[SB18] Niloufar Salehi and Michael S.

Bernstein. Ink: Increasing worker agency to reduce friction in hiring crowd workers. ACM Transactions on Computer-Human Interaction, 25(2):10:1–10:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Singh:2023:WHS

[SBSG23] Nikhil Singh, Guillermo Bernal, Daria Savchenko, and Elena L. Glassman. Where to hide a stolen elephant: Leaps in creative writing with multimodal machine intelligence. ACM Transactions on Computer-Human Interaction, 30(5): 68:1–68:??, October 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3511599.

Schleyer:2012:CAR

[SBSS12] Titus Schleyer, Brian S. Butler, Mei Song, and Heiko Spallek. Conceptualizing and advancing research networking systems. ACM Transactions on Computer-Human Interaction, 19(1):2:1–2:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sun:2002:CMR

[SC02] Chengzheng Sun and David Chen. Consistency maintenance in real-time collaborative graphics editing systems. ACM Transactions on Computer-Human Interaction, 9(1):1–41, March 2002. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Steriadis:2003:DHC

[SC03] Constantine E. Steriadis and Philip Constantinou. Designing human-computer interfaces for quadriplegic people. ACM Transactions on Computer-Human Interaction, 10(2):87–118, 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sobel:2024:ICT

[SDBK24] Kiley Sobel, Maitraye Das, Sara Behbakht, and Julie A. Kientz. Incloodle-Classroom: Technology for inclusive joint media engagement in a neurodiverse kindergarten classroom. ACM Transactions on Computer-Human Interaction, 31(3):41:1–41:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674506.

Sabie:2022:DIM

[SEA22] Dina Sabie, Cansu Ekmekcioglu, and Syed Ishtiaque Ahmed. A decade of international migration research in HCI: Overview, challenges, ethics, impact, and future directions. ACM Transactions on Computer-Human Interaction, 29(4):30:1–30:35, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490555.

Seaborn:2016:APE

[SED^{+}16] Katie Seaborn, Jamal Edey, Gregory Dolinar, Margot Whitfield, Paula Gardner, Carmen Branje, and Deborah I. Fels. Accessible play in everyday spaces: Mixed reality gaming for adult powered chair users. ACM Transactions on Computer-Human Interaction, 23(2):12:1–12:??, May 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Slovak:2015:TDS

[SF15] Petr Slovák and Geraldine Fitzpatrick. Teaching and developing social and emotional skills with technology. ACM Transactions on Computer-Human Interaction, 22(4): 19:1–19:??, July 2015. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Spiel:2019:AAC

[SFKF19] Katta Spiel, Christopher Frauenberger, Os Keyes, and Geraldine Fitzpatrick. Agency of autistic children in technology research — a critical literature review. ACM Transactions on Computer-Human Interaction, 26(6): 38:1–38:??, December 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3344919.

Sidenmark:2020:EHT

[SG20] Ludwig Sidenmark and Hans Gellersen. Eye, head and torso coordination during gaze shifts in virtual reality. ACM Transactions on Computer-Human Interaction, 27(1):4:1– 4:40, January 2020. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3361218.

Spiel:2021:PPH

[SG21] Katta Spiel and Kathrin Gerling. The purpose of play: How HCI games research fails neurodivergent populations. ACM Transactions on Computer-Human Interaction, 28(2): 11:1–11:40, April 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3432245.

Sutcliffe:2006:IIC

[SGFT06] Alistair Sutcliffe, Brian Gault, Terence Fernando, and Kevin Tan. Investigating interaction in CAVE virtual environments. ACM Transactions on Computer-Human Interaction, 13(2):235–267, June 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Song:2009:MFC

[SGL09] Hyunyoung Song, François Guimbretière, and Hod Lipson. The ModelCraft framework: Capturing freehand annotations and edits to facilitate the 3D model design process using a digital pen. ACM Transactions on Computer-Human Interaction, 16(3):14:1–14:??, September 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Svensson:2005:DEK

[SHC05] Martin Svensson, Kristina Höök, and Rickard Cöster. Designing and evaluating Kalas: a social navigation system for food recipes. ACM Transactions on Computer-Human Interaction, 12(3):374–400, September 2005. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sharit:2008:IRK

[SHCP08] Joseph Sharit, Mario A.

Hernández, Sara J. Czaja, and Peter Pirolli. Investigating the roles of knowledge and cognitive abilities in older adult information seeking on the Web. ACM Transactions on Computer-Human Interaction, 15(1):3:1–3:??, May 2008. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Schuessler:2024:GHI

[SHD^{+}24] Martin Schuessler, Luca Hormann, Raimund Dachselt, Andrew Blake, and Carsten Rother. Gazing heads: Investigating gaze perception in video-mediated communication. ACM Transactions on Computer-Human Interaction, 31(3):39:1–39:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660343.

Singh:2024:PSE

[SHG^{+}24] Aneesha Singh, Marusa Hrobat, Suxin Gui, Nadia Bianchi-Berthouze, Judith Ley-Flores, Frederic Bevilacqua, Joaquin R. Diaz Duran, Elena MÁrquez Segura, and Ana Tajadura-JimÉnez. Pushed by sound: Effects of sound and movement direction on body perception, experience quality, and exercise support. ACM Transactions on Computer-Human Interaction, 31(4): 53:1–53:??, August 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3648616.

Sears:2007:ISI

[SHM07] Andrew Sears, Vicki L. Hanson, and Brad Myers. Introduction to special issue on computers and accessibility. ACM Transactions on Computer-Human Interaction, 14(3):11:1–11:??, September 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sevilla:2007:WAI

[SHMA07] Javier Sevilla, Gerardo Herrera, Bibiana Martínez, and Francisco Alcantud. Web accessibility for individuals with cognitive deficits: a comparative study between an existing commercial Web and its cognitively accessible equivalent. ACM Transactions on Computer-Human Interaction, 14(3):12:1–12:??, September 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Shneiderman:2000:CCU

[Shn00] Ben Shneiderman. Creating creativity: user interfaces for supporting innovation. ACM Transactions on Computer-Human Interaction, 7(1):114– 138, March 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
2000-7-1/p114-shneiderman/

.

StAmant:2007:MBE

[SHR07] Robert St. Amant, Thomas E. Horton, and Frank E. Ritter. Model-based evaluation of expert cell phone menu interaction. ACM Transactions on Computer-Human Interaction, 14(1):1:1–1:??, May 2007. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Shen:2024:ESM

[SHS24] Vivian Shen, Chris Harrison, and Craig Shultz. Expressive, scalable, mid-air haptics with synthetic jets. ACM Transactions on Computer-Human Interaction, 31(2):14:1–14:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3635150.

Sanches:2019:ARI

[SHSS19] Pedro Sanches, Kristina Höök, Corina Sas, and Anna Ståhl. Ambiguity as a resource to inform proto-practices: The case of skin conductance. ACM Transactions on Computer-Human Interaction, 26(4):21:1–21:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3318143.

Su:2021:IPR

[SHVH^{+}21] Norman Makoto Su, Jean Hardy, Morgan Vigil-Hayes, Tiffany Veinot, and Rob Comber. Introduction: Performing rurality with computing. ACM Transactions on Computer-Human Interaction, 28(3):16e:1–16e:13, July 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3461832.

Schnadelbach:2012:EPD

[SIK^{+}12] Holger Schnädelbach, Ainojie Irune, David Kirk, Kevin Glover, and Patrick Brundell. ExoBuilding: Physiologically driven adaptive architecture. ACM Transactions on Computer-Human Interaction, 19(4):25:1–25:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Shaer:2009:SPD

[SJ09] Orit Shaer and Robert J. K.

Jacob. A specification paradigm for the design and implementation of tangible user interfaces. ACM Transactions on Computer-Human Interaction, 16(4):20:1–20:??, November 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Shaer:2009:ISI

[SJGL09] Orit Shaer, Robert J. K. Jacob, Mark Green, and Kris Luyten. Introduction to the special issue on UIDL for next-generation user interfaces. ACM Transactions on Computer-Human Interaction, 16(4):16:1–16:??, November 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Schnadelbach:2019:AAP

[SJU19] Holger Schnädelbach, Nils Jäger, and Lachlan Urquhart. Adaptive architecture and personal data. ACM Transactions on Computer-Human Interaction, 26(2):12:1–12:??, April 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301426.

Sun:1998:ACC

[SJZ^{+}98] Chengzheng Sun, Xiaohua Jia, Yanchun Zhang, Yun Yang, and David Chen. Achieving convergence, causality preservation, and intention preservation in real-time cooperative editing systems. ACM Transactions on Computer-Human Interaction, 5(1):63– 108, March 1998. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1998-5-1/p63-sun/.

Spence:2023:MTC

[SKB^{+}23] Jocelyn Spence, Boriana Koleva, Steve Benford, Dimitrios Darzentas, Martin Flintham, Kevin Glover, Hanne Wagner, Rebecca Gibson, and Emily-Clare Thorn. “More than a cliché”: Experiencing hybrid gifting in the wild. ACM Transactions on Computer-Human Interaction, 30(4): 55:1–55:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3577015.

Snow:2024:HWE

[SKDM24] Stephen Snow, Awais Hameed Khan, Kaleb Day, and Ben Matthews. Household Wattch: Exploring opportunities for surveillance and consent through families’ household energy use data. ACM Transactions on Computer-Human Interaction, 31(4):55:1–55:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673228.

Sharma:2024:PGH

[SKN24] Vishal Sharma, Neha Kumar, and Bonnie Nardi. Postgrowth human-computer interaction. ACM Transactions on Computer-Human Interaction, 31(1):9:1–9:??, February 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624981.

Sedig:2001:RIM

[SKW01] Kamran Sedig, Maria Klawe, and Marvin Westrom. Role of interface manipulation style and scaffolding on cognition and concept learning in learnware. ACM Transactions on Computer-Human Interaction, 8(1):34–59, 2001. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2001-8-1/p34-sedig/p34-sedig.
pdf;   http://www.acm.org/
pubs/citations/journals/
tochi/2001-8-1/p34-sedig/

.

Su:2019:DMA

[SLBB19] Norman Makoto Su, Amanda Lazar, Jeffrey Bardzell, and Shaowen Bardzell. Of dolls and men: Anticipating sexual intimacy with robots. ACM Transactions on Computer-Human Interaction, 26(3): 13:1–13:??, June 2019. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3301422.

Smith:2020:CDA

[SLM^{+}20] C. Estelle Smith, Zachary Levonian, Haiwei Ma, Robert Giaquinto, Gemma Lein-Mcdonough, Zixuan Li, Susan O’conner-Von, and Svetlana Yarosh. “I cannot do all of this alone”: Exploring instrumental and prayer support in online health communities. ACM Transactions on Computer-Human Interaction, 27(5):38:1–38:41, October 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3402855.

Sutton:2022:SCA

[SLP22] Jonathan Sutton, Tobias Langlotz, and Alexander Plopski. Seeing colours: Addressing colour vision deficiency with vision augmentations using computational glasses. ACM Transactions on Computer-Human Interaction, 29(3):26:1–26:53, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486899.

Sanchez:1994:HEO

[SLS94] J. Alfredo Sánchez, John J.

Leggett, and John L. Schnase. HyperActive: extending an open hypermedia architecture to support agency. ACM Transactions on Computer-Human Interaction, 1(4):357– 382, December 1994. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-4/p357-sanchez/

.

Sha:2018:ISA

[SLY^{+}18] Long Sha, Patrick Lucey, Yisong Yue, Xinyu Wei, Jennifer Hobbs, Charlie Rohlf, and Sridha Sridharan. Interactive sports analytics: an intelligent interface for utilizing trajectories for interactive sports play retrieval and analytics. ACM Transactions on Computer-Human Interaction, 25(2):13:1–13:??, April 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sease:2011:OHM

[SM11a] Robin Sease and David W. Mc-Donald. The organization of home media. ACM Transactions on Computer-Human Interaction, 18(2):9:1–9:??, June 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Semaan:2011:TMS

[SM11b] Bryan Semaan and Gloria Mark. Technology-mediated social arrangements to resolve breakdowns in infrastructure during ongoing disruption. ACM Transactions on Computer-Human Interaction, 18(4):21:1–21:??, December 2011. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sarcevic:2012:TET

[SMB12] Aleksandra Sarcevic, Ivan Marsic, and Randal S. Burd. Teamwork errors in trauma resuscitation. ACM Transactions on Computer-Human Interaction, 19(2):13:1–13:??, July 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Suhm:2001:MEC

[SMW01] Bernhard Suhm, Brad Myers, and Alex Waibel. Multimodal error correction for speech user interfaces. ACM Transactions on Computer-Human Interaction, 8(1):60– 98, 2001. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2001-8-1/p60-suhm/p60-suhm.
pdf;   http://www.acm.org/
pubs/citations/journals/
tochi/2001-8-1/p60-suhm/.

Sas:2017:EDP

[SN17] Corina Sas and Carman Neustaedter. Exploring DIY practices of complex home technologies. ACM Transactions on Computer-Human Interaction, 24(2):16:1–16:??, May 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Suto:2023:CIC

[SNT^{+}23] Kai Suto, Yuta Noma, Kotaro Tanimichi, Koya Narumi, and Tomohiro Tachi. Crane: an integrated computational design platform for functional, foldable, and fabricable origami products. ACM Transactions on Computer-Human Interaction, 30(4):52:1–52:??, August 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3576856.

Sondergaard:2020:TDD

[Søn20] Marie Louise Juul Søndergaard. Troubling design: a design program for designing with women’s health. ACM Transactions on Computer-Human Interaction, 27(4):24:1–24:36, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3397199.

Schaper:2021:CDT

[SP21] Marie-Monique Schaper and Narcis Pares. Co-design techniques for and with children based on physical theatre practice to promote embodied awareness. ACM Transactions on Computer-Human Interaction, 28(4):22:1–22:42, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3450446.

Saksono:2024:SCF

[SP24] Herman Saksono and Andrea G. Parker. Socio-cognitive framework for personal informatics: a preliminary framework for socially-enabled health technologies. ACM Transactions on Computer-Human Interaction, 31(3): 42:1–42:??, June 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3674504.

Schadenberg:2021:PRA

[SRE^{+}21] Bob R. Schadenberg, Dennis Reidsma, Vanessa Evers, Daniel P. Davison, Jamy J. Li, Dirk K. J. Heylen, Carlos Neves, Paulo Alvito, Jie Shen, Maja Pantić, Björn W. Schuller, Nicholas Cummins, Vlad Olaru, Cristian Sminchisescu, Snezana Babović Dimitrijević, Suncica Petrović, Aurélie Baranger, Alria Williams, Alyssa M. Alcorn, and Elizabeth Pellicano. Predictable robots for autistic childrenvariance in robot behaviour, idiosyncrasies in autistic children’s characteristics, and child-robot engagement. ACM Transactions on Computer-Human Interaction, 28(5): 36:1–36:42, October 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3468849.

Sallnas:2000:SPC

[SRGS00] Eva-Lotta Sallnäs, Kirsten Rassmus-Gröhn, and Calle Sjöström. Supporting presence in collaborative environments by haptic force feedback. ACM Transactions on Computer-Human Interaction, 7(4):461– 476, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-4/p461-sallnas/p461-
sallnas.pdf;   http://www.
acm.org/pubs/citations/journals/

                                        tochi/2000-7-4/p461-sallnas/

.

Sears:1994:SME

[SS94] Andrew Sears and Ben Shneiderman. Split menus: effectively using selection frequency to organize menus. ACM Transactions on Computer-Human Interaction, 1(1):27– 51, March 1994. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1994-1-1/p27-sears/

.

Sawhney:2000:NRS

[SS00] Nitin Sawhney and Chris Schmandt. Nomadic radio: speech and audio interaction for contextual messaging in nomadic environments. ACM Transactions on Computer-Human Interaction, 7(3):353– 383, 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tochi/
                                        2000-7-3/p353-sawhney/p353-
                                        sawhney.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        tochi/2000-7-3/p353-sawhney/

.

Schneider:2016:UME

[SSC^{+}16] Bertrand Schneider, Kshitij Sharma, Sébastien Cuendet, Guillaume Zufferey, Pierre Dillenbourg, and Roy Pea. Using mobile eye-trackers to unpack the perceptual benefits of a tangible user interface for collaborative learning. ACM Transactions on Computer-Human Interaction, 23(6): 39:1–39:??, December 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sauve:2022:PCF

[SSH22] Kim Sauvé, Miriam Sturdee, and Steven Houben. Physecology: a conceptual framework to describe data physicalizations in their real-world context. ACM Transactions on Computer-Human Interaction, 29(3):27:1–27:33, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505590.

Sharma:2023:SCD

[SSHM^{+}23] Adwait Sharma, Christina Salchow-Hömmen, Vimal Suresh Mollyn, Aditya Shekhar Nittala, Michael A. Hedderich, Marion Koelle, Thomas Seel, and Jürgen Steimle. SparseIMU: Computational design of sparse IMU layouts for sensing fine-grained finger microgestures. ACM Transactions on Computer-Human Interaction, 30(3):39:1–39:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569894.

Schwartz:2013:UPM

[SSRW13] Tobias Schwartz, Gunnar Stevens, Leonardo Ramirez, and Volker Wulf. Uncovering practices of making energy consumption accountable: a phenomenological inquiry. ACM Transactions on Computer-Human Interaction, 20(2):12:1–12:??, May 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Salminen:2022:CUP

[SSS^{+}22] Joni Salminen, Sercan Sengün, João M. Santos, Soon-Gyo Jung, and Bernard Jansen. Can unhappy pictures enhance the effect of personas? A user experiment. ACM Transactions on Computer-Human Interaction, 29(2):14:1–14:59, April 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3485872.

Staahl:2021:VRS

[STB21] Anna Ståhl, Vasiliki Tsaknaki, and Madeline Balaam. Validity and rigour in soma designsketching with the soma. ACM Transactions on Computer-Human Interaction, 28(6): 38:1–38:36, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3470132.

Swan:2008:MPC

[STH08] Laurel Swan, Alex S. Taylor, and Richard Harper. Making place for clutter and other ideas of home. ACM Transactions on Computer-Human Interaction, 15(2):9:1–9:??, July 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Shoemaker:2012:TPM

[STKB12] Garth Shoemaker, Takayuki Tsukitani, Yoshifumi Kitamura, and Kellogg S. Booth. Two-part models capture the impact of gain on pointing performance. ACM Transactions on Computer-Human Interaction, 19(4):28:1–28:??, December 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sun:2002:UCI

[Sun02] Chengzheng Sun. Undo as concurrent inverse in group editors. ACM Transactions on Computer-Human Interaction, 9(4):309–361, December 2002. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Slater:1995:TSI

[SUS95] Mel Slater, Martin Usoh, and Anthony Steed. Taking steps: the influence of a walking technique on presence in virtual reality. ACM Transactions on Computer-Human Interaction, 2(3):201– 219, September 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-3/p201-slater/

.

Sutcliffe:2000:EUR

[Sut00] Alistair Sutcliffe. On the effective use and reuse of HCI knowledge. ACM Transactions on Computer-Human Interaction, 7(2):197–221, June 2000. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
2000-7-2/p197-sutcliffe/.

Svanaes:2013:IDL

[Sva13] Dag Svanæs. Interaction design for and with the lived body: Some implications of Merleau-Ponty’s phenomenology. ACM Transactions on Computer-Human Interaction, 20(1):8:1–8:??, March 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Scurto:2021:DDR

[SVCB21] Hugo Scurto, Bavo Van Kerrebroeck, Baptiste Caramiaux, and Frédéric Bevilacqua. Designing deep reinforcement learning for human parameter exploration. ACM Transactions on Computer-Human Interaction, 28(1):1:1– 1:35, February 2021. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3414472.

Soute:2017:DER

[SVDM17] Iris Soute, Tudor Vacaretu, Jan De Wit, and Panos Markopoulos. Design and evaluation of RaPIDO, a platform for rapid prototyping of interactive outdoor games. ACM Transactions on Computer-Human Interaction, 24(4): 28:1–28:??, September 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Stevens:2009:CSA

[SW09] Gunnar Stevens and Volker Wulf. Computer-supported access control. ACM Transactions on Computer-Human Interaction, 16(3):12:1–12:??, September 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sun:2022:MSG

[SWLM22] Zhida Sun, Sitong Wang, Chengzhong Liu, and Xiaojuan Ma. Metaphoraction: Support gesture-based interaction design with metaphorical meanings. ACM Transactions on Computer-Human Interaction, 29(5):45:1–45:??, October 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3511892.

Sonnenwald:2003:ESC

[SWM03] Diane H. Sonnenwald, Mary C. Whitton, and Kelly L. Maglaughlin. Evaluating a scientific collaboratory: Results of a controlled experiment. ACM Transactions on Computer-Human Interaction, 10(2):150– 176, 2003. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Sas:2016:DRL

[SWZ16] Corina Sas, Steve Whittaker, and John Zimmerman. Design for rituals of letting go: an embodiment perspective on disposal practices informed by grief therapy. ACM Transactions on Computer-Human Interaction, 23(4):21:1–21:??, September 2016. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Swearngin:2024:TAA

[SWZ^{+}24] Amanda Swearngin, Jason Wu, Xiaoyi Zhang, Esteban Gomez, Jen Coughenour, Rachel Stukenborg, Bhavya Garg, Greg Hughes, Adriana Hilliard, Jeffrey P. Bigham, and Jeffrey Nichols. Towards automated accessibility report generation for mobile apps. ACM Transactions on Computer-Human Interaction, 31(4):54:1–54:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674967.

Sun:2006:TAS

[SXS^{+}06] Chengzheng Sun, Steven Xia, David Sun, David Chen, Haifeng Shen, and Wentong Cai. Transparent adaptation of single-user applications for multi-user real-time collaboration. ACM Transactions on Computer-Human Interaction, 13(4):531–582, December 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Schmandt:1997:ISI

[SY97] Chris Schmandt and Nichole Yankelovich. Introduction to the special issue on speech as data. ACM Transactions on Computer-Human Interaction, 4(1):1, March 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-1/p1-schmandt/

.

Spiers:2023:BIP

[SYK23] Adam Spiers, Eric Young, and Katherine J. Kuchenbecker. The S-BAN: Insights into the perception of shape-changing haptic interfaces via virtual pedestrian navigation. ACM Transactions on Computer-Human Interaction, 30(1):11:1–11:??, February 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3555046.

Sun:2019:ELO

[SYS19] Ke Sun, Chun Yu, and Yuanchun Shi. Exploring low-occlusion qwerty soft keyboard using spatial landmarks. ACM Transactions on Computer-Human Interaction, 26(4):20:1–20:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3318141.

Schaffer:1996:NHC

[SZG^{+}96] Doug Schaffer, Zhengping Zuo, Saul Greenberg, Lyn Bartram, John Dill, Shelli Dubs, and Mark Roseman. Navigating hierarchically clustered networks through fisheye and full-zoom methods. ACM Transactions on Computer-Human Interaction, 3(2):162– 188, June 1996. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1996-3-2/p162-schaffer/

.

Semertzidis:2023:BCI

[SZM23] Nathan Semertzidis, Fabio Zambetta, and Florian “Floyd” Mueller. Brain-computer integration: a framework for the design of brain-computer interfaces from an integrations perspective. ACM Transactions on Computer-Human Interaction, 30(6):86:1–86:??, December 2023. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603621.

Tang:2007:ALT

[Tan07] John C. Tang. Approaching and leave-taking: Negotiating contact in computer-mediated communication. ACM Transactions on Computer-Human Interaction, 14(1):5:1–5:??, May 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Taylor:2024:MEI

[TB24] Jordan Taylor and Amy Bruckman. Mitigating epistemic injustice: The online construction of a bisexual culture. ACM Transactions on Computer-Human Interaction, 31(4):48:1–48:??, August 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3648614.

Thieme:2020:MLM

[TBD20] Anja Thieme, Danielle Belgrave, and Gavin Doherty. Machine learning in mental health: a systematic review of the HCI literature to support the development of effective and implementable ML systems. ACM Transactions on Computer-Human Interaction, 27(5):34:1–34:53, October 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3398069.

Tomlinson:2013:CIP

[TBN^{+}13] Bill Tomlinson, Eli Blevis, Bonnie Nardi, Donald J. Patterson, M. Six Silberman, and Yue Pan. Collapse informatics and practice: Theory, method, and design. ACM Transactions on Computer-Human Interaction, 20(4):24:1–24:??, September 2013. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Thomas:2001:ACA

[TC01] Bruce H. Thomas and Paul Calder. Applying cartoon animation techniques to graphical user interfaces. ACM Transactions on Computer-Human Interaction, 8(3):198– 222, September 2001. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Torre:2021:EAV

[TCD^{+}21] Ilaria Torre, Emma Carrigan, Katarina Domijan, Rachel Mc-Donnell, and Naomi Harte. The effect of audio-visual smiles on social influence in a cooperative human-agent interaction task. ACM Transactions on Computer-Human Interaction, 28(6):44:1–44:38, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3469232.

Thimbleby:2001:UAM

[TCJ01] Harold Thimbleby, Paul Cairns, and Matt Jones. Usability analysis with Markov models. ACM Transactions on Computer-Human Interaction, 8(2):99–132, 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
tochi/2001-8-2/p99-thimbleby/

.

Teevan:2010:PP

[TDH10] Jaime Teevan, Susan T. Dumais, and Eric Horvitz. Potential for personalization. ACM Transactions on Computer-Human Interaction, 17(1):4:1– 4:??, March 2010. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tuli:2019:GTE

[TDKS19] Anupriya Tuli, Shruti Dalvi, Neha Kumar, and Pushpendra Singh. “It’s a girl thing”: Examining challenges and opportunities around menstrual health education in India. ACM Transactions on Computer-Human Interaction, 26(5):29:1–29:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3325282.

Tashman:2012:WLL

[TE12] Craig Tashman and W. Keith Edwards. WindowScape:

Lessons learned from a taskcentric window manager. ACM Transactions on Computer-Human Interaction, 19(1):8:1– 8:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Terfurth:2024:DRV

[TGG24] Leonie Terfurth, Klaus Gramann, and Lukas Gehrke. Decoding realism of virtual objects: Exploring behavioral and ocular reactions to inaccurate interaction feedback. ACM Transactions on Computer-Human Interaction, 31(3):29:1–29:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660345.

Tan:2006:PLD

[TGSP06] Desney S. Tan, Darren Gergle, Peter Scupelli, and Randy Pausch. Physically large displays improve performance on spatial tasks. ACM Transactions on Computer-Human Interaction, 13(1):71–99, March 2006. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tuch:2015:DHN

[TH15] Alexandre N. Tuch and Kasper Hornbæk. Does Herzberg’s notion of hygienes and motivators apply to user experience? ACM Transactions on Computer-Human Interaction, 22(4):16:1–16:??, July 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Terveen:1999:COV

[THA99] Loren Terveen, Will Hill, and Brian Amento. Constructing, organizing, and visualizing collections of topically related Web resources. ACM Transactions on Computer-Human Interaction, 6(1):67– 94, March 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-1/p67-terveen/.

Thimbleby:2004:UID

[Thi04] Harold Thimbleby. User interface design with matrix algebra. ACM Transactions on Computer-Human Interaction, 11(2):181–236, June 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Thieme:2023:DHC

[THL^{+}23] Anja Thieme, Maryann Hanratty, Maria Lyons, Jorge Palacios, Rita Faia Marques, Cecily Morrison, and Gavin Doherty. Designing human-centered AI for mental health: Developing clinically relevant applications for online CBT treatment. ACM Transactions on Computer-Human Interaction, 30(2):27:1–27:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3564752.

Tanaka-Ishii:2009:KLL

[TIG09] Kumiko Tanaka-Ishii and Julian Godon. Kansuke: a logograph look-up interface based on a few modified stroke prototypes. ACM Transactions on Computer-Human Interaction, 16(2):11:1–11:17, June 2009. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tekinbas:2021:DYC

[TJLS21] Katie Salen Tekinbas, Krithika Jagannath, Ulrik Lyngs, and Petr Slovák. Designing for youth-centered moderation and community governance in Minecraft. ACM Transactions on Computer-Human Interaction, 28(4):24:1–24:41, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3450290.

Toups:2011:TCG

[TKH11] Zachary O. Toups, Andruid Kerne, and William A. Hamilton. The team coordination game: Zero-fidelity simulation abstracted from fire emergency response practice. ACM Transactions on Computer-Human Interaction, 18(4):23:1–23:??, December 2011. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tan:1999:PTT

[TkWSR99] Bernard C. Y. Tan, Kwok kee Wei, Choon-Ling Sia, and Krishnamurthy S. Raman. A partial test of the task-medium fit proposition in a group support system environment. ACM Transactions on Computer-Human Interaction, 6(1):47– 66, March 1999. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
1999-6-1/p47-tan/.

Toups:2019:MMA

[TLA^{+}19] Z. O. Toups, Nicolas Lalone, Sultan A. Alharthi, Hitesh Nidhi Sharma, and Andrew M. Webb. Making maps available for play: Analyzing the design of game cartography interfaces. ACM Transactions on Computer-Human Interaction, 26(5):30:1–30:??, September 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3336144.

Tran:2024:CPS

[TLY^{+}24] Tanh Quang Tran, Tobias Langlotz, Jacob Young, Thomas W. Schubert, and Holger Regenbrecht. Classifying presence scores: Insights and analysis from two decades of the igroup presence questionnaire (IPQ). ACM Transactions on Computer-Human Interaction, 31(5):61:1–61:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3689046.

Thomas:2002:INU

[TM02] Peter Thomas and Robert D. Macredie. Introduction to the new usability. ACM Transactions on Computer-Human Interaction, 9(2):69–73, June 2002. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Terveen:2005:SMF

[TM05] Loren Terveen and David W. McDonald. Social matching: a framework and research agenda. ACM Transactions on Computer-Human Interaction, 12(3):401–434, September 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tyack:2024:SDT

[TM24] April Tyack and Elisa D. Mekler. Self-determination theory and HCI games research: Unfulfilled promises and unquestioned paradigms. ACM Transactions on Computer-Human Interaction, 31(3): 40:1–40:??, June 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3673230.

Taylor:1995:CSA

[TNB^{+}95] Richard N. Taylor, Kari A. Nies, Gregory Alan Bolcer, Craig A. MacFarlane, Kenneth M. Anderson, and Gregory F. Johnson. Chiron-1: a software architecture for user interface development, maintenance, and run-time support. ACM Transactions on Computer-Human Interaction, 2(2):105–144, June 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-2/p105-taylor/

.

Tajima:2022:WTU

[TNLK22] Daisuke Tajima, Jun Nishida, Pedro Lopes, and Shunichi Kasahara. Whose touch is this?: Understanding the agency trade-off between userdriven touch vs. computerdriven touch. ACM Transactions on Computer-Human Interaction, 29(3):24:1–24:27, June 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3489608.

Taranta:2021:MEE

[TPM^{+}21] Eugene M. Taranta II, Corey R.

Pittman, Mehran Maghoumi, Mykola Maslych, Yasmine M. Moolenaar, and Joseph J. Laviola Jr. Machete: Easy, efficient, and precise continuous custom gesture segmentation. ACM Transactions on Computer-Human Interaction, 28(1):5:1–5:46, February 2021. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428068.

TOCHI-staff:2004:R

[Tra04] Transactions on Computer-Human Interaction staff. 2003 reviewers. ACM Transactions on Computer-Human Interaction, 11(1):120, March 2004. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tu:2015:DSB

[TRZ15] Huawei Tu, Xiangshi Ren, and Shumin Zhai. Differences and similarities between finger and pen stroke gestures on stationary and mobile devices. ACM Transactions on Computer-Human Interaction, 22(5):22:1–22:??, October 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tsandilas:2018:FAC

[Tsa18] Theophanis Tsandilas. Fallacies of agreement: a critical review of consensus assessment methods for gesture elicitation. ACM Transactions on Computer-Human Interaction, 25(3):18:1–18:??, June 2018. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Takagi:2007:ANW

[TSFA07] Hironobu Takagi, Shin Saito, Kentarou Fukuda, and Chieko Asakawa. Analysis of navigability of Web applications for improving blind usability. ACM Transactions on Computer-Human Interaction, 14(3):13:1–13:??, September 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Truschin:2014:DIM

[TSGK14] Sergej Truschin, Michael Schermann, Suparna Goswami, and Helmut Krcmar. Designing interfaces for multiple-goal environments: Experimental insights from in-vehicle speech interfaces. ACM Transactions on Computer-Human Interaction, 21(1):7:1–7:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Thebault-Spieker:2017:TGU

[TSTH17] Jacob Thebault-Spieker, Loren Terveen, and Brent Hecht. Toward a geographic understanding of the sharing economy: Systemic biases in UberX and TaskRabbit. ACM Transactions on Computer-Human Interaction, 24(3):21:1–21:??, July 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Tuch:2016:LWG

[TVH16] Alexandre N. Tuch, Paul Van Schaik, and Kasper Hornbæk. Leisure and work, good and bad: The role of activity domain and valence in modeling user experience. ACM Transactions on Computer-Human Interaction, 23(6):35:1–35:??, December 2016. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Uzor:2019:RRE

[UB19] Stephen Uzor and Lynne Baillie. Recov-R: Evaluation of a home-based tailored exergame system to reduce fall risk in seniors. ACM Transactions on Computer-Human Interaction, 26(4):23:1–23:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3325280.

Ullmer:2005:TCS

[UIJ05] Brygg Ullmer, Hiroshi Ishii, and Robert J. K. Jacob. Token + constraint systems for tangible interaction with digital information. ACM Transactions on Computer-Human Interaction, 12(1):81–118, March 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Uzor:2021:EFC

[UK21] Stephen Uzor and Per Ola Kristensson. An exploration of freehand crossing selection in head-mounted augmented reality. ACM Transactions on Computer-Human Interaction, 28(5):33:1–33:27, October 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462546.

Uchidiuno:2023:WYN

[USK^{+}23] Judith Odili Uchidiuno, Jaemarie Solyst, Jonaya Kemper, Erik Harpstead, Ross Higashi, and Jessica Hammer. “What’s Your Name Again?”: How race and gender dynamics impact codesign processes and output. ACM Transactions on Computer-Human Interaction, 30(6):85:1–85:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603624.

Vazquez-Alvarez:2016:DIM

[VAAB^{+}16] Yolanda Vazquez-Alvarez, Matthew P.

Aylett, Stephen A. Brewster, Rocio Von Jungenfeld, and Antti Virolainen. Designing interactions with multilevel auditory displays in mobile audio-augmented reality. ACM Transactions on Computer-Human Interaction, 23(1):3:1– 3:??, February 2016. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Vinnikov:2017:GCA

[VAF17] Margarita Vinnikov, Robert S.

Allison, and Suzette Fernandes. Gaze-contingent auditory displays for improved spatial attention in virtual reality. ACM Transactions on Computer-Human Interaction, 24(3):19:1–19:??, July 2017.

CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

vanBerkel:2023:MAP

[vBBSB23] Niels van Berkel, Maura Bellio, Mikael B. Skov, and Ann Blandford. Measurements, algorithms, and presentations of reality: Framing interactions with AI-Enabled decision support. ACM Transactions on Computer-Human Interaction, 30(2):32:1–32:??, April 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571815.

VanVugt:2010:EFS

[VBHK10] Henriette C. Van Vugt, Jeremy N. Bailenson, Johan F. Hoorn, and Elly A. Konijn. Effects of facial similarity on user responses to embodied agents. ACM Transactions on Computer-Human Interaction, 17(2):7:1–7:??, May 2010. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Velt:2020:TBG

[VBR20] Raphael Velt, Steve Benford, and Stuart Reeves. Translations and boundaries in the gap between HCI theory and design practice. ACM Transactions on Computer-Human Interaction, 27(4):29:1–29:28, September 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3386247.

Velloso:2017:MCS

[VCN^{+}17] Eduardo Velloso, Marcus Carter, Joshua Newn, Augusto Esteves, Christopher Clarke, and Hans Gellersen. Motion correlation: Selecting objects by matching their movement. ACM Transactions on Computer-Human Interaction, 24(3):22:1–22:??, July 2017. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

VanderZanden:2001:UMD

[VH01] Bradley T. Vander Zanden and Richard Halterman. Using model dataflow graphs to reduce the storage requirements of constraints. ACM Transactions on Computer-Human Interaction, 8(3):223– 265, September 2001. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Vicente:2000:HGK

[Vic00] Kim J. Vicente. HCI in the global knowledge-based economy: designing to support worker adaptation. ACM Transactions on Computer-Human Interaction, 7(2):263– 280, June 2000. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tochi/
2000-7-2/p263-vicente/.

Vuong:2024:NDB

[VJR24] Tung Vuong, Giulio Jacucci, and Tuukka Ruotsalo. Naturalistic digital behavior predicts cognitive abilities. ACM Transactions on Computer-Human Interaction, 31(3): 36:1–36:??, June 2024. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3660341.

Vertanen:2014:CTE

[VK14] Keith Vertanen and Per Ola Kristensson. Complementing text entry evaluations with a composition task. ACM Transactions on Computer-Human Interaction, 21(2):8:1– 8:??, April 2014. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

VanSchaik:2007:DPR

[VL07] Paul Van Schaik and Jonathan Ling. Design parameters of rating scales for Web sites. ACM Transactions on Computer-Human Interaction, 14(1):4:1–4:??, May 2007. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

VanderZanden:1995:DCB

[VM95] Brad Vander Zanden and Brad A. Myers. Demonstrational and constraintbased techniques for pictorially specifying application objects and behaviors. ACM Transactions on Computer-Human Interaction, 2(4):308– 356, December 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-4/p308-vander_
zanden/.

VanderZanden:1994:IPV

[VMGS94] Brad Vander Zanden, Brad A.

Myers, Dario A. Giuse, and Pedro Szekely. Integrating pointer variables into oneway constraint models. ACM Transactions on Computer-Human Interaction, 1(2):161– 213, June 1994. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-2/p161-vander_
zanden/.

VanSchaik:2015:ACC

[VML15] Paul Van Schaik, Raza Habib Muzahir, and Mike Lockyer. Automated computational cognitive-modeling: Goal-specific analysis for large websites. ACM Transactions on Computer-Human Interaction, 22(3):14:1–14:??, June 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

VanBerkel:2024:CBA

[VP24] Niels Van Berkel and Henning Pohl. Collaborating with bots and automation on Open-StreetMap. ACM Transactions on Computer-Human Interaction, 31(3):38:1–38:??, June 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665326.

Vines:2015:AOP

[VPW^{+}15] John Vines, Gary Pritchard, Peter Wright, Patrick Olivier, and Katie Brittain. An ageold problem: Examining the discourses of ageing in HCI and strategies for future research. ACM Transactions on Computer-Human Interaction, 22(1):2:1–2:??, March 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Vazquez:2014:APF

[VS14] Marynel Vázquez and Aaron Steinfeld. An assisted photography framework to help visually impaired users properly aim a camera. ACM Transactions on Computer-Human Interaction, 21(5):25:1–25:??, November 2014. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

vanSchaik:2012:UEI

[vSHL12] Paul van Schaik, Marc Hassenzahl, and Jonathan Ling. User-experience from an inference perspective. ACM Transactions on Computer-Human Interaction, 19(2):11:1–11:??, July 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Mechelen:2023:ETK

[VSS^{+}23] Maarten Van Mechelen, Rachel Charlotte Smith, Marie-Monique Schaper, Mariana Tamashiro, Karl-Emil Bilstrup, Mille Lunding, Marianne Graves Petersen, and Ole Sejer Iversen. Emerging technologies in K– 12 education: a future HCI research agenda. ACM Transactions on Computer-Human Interaction, 30(3):47:1–47:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569897.

Vidal:2023:IBM

[VSW23] Laia Turmo Vidal, Elena Márquez Segura, and Annika Waern. Intercorporeal biofeedback for movement learning. ACM Transactions on Computer-Human Interaction, 30(3): 43:1–43:??, June 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3582428.

Velez:2004:WCH

[VTS^{+}04] Maria Velez, Marilyn Mantei Tremaine, Aleksandra Sarcevic, Bogdan Dorohonceanu, Allan Krebs, and Ivan Marsic. “who’s in charge here?”: Communicating across unequal computer platforms. ACM Transactions on Computer-Human Interaction, 11(4):407– 444, December 2004. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Vatavu:2022:CAC

[VW22] Radu-Daniel Vatavu and Jacob O. Wobbrock. Clarifying agreement calculations and analysis for end-user elicitation studies. ACM Transactions on Computer-Human Interaction, 29(1):5:1–5:70, February 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3476101.

Venn-Wycherley:2024:REE

[VWKL^{+}24] Megan Venn-Wycherley, Ahmed Kharrufa, Susan Lechelt, Rebecca Nicholson, Kate Howland, Abrar Almjally, Anthony Trory, and Vidya Sarangapani. The realities of evaluating educational technology in school settings. ACM Transactions on Computer-Human Interaction, 31(2):26:1–26:??, April 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3635146.

Wang:2021:LCL

[Wan21] Yi Wang. Living in a city, living a rural life: Understanding second generation Mingongs’ experiences with technologies in China. ACM Transactions on Computer-Human Interaction, 28(3):21:1–21:29, July 2021. CODEN ATCIF4.

ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3448979.

Ware:1994:ROV

[WB94] Colin Ware and Ravin Balakrishnan. Reaching for objects in VR displays: lag and frame rate. ACM Transactions on Computer-Human Interaction, 1(4):331– 356, December 1994. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1994-1-4/p331-ware/

.

Wallace:2022:TIR

[WBD^{+}22] Shaun Wallace, Zoya Bylinskii, Jonathan Dobres, Bernard Kerr, Sam Berlow, Rick Treitman, Nirmal Kumawat, Kathleen Arpin, Dave B. Miller, Jeff Huang, and Ben D. Sawyer. Towards individuated reading experiences: Different fonts increase reading speed for different individuals. ACM Transactions on Computer-Human Interaction, 29(4): 38:1–38:56, August 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3502222.

Wirfs-Brock:2021:ENS

[WBFDK21] Jordan Wirfs-Brock, Alli Fam, Laura Devendorf, and Brian Keegan. Examining narrative sonification: Using firstperson retrospection methods to translate radio production to interaction design. ACM Transactions on Computer-Human Interaction, 28(6):41:1–41:34, December 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3461762.

Wiesen:2023:AHR

[WBW^{+}23] Carina Wiesen, Steffen Becker, René Walendy, Christof Paar, and Nikol Rummel. The anatomy of hardware reverse engineering: an exploration of human factors during problem solving. ACM Transactions on Computer-Human Interaction, 30(4):62:1–62:??, August 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3577198.

Wakkary:2013:SDF

[WDHM13] Ron Wakkary, Audrey Desjardins, Sabrina Hauser, and Leah Maestri. A sustainable design fiction: Green practices. ACM Transactions on Computer-Human Interaction, 20(4):23:1–23:??, September 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wexelblat:1995:ANG

[Wex95]  Alan Wexelblat.  An ap-

proach to natural gesture in virtual environments. ACM Transactions on Computer-Human Interaction, 2(3):179– 200, September 1995. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-3/p179-wexelblat/

.

Wildemuth:1998:HVB

[WFD98] Barbara M. Wildemuth, Charles P.

Friedman, and Stephen M. Downs. Hypertext versus Boolean access to biomedical information: a comparison of effectiveness, efficiency, and user preferences. ACM Transactions on Computer-Human Interaction, 5(2):156– 183, June 1998. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1998-5-2/p156-wildemuth/

.

Whittaker:2001:CVM

[WH01] Steve Whittaker and Julia Hirschberg. The character, value, and management of personal paper archives. ACM Transactions on Computer-Human Interaction, 8(2):150– 170, 2001. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2001-8-2/p150-whittaker/
p150-whittaker.pdf;   http:
//www.acm.org/pubs/citations/

                                        journals/tochi/2001-8-2/
                                        p150-whittaker/.

Williams:2023:MTA

[WIKW23] Alex C. Williams, Shamsi Iqbal, Julia Kiseleva, and Ryen W. White. Managing tasks across the worklife boundary: Opportunities, challenges, and directions. ACM Transactions on Computer-Human Interaction, 30(3):48:1–48:??, June 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3582429.

Whittaker:2004:COC

[WJN^{+}04] Steve Whittaker, Quentin Jones, Bonnie Nardi, Mike Creech, Loren Terveen, Ellen Isaacs, and John Hainsworth. ContactMap: Organizing communication in a social desktop. ACM Transactions on Computer-Human Interaction, 11(4):445–471, December 2004. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wisiecka:2023:SCD

[WKK^{+}23] Katarzyna Wisiecka, Yuumi Konishi, Krzysztof Krejtz, Mahshid Zolfaghari, Birgit Kopainsky, Izabela Krejtz, Hideki Koike, and Morten Fjeld. Supporting complex decision-making: Evidence from an eye tracking study on in-person and remote collaboration. ACM Transactions on Computer-Human Interaction, 30(5):78:1–78:??, October 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3581787.

Ware:1997:SUO

[WL97] Colin Ware and Kathy Lowther.

Selection using a one-eyed cursor in a fish tank VR environment. ACM Transactions on Computer-Human Interaction, 4(4):309–322, December 1997. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-4/p309-ware/

.

Wu:2008:QNM

[WL08] Changxu Wu and Yili Liu.

Queuing network modeling of transcription typing. ACM Transactions on Computer-Human Interaction, 15(1):6:1– 6:??, May 2008. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wiese:2021:IMI

[WL21] Eliane S. Wiese and Marcia C. Linn. “It Must Include Rules”: Middle school students’ computational thinking with computer models in science. ACM Transactions on Computer-Human Interaction, 28(2):10:1–10:41, April 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3415582.

Wingrave:2009:NTE

[WLB09] Chadwick A. Wingrave, Joseph J. Laviola, Jr., and Doug A. Bowman. A natural, tiered and executable UIDL for 3D user interfaces based on Concept-Oriented Design. ACM Transactions on Computer-Human Interaction, 16(4):21:1–21:??, November 2009. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Watson:2015:MUP

[WLB15] Jason Watson, Heather Richter Lipford, and Andrew Besmer. Mapping user preference to privacy default settings. ACM Transactions on Computer-Human Interaction, 22(6):32:1–32:??, December 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wobbrock:2006:AIS

[WM06] Jacob O. Wobbrock and Brad A. Myers. Analyzing the input stream for characterlevel errors in unconstrained text entry evaluations. ACM Transactions on Computer-Human Interaction, 13(4):458– 489, December 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wang:2015:ICC

[WM15] Dakuo Wang and Gloria Mark. Internet censorship in China:

Examining user awareness and attitudes. ACM Transactions on Computer-Human Interaction, 22(6):31:1–31:??, December 2015. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wybrow:2008:CUO

[WMMS08] Michael Wybrow, Kim Marriott, Linda Mciver, and Peter J. Stuckey. Comparing usability of one-way and multiway constraints for diagram editing. ACM Transactions on Computer-Human Interaction, 14(4):19:1–19:??, January 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wan:2016:DGM

[WMRW16] Lin Wan, Claudia Müller, Dave Randall, and Volker Wulf. Design of a GPS monitoring system for dementia care and its challenges in academia-industry project. ACM Transactions on Computer-Human Interaction, 23(5):31:1–31:??, November 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wakkary:2022:TYM

[WOB22] Ron Wakkary, Doenja Oogjes, and Armi Behzad. Two years or more of co-speculation: Polylogues of philosophers, designers, and a tilting bowl. ACM Transactions on Computer-Human Interaction, 29(5):

47:1–47:??, October 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514235.

Wolber:1997:PIB

[Wol97] David Wolber. Pavlov: an interface builder for designing animated interfaces. ACM Transactions on Computer-Human Interaction, 4(4):347– 386, December 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-4/p347-wolber/

.

Ware:1999:RVO

[WR99] Colin Ware and Jeff Rose. Rotating virtual objects with real handles. ACM Transactions on Computer-Human Interaction, 6(2):162–180, June 1999. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
tochi/1999-6-2/p162-ware/

.

Whittaker:1997:TML

[WSKS97] Steve Whittaker, Jerry Swanson, Jakov Kucan, and Candy Sidner. TeleNotes: managing lightweight interactions in the desktop. ACM Transactions on Computer-Human Interaction, 4(2):137– 168, June 1997. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-2/p137-whittaker/

.

Wyche:2016:MPA

[WSO16] Susan Wyche, Nightingale Simiyu, and Martha E. Othieno. Mobile phones as amplifiers of social inequality among rural Kenyan women. ACM Transactions on Computer-Human Interaction, 23(3): 14:1–14:??, July 2016. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Walmsley:2014:DII

[WST14] William S. Walmsley, W. Xavier Snelgrove, and Khai N. Truong. Disambiguation of imprecise input with onedimensional rotational text entry. ACM Transactions on Computer-Human Interaction, 21(1):4:1–4:??, February 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wu:2023:MSS

[WSWL23] Qunfang Wu, Yisi Sang, Dakuo Wang, and Zhicong Lu. Malicious selling strategies in livestream e-commerce: a case study of Alibaba’s Taobao and ByteDance’s TikTok. ACM Transactions on Computer-Human Interaction, 30(3): 35:1–35:??, June 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3577199.

Wiberg:2005:MAS

[WW05] Mikael Wiberg and Steve Whittaker. Managing availability: Supporting lightweight negotiations to handle interruptions. ACM Transactions on Computer-Human Interaction, 12(4):356–387, December 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wang:2022:DMH

[WWD^{+}22] April Yi Wang, Dakuo Wang, Jaimie Drozdal, Michael Muller, Soya Park, Justin D. Weisz, Xuye Liu, Lingfei Wu, and Casey Dugan. Documentation matters: Human-centered AI system to assist data science code documentation in computational notebooks. ACM Transactions on Computer-Human Interaction, 29(2): 17:1–17:33, April 2022. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3489465.

Wu:2019:LDP

[WWH19] Tongshuang Wu, Daniel S. Weld, and Jeffrey Heer. Local decision pitfalls in interactive machine learning: an investigation into feature selection in sentiment analysis. ACM Transactions on Computer-Human Interaction, 26(4):24:1–24:??, July 2019. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3319616.

Watson:1997:MLD

[WWHW97] Benjamin Watson, Neff Walker, Larry F. Hodges, and Aileen Worden. Managing level of detail through peripheral degradation: effects on search performance with a head-mounted display. ACM Transactions on Computer-Human Interaction, 4(4):323– 346, December 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-4/p323-watson/

.

Wright:2008:AEC

[WWM08] Peter Wright, Jayne Wallace, and John McCarthy. Aesthetics and experience-centered design. ACM Transactions on Computer-Human Interaction, 15(4):18:1–18:??, November 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Wiedenbeck:1997:HPL

[WZ97] Susan Wiedenbeck and Patti L.

Zila. Hands-on practice in learning to use software: a comparison of exercise, exploration, and combined formats. ACM Transactions on Computer-Human Interaction, 4(2):169–196, June 1997. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1997-4-2/p169-wiedenbeck/

.

Xia:2022:IDG

[XGA^{+}22] Haijun Xia, Michael Glueck, Michelle Annett, Michael Wang, and Daniel Wigdor. Iteratively designing gesture vocabularies: a survey and analysis of best practices in the HCI literature. ACM Transactions on Computer-Human Interaction, 29(4):37:1–37:54, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503537.

Xambo:2013:LJR

[XHM^{+}13] Anna Xambó, Eva Hornecker, Paul Marshall, Sergi Jordà, Chris Dobbyn, and Robin Laney. Let’s jam the reactable: Peer learning during musical improvisation with a tabletop tangible interface. ACM Transactions on Computer-Human Interaction, 20(6): 36:1–36:??, December 2013. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Xiao:2023:AIH

[XJS23] Sijia Xiao, Shagun Jhaver, and Niloufar Salehi. Addressing interpersonal harm in online gaming communities: The opportunities and challenges for a restorative justice approach. ACM Transactions on Computer-Human Interaction, 30(6):83:1–83:??, December 2023. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603625.

Xu:2012:MEU

[XLC12] Lingling Xu, Julian Lin, and Hock Chuan Chan. The moderating effects of utilitarian and hedonic values on information technology continuance. ACM Transactions on Computer-Human Interaction, 19(2):12:1–12:??, July 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Xu:2023:CWB

[XPL23] Long Xu, Su Jin Park, and Sangwon Lee. Color2Vec: Web-based modeling of wordcolor association with sociocultural contexts. ACM Transactions on Computer-Human Interaction, 30(4): 51:1–51:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571816.

Xie:2022:IDP

[XRL^{+}22] Jingyi Xie, Madison Reddie, Sooyeon Lee, Syed Masum Billah, Zihan Zhou, Chun-Hua Tsai, and John M. Carroll. Iterative design and prototyping of computer vision mediated remote sighted assistance. ACM Transactions on Computer-Human Interaction, 29(4):36:1–36:40, August 2022. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3501298.

Xiao:2020:TMA

[XZL^{+}20] Ziang Xiao, Michelle X. Zhou, Q. Vera Liao, Gloria Mark, Changyan Chi, Wenxi Chen, and Huahai Yang. Tell me about yourself: Using an AIpowered chatbot to conduct conversational surveys with open-ended questions. ACM Transactions on Computer-Human Interaction, 27(3): 15:1–15:37, June 2020. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381804.

Yamamoto:2023:WRW

[YCVV23] Fujiko Robledo Yamamoto, Janghee Cho, Amy Voida, and Stephen Voida. “We are Researchers, but we are also Humans”: Creating a design space for managing graduate student stress. ACM Transactions on Computer-Human Interaction, 30(5): 75:1–75:??, October 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3589956.

Yamada:1995:DEH

[YHS95] Shoji Yamada, Jung-Kook Hong, and Shigeharu Sugita.

Development and evaluation of hypermedia for museum education: validation of metrics. ACM Transactions on Computer-Human Interaction, 2(4):284–307, December 1995. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1995-2-4/p284-yamada/

. See corrigendum [YHS96].

Yamada:1996:CDE

[YHS96] Shoji Yamada, Jung-Kook Hong, and Shigeharu Sugita. Corrigendum [“Development and evaluation of hypermedia for museum education: validation and metrics”, ACM Trans. Human Interact. 2, 4(Dec. 1995) 284–307]. ACM Transactions on Computer-Human Interaction, 3(3):Page 285, September 1996. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tochi/1996-3-3/p285-yamada/

. See [YHS96].

Yannier:2016:APO

[YHWK16] Nesra Yannier, Scott E. Hudson, Eliane Stampfer Wiese, and Kenneth R. Koedinger. Adding physical objects to an interactive game improves learning and enjoyment: Evidence from EarthShake. ACM Transactions on Computer-Human Interaction, 23(4):

26:1–26:??, September 2016. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Yuan:2021:CMH

[YLR21] Haiyue Yuan, Shujun Li, and Patrice Rusconi. CogTool+: Modeling human performance at large scale. ACM Transactions on Computer-Human Interaction, 28(2):15:1–15:38, April 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447534.

Yang:2014:UUC

[YLZ14] Huahai Yang, Yunyao Li, and Michelle X. Zhou. Understand users’ comprehension and preferences for composing information visualizations. ACM Transactions on Computer-Human Interaction, 21(1):6:1– 6:??, February 2014. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Yuill:2012:MCD

[YR12] Nicola Yuill and Yvonne Rogers. Mechanisms for collaboration: a design and evaluation framework for multi-user interfaces. ACM Transactions on Computer-Human Interaction, 19(1):1:1–1:??, March 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Yang:2020:EDE

[YRA20] Mochen Yang, Yuqing Ren, and Gediminas Adomavicius. Engagement by design: an empirical study of the “reactions” feature on Facebook business pages. ACM Transactions on Computer-Human Interaction, 27(6):43:1–43:35, November 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3412844.

Yamanaka:2020:NUD

[YS20] Shota Yamanaka and Wolfgang Stuerzlinger. Necessary and unnecessary distractor avoidance movements affect user behaviors in crossing operations. ACM Transactions on Computer-Human Interaction, 27(6):44:1–44:31, November 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418413.

Yesilada:2007:EDS

[YSHG07] Yeliz Yesilada, Robert Stevens, Simon Harper, and Carole Goble. Evaluating DANTE: Semantic transcoding for visually disabled users. ACM Transactions on Computer-Human Interaction, 14(3): 14:1–14:??, September 2007. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

You:2023:BSD

[YTL^{+}23] Yue You, Chun-Hua Tsai, Yao Li, Fenglong Ma, Christopher Heron, and Xinning Gui. Beyond self-diagnosis: How a chatbot-based symptom checker should respond. ACM Transactions on Computer-Human Interaction, 30(4): 64:1–64:??, August 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3589959.

Zhai:2005:ISB

[ZB05] Shumin Zhai and Victoria Bellotti. Introduction to sensingbased interaction. ACM Transactions on Computer-Human Interaction, 12(1):1–2, March 2005. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zhai:1996:PEU

[ZBM96] Shumin Zhai, William Buxton, and Paul Milgram. The partial-occlusion effect: utilizing semitransparency in 3D human-computer interaction. ACM Transactions on Computer-Human Interaction, 3(3):254–284, September 1996. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tochi/1996-3-3/p254-zhai/

.

Zhang:2024:HPM

[ZCX^{+}24] Hongbo Zhang, Pei Chen, Xuelong Xie, Zhaoqu Jiang, Zihong Zhou, and Lingyun Sun. A hybrid prototype method combining physical models and generative artificial intelligence to support creativity in conceptual design. ACM Transactions on Computer-Human Interaction, 31(5):65:1–65:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3689433.

Zhai:2014:ETT

[Zha14]  Shumin Zhai.  Editorial:

TOCHI turns twenty. ACM Transactions on Computer-Human Interaction, 21(1):1:1– 1:??, February 2014. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zhai:2015:TEC

[Zha15] Shumin Zhai. TOCHI Editorin-Chief transition: Farewell from Shumin Zhai, welcome Ken Hinckley. ACM Transactions on Computer-Human Interaction, 22(6):27:1–27:??, December 2015. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zajac:2023:CFA

[ZLD^{+}23] Hubert D. Zajac, Dana Li, Xiang Dai, Jonathan F. Carlsen, Finn Kensing, and Tariq O. Andersen. Clinician-facing AI in the wild: Taking stock of the sociotechnical challenges and opportunities for HCI. ACM Transactions on Computer-Human Interaction, 30(2): 33:1–33:??, April 2023. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL https://dl.acm.org/
doi/10.1145/3582430.

Zou:2024:EUC

[ZLM^{+}24] Yixin Zou, Khue Le, Peter Mayer, Alessandro Acquisti, Adam J. Aviv, and Florian Schaub. Encouraging users to change breached passwords using the protection motivation theory. ACM Transactions on Computer-Human Interaction, 31(5):63:1–63:??, October 2024. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3689432.

Zhao:2008:DSU

[ZPSL08] Haixia Zhao, Catherine Plaisant, Ben Shneiderman, and Jonathan Lazar. Data sonification for users with visual impairment: a case study with georeferenced data. ACM Transactions on Computer-Human Interaction, 15(1):4:1–4:??, May 2008. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zimmermann:2021:NPM

[ZR21] Verena Zimmermann and Karen Renaud. The nudge puzzle: Matching nudge interventions to cybersecurity decisions. ACM Transactions on Computer-Human Interaction, 28(1):7:1–7:45, February 2021. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429888.

Zambetta:2020:RPW

[ZRT^{+}20] Fabio Zambetta, William Raffe, Marco Tamassia, Florian ’Floyd‘ Mueller, Xiaodong Li, Niels Quinten, Rakesh Patibanda, Daniel Dang, and Jon Satterley. Reducing perceived waiting time in theme park queues via an augmented reality game. ACM Transactions on Computer-Human Interaction, 27(1):3:1–3:30, January 2020. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3361524.

Zaphiris:2006:TSD

[ZS06] Panayiotis Zaphiris and Rifaht Sarwar. Trends, similarities, and differences in the usage of teen and senior public online newsgroups. ACM Transactions on Computer-Human Interaction, 13(3):403– 422, September 2006. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zhou:2015:MDM

[ZSC^{+}15] Jianlong Zhou, Jinjun Sun, Fang Chen, Yang Wang, Ronnie Taib, Ahmad Khawaji, and Zhidong Li. Measurable decision making with GSR and pupillary analysis for intelligent user interface. ACM Transactions on Computer-Human Interaction, 21(6): 33:1–33:??, January 2015. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zancanaro:2012:CNC

[ZSE^{+}12] Massimo Zancanaro, Oliviero Stock, Zvi Eisikovits, Chaya Koren, and Patrice L. Weiss. Co-narrating a conflict: an interactive tabletop to facilitate attitudinal shifts. ACM Transactions on Computer-Human Interaction, 19(3):24:1–24:??, October 2012. CODEN AT-CIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Zoran:2014:HAC

[ZSNP14] Amit Zoran, Roy Shilkrot, Suranga Nanyakkara, and Joseph Paradiso. The hybrid artisans: a case study in smart tools. ACM Transactions on Computer-Human Interaction, 21(3):15:1–15:??, June 2014. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).